A Bibliography of Papers in Lecture Notes in Computer Science (1992)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

22 March 2018 Version 1.21

Title word cross-reference

1 [483]. 1.3n + O(log n) [1885]. 3

[1857, 1722, 1392]. A P_{3T} [618].

ϵn log n + O(1) [1885]. k [1838, 1907, 1839]. P =?NP [1854]. VAM_{1p} [1252]. N [1653, 1305]. n log n [1885].

* [189].

-Cut [1907]. -Cutwidth [1838]. -D [1857, 1722]. -Dimensional [483]. -Edge-Connectivity [1839]. -syntactic [1305].

/ [1644].

1 [1159]. 1-L [1159]. 1990s [1572, 1573].

2 [916, 903, 1380, 87, 1917]. 2- [915]. 2-Connectivity [916]. 2-Spanners [903]. 2-way [1079, 1917]. 2.0 [610].

3 [915, 1856]. 3- [1189]. 3-Connectivity [915]. 3-Sided [1856]. 316 [1505].

8n [1853].

90 [526, 88]. 9000 [1490]. 91 [600]. 9X [382].

= [1367, 879].

A&M [1604]. ABASE [352]. Abductive [112]. Abelian [957]. Abstract [1909, 114, 1832, 166, 996, 1254, 1256, 563, 1251, 748, 284, 875, 1918, 1292, 131, 1903, 1258, 1855, 1191, 1902, 1255, 1945, 550, 1084, 45, 46, 39, 1097]. Abstracting [125]. Abstraction [779, 286, 219, 999, 1477, 1564]. Abstractions [1420, 791]. Academic [1581]. ACC [1878]. Access [1426, 1413, 1160, 836, 1790, 184, 732, 1274, 191, 1358, 1429]. Accuracy [1363]. Accurately [970]. achievements [703]. Acoustical [409]. ACP [1229]. Acquiring [267, 1566, 1127, 1565, 269]. Acquisition [267, 944, 945, 265, 264, 249, 941, 818, 942, 257, 261, 454, 949, 452, 427, 939, 256, 254]. Across [1587]. Action [703, 1626, 522, 1199, 435, 1269, 1227, 1077, 1076, 300, 706, 707]. Actions [843]. Active [1770, 839, 1520, 1519]. Activities [1464]. Activity [836, 949, 462]. Actor [49, 1235, 743]. ACTRESS [1626]. Ada [373, 395, 393, 383, 375, 382, 377, 392, 384, 385, 371, 391, 394, 379, 386, 372, 387, 389, 378, 374, 380]. Ada-9X [382]. ADAPT [666]. Adaptability [513]. Adaptable [729]. ADAPTing [88]. Adaptive [1441, 390, 666, 679, 934, 973, 1383, 367]. Adaptively [1945]. Adder [1410]. Adding [587, 724]. Additive [864]. Address [1569]. adequate [1099]. Adjusting [349, 1845, 323]. Administration [780]. ADTs [1084, 1624]. Advanced [461, 73, 720, 1512, 809, 322, 405, 473]. Advantage [396]. Advantages [692]. Advisory [341]. AFA [1085]. Against [975]. Age [1470, 764]. Agency [1481]. Agent [680, 146, 734]. Agents [859, 1821, 238, 730, 732, 1010, 735, 718]. AI [461, 1513, 819, 824, 813, 810, 801]. Aid [329]. Aide [250]. Aided [346, 680]. Aiding [646]. Aids [328]. aims [703]. Aircraft [440]. Algebra [1364, 1225, 276, 1053, 286, 879, 689, 704, 709, 1096, 1294]. Algebraic [1432, 1052, 33, 1420, 1760, 841, 1272, 1382, 1007, 846, 1278, 1298]. Algebras [277, 111, 564, 293, 841, 1035, 1152]. algèbre [665]. Algorithm [30, 1454, 1714, 1455, 985, 1365, 902, 1440, 1857, 1436, 1401, 530, 1952, 102, 905, 93, 1893, 1356, 1177, 1448, 29, 1360, 1346, 351, 1398, 1219, 989, 1843, 234, 1370, 1349, 1412, 1424, 1727, 971, 31, 1372, 1427, 1404, 100, 1295, 1942]. Algorithmic [113, 764, 1261, 1029, 891]. Algorithmics [1153, 1836]. Algorithms [106, 1394, 1865, 1732, 977, 1447, 1432, 235, 1725, 1943, 1838, 1844, 560, 1653, 310, 1951, 1811, 481, 1442, 1395, 1256, 480, 760, 969, 1458, 535, 897, 906, 1846, 1646, 1842, 1847, 1729, 925, 467, 1723, 980, 1861, 307, 197, 1360, 1433, 1339, 1915, 1918, 107, 1907, 1840, 908, 1189, 1418, 1960, 1366, 1858, 1876, 1964, 1863, 98, 1332, 20, 15]. Alignment [1728]. Allocation [1537, 1633, 93, 1632, 62, 508, 1446]. Almost [1895, 1184]. Alpha [779, 497]. Alpha-Channel [779]. Alphabet [1883]. Alternating [1954]. Alternative [400]. AMAC [1807]. Amalgamation [1825]. Ambiguous [1524, 453]. Amenable [1437]. Among [1866, 1837, 1230]. Amount [853]. AMPATS [394]. AMT [1419]. Analog [1935]. Analogical [429]. analogie [645]. Analogy [1030, 1658, 1662]. Analyses [1265, 1891]. Analysing [407]. Analysis [106, 551, 520, 1714, 672, 448, 73, 1365, 1844, 762, 1403, 995, 167, 1655, 969, 160, 1239, 828, 1692, 1812, 542, 94, 64, 1640, 1815, 29, 816, 1346, 288, 145, 1843, 418, 134, 268, 63, 726, 1245, 1408, 455, 1630, 1902, 570, 1712, 1637, 1966, 1295, 1297, 47, 6, 1939]. Analytic [969, 901]. Analytica [619]. Analyzing [234, 74]. And-Or [510]. AND-Parallel [1237, 614]. And-Parallelism [888]. and/ by [1886]. Andorra [205, 1034]. Angelic [222]. Animation [351]. Answer [840, 1283]. Answering [1061, 649]. Answers [329, 1326, 1900]. antichain [11]. APLUSIX [665]. Application [1892, 1432, 271, 105, 672, 1365, 403, 1056, 937, 1820, 334, 674, 788, 1706, 1841, 406, 404, 811, 371, 1023, 138, 1035, 1382, 1752, 485, 165, 18, 709, 16, 6]. Application-Specific [485]. Applications [786, 1498, 571, 1342, 178, 771, 193, 819, 1779, 773, 1720, 1496, 789, 1621, 1682, 767, 1034, 905, 1730, 1402, 1310, 385, 874, 812, 1344, 1114, 248, 506, 405, 787, 1047, 1881, 801, 1152, 1317, 1946, 444, 1941, 17]. Applicative [306, 1526]. Applied [1432, 421, 1123]. Applying [1489, 354, 433]. Apport [656]. apprenant [678, 681]. apprentissage [681]. Approach [330, 448, 1948, 523, 627, 556, 1633, 68, 204, 171, 160, 1501, 1656, 535, 252, 1745, 927, 451, 250, 521, 1124, 348, 409, 209, 1900, 1660, 263, 474, 877, 130, 1004, 680, 901, 1747, 1602, 211, 1713, 159, 1875, 417, 455, 458, 766, 1361, 1084]. Approaches [808, 1254, 217, 215, 218, 359, 940, 953]. approches [665]. Approximate [1726, 1725, 45, 1723, 1764, 1717, 1894, 1907]. Approximating [3, 1848, 1864, 10]. Approximation [1874, 980, 307, 974, 1908]. APSE [388]. Architectural [793]. Architecture [731, 1415, 478, 537, 643, 646, 1334, 700, 337, 1378, 341, 1812, 155, 770, 514, 1635, 542, 437, 1416, 1344, 496, 508, 1350, 1375, 831, 888, 1743, 1428, 662, 735, 1414, 369, 1541, 1373, 647, 659]. Architecture-Independent [1344]. Architectures [1943, 1362, 820, 1574, 1399, 722, 493, 1522, 747, 769, 88, 494, 1407, 453, 772]. Area [423, 198, 953]. Argument [508]. Argument-Fetching [508]. Argumentation [691]. Arguments [1317]. ARIES [1774]. Arithmetic [383, 1167, 861, 362]. Armour [425]. Arrangements [1892]. Array [1357, 64, 1352, 61, 88, 1370, 1958, 1966, 1427]. Arrays [1158, 510, 1359, 1552, 1444, 1358, 539, 24]. Arrow [1306]. Arrows [1306]. Art [1511, 12]. Article [636]. Artificial [1566, 407, 721, 802, 396, 138, 413]. artificielle [664]. Artists [336]. ASAX [1812]. Asking [1657]. Aspects [33, 34, 826, 817, 793, 457, 841, 1090]. Aspectual [1124]. Assemblies [434]. Assembly [1046, 233]. Assessing [328]. Assessment [1609, 456, 388, 953]. Assignment [1869, 502, 540, 1872, 1881, 1645]. Assignments [759]. Assistant [362]. Assisted [631]. Assisting [401]. Associating [1798]. Association [838]. Associative [1956]. Associativity [518]. Assumptions [844, 1835]. Assurance [1802]. asymptotic [1109]. Asymptotical [318]. Asymptotics [901]. Asynchronous [1369, 1001, 477, 738, 1912, 1003, 1373, 708]. Asynchronously [859]. At-Most-Once [1219]. ATC [377]. ATM [799]. Atomic [1771, 1534, 1772, 706]. Atomicity [1773]. ATP [1056, 294]. Attention [676]. Attribute [1620, 1621, 1619, 1244, 1245, 1273]. Attribute-Directed [1619]. Attributed [9, 1616, 1253]. Attributes [1763]. Audio [786, 1753, 768, 797]. Audio/Video [786, 1753]. Audit [1812]. Augmentation [980, 1839]. Augmented [1845]. Augmenting [166]. Aurora [1539]. Authentication [1803, 1796, 1802, 1804, 1797]. Author [347]. Authoring [352, 357, 359, 361]. Authoring/Learning [359]. Autoepistemic [1815, 1105]. Automata [274, 273, 1218, 1242, 1889, 1168, 1001, 1215, 964, 1175, 39, 1900, 1180, 1958, 881, 1187, 1639, 1890, 1917, 1192, 1220, 1227, 1079, 1107]. Automata-Theoretic [1900, 1890]. Automated [691, 1031, 609, 580, 599, 864, 666, 170, 1792, 1023, 566, 624, 689, 610, 613, 1938]. Automatic [551, 1638, 1513, 1055, 53, 249, 818, 1458, 1015, 148, 1402, 582, 1445]. Automatically [1566, 1829]. Automation [550]. Automatization [399]. Automaton [1234, 1182]. Automobile [1494]. Autonomous [1934, 1801, 718]. Autonomy [1129]. Average [1346, 971, 1896]. Average-Case [1346]. Avoidable [957]. Avoidance [385]. Axiomatic [535]. Axiomatization [758]. Axiomatizing [1229, 1299]. Axioms [549, 1023, 1044].

B [1388, 1787]. B- [1787]. Back [524, 41, 1412]. Back-Propagation [1412]. Background [1647]. Backpropagation [443]. Backtracking [544, 215, 1259]. Backward [289, 1411]. Balanced [868, 195, 67]. Balancing [107, 1382, 1383, 1434, 1384]. Band [1367]. Banded [1360]. Bandwidth [777]. Bandwidth-Constrained [777]. Banishing [867]. Banking [1499]. Barbed [1009]. Barcan [135]. Barriers [1795]. Base [1054, 420, 1018, 1127, 348, 266, 1314, 653]. Based [691, 326, 947, 1457, 1754, 576, 618, 1498, 430, 1342, 731, 632, 1321, 441, 447, 578, 534, 643, 80, 265, 590, 403, 1119, 673, 278, 820, 1056, 1557, 264, 1058, 1262, 1756, 1060, 1501, 1746, 1745, 1322, 1487, 341, 1621, 1708, 1812, 622, 1333, 1632, 94, 154, 942, 722, 755, 1955, 406, 1377, 404, 468, 652, 531, 440, 1774, 1402, 658, 415, 1344, 352, 876, 431, 289, 1452, 357, 469, 418, 463, 471, 822, 1533, 365, 419, 35, 414, 464, 136, 268, 421, 1721, 949, 270, 417, 141, 741, 640, 429, 1963, 1434, 1964, 1417, 610, 402]. Based [597, 1800, 98, 749, 446, 325, 401, 703, 214, 1252, 19, 717, 1304, 1313, 1435, 1251, 888, 376, 797]. Bases [116, 984, 723, 736, 1718, 1017]. Basic [584, 703, 180, 714, 709, 1292]. basic-net [709]. Basis [1065, 1393, 952]. Bayesian [682]. BBN [509, 1364, 79]. BC [1367]. BCK [868]. BCK-Minimal [868]. Be [1867, 313, 1648]. Beam [446]. Beginning [692]. Behavior [940, 1090]. Behavioral [1549, 999]. Behaviour [1466, 803, 125, 408, 417, 1075]. Behaviour-Based [417]. Belief [1813, 815, 1314, 682, 1313]. Benchmark [483, 623, 1355, 53]. Benchmarking [1406]. Bendix [35]. Benefits [364]. Best [1388]. Best-First [1388]. Between [626, 1560, 1883, 1006, 1010, 987, 936, 1925, 1285, 1139, 1144]. Beyond [1546]. Bi [456, 1184]. Bi-Directional [456]. Bias [1647]. Bibliographic [1671, 1675, 1679, 1684, 1689, 1698, 1704, 1709]. Bibliography [1711]. bicentric [1932]. Biharmonic [1437]. Bin [197]. Binary [277, 913, 119, 1851, 5]. Binding [1413, 1238, 510]. BinProlog [1270]. Biology [103]. Bipartite [484]. Bisectorsup [189]. Bisimularity [1208]. Bisimulation [1911, 1207, 1009, 1200, 1077, 1085]. Bisimulations [524]. BLAS [1361]. Block [1447, 1333, 1372]. Boards [448]. Bono [262]. Boolean [303, 196, 1855]. Bootstrapping [660]. Bother [669]. Bottom [1257, 1271]. Bottom-Up [1257, 1271]. Bound [196, 288, 22]. Boundary [929]. Bounded [1897, 1413, 59, 956, 922, 893, 144, 1086, 1087, 17]. Boundedness [1173, 1701, 1703]. Bounds [302, 990, 1024, 1728, 314, 715, 12]. box [704]. Braid [40]. Brains [1543]. Branches [1027]. Branching [1911]. BRAVE [138]. Breadth [187]. Breadth-First [187]. Breaking [1795]. bridgeless [27]. Bringing [764]. Broadband [785, 784]. Broadcast [307, 25]. Broadcasting [1455, 1851, 1450]. Brokers [505]. Browser [162]. Bruijn [1425]. Buchberger [985, 102]. Buchi [1889]. Buffer [187]. Build [1420]. Building [1030, 403, 160, 468, 725, 325]. built [1282]. built-in [1282]. Bulk [897, 527]. Bulk-Synchronous [897]. Burnt [409]. Buses [1852, 1351]. Butterfly [509, 25].

C [1638, 1403, 1529, 1344, 1894, 746, 1530, 1923, 83]. C-Sensitive [1894]. Cabin [440]. Cable [425]. Cache [68, 69, 309, 70, 517, 75]. Cached [1244]. Caches [486, 518]. Caching [567]. CAD [471, 413]. CAD-Environment [471]. CAD-System [413]. cadre [665]. CAI [343, 358]. CAIS [388]. CAIS-A [388]. CAL [347, 366, 369]. CAL-Courseware [369]. CALC [321]. Calculation [642]. Calculations [1364]. Calculi [1910, 555, 135, 885]. Calculus [1022, 1053, 571, 279, 203, 220, 228, 207, 281, 873, 1197, 321, 737, 993, 739, 1029, 884, 1308, 140, 889, 588, 758, 1323, 1277, 704, 1299]. Calibration [436]. Call [220, 1283, 1928, 354, 370]. Call-by-Value [220]. CALLGUIDE [370]. CAMEL [1258]. Camera [436]. Camera-Guided [436]. Can [313, 1867, 1648]. Candela [1422]. Canonical [1049]. Capabilities [687, 410]. Capacities [1078]. Capacitors [329]. CAPP [426]. Capture [828, 971]. Cardinality [1736, 1737]. Carlo [1423]. Cartesian [231]. Cascading [914]. Case [691, 947, 631, 1544, 1904, 942, 406, 1346, 1219, 1885, 1896, 401, 1645, 1098, 175, 155, 177, 176]. Case-Based [691, 947, 942, 406, 401]. Cases [944, 945, 946]. Cat [299, 765]. Catalogue [834]. Catalyst [270]. Categorical [871, 1258, 997, 1765]. Categorically [1000]. Categories [862, 231]. Category [229, 239]. Caught [626]. Causal [265, 652, 704]. Caused [671]. Caylay [1453]. CBR [801]. CC [1178]. CC-Paradigm [1178]. CCITT [796]. CCS [1912]. CCSDS [1429]. CD [793]. CD-I [793]. Cellular [1001, 1182, 1958, 1187]. Center [906, 782]. Centered [1471]. Central [553]. Centred [177, 725]. Ceramic [409]. CERN [92]. CERN-MPPC [92]. Certification [1798]. Certified [1915]. CG [1454]. chain [11]. Chaining [590, 1411, 733]. Challenge [568, 624, 389, 1933]. Challenges [827, 1939]. Change [1835, 1314, 1484]. Changes [1166, 1555]. Changing [1548]. Channel [780, 779, 1792, 531, 1429]. Channel-Based [531]. Channels [511]. Characteristic [98, 1195]. Characteristic-Set-Based [98]. Characterization [992, 859, 1677, 1953, 1175, 1661, 993]. Characterizations [1190, 893]. Characterizing [1859, 1857, 502, 1194]. Check [1240, 1027]. Checker [608]. Checking [1206, 266, 882, 1010, 1246, 1081]. Chemical [701]. CHEMPROF [701]. Chew [1868]. Children [356, 363]. Chinese [581, 1791]. CHIP [1071]. Chips [504]. Choice [531, 1531, 706]. Chromatic [910]. Chronological [1307]. Church [1107]. Circle [931]. Circuit [448, 1855, 955, 1928]. Circuits [1878, 672, 1895, 280, 900, 198, 1389]. Circular [1731, 1962]. Class [1759, 1208, 1857, 1846, 1659, 1155, 10]. Classes [1650, 1652, 992, 1848, 829, 893, 714]. Classical [578, 1014, 756, 1309, 757, 134, 1029, 860, 1103]. Classification [1656, 263, 269, 469, 455, 1135]. Classifying [1742]. Clausal [1036]. Clauses [585]. Cleanroom [1610]. Clinical [1500]. Clock [1349]. Clocks [998]. Close [1401]. Closed [844, 239, 231, 1438]. Closest [927]. Closure [856, 1676, 1079, 1089, 1109]. CLP [1276, 1238, 1300, 1833]. Clues [1561]. Cluster [1331]. Clustering [1739]. CNC [466]. CNET [1343, 1454]. Coarse [732]. Coarse-Grained [732]. COCA [686]. Code [580, 84, 1611, 375, 1631, 160, 1792, 1641, 1387, 378, 376, 51]. Codes [1444]. Coding [1438]. cognitif [681]. COGNITIO [655]. Cognition [655, 1560, 1658, 936, 953]. Cognitive [331, 672, 657, 950, 355, 951, 948, 952, 656, 659]. Cograph [905]. Coherence [70, 75]. Coincidence [1628]. Collaborative [699].

Collection [1515, 1542, 1238, 1529, 1535, 1517, 1534, 1536, 1522, 1540, 1533, 1531, 1520, 1530, 1523, 1519, 1539, 1514, 1541]. Collection-Cooperative [1530]. Collections [1540]. Collector [1524, 1528, 1518, 1538]. Collision [971]. Color [1730]. Coloring [920, 1842, 15]. Colour [355]. Colouring [319]. Combination [1354, 184, 1035]. Combinatorial [1760, 1150]. Combinatorics [761]. Combinators [561, 1097]. Combinatory [556, 220]. Combine [913]. Combined [1745]. Combining [554, 460, 1627, 512, 143]. Commercial [580]. Common [398, 252, 1883, 1729, 722, 101, 1718, 1541]. Common-Sense [398]. Communicating [1869, 859, 532, 1431, 1224, 547, 234, 1006, 1191, 505]. Communication [785, 1334, 480, 779, 788, 796, 738, 60, 911, 531, 519, 428, 494, 423, 1728, 1467, 1003, 781, 704, 708]. Communication-Free [60]. Commutations [963, 1186]. Commutative [1436]. Compact [232, 318]. Companion [698]. Company [1602]. Comparative [528, 953, 1100]. Compare [255]. Comparing [1254, 1635, 1355, 1108]. Comparison [1374, 996, 909, 1360, 356, 1721, 713]. Comparisons [1725, 1181, 1885]. Compatible [1691, 1700, 1695, 1690, 1699]. Competence [354]. Competition [1032]. Competitive [1844, 1843]. Compilation [55, 122, 66, 1385, 1064, 1026, 87, 653, 142, 389]. Compile [73, 1527, 1431, 1640, 1258, 1637]. Compile-Time [1527, 1637, 73]. Compiler [1638, 1626, 1629, 1820, 70, 1252, 1639, 1624]. Compilers [133, 1627]. Compiling [128, 1642]. Complete [1826, 484, 1251, 35, 885, 1308, 21, 5, 1841]. Completely [36]. Completeness [1909, 1631, 281, 867, 1012, 1292, 1106, 1110]. Completion [986, 35, 1049, 558, 1304].

Complex [1783, 420, 646, 450, 1634, 1170, 189, 312, 452, 763, 1737, 165]. Complexity [1862, 1158, 1854, 302, 1161, 990, 1162, 560, 1167, 995, 1860, 911, 872, 196, 1915, 1151, 1855, 1885, 1896, 1105, 955, 1154, 766, 1156, 308, 1456, 1085, 10, 11, 22]. Component [161]. Components [488, 982, 268, 489, 1627]. Composite [163, 1860]. Composition [1763, 1008, 222, 529, 1074]. Compositional [523, 283, 1212, 1228, 1234]. Compositions [1234]. Comprehensive [1517, 1491]. Compressed [1719]. Compression [1727]. Computability [766]. Computation [1235, 1160, 1161, 1769, 1084, 101, 314, 1439, 1033, 1152, 97, 1195, 1368, 1927]. Computational [655, 1168, 182, 206, 184, 1879, 1449, 188, 670, 103, 1273]. Computations [1181, 1397, 134, 1382]. Compute [89, 987]. computed [1283]. Computer [326, 329, 1340, 1053, 1511, 482, 673, 1795, 699, 1, 346, 1596, 991, 354, 809, 473, 1597, 1134, 1594, 1930, 1925, 124, 1919]. Computer-Based [326, 673]. Computers [1374, 1644, 1421, 1410, 1386, 356]. Computing [1766, 827, 487, 902, 1326, 784, 832, 569, 1351, 355, 1731, 624, 772, 741, 1643, 1044, 1935, 1920, 1940]. Concave [1717]. Concept [334, 730, 1647, 1132, 1042, 18]. Concepts [1430, 1459, 1544, 1670, 410]. Conceptual [178, 1742, 179, 169, 1756, 249, 153, 250, 170, 1751, 259, 440, 258, 1564]. Conceptualization [1316]. Concerning [964]. Concise [1449]. Conclog [96]. Conclusion [1655, 1710]. Concrete [235]. Concurrency [237, 1345, 236, 836, 1185, 1186, 746, 52, 749, 1926, 1100, 712]. Concurrency-Related [1345]. Concurrent [534, 1527, 525, 316, 95, 1204, 384, 870, 504, 747, 515, 739, 499, 1010, 1538, 1230, 741, 1298, 1297]. Condensed [422, 566]. Conditional [847, 1146]. Conditions [58, 1706, 1412, 1281]. Conference [1919, 1800]. Configuration [441, 440, 442, 258]. Configurations [1216]. Conflicts [1731]. Confluence [38, 1196]. Confluent [37]. Confrontation [174]. Congruence [1201]. Congruent [638]. Congruity [1865]. Conjecture [1898]. Conjectures [575]. Conjugate [1371]. Conjunctions [872]. Conjunctive [1897, 872]. connaissances [675, 651]. Connected [1784, 1450, 23]. Connecting [751]. Connection [1254, 637, 1332, 1942]. Connectionism [802, 431]. Connectionist [408, 444]. Connections [244, 1925]. Connectives [1124]. Connectivity [1849, 1839, 1156, 915, 916]. Conquer [1458]. Consecutive [1882]. Conservative [1524]. Considerations [782, 387, 143]. Considered [1405]. Consistency [1045, 304, 266, 769, 129, 1782, 1294]. Consistent [1858]. Constant [1850, 1954]. Constants [1035]. Constrained [1866, 777, 585]. Constraint [1016, 174, 837, 400, 1493, 805, 1807, 1324, 432, 449, 214, 1289, 1290]. constraint-based [214]. Constraints [1899, 438, 986, 840, 1558, 170, 1955, 440, 1736, 894, 562, 1737, 1289]. Constructing [317]. Construction [1419, 723, 909, 1015, 232, 1250, 617, 324, 664]. Constructions [660, 758]. Constructive [1031, 762, 761, 1322, 866, 751, 875, 130, 552, 753, 256, 1294, 1108]. Constructivity [760, 1328, 766]. Constructors [594, 1244]. Consumption [120]. Content [967]. Contention [512]. Context [960, 1206, 1911, 1207, 965, 1378, 629, 1673, 1253, 1175, 958, 1149, 1957, 1772, 1662]. Context-Free [1206, 1911, 1207, 965, 1175, 958, 1149]. Context-Freeness [1673]. Context-Sensitive [960]. Context-Sensitivity [1957]. Context-Specific [1772]. Continuation [1270, 1532]. Continuations [498]. Continuing [1601, 1592, 1581]. Continuity [892]. Continuous [235, 862, 238, 1308, 791, 783, 781, 1144]. Continuous-Media [791]. Contour [1964]. Contraction [1314]. Contracts [1537]. Contradict [677]. contributions [1137]. Control [646, 773, 1236, 573, 836, 440, 209, 1416, 746, 427, 141, 1928, 1963, 1107, 834, 445, 416, 1829, 1934, 1936, 7, 1937]. Controlled [1947]. Controlling [1020]. Controls [390, 1790]. Conventional [507, 64, 1116]. Convergence [408]. Convergent [593, 1867]. Conversation [731, 366]. Conversational [342]. conversion [1296]. Convex [930, 1893, 1864, 1189, 190, 1858]. Cookbook [1488]. coopératif [633]. Cooperation [719]. Cooperative [333, 1061, 730, 359, 365, 888, 1530, 270, 725, 735, 729, 734]. Coordinated [1466]. coordinates [1145]. Coordinating [1938, 718]. Coordination [1479, 734]. Coping [534, 291]. Copying [1533]. Coroutines [1525]. Corporate [1607, 1609]. Correct [1629]. Correction [799, 1955]. Correctness [1080, 242, 580, 535, 887, 233]. Cost [1413, 1433, 1728, 1382]. Costs [364]. Count [970]. Counter [1561]. Counter-Intuitive [1561]. Counterexamples [1292]. Counters [1767]. Counting [1162]. Coupled [721, 517, 1232]. Course [1577, 1609, 1056, 1574, 1605, 350, 321]. Courses [1576, 1578]. Courseware [337, 360, 369]. Cover [581, 1805]. CRCW [899]. Create [364]. Creating [1930, 1608]. Creation [1781, 1627]. Credit [1645]. CREW [899]. CRI [971]. CRI-Length [971]. Criteria [1036, 592]. Critical [767, 1887, 787, 295]. Critics [1028]. Crossing [1876, 192, 23]. Cryptographic [1811]. Cryptology [1793]. CSG [1396]. CSP [1222, 1476, 298]. Cube [1450, 23]. Cube-Connected-Cycles [1450]. cubic [27]. Cultivate [1546]. Cumulative [847].

Current [245, 1613, 810]. Curriculum [643, 1584, 349, 1597]. Curry [1030]. Curves [1874]. Cut [1846, 538, 1907, 850, 895, 1102]. Cut-Elimination [895, 1102]. Cuts [983]. cutting [1087]. Cutwidth [1838]. Cycle [557, 154, 500]. Cycles [1450, 26, 27, 23]. Cyclic [1632].

D [1857, 1722, 50, 231, 1392, 143]. D-Objects [1392]. d’acquisition [675]. DAMP [1337]. dans [665, 644, 645, 675, 651, 685, 678]. DAP [1419, 1435, 1355]. DAP-Based [1435]. d’apprentissage [674]. DASH [521]. Data [692, 1651, 1783, 1342, 492, 1753, 84, 235, 857, 536, 837, 913, 720, 1784, 1386, 1799, 1550, 182, 535, 1746, 721, 58, 382, 1322, 1239, 72, 1751, 1565, 70, 538, 816, 1774, 527, 158, 875, 1385, 77, 775, 1772, 1339, 515, 1748, 1387, 1916, 1376, 1424, 1747, 63, 1624, 1743, 1630, 835, 85, 781, 1358, 1429, 165, 1084]. Data-Oriented [382]. Data-Parallel [1385, 1387]. Data/Control [58]. Database [731, 147, 1018, 526, 720, 1471, 1771, 840, 1019, 1745, 721, 522, 1780, 155, 1760, 1790, 184, 185, 199, 491, 130, 463, 1320, 1066, 833, 838, 726, 1752, 1762, 834, 724, 151, 1924]. Databases [1770, 152, 1775, 1785, 397, 806, 1041, 1761, 171, 153, 1558, 1758, 722, 1806, 1900, 829, 1782, 191, 1545, 1757, 1831, 1497, 1765, 1091]. Dataflow [1415, 995, 71, 1172, 1416, 508, 1414]. Datalog [1778, 1769, 1779, 1749, 1173, 139]. DATALOGsup [1173]. DATE [439]. DBC [526, 833]. DBC/1012 [526, 833]. DCE [341]. DDsup [1367]. DDucks [271]. d’E.I.A.O [651]. Deadlock [1441, 1823, 544, 485]. Deadlock-Free [1441]. Deadlocking [1074]. Deal [64]. Debate [1546]. Debugger [78, 1336]. Debuggers [387]. Debugging [1512, 1059, 666, 1261, 139, 1246].

Decidability [1779]. Decidable [1207, 1208, 593]. Deciding [29]. Decimal [383]. Decision [554, 411, 451, 1646, 1693, 1696, 303, 1187, 1890, 724, 1082]. Decision-Making [411]. Declarative [507, 1785, 178, 157, 1250, 126, 1647, 1320, 1263, 139, 1088]. decomposable [4]. Decomposing [929]. Decomposition [904, 1763, 1202, 230, 541, 186, 1941]. Decompositional [474]. Decompositions [930, 1765]. Decoupled [1375]. Decreasing [1891]. Deduction [1031, 609, 1056, 599, 1814, 864, 1816, 1023, 566, 132, 134, 1029, 601, 1301]. Deductions [1012]. Deductive [1775, 1318, 1785, 179, 1041, 1761, 1019, 153, 1013, 1900, 589, 1831, 1497]. Deep [114]. Default [1315, 1133, 1777, 1325, 1105]. default-and [1105]. Defeasible [1317]. Definable [1818, 1109]. Defined [183]. Defining [40]. Definite [1367, 119, 966, 142]. Definition [561, 1137]. Definitional [1287, 209, 850]. Definitions [223, 203, 208, 873, 1181]. Degree [1159, 918, 1908, 1087]. Delay [773]. Delays [911]. Deletion [1176]. Delivery [1219]. DELPHI [459]. DELPHI-EXPERT [459]. DELTA [1337]. DELTA-T [1337]. Demo [1827]. DEMON [703]. Demonstration [539]. Denotational [1088, 221]. Denotations [860]. Densities [1194]. density [1146]. dependability [1921]. Dependable [1920]. Dependence [384, 1359, 63]. Dependencies [291]. Dependency [1764]. Dependent [857, 511, 104, 752, 877, 224, 971]. Depth [1853, 900]. Deques [301]. Derivation [1674, 1255]. Derivations [1672, 1302, 1291]. Deriving [1311]. Describing [1486, 1948]. Description [178, 278, 621, 1557, 606, 748, 1475, 1488, 1422, 1154, 449, 1136, 1141]. Descriptions [1862, 1139]. Descriptive [1550, 269]. Desiderata [1653]. Design [152, 327, 439, 488, 425, 448, 166, 447, 523, 646, 1483, 1811, 264, 334, 773, 1436, 1745, 1584, 782, 696, 155, 346, 467, 468, 350, 77, 1563, 474, 472, 661, 360, 693, 362, 173, 1747, 1308, 268, 432, 399, 168, 1508, 653, 939, 1575, 1548, 1757, 1417, 402, 749, 401, 376, 703, 1930]. Designed [371, 1102]. Designing [381, 336, 484, 648, 1359, 1908]. Designs [1745, 1333, 496, 1739]. Desk [727]. Detachment [566]. Detecting [690]. Detection [1345, 1342, 1906, 317, 1356, 385, 16]. Detectors [459]. Determinate [318]. Determination [1125]. Determining [1865, 925]. Determinism [1176]. Deterministic [1769, 959]. Develop [1621]. Developed [1275]. Developers [951]. Developing [1054, 353, 270, 389]. Development [947, 373, 441, 1629, 1609, 390, 1466, 530, 345, 1475, 755, 246, 1344, 1132, 149, 268, 1625, 1467, 1007, 1210, 735, 1800, 51]. Developments [1854, 1918, 365]. deviations [1936]. Devices [1361]. d’exercices [665]. DFA’s [1719]. Diagnosis [459, 430, 672, 657, 466, 679, 822, 464, 470, 465, 689, 429, 458]. Diagnostic [408, 423, 681]. Diagnostics [455]. Diagram [154]. Diagrams [932, 692, 1744, 13]. Dialect [138]. Dialogue [825, 1505, 684, 633]. Diameter [904]. DIANA [376]. DIANA-Based [376]. Dictionaries [1787]. Dictionary [1732, 1157, 1391, 1733, 1884, 1132, 1747]. Difference [572]. Differences [1880, 252]. Differencial [1353]. Different [389]. Differential [1432, 823, 104, 1937]. Differential-Algebraic [1432]. Differential-geometric [1937]. Differentiating [253]. Digit [1410, 1404]. Digital [1566, 798, 998, 768, 794, 472, 1933, 795, 797, 1935]. DIGITEF [644]. Digraph [919, 18]. Dilemmas [1500]. Dimensional [932, 1951, 913, 86, 483, 923, 1215, 1954, 1189, 1193, 1966, 928]. Dimensions [1865].

Direct [309, 897, 1636, 41]. Directed [1626, 1536, 1618, 1619]. Direction [56]. Directional [456]. Directive [254, 674]. Discipline [1570]. Discourse [169, 342]. Discovery [637]. Discrete [1225, 631]. discretized [1144]. Discussing [1491]. Discussion [1580, 265, 1588, 1586, 1600]. Disjoint [554, 26, 27]. Disjunctive [1039, 1775, 1041, 1761, 1326, 1328, 211, 976]. Display [1731]. Displaying [1396]. Disproving [575]. Dissection [1362]. Dissemination [480]. Distance [1395, 56, 1587, 1721]. Distributed [1515, 49, 1457, 395, 827, 698, 80, 1362, 1208, 771, 278, 819, 1799, 194, 339, 1420, 1745, 1906, 95, 1331, 748, 102, 317, 1517, 538, 1790, 531, 1804, 732, 1346, 775, 396, 371, 1348, 88, 187, 1381, 494, 1518, 1407, 1520, 499, 495, 1347, 1408, 1384, 735, 1519, 1800, 83, 712]. Distributing [533]. Distribution [393, 978, 121, 1745, 782, 1797, 719, 190, 1145, 1800]. Distributions [1861]. divergence [1285]. Divide [1458]. Divide-and-Conquer [1458]. Divisors [101]. DNA [925, 1962, 1664]. Do [1471, 1544, 323]. Document [1809]. Documentation [148, 270]. Documents [795]. Doing [346]. Domain [226, 1644, 232, 129, 261, 452, 1941]. Domain-Driven [261]. Domains [334, 1495, 231, 230, 1801, 976, 165]. Douglas [1602]. Down [1242, 1619]. Draw [919, 336]. Drawing [1952, 414]. Drawings [1857]. Driven [634, 538, 1113, 261, 256]. Dryer [424]. DSDL [834]. Dual [1899]. Duality [1718, 1214]. Dumb [695]. d’un [656, 678, 659]. Duration [281]. During [149, 1531]. Dynamic [744, 1516, 823, 525, 1319, 20, 502, 915, 927, 522, 93, 42, 916, 1733, 1260, 1425, 1531, 831, 1355, 891, 1451, 1384, 1964, 369, 1446, 711]. Dynamicity [784]. Dynamics [1476, 1448]. dynamique [633]. Dynamo [1384]. Dyslexic [363]. DZ [1727].

E-Narrowing [592]. E-Unification [556]. E3 [1472]. Easy [994]. Eco [940]. Eco-Pragmatic [940]. Ecological [1532]. Economic [1547]. Economics [444]. EDDI [678]. Edge [17, 1839]. Edit [1721]. Editor [1457, 176]. Edo [1548]. Edo/Tokyo [1548]. EDS [473]. Educating [1590]. Education [1601, 1609, 338, 1572, 1573, 764, 1592, 344, 1596, 1581, 1579]. Educational [330, 350, 365, 324, 697, 367]. EER [1749]. Effect [1128, 710]. Effective [1409, 892]. Effectiveness [69, 364]. Effects [355, 1663, 971]. Efficiency [511, 1379]. Efficient [1732, 1357, 910, 536, 608, 481, 1440, 1857, 923, 1952, 1692, 467, 475, 1346, 474, 1182, 1525, 189, 1397, 1324, 198, 100, 1942]. Efficiently [1648]. Ehrenfeucht [856]. Ehrenfeucht-Fraisse [856]. Eigensolution [1427]. Eigenvalue [1365]. Eigenvalues [1368]. EKS [1070]. EKS-VI [1070]. Elaborating [163]. Elaboration [385]. Elaboration-Time [385]. Elastic [107]. Electrical [1432, 672]. Electron [446]. Electronic [1810, 1602, 539]. Electronics [329]. Element [425]. Elementary [1024, 648, 901, 9, 1279]. Elevation [1562]. eleve [656]. Elf [210]. Elicitation [262]. Eliciting [171]. Elimination [567, 422, 1163, 1820, 850, 895, 3, 1102]. ElipSys. [1058]. Elliptic [1372]. else [241]. Embedded [1768, 1019, 379]. Embedding [476, 579, 1743, 1332, 5]. Embeddings [1425, 1230]. EMC [448]. Empirical [1725]. Empirically [949]. Employing [840, 1331]. Emptiness [1179, 1107]. Empty [596]. Enactment [1469]. Encoding [1559]. encryption [19]. End [774, 728, 463, 176]. End-to-End [774]. Energy [459, 1955, 92]. Enforcement [495, 1801]. Engagement [324]. Engine [727, 1270, 1532]. Engineering [1615, 1511, 626, 1577, 1609, 1599, 1462, 334, 1595, 1500, 247, 1572, 1573, 1606, 1476, 1571, 1596, 1570, 1614, 545, 1605, 1585, 1610, 950, 1604, 1581, 1587, 1583, 817, 410, 260, 365, 419, 1578, 149, 1608, 940, 1582, 1598, 159, 936, 156, 948, 1593, 1597, 374, 953, 952, 1594]. Engineers [1584, 1589]. Enhanced [126]. Enhancing [69, 687, 410]. Enrichments [1045]. Entailment [578]. Enterprise [398]. Entities [1756]. Entity [1745, 1740, 126, 1750, 1739]. Entity-Relationship [1740, 1750]. Entity-Relationship-Based [1745]. Entrenchment [1314]. Enumeration [983]. Enumerative [197]. Envelope [1873]. Envelopes [1383]. Environment [1343, 1454, 439, 632, 147, 1057, 1059, 634, 789, 1621, 1250, 346, 1472, 538, 1749, 351, 1344, 251, 471, 680, 177, 546, 838, 670, 139, 640, 641, 1451, 445, 1643, 51, 1456]. Environmental [1560, 1551, 366, 1568]. Environments [326, 628, 390, 1471, 309, 95, 730, 775, 322, 177, 1627]. environnement [674]. Epistemic [1314, 896]. Epistemology [637]. Equal [1433, 1963]. Equality [607, 587, 962, 623, 889, 271]. Equation [1437]. Equational [1265, 1045, 554, 1163, 1267, 1097, 35, 1035, 894, 1305, 1278]. Equations [823, 1943, 1431, 104, 1353, 39, 29, 28, 241, 1424, 31, 595]. Equipment [796]. Equivalence [223, 1207, 1779, 1199, 1212, 1299, 1099, 717]. Equivalences [1198, 1213, 1010]. Era [324]. Erasing [1175]. EREW [899, 979]. Ergonomic [447]. Error [459, 799, 657, 1955, 1728]. Errors [1345, 671]. Esprit [703]. Essential [1405]. ESTEAM [1505]. ESTEAM-316 [1505]. Estimate [989]. Estimating [1409, 69]. Estimation [1456]. Estimations [471]. Eta [1296]. Eta-conversion [1296]. Ethernet [1223]. étudiant [675, 659]. EUODHILOS [1069]. Europe [247]. European [1481]. Evaluating [689, 1102]. Evaluation [1638, 1241, 167, 1778, 1256, 1485, 1259, 317, 518, 517, 658, 491, 362, 1244, 1424, 1713, 595, 1271, 378, 1142, 1139, 1404, 1737]. Evaluation-Interpolation [1424]. evaluations [4]. Even [1963]. Event [524, 239]. Events [304, 1188]. Every [1184]. Everything [1126]. Eves [621]. Evidence [1559, 1124]. Evolution [1759, 1483, 89, 1752]. Evolving [1486, 1472]. Exact [1878]. Exact-Threshold [1878]. Example [332, 178, 1510, 1484]. Examples [691, 1814, 535, 946, 1664, 763, 1317]. Exchange [1169, 1606, 1335]. Exclusion [1346]. Executable [56]. Executing [790, 1043, 292]. Execution [49, 1342, 48, 124, 1336]. Executives [380]. Exegesis [526]. Existence [100]. Exists [1471]. Expanding [1327]. Expansion [1910, 1150]. Exper [240]. Experience [1607, 53, 1574, 1488, 1518]. Experiences [1611, 1603]. Experiment [1824, 590, 333]. Experimental [1256, 641, 1923]. Experimentation [332]. Experimenting [1454, 1472]. Experiments [1649, 265, 566]. Expert [947, 425, 459, 448, 1499, 264, 400, 696, 466, 263, 431, 940, 1502, 951, 943, 258, 1411, 458, 694, 416, 401, 459]. Expertise [648, 654, 697, 685]. Experts [454]. EXPITS [641]. Explainable [264]. Explanation [635, 429, 694]. Explanation-Based [429]. Explanations [669]. Explicit [492, 1325, 140, 74, 85, 1296]. Exploit [86]. Exploiting [527, 402]. Exploration [694]. Exploring [368]. Exponent [1023]. Exponential [1895, 1184]. Expressed [1755]. Expressibility [886]. Expressing [163, 1558]. Expression [492, 1717, 1404]. Expressions [1719, 963, 1659, 740]. Expressive [890, 955, 1079]. Extended [1909, 201, 655, 207, 283, 225, 1292, 1541, 39]. Extending [390, 1019, 64, 432]. Extensibility [489]. Extensible [1253, 185]. Extension [1825, 1554, 1258, 35, 499].

Extensions [484, 1065, 506, 1777, 834]. External [1066]. Extrapolation [104]. Extremely [1867]. Eye [436]. Eye-In-Hand [436].

F [1387, 143]. F-Code [1387]. Facilities [1822]. Factoring [311]. Factorization [1367, 312]. Factory [1495]. Facts [1561, 939]. Failure [1039, 1733, 579, 1027, 604]. Failure-Resistant [604]. Failures [1605]. Fair [1222]. Faithful [1781]. falsity [1087]. Families [183]. Family [1627]. Fanout [484]. Fantasies [939]. Farm [1368]. Farming [1337]. Fashioned [272]. Fast [1726, 979, 1715, 1723, 15, 1252, 1907, 1840, 981, 1295]. Fault [1799, 484, 1851, 466, 464, 1518, 465, 1444, 18]. Fault-Tolerant [484, 1851, 1518, 1444]. Faults [1329, 406]. Faulty [303, 1450]. FAUST [620]. FAUST-Prover [620]. Fawlty [636]. FCG [1641]. FDDI [798, 797]. FDDI-Based [797]. Feasibility [857]. Feasible [1155]. Feature [1421, 658, 415, 141]. Feature-Based [415]. Federation [1789]. Federative [730]. Fetching [508]. Few [930]. FFT [483]. Fibre [425]. FIDO [129]. Field [648]. Fields [1546]. fifo [708]. fifo-net [708]. Figure [1708]. Figure-Generating [1708]. figures [664]. File [1526]. Filter [1694]. Filtering [217, 216, 1353]. final [1294]. Finalization [1528]. Finding [311, 1777, 1876]. Fine [62, 1923]. Fine-Grain [62, 1923]. Finger [1886]. Finitary [203]. Finite [533, 857, 1168, 1495, 964, 39, 1180, 129, 1661, 1639, 1010, 1917, 891, 1323, 1654]. finiteness [1302]. First [1386, 577, 1388, 1544, 187, 885, 136, 140, 889, 955, 305, 1282, 1303, 1083, 1099, 21]. First-Order [577, 1544, 955, 1282, 1303]. Fixed [1899, 1889, 1678, 891, 45]. Fixed-Point [1678]. Fixpoint [1283, 1308].

Fixpoints [1201]. Flang [128, 1642]. Flat [535, 95]. Flexible [155, 440, 185, 364]. Floating [1410, 1439]. Floating-Point [1410, 1439]. Flow [692, 72, 64, 1918, 63, 1630]. Flows [310, 917]. Fluid [1355]. fo [514]. Focal [953]. Focus [676]. Fold [1249]. Fold-Unfold [1249]. Folding [1247]. FOLON [1250]. fondé [633]. FOOPS [740]. Forecasting [444]. Forest [1394]. Foreword [935]. Form [1825, 1779, 577, 872, 1360, 668, 1295, 965]. Formal [742, 1803, 1925, 278, 163, 1510, 338, 1584, 255, 1503, 1792, 1794, 547, 751, 1583, 1582, 1602, 1474, 831, 1007, 1136, 613, 1223, 51]. Formalisation [259]. Formalism [1477]. former [583]. Forms [1655, 1098]. Formula [135]. Formulae [1163, 577, 880]. Formulas [868, 955, 849, 1083]. Formulas-as-types [849]. Forth [524]. Fortran [50, 88]. Fortune [1542]. Forward [799, 590, 1573, 289, 733]. Forward-Chaining [733]. Found [578]. Foundation [73, 136, 1613]. Foundations [1785, 874, 318, 1007]. Founded [1776, 1325, 1012, 1190, 1301]. Four [53]. Fractional [914]. fragment [1099]. Fragmentary [1834]. Fragmented [1799]. Fragments [1014]. Fraisse [856]. Frame [1013, 424]. Framework [1051, 1803, 115, 175, 683, 1832, 1459, 1119, 788, 1315, 255, 657, 153, 1632, 1596, 158, 1307, 472, 442, 149, 739, 934, 688, 150, 595, 529, 610, 729, 733]. Frameworks [940, 939]. FRAPPS [610]. Free [1441, 587, 1016, 1206, 1911, 1207, 544, 965, 1436, 60, 1175, 958, 594, 29, 1149, 28, 134, 485]. Freedom [372]. Freeness [1673, 1902]. Frequency [1353, 668]. Frequency-Filtering [1353]. FRESCO [730]. Freshness [1802]. FRI [604]. FRIENDLY [1064]. FRIENDLY-WAM [1064]. Fringe [913]. Front [728, 463, 176]. Front-End [728].

Front-End-To-Database [463]. frontiers [1924]. Full [1909, 510, 1375, 888]. Fully [147, 1410, 916, 1263, 1191, 1404]. Function [871, 1146, 1142]. Functional [201, 1018, 503, 1236, 1641, 1264, 498, 1258, 1261, 419, 1424, 1266, 835, 1271, 705, 1290, 1088]. Functional/Logic [1018, 1258]. Functionalities [402]. Functionality [649]. Functionals [224]. Functions [1650, 235, 302, 972, 511, 865, 1898, 1700, 1699, 815, 1733, 196, 1062, 639, 1840, 1244, 1272, 305, 24, 1091]. Functorial [1185]. fundamental [1925]. Fundamentals [1932, 1736]. Funding [1612, 1613]. Funnel [1859]. Funnel-Shaped [1859]. FUNSOFT [1485, 1487]. Fusion [469, 470]. Future [1043, 1890, 804]. Fuzzy [874, 431, 433].

Gaining [396]. Galactic [879]. Galerkin [1941]. gallery [12]. Gallina [1904]. Galois [1254]. Game [194, 1870, 653]. Games [1909, 856, 1220]. Gap [1717]. Garbage [1515, 1542, 1524, 1238, 1529, 1517, 1534, 1536, 1522, 1540, 1531, 1518, 1520, 1538, 1530, 1523, 1519, 1539, 1514, 1541]. Gate [1878]. GAZER [611]. gcd [1393]. GCLA [200, 209]. GDoM [1809]. General [1040, 1262, 1233, 1665, 1691, 1700, 1914, 1128, 966, 1069, 1743, 1290]. Generalised [254]. Generality [626]. Generalization [1318, 1653, 31]. Generalization-to- [1653]. Generalized [1039, 1365, 310, 480, 1281, 1872, 1335, 132, 955, 1712, 19]. Generalizing [1313]. Generate [1871]. Generated [1049, 1829]. générateur [651]. Generating [944, 536, 839, 1617, 1708, 1685, 1686, 1687, 1688, 903, 1867, 472, 258, 1429]. Generation [691, 1638, 1810, 84, 820, 169, 249, 840, 622, 579, 148, 1958, 719, 1423, 783, 369, 665]. Generational [1538, 1523]. Generative [1229]. Generator [1626, 1641]. Generators [375]. Generic [267, 175, 1057, 1256, 357, 1840, 887, 8].

Genericity [1782]. générique [678]. Genetic [1646, 197, 667]. Genome [1721]. Gentle [1497]. Gentzen [135, 895]. Gentzen-Style [895]. Gentzen-type [135]. genus [17]. GeoBlock [345]. Geographic [1558, 184, 1561, 1568]. Geometric [181, 537, 924, 1866, 189, 1662, 191, 449, 1937]. Geometrical [1865]. Geometrically [1857]. géometriques [645, 664]. Geometry [182, 1, 345, 184, 1551, 188, 603, 552]. Geospatial [1567]. GFLOPS [542]. GIS [1546, 1563]. GKD [913]. GKD-Trees [913]. Global [1520, 1875]. goal [1304]. Godel [1824]. Going [630, 1929]. Good [998, 1211]. Grace [702]. Gradated [1956]. Graded [896]. Gradient [1371]. Gradual [1555]. Graduate [344]. Grain [66, 62, 1923]. Grained [732]. Grammar [1018, 559]. Grammars [1947, 1620, 965, 1621, 1708, 1685, 1668, 1705, 1706, 1707, 1687, 1957, 1245, 1246, 1959, 1273, 1966, 9]. Grammatical [1724]. Granularity [629]. Graph [1947, 904, 1148, 965, 760, 65, 477, 1692, 1685, 1686, 1693, 1695, 1690, 1699, 1687, 1688, 1696, 1842, 183, 1377, 980, 540, 1359, 6, 1538, 1523, 1849, 1156, 100, 8, 18, 9, 16, 7]. Graph-Generating [1685, 1686]. Graph-Theoretic [1693, 1695, 1690, 1699, 1696]. graph-transformations [8]. Graphic [1621]. Graphical [1457, 1755, 608, 837, 347]. Graphics [327, 1]. Graphs [1859, 481, 1334, 922, 920, 153, 915, 58, 304, 1453, 1632, 916, 183, 1841, 925, 918, 1848, 982, 1181, 294, 1425, 988, 1840, 1876, 1839, 449, 27, 4, 20, 15, 19, 21, 17]. Grasping [417, 436]. Greater [344]. Greatest [101]. Greedy [1453]. Greibach [965]. Grid [1354, 1402, 1864, 1344, 1856]. Grid-Based [1402]. Grids [1344, 465, 5]. Gröbner [984]. Ground [252, 127, 1295]. Groundness [1902]. Group [28, 40]. Groups [1436, 29, 1023]. Growing [960]. Grows [1945]. GTS [67]. Guaranteed [1728]. Guided [637, 661, 436, 601]. Guidon [627]. Guidon-Manage [627].

Hamiltonian [1169]. Hand [436]. Handler [1823]. Handling [913, 1550, 671]. Hard [381, 1867, 845, 380, 1880]. Hardware [1943, 1811, 521, 280, 1331, 1067, 1428]. Harrop [880]. Hash [972]. Hashing [193]. HDTV [1933]. Heap [909]. Heaviest [1716]. Heidelberg [770]. height [3]. HeiTS [770]. Help [727, 661, 662]. Help-Desk [727]. Herbrand [1825, 883]. Hereditary [880]. Herky [605]. Heterogeneous [731, 838]. Heuristic [121, 93, 648, 1663]. Heuristics [445, 1140]. hierarchic [1303]. Hierarchical [1557, 1559, 65, 1828, 1632, 1438, 52]. Hierarchically [183]. Hierarchies [81, 968, 1104]. Hierarchy [1895, 1954, 440, 1781, 849]. High [625, 459, 1374, 78, 839, 193, 1059, 1278, 605, 770, 1376, 1622, 968, 92, 1643, 781]. High-Level [78, 839, 1622, 1643]. High-level-replacement [1278]. High-Speed [770, 781]. Higher [932, 1040, 852, 1869, 556, 204, 338, 1215, 1257, 755, 1191, 1272, 1193, 305, 233, 1277]. Higher-Order [1040, 556, 204, 1257, 1272, 233]. Highley [392]. Highly [478, 93]. highway [1938]. Hilbert [1718]. Hoare [283, 871]. HOL [615]. Homology [1215]. Homomorphic [225]. Homotopy [1365]. Honest [1863]. Honours [1585]. Horizons [1927]. Horn [1040, 844, 211]. Howard [1030]. HRG’s [1686, 1669, 1688, 1680]. HRL [1683]. HRL’s [1694, 1681]. HSP [229]. Hull [1189]. Hulls [549]. humaine [685]. Human [699, 677, 943, 725, 1930]. Human-Centred [725]. Human-Computer [699, 1930]. Hybrid [1415, 1773, 451, 290, 294, 1209]. Hydrodynamic [1365]. Hyper [607, 568].

Hyper-Linking [607]. Hyper-Resolution [568]. Hypercard [333, 364]. Hypercube [1380, 515, 453, 23]. Hypercubes [1330]. Hypercycles [544]. Hyperedge [1708, 1668, 1707]. Hyperedge-Replacement [1668, 1707]. Hyperedges [1666, 1667]. Hypergraph [1705, 1706, 1452]. Hypergraph-Based [1452]. Hypergraph-Replacement [1705, 1706]. Hypergraphs [1666, 1667]. Hypermedia [330, 352, 320, 367]. Hypermedia-Based [352]. HyperMIR [340]. Hypernode [830]. Hyperplane [60]. Hypertext [370]. Hypothesis [1663]. Hypothetical [1955].

I-Categories [862]. I/O [1266, 1220]. i860 [1361]. IBM [1610]. ICAI [358]. ICL [615]. Iconic [634]. Id [55]. idea [707]. Ideas [368]. Identifiable [1650]. Identification [1792, 269, 1661, 1622, 1782]. Identifying [1720, 1664]. Identity [1800]. Identity-Based [1800]. IDRIS [173]. if [241]. if-then-else [241]. Ignorance [1307]. II [209, 200, 110, 111, 349, 845, 1531]. II. [1084]. II/III [1531]. III [1275, 1363]. IKEA [725]. ILISCE [445]. Image [482, 1435, 1438, 1449, 418, 107, 187, 1397, 399, 1960, 1964, 1146, 1141]. Images [1959]. Imaging [1350]. Immune [1184]. Impact [1479, 778, 62]. Impacts [801]. Imperative [536]. Implement [251, 1083]. Implementation [30, 520, 1430, 1052, 111, 105, 1629, 1483, 503, 1435, 1537, 1421, 119, 1298, 95, 280, 1526, 770, 1243, 547, 1749, 1237, 1525, 1132, 291, 1269, 360, 133, 831, 1624, 1793, 1643, 1361, 1942]. Implementations [1447, 534, 1427]. Implemented [371, 614]. Implementing [762, 204, 803, 280, 207, 770, 538, 1224, 589, 1622, 1272, 1273]. Implicates [569]. Implication [1310, 1118]. Implications [1019, 802]. Implicit [1892, 1884]. Important [174, 320]. Impossibility [550].

Impressionistic [336]. Improve [1386, 1536]. Improved [587, 1390, 1448]. Improvement [1601]. Improvements [924, 31]. Improving [511, 1891, 696, 1236, 1640]. IMPS [606]. In-Place [1885]. Inaccurate [1651]. Inaugural [1919]. Including [801]. Incomplete [586, 1330]. Inconsistency [1265, 677]. Inconsistent [1320]. Increase [198]. Increasing [1716]. Increasing/ Common [1716]. Incremental [1265, 1241, 1161, 1778, 1769, 1291, 1535, 1534, 1522, 882, 1533, 1244, 981]. Incrementally [569]. Indefinite [1775]. Independent [925, 1237, 1344, 14]. Index [185]. Indexing [1251]. indicator [24]. Indirect [61]. Indistinguishabilities [1879]. Individual [1606, 1614, 705]. Induction [581, 123, 256, 1044, 1567, 604]. Inductive [808, 203, 1060, 1656, 898, 208, 1028, 1647, 1663, 412]. Industrial [441, 526, 700, 1034]. Industry [1577, 1576, 1581, 1494]. Industry-Oriented [1576, 1577]. Inequalities [312]. Inference [1120, 1656, 898, 1764, 1311, 1275, 1092]. Inferences [429]. Inferred [1268]. Inferring [1183]. Infinitary [1766, 991, 1109]. Infinite [1016, 857, 1208, 1001, 1027, 1047, 1049, 716, 1098]. infinite-valued [1098]. Infinity [931]. Influence [1614]. Informal [1602]. Informant [1660]. Information [967, 1775, 226, 1127, 179, 1784, 480, 899, 1771, 1559, 1553, 154, 148, 1563, 1877, 938, 469, 405, 149, 173, 1568, 719, 845, 1214, 168, 1743, 473, 725, 172, 729, 1138, 733, 734, 1648, 1144]. Informed [1870]. Inheritance [223, 1835, 1275]. Initial [1472]. Inner [1351]. Innermost [1037]. Input [1861, 570]. Inputs [358]. Insertion [1176]. Inspection [1611]. Inspired [1639]. Instruction [1634, 679, 1582, 63, 650]. Instructional [461, 327, 328, 676, 687, 654].

Instructions [1636]. Integer [979]. Integrate [727]. Integrated [147, 1633, 337, 1745, 1621, 792, 405, 1743]. Integrating [377, 836, 350, 1336, 1578, 146, 772, 1474, 1597, 57, 1594]. Integration [439, 731, 174, 406, 793, 1578, 838, 168, 1741, 835, 389, 172]. Integrity [116, 840, 1558, 1738]. Intelligence [1566, 721, 802, 396, 719, 138, 413, 664]. Intelligence-Database [721]. Intelligent [625, 695, 683, 643, 646, 397, 825, 390, 339, 340, 696, 652, 415, 630, 353, 686, 687, 667, 362, 434, 270, 427, 725, 462, 662, 416, 671, 1946, 734, 642, 633, 659]. intelligents [664]. Intensional [1814, 840]. Intensionally [865]. Intensive [1114]. Intention [634]. Intention-Driven [634]. Interaction [1754, 1164, 1125, 1259, 51, 1930]. Interactions [277]. Interactive [152, 1052, 1611, 340, 1064, 173, 1266, 256]. Interargument [1833]. Interconnection [1394, 1333, 476, 1452, 1375]. Interconnects [539]. Interface [608, 1203, 1528, 1563, 548, 462, 1593, 388]. Interfaces [1504, 1006, 797]. Interfacing [1264]. Interfering [917]. Interleaving [858, 706]. Intermediate [482, 261]. International [372]. Internetworks [773]. Interoperability [397]. Interpolation [974, 1424]. Interpret [1744]. Interpretation [114, 1832, 1254, 1256, 871, 1903, 1029, 414, 1902, 1255, 45]. Interpretations [560, 996, 255, 752, 883, 46, 1279]. Interpretative [941]. Interpreter [1276, 1562, 1257]. Interpreters [133, 888]. Interpreting [462]. Interprocedural [1260, 1628, 44, 1630]. Intersection [1893, 989]. Intersections [1179]. Interval [576, 481, 304, 1632, 925, 876, 1890, 1076, 24]. Interval-Based [576, 876]. Intervals [319]. Introduction [984, 478, 824, 1025, 1468, 212, 1665, 814, 1816, 805, 807, 719, 1112, 1473, 1463, 821, 801, 1497, 1072, 1482]. introductory [214]. Intrusion [1799]. Intuitionistic [1321, 588, 849]. Intuitions [1828, 751]. Intuitive [1561]. Invariance [854, 1347]. Invariant [1956]. Invariants [206]. Inverse [597]. Inversion [1370, 1444]. Invertibility [1877]. Investigations [28]. Invited [1225, 1254, 897, 923, 1236, 1219, 1197, 1209, 1214, 900, 926, 1277, 1300, 1288]. involving [21]. IPSC [1360, 1402, 1380]. iPSC/2 [1380]. IPSC/860 [1360, 1402]. Isabelle [600]. Isabelle-91 [600]. ISAR [1052]. ISO [1490]. Isomorphism [1030, 1898]. Issues [487, 760, 1589, 784, 954, 749]. Iterated [1167]. Iteration [242, 1369, 68, 1863]. Iterations [1891, 18]. Iterative [1158, 1418]. ITS-Shell [369]. ITSs [700].

Jacobi [1447, 1369, 1427]. Jade [52]. Jets [407]. Jitter [773]. Join [490]. Joint [1595]. Joyce [1345]. Jumping [1655, 1851].

K-Colouring [319]. K-Nesting [319]. K-Project [1386]. KADS [252, 259, 246]. Kanji [343]. kappa [921]. kappa-Trees [921]. KCM [727, 1067]. KCM/ [727]. KEEP [260]. Kernel [205, 1034, 514, 768]. Key [1797, 1800]. Keynote [1569]. Keyword [1715]. Kind [1620]. Kinematic [1644]. KITSS [419]. KIV [617]. KIWIS [180]. KL1 [614]. Kleene [881, 1677, 1096]. Kleene-Type [1677]. Knapsack [1357]. Knee [1875]. Knock [1875]. Knock-Knee [1875]. Know [1836]. Knowledge [267, 1486, 1566, 1498, 430, 944, 331, 727, 1054, 420, 945, 441, 460, 683, 116, 1018, 447, 1117, 1127, 643, 80, 265, 590, 403, 820, 450, 1129, 956, 723, 264, 1130, 249, 941, 818, 1501, 247, 1125, 803, 451, 341, 696, 1834, 648, 1565, 154, 942, 722, 814, 348, 736, 468, 652, 1647, 950, 1307, 269, 817, 469, 1132, 266, 1121, 418, 1114, 463, 471, 1551, 257, 260, 419, 261, 841, 414, 940, 1507, 268, 262, 454, 653, 421, 949, 270, 936, 452, 427, 258, 939, 156, 455, 948, 256, 682, 446, 953, 1567, 1123, 952]. Knowledge [325, 1017, 254, 1929]. Knowledge-Acquisition [249]. Knowledge-Base [1018]. Knowledge-Based [430, 441, 447, 643, 80, 590, 403, 264, 1501, 341, 722, 469, 463, 419, 414, 268, 421, 270, 325]. Knuth [35]. KOAN [1380]. KolmogorovComplexity [954]. KPROP [614]. Kripke [752]. KryptoKnight [1797]. KUSET [446].

L [1498, 1159, 976]. L-Domains [976]. Lab [1513]. label [704]. Labelled [1771]. Laboratory [1611, 1603, 1606, 1604, 321, 1625]. Lagrangian [1899]. Lambda [1321, 220, 228, 555, 993, 207, 873, 1299]. lambda-Calculus [207, 873, 1299]. Lambda-Terms [1321]. Lambda-Value [220]. lambdamu [1029]. lambdamu-Calculus [1029]. Language [1755, 84, 1018, 534, 178, 523, 1046, 450, 1504, 169, 86, 1756, 205, 657, 1812, 826, 342, 1034, 1503, 1257, 1904, 347, 1749, 1422, 1660, 1449, 668, 164, 878, 1063, 222, 1748, 498, 1507, 1324, 1713, 1272, 1625, 159, 227, 653, 141, 1548, 1271, 1042, 233, 671, 151, 8, 1483]. Languages [742, 1516, 331, 1768, 223, 960, 1117, 1524, 1826, 306, 1019, 1767, 207, 1236, 1953, 1175, 493, 958, 1179, 1149, 1641, 1958, 1387, 360, 1261, 1006, 739, 966, 1231, 453, 1727, 1230, 1961, 1630, 1194, 1963, 716, 1296, 44, 712]. Lapse [315]. Large [1365, 375, 1378, 736, 1020, 1452, 408, 1905, 834, 1427, 1936]. largest [1085]. L’assistance [644]. Last [329]. LaToKi [1271]. Lattice [1644, 1840, 1393]. Lattices [883]. Law [1593]. Layering [300]. Layers [741]. Layout [1148, 70]. Layouts [1875]. Lazy [1239, 1259, 622, 1641, 1261, 993, 1266, 595, 1523]. LDL [1625]. Leaf [1954]. Leaf-Size [1954]. Learner [690]. Learners [1166]. Learning [326, 1649, 1651, 852, 1053, 632, 1652, 698, 333, 673, 1654, 699, 160, 629, 1646, 343, 345, 346, 348, 1659, 807, 1660, 1647, 428, 630, 431, 638, 471, 248, 809, 322, 257, 680, 364, 366, 323, 688, 670, 412, 190, 660, 640, 1961, 429, 1963, 694, 445, 1648, 1645]. Least [1825]. Lecture [1254, 1236, 1197]. Legacy [827]. Legislation [1506]. LEI [1562]. Leitmotiv [1243]. Lemma [1682, 1673, 1681]. Lemmaizing [567]. Length [1894, 1876, 1146, 971]. Lengths [1024]. Less [1212]. Lesson [358]. Letters [957]. Level [1018, 78, 839, 1130, 1059, 1595, 779, 1124, 1790, 518, 1122, 357, 1608, 940, 1622, 63, 831, 1246, 1414, 1643, 1278]. Levels [1479, 482, 1564]. Lexical [1120, 1127, 1119, 1129, 1130, 1125, 1132, 1121, 1133, 1113, 1118, 1128, 1123]. lexicographic [21]. Lexicon [1744, 1126, 1116, 1131, 1115]. library [1922]. Licensing [1589]. Life [154]. Lighting [447]. LIKE [1744]. LIM [568]. Liminf [243]. Limited [1825, 898, 872, 1180]. Linda [73]. Lindenbaum [846]. Lindenbaum-Algebraic [846]. Line [1892, 1874, 1844, 18, 1410, 1952, 1842, 1786, 401, 1404, 1146]. Linear [1909, 1364, 1357, 1051, 851, 1838, 516, 1779, 1312, 864, 1873, 1014, 869, 1212, 1177, 1370, 639, 1359, 885, 136, 140, 889, 961, 1366, 570, 601, 1255, 926, 1289]. Linear-Input [570]. Lines [1873, 930, 5]. Linguistic [1120, 1562, 826, 1118]. linguistique [656]. Link [906, 539]. Linked [639]. Linking [607, 251]. Links [226]. LIPS [1264]. LISP [1640, 65, 94, 1518, 1541]. List [861, 912]. lists [47]. Literacy [701]. Literals [1831]. Literate [158]. Little [591]. Live [768]. Liveness [1913, 1209]. Load [513, 107, 1382, 1383, 1434, 1384]. Load-Balancing [107]. Loading [1358]. Local [68, 1891, 467, 1738, 1181, 993, 1908, 1721, 1383, 1083]. Localities [1205, 1004]. Locality [1562, 1536]. Location [1892, 928, 1142, 1143]. Locations [1912]. Löf [754]. Log [1822, 1853]. Logic [1766, 1909, 1050, 1249, 1265, 551, 1051, 851, 1788, 1040, 1039, 576, 1158, 1516, 109, 854, 1321, 33, 201, 114, 1016, 808, 1819, 115, 202, 855, 1832, 1509, 578, 1822, 596, 1046, 1056, 24, 118, 1057, 1817, 995, 1058, 556, 1527, 1510, 806, 204, 1043, 1500, 1312, 1493, 1326, 1038, 864, 220, 803, 1267, 207, 1236, 1014, 1250, 123, 1328, 848, 1257, 869, 125, 1212, 870, 541, 871, 991, 1815, 1506, 1492, 1887, 805, 994, 350, 875, 876, 1264, 1300, 129, 130, 131, 292, 1309, 1067, 134, 845, 1747, 1263, 135, 885, 1308, 136, 1888, 137]. Logic [1231, 1890, 211, 140, 889, 842, 614, 891, 1902, 835, 433, 598, 1042, 144, 846, 849, 1829, 1327, 896, 143, 1301, 1302, 2, 1925, 1089, 1289, 1093, 1297, 1095, 1290, 1288, 1098, 1099, 1094, 1105, 1109]. Logic-plus-Functional [201]. Logical [110, 859, 1826, 1745, 206, 1092, 1274, 1309, 136, 141, 1115, 893, 759, 550, 1101, 1085]. Logical-Based [141]. Logician [1068]. Logics [275, 1021, 618, 855, 578, 863, 1315, 847, 1828, 1213, 1311, 1069, 890, 850, 1306, 597, 1323]. Logspace [893, 21]. logsub [1853]. Long [355]. Long-Term [355]. Longest [1883, 1729, 1718]. Loop [71, 501, 70, 1438, 67, 508]. Loops [1401, 60, 61]. loss [710]. LOTOS [1224]. Low [967, 904, 1413, 482, 968, 192]. Low-Diameter [904]. Lower [302, 990, 1024, 196]. Lowlevel [1820]. Lowness [1862]. LR [1617, 1618]. LUSTRE [280].

Machine [852, 580, 990, 1435, 1130, 160, 1377, 807, 428, 415, 180, 356, 472, 515, 248, 809, 1258, 257, 833, 433, 834, 416, 1567, 1155, 1332, 1808, 1083, 1942]. Machines [441, 526, 1391, 898, 71, 1952, 1251, 72, 1954, 1224, 547, 466, 1543, 103, 891, 893, 1155]. Macintoshes [603]. Made [316, 930, 994].

Maintainable [723]. Maintaining [931, 1787, 1193]. Maintenance [403, 404, 811, 246, 982, 266, 150]. Major [623]. Makanin [30, 29, 31]. Make [251]. Making [411, 937, 451, 638]. Malign [1861]. Manage [627]. Management [1516, 108, 1054, 271, 403, 837, 825, 65, 514, 1377, 246, 1521, 1114, 405, 1532, 1752, 1541, 724]. Manager [1809, 1505, 185, 1774, 830]. Managing [373, 1486, 785]. Manifold [1430]. Manipulate [1546]. Manipulation [1748]. Manipulations [1949, 1386]. Manufacture [425]. Manufacturing [426]. Many [1021, 596, 863, 1883, 1121, 1323]. Many-Valued [1021, 863, 1323]. Map [1421, 931]. Mapped [309]. Mapping [1943, 1362, 1334, 309, 1339, 1445, 940]. mappings [1139]. Maps [1566]. Markedness [668]. Market [443]. Marshall [1596]. Martin [754]. Martin-Löf [754]. MASCOT [371]. Massive [89]. Massively [1340, 1410, 194, 493, 1371, 103, 834, 92, 1541]. Master [1595]. Mastering [356]. Mastery [673]. Match [1639]. Matching [1732, 1788, 1714, 1726, 572, 1725, 924, 1951, 1440, 593, 1722, 1616, 1730, 1733, 1723, 1724, 1717, 1914, 1960, 973, 161, 1134, 1135, 1136]. Materializing [1786]. Mathematica [619]. Mathematical [551, 111, 1419, 878, 1919]. Mathematics [625, 631, 764, 1108]. Mathpert [1053]. Matrices [1367, 420, 1360, 1368, 1427]. Matricial [242]. Matrix [516, 902, 1370, 1718, 1424, 1407, 1444, 1963]. Matter [335]. Mature [1535]. Max [1846]. Max-Cut [1846]. Maxima [1850]. Maximal [1440, 14]. Maximality [707]. Maximum [925, 918, 1918]. May [1836]. Mazurkiewicz [236]. McDonnell [1602]. Meaning [1125, 1556, 1131, 1118]. Means [416]. Measurements [1337]. Measures [243, 813]. Measuring [120, 688]. mécanisme [644]. Mechanical [455].

Mechanism [439, 1262, 94, 1336, 746]. Mechanisms [649]. Mechanizing [1028]. Media [791, 783, 781]. Mediating [261]. Medical [1350]. Medicine [340]. Medium [1222, 336]. MegaLog [1054]. Meiko [832, 1355]. Membership [986]. memoization [1286]. Memories [517]. Memory [106, 1516, 108, 1426, 1409, 1374, 536, 80, 1362, 193, 1527, 68, 898, 81, 1331, 102, 1379, 1956, 518, 1380, 1521, 1522, 88, 1916, 1381, 1376, 494, 1407, 842, 1532, 1384, 75, 1541, 1083]. Memory-Efficient [536]. Memory-Reuse [1527]. Mental [632, 660]. Merging [301, 979, 1174, 475, 1886, 838]. Mesh [1455, 1852]. Meshes [195, 1397]. Message [1338, 956, 1399, 1887, 1219, 505]. Message-Brokers [505]. Meta [1276, 1819, 1822, 1826, 1825, 1814, 1828, 210, 589, 838, 1831, 1752, 177]. Meta-CASE [177]. Meta-Integration [838]. Meta-Interpreter [1276]. Meta-Logics [1828]. Meta-Programming [1822, 1826, 1825, 1814, 1831]. Meta-Theory [210, 589]. Metalevel [1823, 117, 1820, 1815, 133, 1818]. Metalogic [1821]. Metamodeling [176]. Metaphor [1116]. Metaprograms [1835, 142]. Metastructures [1253]. Metatheorem [1701, 1693]. Method [587, 1826, 179, 169, 1462, 104, 1353, 1371, 1739, 1474, 427, 1428, 1417, 1400, 597]. méthodologie [674]. Methodologies [1735]. Methodology [200, 425, 105, 167, 1745, 728, 365, 1094]. Methods [175, 1584, 1792, 814, 184, 1583, 257, 1582, 1407, 198, 253, 1153, 1836, 703, 1941, 1937, 1135, 1140]. Metric [987]. Metrics [1798, 374, 376]. MGTP [622]. Micro [345]. Micro-World [345]. Microlog [528]. Microprocessor [580]. Microsystems [792]. Middle [573]. Middle-Out [573]. MIDI [336]. MIMD [1409, 507, 1391, 1448, 1383]. Min [1846, 1907]. Min-Cut [1846]. Mind [1166].

Minimal [1441, 880, 1103, 868]. Minimise [494]. Minimization [1221, 1955]. Minimizing [1775, 918, 508]. Minimum [1394, 1874, 906, 933, 307, 1876, 1450, 3]. Mining [1502]. MinMax [1894]. Minors [866]. Mise [678]. Mismatches [1714]. Misrepresentation [684]. Mixed [513]. Mixed-Mode [513]. Mixing [441]. Mixing-Machines-Development [441]. MIZ [1062]. MIZ-PR [1062]. ML [997]. Mobile [1198, 477, 436]. Mobility [1178]. Modal [851, 863, 876, 841, 135, 890, 835, 1306]. Modalities [896]. Modality [1307, 143]. Mode [513]. Model [567, 1754, 488, 1040, 1478, 331, 1081, 632, 34, 1827, 302, 1206, 438, 265, 858, 1811, 169, 720, 1557, 909, 398, 502, 1746, 721, 1233, 1125, 1606, 1558, 1740, 622, 1751, 1547, 579, 942, 406, 547, 836, 404, 1491, 1791, 652, 1774, 199, 679, 829, 822, 548, 260, 667, 508, 1662, 240, 1007, 429, 835, 1490, 1564, 830, 402, 172, 1965, 708, 1141]. Model-Based [632, 265, 942, 406, 404, 822, 429, 402]. Model-Theoretic [34, 1007]. modèle [633, 674, 678]. Modeling [673, 700, 1550, 677, 424, 251, 358, 1735, 680, 1407, 697]. modelisation [656, 675, 685, 659]. Modelling [1461, 1479, 1753, 178, 1483, 837, 1481, 1489, 1796, 154, 1488, 711, 658, 415, 158, 668, 661, 405, 261, 1186, 1743, 949, 427, 1460, 165, 705]. Modellings [1314]. Models [275, 439, 1783, 1374, 79, 179, 249, 250, 1751, 869, 259, 126, 1449, 354, 1563, 408, 993, 1598, 464, 640, 682, 1273, 1155, 308, 254, 1396, 714, 715, 1083, 712]. Modes [480]. Modification [840]. Modifications [1669, 1787]. MODULA [86, 87]. MODULA-2 [87]. MODULA-S [86]. Modular [1620, 705, 1334, 1037, 101, 1359, 714, 1281]. modularisation [1101]. Modularity [886]. Modularizing [1819]. Modulation [433]. Modules [887]. Modulo [1201]. Molecular [1448]. Molecules [1722]. MOLTKE [266]. MOLTKE/HyDi [266]. Monadic [4]. Monitor [1337]. Monitoring [448, 1568]. Monochrome [1438]. Monoidal [239]. Monotonic [863, 1316, 1660, 112]. Monotonicity [855]. Monotonous [189]. Monte [1423]. Morphisms [1167]. Mosaic [1923]. Mosca [386]. Most [1219]. Motion [1892, 182, 1451]. Motivational [690]. Mouse [356, 299]. Movement [1916]. Movie [794]. Movies [794]. Moving [932, 931, 950, 417, 13]. MPPC [92]. MSL [878]. Msup [1808]. Mu [279]. Mu-Calculus [279]. Much [1648]. Mulmuley [989]. Multi [1948, 913, 264, 840, 1331, 66, 907, 101, 424, 385, 1522, 394, 680, 734, 704, 1337, 95]. Multi-Agent [680, 734]. Multi-Database [840]. Multi-Dimensional [913]. Multi-Expert [264]. Multi-Grain [66]. multi-label [704]. Multi-Packet [907]. Multi-Processor [1331, 394]. Multi-Resolution [1948]. Multi-Threaded [101, 1522]. Multi-Transputer [1337, 95]. Multi-Unit [385]. Multi-Zone [424]. Multiagent [744, 647]. Multicast [1847]. Multicomputer [1375, 1923]. Multicomputers [513, 1452, 1384, 512, 485]. Multidimensional [492, 1960, 1959]. Multigrid [104, 1437]. multihead [1079]. Multilayered [1834]. Multilevel [1809, 1806, 652, 1807, 1808]. Multimedia [790, 776, 785, 771, 788, 789, 782, 792, 778, 775, 777, 793, 359, 322, 795]. Multipacket [91]. Multiparadigm [1475, 1288]. Multiple [1651, 618, 727, 1415, 1844, 1715, 435, 358, 639, 423, 1728, 1314, 640, 1033, 1564, 650, 539]. Multiple-Quantum-Well [539]. Multiple-Valued [618]. Multiplication [99]. Multiplications [902]. Multiplicative [1909, 864]. Multiplicities [962]. Multiplicity [971]. Multiplicity-Dependent [971]. Multiport [1376]. Multiprocessing [1438]. Multiprocessor [1443, 521, 1341, 517, 1434, 90, 1417, 1396]. Multiprocessors [1441, 106, 1409, 507, 486, 81, 102, 1448, 518, 75]. Multiscale [1398]. Multisets [933]. Multithreaded [1378]. Multiuser [786]. Multivalued [576]. Multivariate [314]. Multiway [1232]. Multiwindow [1341]. Muse [509]. Musicians [336]. Mutual [1346, 1144].

n [1853]. Nagging [329]. Naming [1818]. Narrowing [592, 1254, 1060, 1237, 1272, 1293, 1291, 1292]. National [1613]. Natural [1241, 609, 1755, 1056, 599, 1504, 169, 826, 1503, 210, 1029, 1507, 1727, 159]. Naturally [1768]. Nature [1188]. Navigational [367]. NC [904]. NCR [833]. NCR/Teradata [833]. Near [978, 902, 211]. Near-Horn [211]. Near-Optimal [902]. Near-Perfect [978]. Nearest [422, 988]. Nearest-Neighbor [988]. Nearly [1850]. Nearly-Constant [1850]. Necessary [754]. Need [1462, 940, 1836]. Needs [1114]. neg [1173]. Negation [1039, 1163, 117, 1038, 1328, 579, 96, 880, 845, 1325, 1829]. Negations [872]. Negative [845, 660]. négociation [633]. Negotiation [734]. Neighbor [988]. Neighbour [422, 1556]. Neighbourhood [1434]. Neighbourhood-Based [1434]. NelleN [158]. Nested [1362, 1773, 60, 199]. Nesting [319, 1748]. NESTOR [359]. Net [1457, 384, 424, 1417, 714, 715, 708, 709, 717]. Netlog [1042]. Nets [1000, 608, 1015, 1172, 1485, 1487, 1002, 1199, 428, 1217, 886, 1543, 1003, 1073, 1078, 444, 703, 1081, 705]. Network [1432, 422, 310, 1442, 1169, 779, 782, 784, 1331, 778, 542, 1871, 1955, 1452, 917, 405, 667, 465, 795, 783, 1945, 1404, 308, 1446, 6, 25]. Networks [1394, 774, 407, 1329, 799, 1016, 503, 1410, 1431, 484, 443, 1944, 1354, 1851, 1333, 476, 1853, 404, 1445, 1335, 1151, 423, 1908, 682, 512, 781, 1922]. Neural [407, 1944, 1956, 424, 1151, 465, 1543, 1965, 1945, 1940]. Neurocomputing [824]. Next [1735, 719, 783]. Nexttime [1212]. Nexttime-Less [1212]. NFA’s [1719]. niveau [675]. no [1520]. Node [913]. Nodes [303]. noisy [1144]. Non [1036, 1405, 112, 855, 578, 1769, 863, 844, 1125, 1695, 869, 1289, 1316, 917, 886, 1721, 1876, 597, 674]. Non-Clausal [1036]. Non-Compatible [1695]. Non-Crossing [1876]. Non-Deterministic [1769]. Non-Essential [1405]. Non-Horn [844]. Non-Interfering [917]. Non-Lexical [1125]. Non-linear [1289]. Non-Local [1721]. Non-Monotonic [863, 1316, 112]. Non-Monotonicity [855]. Non-Stable [869]. Non-Standard [597]. Nonconvex [929]. Nondeterminism [237, 1180, 222]. Nondeterministic [7]. Nonlinear [1026, 1868]. Nonmonotonic [848]. Nonwellfounded [227]. Normal [1039, 1830, 965, 1779, 577, 872, 1098, 1295]. Normal-Form [1779]. Normalization [1165]. Normalizing [1868]. Normative [1794]. Norms [1372]. Note [37, 1951, 1671, 1675, 1679, 1684, 1689, 1698, 1704, 1709]. Notes [1509, 810]. Nothing [491]. notions [717]. Novel [1352]. Novices [628]. Novikov [895]. Novikov-Style [895]. NP [1880, 1841, 1106]. NP-Complete [1841]. NP-Hard [1880]. NREVERSAL [1542]. NUDIST [726]. Number [1899, 1874, 872, 1444, 983, 23]. Numbers [1167, 1860, 660, 763, 192]. Numeration [961]. Numeric [478]. Numerical [1419, 1369, 1442, 530, 470, 1428]. Nuprl [590].

O [1853, 1266, 1220]. OASIS [164]. Object [742, 373, 152, 750, 1754, 1478, 330, 1783, 1824, 147, 1759, 223, 534, 438, 1164, 1799, 1443, 1420, 721, 153, 828, 1758, 748, 1751, 154, 493, 1790, 186, 1536, 164, 145, 472, 728, 747, 829, 410, 737, 739, 838, 89, 146, 1782, 741, 1630, 1757, 1597, 489, 1752, 1637, 1966, 749, 1945, 151, 1277, 44, 1142, 1143]. Object-Based [1754, 534, 741, 749]. Object-Oriented [742, 373, 152, 330, 1824, 147, 1759, 223, 438, 1799, 1443, 1420, 721, 153, 828, 1758, 1790, 164, 145, 728, 747, 829, 410, 739, 838, 1782, 1630, 1757, 1597, 489, 1637, 151, 1277, 44]. Objective [145]. Objectives [650]. ObjectMath [1643]. OBJECTOR [361]. Objects [1051, 115, 1948, 1546, 1857, 1835, 778, 1535, 1392, 728, 189, 1520, 831, 746, 417, 455, 1519, 449]. Observation [1201, 1099, 1144]. Observations [1917]. Observed [1185]. Observer [1554]. Obstacles [1837]. OCCAM [709, 1429]. Occupancy [465]. Occur [1240, 1027]. Occur-Check [1240, 1027]. Occurrences [1720]. Octrees [1949]. oeuvre [678]. Office [213]. Offs [236]. OIKOS [1470]. OL [959]. Old [272]. Old-Fashioned [272]. Omega [1031, 1107, 883]. omega-automata [1107]. omega-Herbrand [883]. Omega-Rule [1031]. Omniscience [1309]. On-Line [1844, 1410, 1842, 1786, 1404]. Once [1219]. One [956, 1898, 1177, 598, 38, 1196, 1098]. One-Message [956]. One-Rule [38, 1196]. One-State [1177]. One-Way [1898]. Ones [1882]. Online [344]. Only [1916]. Ontic [590]. onto [1943, 1445, 940]. Ontology [1113, 1131]. Ontology-Driven [1113]. OO [1734]. OODB [1274]. OODBs [1781]. Open [1051, 202, 1789, 337, 343, 832, 775]. Operating [778, 783, 834]. Operation [1832, 720, 1635]. Operation-Triggered [1635]. Operational [1218, 110, 157, 246, 251, 1234, 498, 221]. Operationalize [250]. Operationalizing [412]. Operations [1119, 527, 908, 1721]. Operator [276, 277, 1177]. Operators [113, 1950, 497, 883, 892]. Opium [1059, 1512]. Opportunistic [512]. Opportunities [1613, 1939]. Optic [425].

Optical [487, 1375, 539]. Optimal [977, 1951, 902, 484, 480, 309, 72, 319, 5, 1847, 1893, 16, 907, 1433, 1412, 1960, 1400, 1937]. Optimality [1908]. Optimisation [490, 1494]. Optimised [403]. Optimization [1256, 71, 58, 218, 1871, 197, 582, 1339, 57, 1940]. Optimizations [1820, 1779, 870]. Optimizing [1770, 1769, 85]. Optimum [1401, 1916]. Opto [539]. Opto-Electronic [539]. Or-Parallel [121, 1413, 1238, 1032, 538, 138]. Or-Parallelism [110]. Oracle [1284, 893, 832]. Oracles [1895]. Orca [828]. Order [1040, 852, 1869, 564, 902, 556, 204, 577, 1544, 1685, 1687, 1257, 755, 594, 1048, 879, 885, 136, 1324, 1191, 1272, 140, 889, 955, 1078, 305, 233, 1282, 1303, 4, 1083, 1280, 1099]. Order-Sorted [564, 594, 1048, 879, 1324, 1280]. Ordered [1022, 866, 1323]. Ordering [870, 585, 16]. Orders [304]. Ordinal [861, 1100]. ordinary [1084]. Organization [1608, 726]. Organizational [168, 724]. Organizing [1421, 1945]. Orientation [1553, 1488, 188]. Oriented [742, 373, 152, 1242, 330, 1824, 147, 1759, 223, 438, 1576, 1799, 1443, 1420, 721, 153, 382, 828, 1758, 493, 1790, 164, 145, 472, 728, 747, 829, 410, 739, 838, 1782, 1630, 1757, 1597, 489, 1637, 1042, 151, 1277, 1577, 44, 1304]. Orthoconvex [1856]. Orthogonal [594]. Orthogonalization [1372]. OSCAR [66]. Other [1854]. Our [1264]. Outline [1277]. Outlines [297]. Output [1878]. Overhead [70]. Overloading [564]. Overview [1343, 205, 50, 822, 293, 610, 1539].

P [526, 1880, 1344, 1184]. P-90 [526]. P-Bi-Immune [1184]. Package [364, 433]. Packet [774, 773, 907, 1377]. Packet-Based [1377]. Packet-Switched [774]. Packet-Switching [773]. Packing [197].

Paged [510]. Pairs [927, 22]. Palindromes [977]. Pancake [1394]. Pandora [1823]. Panel [750, 1580, 1588, 1613, 1586, 1600]. PANTER [418]. Paper [479]. Papers [1491]. Parabolic [1353]. Paradigm [395, 941, 545, 1583, 423, 323, 1178]. Paradigms [1614, 801]. Paradox [1780]. Paradoxes [565]. Paragon [520]. Paralation [502]. Parallel [106, 1394, 742, 1457, 977, 1432, 488, 1367, 1157, 1340, 1338, 1342, 1850, 919, 84, 235, 78, 478, 537, 523, 1943, 1362, 1949, 48, 1390, 301, 1365, 516, 1951, 1435, 902, 1167, 481, 1440, 1334, 1421, 1857, 909, 1058, 1410, 1395, 1386, 68, 194, 121, 535, 1436, 1420, 1401, 1388, 1399, 897, 483, 906, 530, 1952, 522, 1354, 1202, 217, 1672, 1707, 1174, 1842, 1331, 622, 905, 93, 648, 514, 832, 104, 1353, 542, 392, 94, 124, 493, 545, 1356, 1723, 1955, 475, 1377, 1392, 1448, 1640, 197, 1360, 1149, 1521, 54, 1449, 1398, 1385, 77, 99, 474]. Parallel [540, 391, 612, 1433, 1348, 1339, 491, 1445, 222, 498, 1958, 506, 1387, 833, 1907, 1412, 546, 1371, 1424, 908, 1189, 1418, 453, 89, 1393, 1538, 314, 1960, 1959, 1408, 103, 1428, 595, 529, 473, 1366, 1033, 1451, 1964, 1411, 1417, 834, 92, 1153, 1836, 1074, 97, 543, 83, 51, 1427, 1155, 1404, 558, 1541, 100, 1456, 539, 15, 1927, 1920, 14, 1413, 1238, 1032, 538, 1237, 138, 614, 1363]. Parallel-III [1363]. Parallelism [492, 110, 86, 65, 510, 1416, 527, 62, 63, 888, 1423, 490, 74, 1414, 85]. Parallelization [1409, 105, 80, 53, 1458, 81, 1402, 57, 98]. Parallelizing [59, 1403, 66, 541, 61]. Parametric [922, 219, 1004]. Paramodulation [584, 1323]. Parikh [1683]. Parlog [110, 48]. ParMod [83]. ParMod-C [83]. Parsers [1617]. ParSim [1408]. Parsing [1618, 1958, 1619, 453]. ’Part [1758, 110, 111, 219]. Partial [1638, 1301, 203, 208, 104, 1353, 871, 1816, 879, 1012, 1078, 713, 1283]. Partially [1436]. Participation [693]. Particle [407].

Partition [68, 1840, 11]. Partitioning [1362, 1334, 72, 60, 67, 540, 189, 1942]. Partitions [1359, 1840]. Parts [344, 468, 415]. Passenger [440]. Passing [1338, 1869, 1399, 1887, 1270, 1532]. Past [1043]. Patent [1593]. Path [1947, 1169, 1239, 1346, 22]. Path-Controlled [1947]. Paths [857, 906, 1837, 1876, 1798, 20, 21]. Pathwidth [1848, 3]. Pattern [1947, 1714, 924, 1951, 1944, 1616, 1717, 1956, 409, 366, 1639, 1960, 973, 1946, 1965, 1966]. Pattern-Match [1639]. Patterns [181, 1560, 1956, 1548, 1283]. PB [1276]. PCP [79]. PCP/PFP [79]. PDEs [1354]. PEACE [1486]. Pedagogic [684]. Pedagogical [646]. pédagogique [674, 685]. Penalties [1717]. [1644, 1254, 390, 727]. 1012 [526, 833]. 860 [1402, 1360]. Aspect [1130]. by [1886]. CM [1460]. Common [1716]. Control [58]. Fold [1248]. HyDi [266]. III [1531]. IM [1460]. Learning [359]. Logic [1018, 1258]. Parallel [1966]. PFP [79]. RTTL [295]. Teradata [833]. Tokyo [1548]. Video [786, 1753]. X-Windows [793]. People [1546, 699, 1544]. PEPSIM [1406]. PEPSIM-ST [1406]. Perfect [1157, 975, 978, 1848, 53]. Perfect-Benchmark [53]. Performance [509, 1337, 1409, 799, 1374, 448, 1369, 193, 1401, 605, 542, 285, 1640, 186, 517, 668, 355, 491, 471, 149, 833, 1407, 715, 1143]. Performances [1433]. Performing [482]. Periodic [1720]. Periods [977]. Permutation [1453, 1299]. persistent [1081]. Persisting [1548]. Personal [1477]. Perspective [1576, 287, 1477, 1579, 1765, 645, 214, 1932]. Petri [1457, 1081, 1000, 705, 715, 1172, 1002, 1199, 1217, 1417, 1003, 1073]. Petri-Net-Based [1417]. Phase [1499]. Phases [1388, 1793]. Phenomenon [1121]. Philosophical [802, 287, 954]. Phonological [354]. Phylogeny [975].

Physical [406, 1339, 1429]. Physically [1155]. Physics [459, 92]. pi [1197]. pi-Calculus [1197]. Picture [1953, 1961]. PIE64 [514]. PIM [426]. Pipeline [490]. Pipelines [1634]. Pipelining [1633, 1410, 1635, 1400]. PIPS [1456]. Place [1885]. Placement [1636]. Places [1075]. Planar [915, 916, 1841, 918, 20]. Planarity [981]. Plane [1876, 1087, 13]. Planification [685]. Planning [1892, 1566, 426, 182, 1501, 1013, 1596, 676, 1028, 349, 811, 435, 812, 358, 810, 1451, 172]. Plans [944, 426, 943, 574, 718]. Plants [430]. Platform [1054, 832, 1387]. Platforms [389]. Plausibility [1095]. Play [1220]. Plays [623]. Pluralistic [1166]. plus [201]. PMI [262]. PML [1474]. Point [1892, 985, 1410, 1678, 1554, 950, 1439, 928]. Pointer [990]. Pointers [61]. Points [932, 1889, 1864, 1531, 891, 13, 45, 1145]. Poisson [1437]. Policy [1801]. Pollution [1568]. Polyadic [1197]. Polycyclic [1373]. Polygon [1864]. Polygonal [1874]. Polygons [1859, 930, 1893, 987]. Polyhedra [930, 1858]. polyhedral [1092]. Polyhedron [929]. Polyinstantiation [1805]. Polymorphic [854, 1062, 1903, 882, 1324, 894]. Polymorphism [219]. Polynomial [560, 972, 958, 872, 101, 1877, 312, 1194, 1400]. Polynomials [974, 311, 314, 1901]. Polytopes [1150]. Pomsets [236]. Pool [89]. Popperian [637, 1654]. Portability [54]. Portable [1513, 1435, 1387, 1384, 543]. Porting [379]. posés [664]. Poset [1170]. Posets [229, 1216]. Position [479]. Positions [1794, 950]. Positive [1367, 123, 1901]. Possibilities [245]. Possible [824, 99]. Possibly [596]. Post [324]. Post-TV [324]. Postconditions [843]. pour [659]. Power [1654, 1168, 699, 1686, 1688, 232, 890, 955, 458, 1079]. Powerful [277, 78, 766, 1937]. PPS [1498]. PR [1062]. Practica [1612]. Practical [1726, 395, 1379, 812, 51]. Practice [245, 1391, 1492, 583, 1598, 298]. Practitioner [1593]. Practitioners [1590, 1591]. Pragmatic [940]. Pragmatics [1133]. PRAM [979]. PRAMs [899, 196]. Precompiling [1529]. Preconditioners [1371]. Preconditioning [1354]. Preconditions [843]. Predicate [1128]. Predicates [1906, 1691, 745]. Predicative [296]. Predictable [1118]. Predicting [391]. Prediction [443, 1382, 675]. Predictive [1658]. Preference [855, 834]. Preferences [1316]. Preliminary [762, 753, 860]. PRELUDE [543]. premières [665]. Preorders [1011]. Prescriptive [1550]. Presence [971]. Presentation [825, 778, 794]. Presentations [790]. Presenting [795]. preservation [707]. Preserving [1248, 1212, 1210, 1075]. Presorted [1390]. Presuppositions [1133]. Primal [1035]. Prime [524, 569]. Primitive [527, 224, 1085, 1285]. Primitives [1820, 1544]. Prince [1496]. Principal [1840]. Principle [598, 2]. Principles [1615, 1612, 155, 813]. Printed [448]. priorities [711]. Priority [301, 769, 908]. Private [785, 486]. Probabilism [1654]. Probabilistic [273, 1714, 1229, 1878, 1913, 1011, 1166, 118, 927, 456, 814, 1228, 1349, 1917, 1712, 682, 1017, 1090]. Probabilities [1229, 631, 1109]. Probability [143, 1146]. Problem [1357, 1240, 534, 1838, 701, 1436, 252, 1846, 1883, 635, 1847, 1013, 1730, 501, 730, 962, 197, 1884, 1179, 1918, 1718, 432, 412, 1200, 253, 1137, 1839, 299, 100, 1645, 912, 12, 10, 22, 1107]. problèmes [664, 645]. Problems [1623, 1854, 1160, 438, 1365, 924, 1162, 1148, 568, 486, 922, 906, 1701, 1693, 1703, 1696, 1841, 1871, 1852, 1716, 911, 285, 1448, 385, 263, 623, 938, 1872, 408, 1662, 1187, 926, 11, 14, 21]. Procedural [331, 117, 125]. Procedure [1649, 1015, 35, 1890, 610, 233, 558, 1082, 1102]. Procedures [554]. Process [1486, 1225, 276, 1469, 1478, 109, 1461, 277, 1479, 271, 1601, 166, 534, 1483, 1459, 503, 1462, 1471, 1466, 1481, 1489, 1476, 1487, 154, 1472, 1488, 1491, 286, 1464, 638, 1477, 260, 293, 1662, 177, 739, 1602, 168, 1467, 427, 1575, 1484, 1465, 1460, 416]. Process-Centered [1471]. Process-Centred [177]. Processes [1229, 1869, 411, 1198, 1205, 1206, 1911, 1207, 1208, 1913, 1011, 532, 1431, 1480, 525, 477, 1204, 1475, 1228, 1348, 234, 296, 505, 708, 1100, 713]. Processing [460, 478, 523, 482, 1435, 1799, 68, 1559, 826, 1503, 183, 814, 1806, 127, 816, 186, 1449, 491, 107, 1807, 1507, 473, 92, 444, 1928, 1939]. Processor [1331, 1352, 1445, 1370, 1444, 90, 1368, 394]. Processors [1401, 1331, 476, 1376, 1427]. Produce [330]. Product [1351, 1477, 1598, 47]. Production [89, 401]. Products [409]. Professional [1589]. Professionals [1592]. Profile [1624]. Profiling [1340]. Program [1050, 328, 1409, 242, 114, 1419, 837, 1008, 65, 340, 530, 348, 77, 582, 494, 1608, 1210, 142, 76, 709, 1110, 1610]. Programmable [370]. programmed [9]. Programming [1343, 1454, 1265, 744, 1051, 692, 1788, 200, 1040, 1516, 1824, 331, 632, 507, 201, 147, 808, 1511, 1819, 115, 79, 332, 855, 1268, 1822, 1826, 673, 1821, 1046, 937, 1825, 1814, 400, 1058, 628, 1510, 204, 1443, 1493, 1420, 1233, 789, 1267, 207, 123, 1328, 848, 648, 50, 1257, 493, 1815, 1492, 805, 994, 351, 54, 1344, 1264, 1300, 1063, 129, 131, 661, 360, 548, 1067, 546, 1747, 1263, 136, 137, 1266, 670, 211, 227, 595, 1230, 791, 804, 1964, 1831, 1643, 1042, 144, 83, 926, 51, 1456, 1302, 1289, 1290, 1288, 1088]. Programs [49, 1249, 852, 1039, 109, 1338, 1830, 580, 1832, 1509, 78, 73, 857, 536, 118, 1817, 995, 119, 1527, 53, 1769, 1043, 1326, 1038, 1906, 1399, 1776, 1891, 1173, 280, 206, 1236, 72, 317, 1250, 573, 94, 125, 384, 545, 870, 541, 1506, 474, 1581, 391, 504, 88, 1445, 358, 1320, 1258, 756, 1027, 363, 845, 884, 1007, 1408, 1275, 74, 1902, 1271, 1417, 1637, 846, 1829, 370, 1327, 85, 1301, 1297, 1094, 1102]. Progress [243]. Proj [231]. Project [727, 1744, 1603, 375, 1386, 398, 1496, 828, 1606, 1614, 463, 321, 359, 92, 479]. projection [1092, 1106]. Projections [1554]. Projective [1]. Projects [1605]. projet [665]. Proliferation [534]. Prolog [1256, 804, 410, 1282, 1284, 108, 727, 1413, 1511, 1820, 119, 120, 1262, 1510, 121, 666, 95, 510, 538, 348, 127, 1252, 1064, 1065, 506, 133, 1066, 1531, 138, 888, 1275, 1270, 1532, 505]. Prolog-Written [133]. Promise [827, 1160]. Promises [487, 285]. Proof [608, 1015, 1828, 866, 1316, 1294, 582, 583, 297, 588, 610, 574, 1080, 1087, 42]. Proof-Theory [1828]. Proofs [551, 1052, 1030, 580, 578, 956, 1025, 1312, 868, 1028, 1063, 756, 860, 1102, 1103]. Propagation [1832, 1412, 44]. Properties [1248, 113, 1913, 561, 1170, 1676, 1692, 1695, 1690, 1118, 884, 1209, 1347, 1210, 1109]. Property [992, 1882, 1868]. Propia [1071]. Proportional [1662]. Proposal [338]. Proposed [997]. Propositional [571, 1213, 614, 891, 849]. ProPre. [1063]. ProQuery [1274]. Protection [425]. Protein [1722, 1720]. Protocol [780, 771, 798, 1222, 255, 796, 531, 1585, 769, 781, 1429]. Protocols [1796, 1802, 529]. PROTOS [1498]. PROTOS-L [1498]. Prototype [616, 414]. Prototyping [1749, 1298]. Provability [1082]. Provably [1629]. Proved [1005]. Prover [611, 618, 1055, 619, 1060, 622, 579, 1062, 612, 614, 603, 620]. Provers [567, 624]. Provided [470]. Providing [635]. Proving [607, 1513, 1321, 590, 599, 1032, 581, 1833, 585, 884, 552, 1901, 597, 1111, 1303, 1094]. Proximity [1193]. Pruning [1027]. Pseudo [137, 1856]. Pseudo-Resolution [137]. Psub [143]. PTAH [478]. Pumping [1682, 1681]. Pundit [1504]. Pure [997, 1781]. Purposes [697]. Push [1242]. Push-Down [1242]. Putting [1599, 297]. Puzzle [1957]. Puzzles [565]. PVS [616]. Pyramid [1332].

Q [1833]. QR [1447]. Quadric [1396]. Quadtrees [1949]. Qualified [702, 43]. Qualitative [823, 1553, 1568, 726, 640, 821, 724]. Quality [409]. Quantification [1748]. Quantifier [1025, 1104]. Quantifiers [144, 2]. Quantifying [853]. Quantitative [672]. Quantity [1331]. Quantizer [1438]. Quantum [539]. Quasi [866, 959]. Quasi-Deterministic [959]. Queries [1755, 1778, 1786, 1659, 490, 649, 1091, 1109]. Query [1768, 1018, 523, 1262, 1756, 840, 1767, 186, 1900, 180, 1563, 893]. Quest [240]. Questions [329, 1657, 285, 951]. Queue [908]. Queues [301, 1383]. Quickly [970]. QUIZ [339].

R [1276]. Race [1342]. Radar [1356].

Radius [1893]. Raleigh [836]. RAM [302]. Random [1032, 920, 15, 19]. randomization [1920]. Randomized [1732, 1897, 1714, 1850, 1379, 974, 1907]. Randomizing [1863]. Range [1147, 1850, 990]. Range-Maxima [1850]. Ranking [1149]. Raster [1546]. Raster-Vector [1546]. Rational [1162, 963, 962, 1575]. Ray [187]. Ray-Z-Buffer [187]. RDBMSs [728]. Reacting [690]. Reactive [390, 1203, 280, 284, 173, 1209]. Reactive/ [390]. Read [1916]. Read-Only [1916]. Real [272, 275, 273, 276, 780, 108, 944, 460, 381, 858, 762, 279, 779, 1910, 1034, 283, 285, 286, 287, 385, 288, 396, 938, 379, 291, 292, 506, 372, 465, 1444, 763, 1368, 1223, 781, 413, 380, 1289]. Real-Number [1444]. Real-Time [273, 780, 108, 460, 381, 858, 279, 779, 1910, 1034, 283, 285, 385, 288, 396, 379, 291, 292, 506, 465, 1223, 781, 380].

Realistic [302, 1871]. Realizability [752]. Realization [1428]. Reason [811]. Reasoning [947, 112, 1509, 1913, 179, 1555, 1553, 1562, 208, 848, 573, 942, 406, 404, 1815, 1561, 1133, 638, 457, 1552, 845, 1069, 821, 1317, 550, 401, 724, 650, 214]. Reasoning-Congruent [638]. Rebalancing [910]. Recipe [272]. Recognition [1947, 1944, 905, 1956, 409, 1149, 418, 1959, 1964, 1946, 1965, 1945]. Recognizable [1953, 1961, 716]. Recognizing [59, 1859, 1873]. Recommendation [1429]. Recommendations [1605]. Reconciling [157]. Reconfigurability [1442]. Reconfigurable [1334, 1442, 1351, 1397, 1404]. Reconfiguring [322, 308]. Recording [1491]. Recovery [1955, 1450, 1358]. Rectangles [1645]. Rectangular [928, 12]. Rectilinear [1847, 1837]. Recurrence [1431, 1366]. Recurrences [59, 961]. Recursion [632, 1026, 881, 1285, 1096]. Recursive [1650, 530, 573, 541, 873, 424, 1062, 224, 1085]. Redefining [1122]. Redescription [1662]. Reduce [512]. Reducer [1538]. Reducing [1331, 1360]. Reduction [306, 207, 70, 1377, 497, 555, 1393, 1523]. Reductions [1897, 967, 1159]. Redundancy [1036, 422, 1820, 1020]. Reference [1459, 667]. References [64, 61]. Referential [1738]. Refinement [532, 1203, 1199, 257, 1888, 1003, 1077, 1076, 1075, 300, 706]. refinements [707]. Refitting [943]. Reflection [850]. Reflective [1821, 747, 753]. Reflex [1950]. Reform [1026]. Refutations [1024]. Regard [387]. Regionalization [1549]. Regions [1537]. Register [1633, 1632, 62]. Regular [36, 1719, 1659, 127, 1717, 1179, 1359, 884, 1959, 961, 1194]. Related [1345, 1670, 1697, 1702, 1858, 1154, 926]. Relating [237, 1037]. Relation [1823, 199, 586, 311, 1914, 841, 1285]. Relation-Finding [311]. Relational [1268, 171, 841, 1141, 1091, 1139, 1136]. Relations [306, 1758, 1760, 1764, 225, 40, 728, 1833, 886, 936, 1741, 1255, 1818, 1104]. Relationship [1560, 1745, 1740, 126, 1750]. Relationships [1555, 1866, 1275]. Relativizations [1854]. Relativized [968]. Relevance [1313]. Reliability [198, 6]. Reliable [972, 519]. Remainder [581]. Remarks [1158]. Remediation [363]. Remote [732]. Renaming [62]. Rendez [395]. Rendez-Vous [395]. Reordering [1831]. Reorganization [913]. Repeated [516]. Replacement [1830, 1708, 1668, 1705, 1706, 1707, 1667, 1278]. Replay [1342, 1336]. Replicated [1806]. Replication [1533]. Replication-Based [1533]. Report [615, 762, 753]. Reporting [990]. Represent [721]. Representation [425, 1117, 1130, 65, 451, 1834, 384, 814, 56, 652, 841, 1128, 1507, 651, 1929]. Representations [166, 1268, 1386, 638, 639, 261, 233, 1108]. Representing [1127, 1644, 450, 803, 449]. Requirements [167, 163, 828, 1124, 1472, 1794, 1581, 159, 783]. Research [703, 1697, 1702, 1578, 688, 89, 1924]. Resemblance [181]. residue [710]. Resistant [604]. Resolution [1021, 1036, 1022, 1948, 1826, 568, 1014, 137, 1324, 971, 610, 598, 1323, 2, 1090, 645]. Resolution-Based [610]. Resource [1377, 288]. Resource-Bound [288]. Restatement [1755]. Restricted [1659, 188, 1881, 2]. Restricting [1663]. Restructuring [501, 76]. Resultant [312]. Results [1869, 1891, 964, 1614, 1760, 926, 1332, 860, 1292, 1105]. Retraining [422]. Retrieval [899, 912]. Return [1869]. Reusable [488, 268, 1627, 380]. Reuse [162, 1576, 1527, 1578, 412, 1579, 161, 378, 489]. Reusing [1742, 1480, 389]. réutilisable [644]. Reversal [1346]. Reverse [1511, 966].

Reversibility [1176]. Reversible [1958]. Reversing [46]. Revision [1813, 1531, 1313]. Revisited [698, 627, 1884, 1868, 892]. Rewrite [986, 1279]. Rewriter [985]. Rewriting [204, 477, 1037, 605, 594, 559, 1868, 1049, 1196, 558, 1280, 1281, 1291, 1286, 6, 7, 1111]. RHODOS [1800]. Right [437]. Rigidity [1907, 1840]. Ring [798]. Rings [476, 91]. RISC [1361]. Robin [1843]. Robot [436, 1451]. Robotic [434]. Robots [435, 1931]. Robust [867, 1517]. Role [1478, 824, 813, 623, 948]. Roles [1756]. ROO [612]. Roots [1524]. Rough [455]. Round [1843]. Routing [1441, 544, 1453, 1852, 907, 195, 91, 485, 18]. RRL [605, 604]. RUE [568]. Rule [1031, 1812, 431, 38, 1196]. Rule-Based [1812, 431]. Rules [839, 586, 1118, 1311]. Run [65]. Run-Time [65]. Runtime [382, 1627]. Russia [1762]. RVF [613].

S [86, 1808]. S4 [135]. sa [685, 1460]. SA/ CM/IM [1460]. Safe [1247, 123, 1091]. Safety [1913, 295]. Sample [1871]. Satisfaction [438, 432]. Satisfiability [872]. Saving [1787]. Scaffolding [638]. Scalability [486]. Scalable [832, 1381, 75, 1519]. Scalar [1633, 57]. Scale [408]. Scales [1561]. Scaling [402]. Scaling-Up [402]. Scattered [899]. Scenarios [447]. Scenes [189]. Scheduler [392]. Scheduling [1634, 1495, 1401, 1399, 522, 1636, 778, 911, 811, 777, 1433, 745, 90, 445, 1404, 1494]. Schema [170, 838, 831, 1741]. Schemas [1742, 1740]. Schemata [1807]. Schematic [877]. schematization [1285]. Scheme [1527, 773, 1251, 122, 66, 1416, 499, 888, 1439, 1902, 19, 1290]. Schemes [1515, 648, 1863]. School [625]. Science [626, 1596, 991, 1613, 950, 948, 1597, 1594, 1930, 1925]. sciences [1919]. Scientific [1491, 1643]. Scientists [951]. scope [703].

Scoping [1623]. Screen [355, 356]. SDEs [167]. Sealed [1264]. Search [910, 194, 467, 1845, 1020, 1023, 583, 132, 1663, 588, 1143, 1140]. Searching [1147, 923, 1841, 1715, 1392, 1870, 1713, 1856]. Second [1810, 820, 657, 668, 4]. second-order [4]. SECReTS [1499]. Secure [1785, 1806, 1807]. Security [1789, 1795, 1794, 1807, 1808]. Segment [989]. Segments [1892, 1874]. Selected [182, 953, 1138]. Selecting [437]. Selection [1036, 511, 1853, 736, 1916]. Selective [1880]. Selector [1631]. Self [267, 1421, 696, 349, 1845, 1945]. Self-Adjusting [349, 1845]. Self-Improving [696]. Self-Organizing [1421, 1945]. Self-Study [267]. Selfspell [363]. Semantic [587, 223, 1117, 578, 1127, 1403, 1784, 1212, 741, 1741]. Semantically [601]. Semantics [1241, 1040, 1039, 109, 1827, 110, 1626, 117, 118, 528, 1825, 120, 1130, 1769, 1761, 1319, 171, 1038, 1776, 1828, 1172, 281, 848, 1124, 738, 1002, 284, 751, 994, 1385, 752, 1234, 1217, 210, 1320, 498, 1269, 1185, 1113, 739, 1325, 997, 1012, 1190, 1191, 227, 221, 753, 743, 1073, 1078, 846, 740, 1301, 1284, 706, 1283, 709, 1294, 1088, 1100, 1101]. Semi [963, 1776, 1186]. Semi-Commutations [963, 1186]. Semi-Strict [1776]. Semialgebraic [1147]. Semicomputable [763]. Semigroups [36, 1187]. Semiring [962]. Sense [398, 251]. Sensible [1825]. Sensing [1931]. Sensitive [960, 1894]. Sensitivity [1957]. Sensor [1565]. Separating [1793]. Separation [1652, 968]. Separators [1841, 17]. SEPIA [1065]. Sequence [925]. Sequences [1390, 1423]. Sequencing [683]. Sequent [140, 889, 588]. Sequential [1516, 493, 1907, 546, 884, 885, 1308, 1966]. Sequential/Parallel [1966]. Sequentiality [594]. Serial [1723, 1360, 1376]. Serialisation [545, 1297]. Series [919, 962, 574, 444]. Series-Parallel [919].

Servers [89]. Service [781]. Session [1607, 1580, 1603, 1576, 1590, 1612, 1588, 1572, 1586, 1600, 1582, 1598, 1594]. Sessions [785]. Set [1788, 581, 1730, 1893, 925, 39, 518, 874, 394, 1184, 1508, 388, 562, 1781, 98, 1085, 1937]. Set-Associativity [518]. Set-Term [1788]. Sets [1147, 1897, 967, 1880, 232, 958, 227, 1047, 190, 961, 455, 1033, 1963, 763, 1195, 912, 14]. Sets-of-Support [1033]. Setting [576]. Shallow [1129]. Shape [1955]. Shaped [1859]. Shared [106, 1426, 1409, 193, 1379, 518, 1380, 1521, 1522, 1344, 491, 1381, 1035, 75]. Shared-Memory [106]. Sharing [1902, 1286]. Sheaf [740]. Shell [252, 250, 686, 176, 370, 369]. Shells [694]. Sherlock [349]. Shift [1351, 1118, 323]. Shifting [950]. Short [27, 713]. Shortening [1025]. shortest [20, 22]. Should [669]. Shrinking [500]. Shuffle [1169, 1335]. Shuffle-Exchange [1335]. SHUNYATA [602]. SICStus [1066]. Sided [1856]. Sigma [1609]. Signal [444, 1144, 1939]. Signals [1356]. Silent [1226]. SIMCDL [546]. SIMD [84, 482, 124, 1350, 1541]. Similarity [925, 1741]. Simple [672, 1401, 316, 216, 1882, 222, 133, 1663, 1190, 766, 54]. Simplex [990, 1428]. Simplex-Method [1428]. Simplifications [1680]. Simplified [1307, 914]. Simplifying [1739]. Simulating [546]. Simulation [1432, 823, 105, 1410, 1443, 350, 639, 386, 1232, 464, 1423]. Simulations [193, 289]. Simulator [1406]. Simultaneous [228]. Since [885]. Single [1023]. Situation [342]. Six [1609]. Size [913, 1617, 1883, 1730, 1954, 1853, 958, 1255]. Skeletal [944, 426]. Skeleton [731]. Slicing [1260]. Small [1167, 1617, 1685, 1687, 1883, 1177, 1867, 1908, 900, 1154]. Snapshots [315]. Social [1464, 1465]. Socio [627]. Socio-Technical [627]. SODA [346]. Software [461, 1486, 488, 1337, 1478, 695, 330, 1461, 1615, 1419, 1601, 1577, 1609, 1483, 162, 1576, 1599, 1510, 1466, 1595, 1480, 160, 1574, 1584, 1572, 1589, 1573, 1606, 1487, 521, 1812, 1592, 1571, 346, 1635, 1596, 1570, 1472, 1614, 1605, 1610, 1604, 1464, 1591, 1581, 1587, 379, 1387, 410, 365, 419, 1578, 177, 1608, 1598, 150, 1602, 617, 1905, 412, 1467, 156, 1579, 1465, 161, 1593, 1597, 389, 735, 1400, 97, 374, 543, 1361, 1594, 1925, 1928]. Software-Monitor [1337]. SOL [151]. Solid [1856]. Solution [1623, 516, 1354, 1353, 1448, 263, 389, 1645]. Solvability [29]. Solve [1436, 1841, 938, 14]. Solver [1437]. Solvers [313]. Solving [1899, 438, 252, 730, 1424, 432, 253, 1366, 894, 1363]. Some [1649, 1337, 1158, 1854, 592, 1814, 1690, 1699, 802, 1090, 1841, 285, 817, 210, 1118, 1211, 901, 1917, 143, 1173, 1848, 11]. Sometimes [1211]. Sorted [564, 596, 563, 594, 1048, 879, 1324, 562, 1280]. Sorting [1390, 1174, 933, 1352, 1886, 1916, 934, 1885]. Sorts [596]. Soundness [1012]. Source [375, 1792, 378]. Sources [1651, 727]. Space [1208, 956, 120, 1481, 1547, 231, 475, 933, 435, 322, 1787, 1189, 1663, 372, 682]. Space-Bounded [956]. Space-Efficient [475]. Space-Saving [1787]. Spaces [932, 1020, 1193]. SPADE [1469]. Spanners [1849, 903]. Spanning [1394, 192]. Sparse [1897, 1862, 1895, 1365, 1354, 903, 974, 1371, 1407]. Sparse-Grid-Preconditioning [1354]. Spatial [1644, 1550, 1559, 1553, 1560, 1562, 1544, 184, 185, 186, 1561, 1552, 191]. Spatial-Linguistic [1562]. Spatio [1565]. Spatio-Temporal [1565]. Special [586]. Special-Relation [586]. Specific [1772, 1210, 485]. Specification [837, 1510, 1466, 1746, 1904, 124, 1794, 285, 288, 876, 164, 878, 1888, 1272, 1474, 296, 141, 1793, 1223, 299, 1097]. Specification-Based [1746]. Specifications [273, 520, 1241, 1218, 839, 1631, 255, 157, 1224, 547, 875, 292, 386, 1622, 1007, 258, 1278, 1298]. Specifying [790, 411, 1887]. spectral [1942]. Speed [770, 1376, 781]. Speeding [1067, 1195]. Speeding-up [1067, 1195]. Speedup [391, 1412]. SPELL [1483]. Spelling [363]. Spes [1050]. sphere [1145]. Spheres [931]. SPIKE [1055]. SPL [1490]. Splicing [1962, 1664]. Split [1770]. Splitting [301]. spread [18]. SQL [834]. Squares [977, 957]. Squaring [516]. ST [707, 1406]. ST-idea [707]. Stability [1365]. Stabilization [1941]. Stable [865, 1776, 869, 1093, 313, 1428]. Stably [933]. Stacks [1844]. Stance [950]. Standard [1510, 597]. Standardized [337]. Stanford [521]. Star [1394]. State [276, 1208, 278, 690, 1224, 547, 1177, 891]. State-Space [1208]. Stategies [544]. Statement [1385]. Statements [552]. Static [1524, 503, 1399, 1891, 1418]. Station [372]. Stationary [1358]. Statistical [956, 1740, 1879, 1349]. Statistics [364, 689, 444]. Steiner [10]. Step [174, 1386, 1038, 1226, 804]. STI [685]. Stochastic [1936, 715, 1940]. Stock [443]. Storage [71, 508, 830]. Stories [1805]. Storing [368]. Story [322]. STRAda [393]. Strategic [1501, 396]. stratégie [674]. Strategies [1646, 341, 155, 207, 195, 668, 357, 688, 670, 421, 324, 1293]. Strategy [770, 553, 1841, 355, 1411, 1143, 1195, 1304]. Stream [1152]. Streams [775]. Strengthening [1663]. Stress [106]. Strict [306, 1776, 437]. Strictness [854, 1245]. Strikes [975]. String [1726, 1725, 1687, 1688, 1730, 1731, 1914, 1713, 1049]. String-Graph-Generating [1687, 1688]. Strings [1174, 1883, 1729]. Strive [553]. Strong [1652, 1165, 252, 1328, 845]. Strong-Problem-Solving-Shell [252]. Strongly [232, 313]. Structural [1218, 1854, 715, 432, 421, 1896]. Structurally [1790]. Structure [520, 1159, 1720, 1331, 765, 993, 1418].

Structured [1046, 343, 1344]. Structures [1120, 235, 536, 524, 861, 182, 1239, 239, 1866, 515, 322, 191, 1115, 85, 7]. Structuring [163, 335]. Stub [1801]. Student [673, 677, 679, 358, 680, 669, 670, 649, 682]. STUDIA [633]. Study [267, 571, 799, 111, 1442, 1904, 1219, 1133, 833]. Studying [766]. Style [342, 126, 1270, 895, 656, 41]. Styles [326]. Subdivisions [928]. Subinstruction [1414]. Subject [335, 355]. Sublogics [849]. Submarine [425]. Submodular [1840]. Suboptimal [983]. Subscripted [1633]. Subsequence [1883, 1729, 1716, 1718]. Subset [1659, 570]. Subsets [754]. Substitution [228]. Substitutions [1165, 1283, 1296]. Substructure [1722]. Subsumption [1242, 1783, 1117]. Subsumption-Oriented [1242]. Subsystem [1341]. Subtrack [1595]. Subtraction [658]. Subtypes [225, 555]. Success [1609, 702]. sufficient [1281]. Suffix [1712]. Suitable [1471]. Sum [344, 574]. SUMIT [362]. Summary [1269, 1946]. SUN [792]. SUPER [1754]. SUPER-Visual [1754]. Supercomputer [526]. Superfiniteness [1900]. Superperfection [921]. Superpolynomial [1895]. Superposable [1192]. Superposition [584, 1036]. Superscalar [1439]. Supervisory [646]. Support [786, 1053, 771, 779, 382, 767, 778, 768, 506, 787, 1067, 455, 1033, 388, 446, 724, 142]. Supported [1479]. Supporting [448, 1464, 222, 429, 161]. SUPRENUM [483, 1355]. Surface [832]. Survey [275, 714, 716, 717]. SWG [388]. Switch [1385]. Switched [774]. switches [1928]. Switching [1413, 773, 1351]. Symbol [1623]. Symbolic [1554, 1336, 1552, 97, 642]. symbolique [681]. Symmetric [1447, 1367, 1880, 1360, 313, 1368]. Symmetrical [885]. Symmetries [181, 571]. Symmetry [1877, 16]. SYMPATIX [482].

Synchronisation [1520]. Synchronization [774, 1329, 776, 1388, 1233, 382, 775, 1348, 1772, 1349, 1232, 489, 710]. Synchronized [790, 964]. Synchronous [897, 1234]. Syntactic [1947, 1948, 470, 1305]. Syntax [1618, 131, 1113, 1313, 42]. Syntax-based [1313]. Syntax-Driven [1113]. Synthesis [1321, 573, 187, 293, 156, 1110]. System [1050, 947, 602, 1051, 1337, 1052, 425, 393, 459, 430, 823, 441, 1268, 698, 1499, 785, 621, 599, 699, 1058, 1512, 606, 194, 339, 1061, 803, 382, 1322, 340, 1476, 522, 341, 696, 1267, 343, 1526, 50, 778, 1475, 285, 1517, 406, 652, 246, 1438, 415, 180, 352, 876, 40, 472, 357, 469, 418, 291, 1797, 506, 361, 787, 1067, 149, 616, 423, 1518, 1888, 268, 951, 838, 1921, 617, 168, 726, 1069, 660, 473, 1275, 783, 601, 1411, 1417, 834, 458, 694, 445, 1070, 1800, 671, 613, 543, 642, 1541, 1396, 1087, 1924, 770, 470, 413]. Systematic [1943, 617]. système [633, 678, 659]. systèmes [651]. Systemic [680]. Systems [1221, 273, 1457, 328, 1432, 395, 1054, 112, 37, 827, 448, 1018, 1789, 381, 533, 643, 80, 858, 1844, 646, 627, 403, 516, 986, 179, 278, 820, 720, 593, 264, 400, 1238, 163, 1431, 390, 1799, 1319, 1771, 840, 1574, 721, 1203, 1037, 377, 1202, 748, 93, 282, 832, 283, 284, 154, 722, 148, 39, 811, 1790, 594, 468, 1887, 184, 185, 517, 466, 1804, 732, 288, 431, 353, 1264, 251, 289, 1867, 686, 1348, 290, 491, 320, 661, 687, 292, 1912, 667, 294, 959, 1868, 1006, 173, 295, 719, 589, 1502, 1209, 1602, 1010, 399, 1191, 1905, 421]. Systems [1049, 1743, 1347, 270, 427, 258, 961, 1962, 725, 1366, 1664, 662, 90, 1363, 610, 743, 1519, 416, 729, 734, 38, 1196, 401, 558, 976, 300, 539, 380, 1818, 1925, 1278, 1298, 1942, 711, 1279, 1294, 6, 717, 7, 1110, 1111]. Systolic [1357, 536, 496, 1359, 1358].

t [1367, 1337]. Table [1897]. Tableau [618, 1321]. Tableau-Based [618, 1321]. Tableaux [587]. Tables [1623, 624]. Tactic [590]. Tactic-Based [590]. Tagging [216].

Tail [573, 989]. Tail-Recursive [573]. Tailoring [461]. Tale [772]. Talk [1225, 1219, 1209, 1214, 1277, 1300, 1288]. Tally [1862]. Task [1413, 683, 121, 65, 255, 58, 93, 392, 384, 540, 355, 940, 1404]. Task-Level [940]. Task-Switching [1413]. Tasks [267, 105, 1521, 812, 1433, 661]. TBLL [1082]. TC2000 [509, 1364, 79]. TCB [1792]. TCCS [999]. Teacher [693]. Teaching [329, 1577, 1056, 345, 1585, 1591, 1587, 360, 323, 1575]. Team [1614]. Technical [627, 455, 1933]. Technique [1770, 1249, 1354, 154]. Techniques [742, 200, 1566, 460, 278, 1238, 1256, 338, 501, 186, 196, 129, 809, 500, 1514, 715]. Technology [1610, 938, 465, 1922, 76, 413]. Telecommunication [404]. Telecommunications [772]. Teleconferencing [796]. Template [1720, 1960]. Temporal [576, 438, 992, 1756, 1043, 577, 1213, 304, 1740, 1124, 1565, 1212, 1887, 875, 457, 885, 1308, 136, 1888, 140, 889, 1545]. Ten [1735]. Tense [1130]. Tense/Aspect [1130]. tensor [47]. tentative [583]. Tenter [424]. Term [1788, 985, 1117, 204, 1037, 1286, 355, 1868, 1639, 558, 1111]. Termination [1248, 592, 1817, 1280, 1037, 1279, 495, 1281, 1094, 1111]. Terminological [450, 208, 1311]. Terminology [1114]. Terminology-Intensive [1114]. Terms [1321, 1322, 232, 127, 132, 1047]. Terrain [1557]. Test [1158, 1882, 958, 409, 388]. Testclasses [844]. Testing [1198, 1011, 1631, 340, 1204, 1199, 921, 318, 419, 421, 689, 1800, 446, 981]. Tests [1389]. Texas [1604]. Text [327, 322, 1131, 1727, 452]. Their [118, 1256, 868, 874, 62, 1187, 983, 1073, 6, 1109]. Them [1836]. Theorem [567, 1303, 1513, 611, 618, 1321, 1055, 590, 619, 599, 1032, 1060, 1678, 1694, 1683, 581, 622, 579, 1628, 1062, 612, 585, 1868, 624, 614, 603, 597]. Theorems [607, 229, 1910]. Theoretic [1040, 731, 34, 1827, 1693, 1695, 1690, 1699, 1696, 1900, 874, 1890, 1007, 100]. Theoretical [571, 1725, 1442, 442, 1123]. Theories [554, 242, 202, 591, 844, 993, 1777, 1035, 1303, 1305, 1925, 1086]. Theory [1235, 274, 32, 226, 1022, 1854, 1205, 655, 1391, 1828, 1780, 1549, 1316, 755, 1794, 39, 1048, 1492, 1310, 874, 1563, 1750, 752, 318, 877, 210, 130, 583, 661, 881, 757, 1151, 1325, 1639, 589, 1543, 298, 753, 754, 1762, 1152, 1306, 682, 305, 1282, 18, 1085, 42, 43, 1108, 1138]. Theory-Guided [661]. There [1560]. Thermodynamics [1542]. thesis [1137]. Thinking [262]. Thinning [1398]. Thrashing [68, 501]. Threaded [101, 1522]. Threads [1414]. Three [1865, 1470, 1729, 831, 1966, 928]. Three-Dimensional [1966, 928]. Three-Level [831]. Threshold [1878, 900]. Thue [37, 38]. Tight [12]. Tightly [517]. TIM [295]. TIM/RTTL [295]. Time [272, 275, 273, 1225, 276, 780, 108, 1850, 1413, 460, 381, 858, 1838, 278, 511, 1527, 315, 279, 779, 65, 1910, 1873, 767, 1034, 283, 104, 1353, 1547, 284, 285, 238, 286, 1226, 1640, 307, 287, 1887, 385, 288, 396, 1877, 1926, 379, 1184, 291, 292, 506, 322, 787, 1214, 372, 465, 140, 297, 783, 1400, 1637, 1223, 781, 1450, 444, 380, 73, 25]. Time-Bounded [1413]. Time-Critical [767, 1887, 787]. Time-Dependent [511, 104]. Time-Lapse [315]. Time-Parallel [104, 1353]. Timed [1221, 274, 277, 1222, 282, 290, 294, 293, 296, 298, 1227, 1223, 300]. Timely [1804]. Timestamping [316]. Timing [1000, 1792, 289, 1209]. Timing-Based [289]. Timing-Channel [1792]. TK [123]. Toeplitz [313]. Token [978, 798]. Tokens [1415, 705]. Tolerance [1799, 18]. Tolerant [484, 1851, 1518, 1444]. Tolerating [1329]. Too [1648]. Tool [818, 160, 1406, 155, 317, 1792, 1336, 353, 82, 394, 1064, 189, 176, 1408, 368, 1384, 694, 376]. Toolkit [786, 1271]. Tools [790, 265, 249, 1476, 630, 1114, 693, 1735, 262, 470, 1508, 891, 172, 367, 1937]. Toolset [388, 1298]. Top [1619]. Top-Down [1619]. Topics [182, 1154, 1138]. TOPMUSS [470]. TOPMUSS-System [470]. Topological [1555, 1558]. Topologically [1858]. Topology [761, 1422]. Topsy [479]. Torus [907]. Total [1650, 1876]. Touch [356]. TOVE [398]. tracage [674]. Trace [1218, 1262, 1002, 94, 82, 1231, 1196]. Trace-Rewriting [1196]. Traces [236, 1001, 1170, 743, 716]. Tracing [187]. Track [1881]. Tracking [469, 417]. Trade [236]. Trade-Offs [236]. Tradeoffs [1584]. Traditional [1795]. Traffic [777]. Trail [1812]. Training [1601, 1608, 1602, 689]. Transaction [1806, 491]. Transactions [1773]. Transducers [1227]. Transductions [1162]. Transfer [1610, 671]. Transformation [1050, 393, 114, 1419, 1395, 1616, 755, 70, 582, 133, 494, 1418]. transformational [1094]. Transformations [1249, 1248, 166, 142, 8]. Transforming [1249, 1830, 1751, 946]. Transforms [537]. Transition [1221, 533, 1319, 282, 1912]. Transitions [1003, 1075]. Transitive [856, 1079, 1089, 1109]. Transivity [624]. Translating [504]. Translation [1130, 728, 1618]. translations [1106]. Transmission [798, 794, 141]. Transparent [732]. Transport [770, 1635, 781]. Transport-Triggered [1635]. Transports [1885]. Transputer [1337, 95, 888, 1446]. Transputer-Based [888]. Transputers [1429]. Traversal [561]. Treating [126]. Treatment [430, 1568]. Tree [643, 909, 194, 922, 1847, 1723, 1724, 1183, 1425, 966, 1739, 1959, 971, 1140, 1389, 3, 4, 10, 1107, 1143]. tree-decomposable [4]. Tree-Width [922]. Trees [1455, 910, 913, 1005, 1616, 866, 1674, 911, 303, 921, 1845, 1886, 189, 1870, 1787, 1266, 890, 1712, 192, 1287, 5, 913]. Treewidth [1848, 3]. Trellis [1192]. Trends [247, 810]. Treshold [1171]. Triangle [1557]. Triangle-Based [1557]. Triangular [1363]. Triangulating [918]. Triangulation [1894, 1189]. Triangulations [1894]. Triconnected [982]. Tridiagonal [1360]. Trig [1053]. Triggered [1635]. Triples [283]. Tropical [962]. Troubleshooting [402]. True [236]. Truth [1897]. Truth-Table [1897]. Tuning [433]. Tuples [1763]. Turing [867, 1954]. Tuteurs [664]. Tutor [673, 666, 696, 342, 636, 663, 702]. Tutorial [691, 1509, 1817, 340, 684, 214]. tutoriel [633, 659]. tutorielle [645]. Tutoring [625, 695, 643, 646, 339, 341, 696, 652, 637, 352, 353, 686, 357, 687, 667, 362, 671, 642]. Tutors [630, 325]. TV [324]. Two [1865, 1374, 975, 223, 1018, 1951, 86, 1130, 1256, 1038, 923, 1124, 1729, 1954, 518, 197, 1360, 1190, 951, 772, 1246, 31, 928]. Two- [928]. Two-Dimensional [1951, 923, 1954]. Two-Level [1018, 1130, 1124, 518, 1246]. Type [229, 1322, 1677, 755, 1790, 1048, 1536, 752, 882, 765, 757, 552, 753, 754, 1277, 44, 135]. Type-Level [1790]. Typechecking [1524]. Typed [228, 873, 225]. Types [527, 875, 877, 219, 1772, 224, 894, 1080, 1084, 43, 1097, 849]. Typing [1903, 42].

UCG [1267]. UCG-E [1267]. UEPS [1810]. Un-Programming [1511]. Unambiguous [1160]. Uncertainty [721, 813, 816, 457, 598]. Undecidability [1779]. Undecidable [962]. Undergraduate [1606, 1604]. Underground [447]. Understand [1544, 1064]. Understanding [472, 688, 399, 1465]. Unexpected [462]. Unfold [1249, 1248]. Unfold/Fold [1248]. Unidimensional [1182]. Unification [32, 554, 36, 557, 33, 34, 564, 1119, 556, 1008, 563, 1253, 1048, 555, 35, 1035, 1047, 562]. Unification-Based [1119]. Unified [1759].

Uniform [1293, 561, 1037, 190, 1389].

Unifying [1660, 877, 739]. Unimodularity [1405]. Union [554]. Unions [1650]. Uniprocessor [1514]. Unique [277]. Uniquely [1868]. Uniqueness [1201, 1243]. Unit [385, 1429]. Universal [193, 1812, 1177, 199, 1182, 1966]. University [1576, 1596, 1604]. UNIX [776, 793, 663]. UNIX/X [793]. UNIX/ X-Windows [793]. Unlimp [1243]. Unsatisfiability [1171]. Unsolvable [1703, 1696]. Unstable [1906]. Until [885]. Untyped [1825]. Update [882, 912]. Updates [1750]. Updating [1771]. upon [1454]. Upper [232, 22]. Urban [1548]. Urgent [277]. usager [644]. Use [1031, 1337, 395, 632, 1744, 1483, 561, 1606, 684, 1, 1028, 1956, 356, 574, 1429]. Used [436]. User [245, 700, 1563, 462, 446]. Uses [358]. Using [1770, 426, 727, 420, 114, 1719, 1222, 375, 1495, 502, 1553, 1443, 1420, 1476, 377, 1616, 696, 1236, 1034, 573, 283, 1353, 676, 925, 468, 1438, 67, 875, 371, 938, 1958, 1359, 1376, 257, 363, 364, 419, 295, 1552, 1244, 465, 1624, 1959, 14, 490, 1490, 1564, 1411, 834, 562, 1368, 734, 1427, 370, 512, 449, 413, 254, 1794]. USSR [583]. utilisant [674].

Vademecum [1813]. Validating [169, 292]. Validation [1487, 170, 1591, 387]. Value [1869, 220]. Value-Passing [1869]. Valued [1021, 618, 863, 1323, 1098]. Vanilla [1825]. Variable [587, 553, 1098]. Variables [1633, 1253]. Variant [1428]. Variants [1337, 1090]. Varieties [36]. Variety [1069]. Various [1655]. Vector [1546, 1438, 1439, 710]. Vectorial [1376]. vehicles [1938]. Verboseness [853]. Verifiability [1657]. Verification [420, 116, 1202, 1796, 283, 1774, 1228, 496, 1591, 1219, 291, 295, 616, 1905, 891, 613]. Verified [1046, 1232, 617]. Verifier [956]. Verify [1222]. Verifying [273, 1209].

Version [203]. Versioning [829]. Versions [836]. Versus [1657, 1477, 1088, 713]. vertex [16]. Vertices [1864]. Very [1376]. VHDL [1508]. VI [1070]. via [1030, 1008, 1202, 1659, 311, 1106]. Video [1611, 798, 768]. Videophone [796]. View [245, 985, 174, 1799, 563, 1691, 1700, 1554, 1750, 950, 1300, 82, 759, 1593]. Viewpoint [342, 931]. Viewpoints [1123]. Views [1759, 1786, 831, 1765]. Vigilance [355]. Virtual [1426, 511, 1380, 1344, 1928]. virtual-circuit [1928]. Virtually [1522]. Visibility [1859, 931]. Visible [638]. Vision [1950, 809, 416, 1134]. Visions [801]. Visual [1944, 1748, 427, 1754]. Visualization [1457, 328, 1338, 1341, 1336, 82, 1731]. Vital [1209]. VLDC’s [1723]. VLSI [504, 1428, 1437]. VLSI-Hardware-Realization [1428]. Voronoi [932, 13]. Vous [395]. VPL [519]. vs [1734, 1253, 197].

w.r.t [1301]. Wadler [47]. Walks [1183]. Wall [1791]. Wallet [1810]. WAM [1080, 111, 1251, 122, 124, 1064, 142]. WAM-Based [1251]. WAM-Support [142]. Waste [430]. Waste-Water [430]. Watching [626]. Water [430]. Way [1898, 1573, 1079, 1917]. Wayfinding [1564]. Weak [1037, 1310, 1200, 1155]. Weakest [1212]. Weakly [1880]. Weakly-P- [1880]. Weight [983]. Weighted [1168, 1839]. Well [1776, 866, 1325, 1012, 1190, 539, 1301]. Well-Founded [1776, 1012, 1190, 1301]. Well-Quasi-Ordered [866]. WEST [698]. where [1929]. Which [329, 623, 357, 1945]. While [918, 931]. Who [1461, 1836]. Whole [344]. Wide [423]. Widening [1254]. Widening/ [1254]. Width [922]. Winch [639]. Windows [793]. Wire [1587]. Wiring [1875]. Within [1454, 1513, 1031, 425, 1596, 369]. Without [1307]. WIZ [414]. Word [1436, 1125, 1122, 31, 1104]. Words [39, 1176, 1121, 1076, 713]. Workbench [1068, 939]. Working [390]. Workpiece [418]. Workshop [800]. Workstation [1354, 767]. Workstations [793]. World [944, 844, 213, 345, 1919, 938, 1123]. Worlds [146]. Wormhole [485]. Worst [1885]. Worthwhile [937]. Wrap [800]. Wrap-Up [800]. Write [1915]. Writers [336]. Written [133].

X [794]. X-Movie [794]. XFSM [547].

Xpnet [608]. XPRAM [548].

Year [1585, 1919, 1922]. Years [1735].

Young [336].

Z [187]. Zero [956]. Zero-Knowledge [956]. Zone [424].

References

Herman:1992:UPG

[1] I. Herman. The use of projective geometry in computer graphics. Lecture Notes in Computer Science, 564:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burckert:1991:RPL

[2] H. J. (Hans-Jurgen) Burckert. A resolution principle for a logic with restricted quantifiers, volume 568 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 3-540-55034-8 (Berlin), 0-387-55034-8 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 116 pp. LCCN QA76.63 .B8 1991.

Bodlaender:1992:ATP

[3] H. L. Bodlaender, J. Gilbert, H. Hafsteinsson, and T. Kloks. Approximating treewidth, pathwidth and minimum elimination tree height. Lecture Notes in Computer Science, 570:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courcelle:1992:MSO

[4] B. Courcelle and M. Mosbah. Monadic second-order evaluations on treedecomposable graphs. Lecture Notes in Computer Science, 570:13–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heckmann:1992:OEC

[5] R. Heckmann, R. Klasing, B. Monien, and W. Unger. Optimal embedding of complete binary trees into lines and grids. Lecture Notes in Computer Science, 570:25–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okada:1992:GRS

[6] Y. Okada and M. Hayashi. Graph rewriting systems and their application to network reliability analysis. Lecture Notes in Computer Science, 570:36–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schurr:1992:NCS

[7] A. Schurr and A. Zundorf. Nondeterministic control structures for graph rewriting systems. Lecture Notes in Computer Science, 570:48–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andries:1992:LGG

[8] M. Andries and J. Paredaens. A language for generic graph-transformations. Lecture Notes in Computer Science, 570:63–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freund:1992:AEP

[9] R. Freund and B. Haberstroh. Attributed elementary programmed graph grammars. Lecture Notes in Computer Science, 570:75–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ihler:1992:CAC

[10] E. Ihler. The complexity of approximating the class Steiner tree problem. Lecture Notes in Computer Science, 570:85–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lonc:1992:CSC

[11] Z. Lonc. On the complexity of some chain and antichain partition problems. Lecture Notes in Computer Science, 570:97–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czyzowicz:1992:TBR

[12] J. Czyzowicz, E. Rivera-Campo, N. Santoro, J. Urrutia, and J. Zaks. Tight bounds for the rectangular art gallery problem. Lecture Notes in Computer Science, 570:105–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guibas:1992:VDM

[13] L. Guibas, J. Mitchell, and T. Roos. Voronoi diagrams of moving points in the plane. Lecture Notes in Computer Science, 570:113–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shoudai:1992:UMI

[14] T. Shoudai and S. Miyano. Using maximal independent sets to solve problems in parallel. Lecture Notes in Computer Science, 570:126–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kedem:1992:FPA

[15] Z. Kedem, K. Palem, G. Pantziou, P. Spirakis, and C. Zaroliagis. Fast parallel algorithms for coloring random graphs. Lecture Notes in Computer Science, 570:135–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiang:1992:OVO

[16] X. Y. Jiang and H. Bunke. Optimal vertex ordering of a graph and its application to symmetry detection. Lecture Notes in Computer Science, 570: 148–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sykora:1992:ESG

[17] O. Sykora and I. Vrto. Edge separators for graphs of bounded genus with applications. Lecture Notes in Computer Science, 570:159–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Du:1992:LDI

[18] D.-Z. Du, Y.-D. Lyuu, and D.-F. Hsu. Line digraph iterations and the spread concept — with application to graph theory, fault tolerance, and routing. Lecture Notes in Computer Science, 570:169–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kucera:1992:GES

[19] L. Kucera. A generalized encryption scheme based on random graphs. Lecture Notes in Computer Science, 570: 180–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feuerstein:1992:DAS

[20] E. Feuerstein and A. Marchetti-Spaccamela. Dynamic algorithms for shortest paths in planar graphs. Lecture Notes in Computer Science, 570: 187–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stewart:1992:CPL

[21] I. Stewart. Complete problems for logspace involving lexicographic first paths in graphs. Lecture Notes in Computer Science, 570:198–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takaoka:1992:NUB

[22] T. Takaoka. A new upper bound on the complexity of the all pairs shortest path problem. Lecture Notes in Computer Science, 570:209–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sykora:1992:CNH

[23] O. Sykora and I. Vrto. On the crossing number of the hypercube and the cube connected cycles. Lecture Notes in Computer Science, 570:214–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damaschke:1992:LAI

[24] P. Damaschke. Logic arrays for interval indicator functions. Lecture Notes in Computer Science, 570:219–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stohr:1992:BTB

[25] E. Stohr. On the broadcast time of the Butterfly network. Lecture Notes in Computer Science, 570:226–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodlaender:1992:DC

[26] H. L. Bodlaender. On disjoint cycles. Lecture Notes in Computer Science, 570:230–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandstadt:1992:SDC

[27] A. Brandstadt. Short disjoint cycles in cubic bridgeless graphs. Lecture Notes in Computer Science, 570:239– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makanin:1992:IEF

[28] G. S. Makanin. Investigations on equations in a free group. Lecture Notes in Computer Science, 572:1–??, 1992.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koscielski:1992:AMA

[29] A. Koscielski. An analysis of Makanin’s algorithm deciding solvability of equations in free groups. Lecture Notes in Computer Science, 572:12–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdulrab:1992:IMA

[30] H. Abdulrab. Implementation of Makanin’s algorithm. Lecture Notes in Computer Science, 572:61–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulz:1992:MAW

[31] K. U. Schulz. Makanin’s algorithm for word equations — two improvements and a generalization. Lecture Notes in Computer Science, 572:85–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baader:1992:UT

[32] F. Baader. Unification theory. Lecture Notes in Computer Science, 572: 151–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bockmayr:1992:ALA

[33] A. Bockmayr. Algebraic and logic aspects of unification. Lecture Notes in Computer Science, 572:171–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bockmayr:1992:MTA

[34] A. Bockmayr. Model-theoretic aspects of unification. Lecture Notes in Computer Science, 572:181–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohsuga:1992:CEU

[35] A. Ohsuga and K. Sakai. Complete equational unification based on an extension of the Knuth–Bendix completion procedure. Lecture Notes in Computer Science, 572:197–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baader:1992:UVC

[36] F. Baader. Unification in varieties of completely regular semigroups. Lecture Notes in Computer Science, 572:210– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Book:1992:NCT

[37] R. V. Book. A note on confluent Thue systems. Lecture Notes in Computer Science, 572:231–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wrathall:1992:COR

[38] C. Wrathall. Confluence of onerule Thue systems. Lecture Notes in Computer Science, 572:237–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karhumaki:1992:SEF

[39] J. Karhumaki. Systems of equations over a finite set of words and automata theory (extended abstract). Lecture Notes in Computer Science, 572:247– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makanina:1992:NSD

[40] A. G. Makanina. New system of defining relations of the braid group. Lecture Notes in Computer Science, 572: 250–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Danvy:1992:BDS

[41] O. Danvy. Back to direct style. Lecture Notes in Computer Science, 582:183– 196, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henglein:1992:DTS

[42] F. Henglein. Dynamic typing: syntax and proof theory. Lecture Notes in Computer Science, 582:197–230, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1992:TQT

[43] M. P. Jones. A theory of qualified types. Lecture Notes in Computer Science, 582:231–256, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larcheveque:1992:ITP

[44] J. M. Larcheveque. Interprocedural type propagation for object-oriented languages. Lecture Notes in Computer Science, 582:257–282, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hankin:1992:AFP

[45] C. Hankin and S. Hunt. Approximate fixed points in abstract interpretation. Lecture Notes in Computer Science, 582:283–306, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hughes:1992:RAI

[46] J. Hughes and J. Launchbury. Reversing abstract interpretations. Lecture Notes in Computer Science, 582:307– 326, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielson:1992:TPW

[47] F. Nielson and H. R. Nielson. The tensor product in Wadler’s analysis of lists. Lecture Notes in Computer Science, 582:327–354, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheese:1992:PEP

[48] A. Cheese. Parallel execution of parlog. Lecture Notes in Computer Science, 586:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agha:1992:DEA

[49] G. Agha, C. Houck, and R. Panwar. Distributed execution of actor programs. Lecture Notes in Computer Science, 589:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hiranandani:1992:OFD

[50] S. Hiranandani, K. Kennedy, C. Koelbel, U. Kremer, and C.-W. Tseng. An overview of the Fortran D programming system. Lecture Notes in Computer Science, 589:18–34, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Werth:1992:IFP

[51] J. Werth, J. Browne, S. Sobek, T. Lee, P. Newton, and R. Jain. The interaction of the formal and the practical in parallel programming environment development: CODE. Lecture Notes in Computer Science, 589:35–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scales:1992:HCJ

[52] D. Scales, M. Rinard, M. Lam, and J. Anderson. Hierarchical concurrency in Jade. Lecture Notes in Computer Science, 589:50–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eigenmann:1992:EAP

[53] R. Eigenmann, J. Hoeflinger, Z. Li, and D. Padua. Experience in the automatic parallelization of four Perfect-Benchmark programs. Lecture Notes in Computer Science, 589:65–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1992:PSP

[54] J. Lee, C. Lin, and L. Snyder. Programming SIMPLE for parallel portability. Lecture Notes in Computer Science, 589:84–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ariola:1992:CI

[55] Z. Ariola and Arvind. Compilation of id. Lecture Notes in Computer Science, 589:99–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johnson:1992:ERD

[56] R. Johnson, W. Li, and K. Pingali. An executable representation of distance and direction. Lecture Notes in Computer Science, 589:122–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tjiang:1992:ISO

[57] S. Tjiang, M. Wolf, M. Lam, K. Pieper, and J. Hennessy. Integrating scalar optimization and parallelization. Lecture Notes in Computer Science, 589:137– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Girkar:1992:ODC

[58] M. Girkar and C. Polychronopoulos. Optimization of data/control conditions in task graphs. Lecture Notes in Computer Science, 589:152–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Callahan:1992:RPB

[59] D. Callahan. Recognizing and parallelizing bounded recurrences. Lecture Notes in Computer Science, 589:169– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1992:CFH

[60] C.-H. Huang and P. Sadayappan. Communication-free hyperplane partitioning of nested loops. Lecture Notes in Computer Science, 589:186– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1992:PLI

[61] L.-C. Lu and M. Chen. Parallelizing loops with indirect array references or pointers. Lecture Notes in Computer Science, 589:201–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nicolau:1992:RAR

[62] A. Nicolau, R. Potasman, and H. Wang. Register allocation, renaming and their impact on fine-grain parallelism. Lecture Notes in Computer Science, 589:218–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rau:1992:DFD

[63] B. R. Rau. Data flow and dependence analysis for instruction level parallelism. Lecture Notes in Computer Science, 589:236–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kallis:1992:ECF

[64] A. Kallis and D. Klappholz. Extending conventional flow analysis to deal with array references. Lecture Notes in Computer Science, 589:251–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Furnari:1992:RTM

[65] M. Furnari and C. Polychronopoulos. Run-time management of Lisp parallelism and the hierarchical task graph program representation. Lecture Notes in Computer Science, 589:266– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kasahara:1992:MGP

[66] H. Kasahara, H. Honda, A. Mogi, A. Ogura, K. Fujiwara, and S. Narita. A multi-grain parallelizing compilation scheme for OSCAR. Lecture Notes in Computer Science, 589:283–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Labarta:1992:BLP

[67] J. Labarta, E. Ayguade, J. Torres, M. Valero, and J. M. Llaberia. Balanced loop partitioning using GTS. Lecture Notes in Computer Science, 589:298–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fang:1992:IPA

[68] J. Fang and M. Lu. An iteration partition approach for cache or local memory thrashing on parallel processing. Lecture Notes in Computer Science, 589:313–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrante:1992:EEC

[69] J. Ferrante, V. Sarkar, and W. Thrash. On estimating and enhancing cache effectiveness. Lecture Notes in Computer Science, 589:328–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ju:1992:RCC

[70] Y.-J. Ju and H. Dietz. Reduction of cache coherence overhead by compiler data layout and loop transformation. Lecture Notes in Computer Science, 589:344–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gao:1992:LSO

[71] G. Gao and Q. Ning. Loop storage optimization for dataflow machines. Lecture Notes in Computer Science, 589: 359–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hardon:1992:OPP

[72] R. Hardon and S. Pinter. Optimal partitioning of programs for data flow machines. Lecture Notes in Computer Science, 589:374–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carriero:1992:FAC

[73] N. Carriero and D. Gelernter. A foundation for advanced compile-time analysis of Linda programs. Lecture Notes in Computer Science, 589:389– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srinivasan:1992:APE

[74] H. Srinivasan and M. Wolfe. Analyzing programs with explicit parallelism. Lecture Notes in Computer Science, 589:405–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thapar:1992:SCC

[75] M. Thapar, B. A. Delagi, and M. J. Flynn. Scalable cache coherence for shared memory multiprocessors. Lecture Notes in Computer Science, 591: 1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolfe:1992:NPR

[76] M. Wolfe. New program restructuring technology. Lecture Notes in Computer Science, 591:13–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lewis:1992:DPP

[77] T. G. Lewis, R. Currey, and J. Liu. Data parallel program design. Lecture Notes in Computer Science, 591:37–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caerts:1992:PHL

[78] Ch. Caerts, R. Lauwereins, and J. A. Peperstraete. A powerful high-level debugger for parallel programs. Lecture Notes in Computer Science, 591:54–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brooks:1992:PPP

[79] E. D. Brooks, B. C. Gorda, and K. H. Warren. The PCP/PFP programming models on the BBN TC2000. Lecture Notes in Computer Science, 591:65–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chapman:1992:KBP

[80] B. M. Chapman and H. M. Herbeck. Knowledge-based parallelization for distributed memory systems. Lecture Notes in Computer Science, 591: 77–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerndt:1992:PMM

[81] M. Gerndt and H. Moritsch. Parallelization for multiprocessors with memory hierarchies. Lecture Notes in Computer Science, 591:89–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malony:1992:TVT

[82] A. D. Malony, D. H. Hammerslag, and D. J. Jablonowski. Trace view: A trace visualization tool. Lecture Notes in Computer Science, 591:102–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weininger:1992:PDP

[83] A. Weininger, Th. Schnekenburger, and M. Friedrich. Parallel and distributed programming with ParMod-C. Lecture Notes in Computer Science, 591:115– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brezany:1992:CGD

[84] P. Brezany and V. Sipkova. Code generation for a data parallel SIMD language. Lecture Notes in Computer Science, 591:127–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolfe:1992:DSO

[85] M. Wolfe and H. Srinivasan. Data structures for optimizing programs with explicit parallelism. Lecture Notes in Computer Science, 591:139– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diestelkamp:1992:MLE

[86] W. Diestelkamp, H. Bi, and A. Bottcher. MODULA-S: A language to exploit two dimensional parallelism. Lecture Notes in Computer Science, 591: 157–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Philippsen:1992:MC

[87] M. Philippsen and W. F. Tichy. MODULA-2 and its compilation. Lecture Notes in Computer Science, 591: 169–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merlin:1992:AFA

[88] J. H. Merlin. ADAPTing Fortran 90 array programs for distributed memory architectures. Lecture Notes in Computer Science, 591:184–200, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reymond:1992:EMP

[89] M. H. Reymond. Evolution of massive parallel compute servers from a research object to a production pool. Lecture Notes in Computer Science, 591: 201–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tripathi:1992:PSM

[90] S. K. Tripathi, G. Serazzi, and D. Ghosal. Processor scheduling in multiprocessor systems. Lecture Notes in Computer Science, 591:208– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makedon:1992:MRR

[91] F. Makedon and A. Simvonis. Multipacket routing on rings. Lecture Notes in Computer Science, 591:226– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vesztergombi:1992:MPP

[92] G. Vesztergombi and F. Rohrbach. Massively parallel processing in high energy physics: The CERN-MPPC project. Lecture Notes in Computer Science, 591:238–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heiss:1992:HAD

[93] H.-U. Heiss and R. Wiesenfarth. A heuristic algorithm for dynamic task allocation in highly parallel systems. Lecture Notes in Computer Science, 591: 252–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ilmberger:1992:APL

[94] H. Ilmberger and S. Thurmel. Analysis of parallel Lisp programs based on a trace mechanism. Lecture Notes in Computer Science, 591:266–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glasser:1992:DIF

[95] U. Glasser, G. Hannesen, M. Karcher, and G. Lehrenfeld. A distributed implementation of flat concurrent Prolog on Multi-Transputer environments. Lecture Notes in Computer Science, 591:277–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacquet:1992:NC

[96] J.-M Jacquet. Negation in conclog. Lecture Notes in Computer Science, 591: 289–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1992:SCP

[97] P. S. Wang. Symbolic computation and parallel software. Lecture Notes in Computer Science, 591:316–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1992:PCS

[98] D. Wang. On the parallelization of characteristic-set-based algorithms. Lecture Notes in Computer Science, 591:338–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lippitsch:1992:MPP

[99] P. Lippitsch, K. C. Posch, and R. Posch. Multiplication as parallel as possible. Lecture Notes in Computer Science, 591:350–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zerovnik:1992:EEP

[100] J. Zerovnik. On the existence of an efficient parallel algorithm for a graph theoretic problem. Lecture Notes in Computer Science, 591:359–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuchlin:1992:MTC

[101] W. Kuchlin. On the multi-threaded computation of modular polynomial greatest common divisors. Lecture Notes in Computer Science, 591:369– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hawley:1992:BAD

[102] D. J. Hawley. A Buchberger algorithm for distributed memory multiprocessors. Lecture Notes in Computer Science, 591:385–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulten:1992:CBM

[103] K. Schulten. Computational biology on massively parallel machines. Lecture Notes in Computer Science, 591: 391–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horton:1992:TPM

[104] G. Horton and R. Knirsch. Timeparallel multigrid in an extrapolation method for time-dependent partial differential equations. Lecture Notes in Computer Science, 591:401–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breitenecker:1992:PST

[105] F. Breitenecker, G. Schuster, I. Husinsky, and J. Fritscher. Parallelization of simulation tasks: Methodology — implementation — application. Lecture Notes in Computer Science, 591:412– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adeli:1992:PAS

[106] H. Adeli and O. Kamal. Parallel algorithms for stress analysis on shared-memory multiprocessors. Lecture Notes in Computer Science, 591: 426–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miguet:1992:ELB

[107] S. Miguet and Y. Robert. Elastic loadbalancing for image processing algorithms. Lecture Notes in Computer Science, 591:438–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bekkers:1992:RTM

[108] Yves Bekkers and Lucien Ungaro. Realtime memory management for Prolog. Lecture Notes in Computer Science, 592:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belmesk:1992:PSL

[109] M. Belmesk. A process semantics of logic programs. Lecture Notes in Computer Science, 592:13–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borger:1992:LOS

[110] Egon Borger and Elvinia Riccobene. Logical operational semantics of parlog: Part II: Or-parallelism. Lecture Notes in Computer Science, 592:27–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borger:1992:WAM

[111] Egon Borger and Dean Rosenzweig. WAM algebras — A mathematical study of implementation: Part II. Lecture Notes in Computer Science, 592: 35–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bondarenko:1992:ASN

[112] A. G. Bondarenko. Abductive systems for non-monotonic reasoning. Lecture Notes in Computer Science, 592:55–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borshchev:1992:PAO

[113] Vladimir B. Borshchev. Properties of algorithmic operators. Lecture Notes in Computer Science, 592:66–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boulanger:1992:DLP

[114] Dmitri Yu. Boulanger. Deep logic program transformation using abstract interpretation. Lecture Notes in Computer Science, 592:79–??, 1992. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brogi:1992:OLP

[115] Antonio Brogi, Evelina Lamma, and Paola Mello. Objects in a logic programming framework. Lecture Notes in Computer Science, 592:102–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bry:1992:IVK

[116] Francois Bry, Rainer Manthey, and Bern Martens. Integrity verification in knowledge bases. Lecture Notes in Computer Science, 592:114–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Costantini:1992:PSM

[117] Stefania Costantini and Gaetano Lanzarone. On procedural semantics of metalevel negation. Lecture Notes in Computer Science, 592:140–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dantsin:1992:PLP

[118] Eugene Dantsin. Probabilistic logic programs and their semantics. Lecture Notes in Computer Science, 592:152– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demoen:1992:IPB

[119] Bart Demoen and Andre Marien. Implementation of Prolog as binary definite programs. Lecture Notes in Computer Science, 592:165–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dikowski:1992:PSM

[120] A. Ja. Dikowski. Prolog semantics for measuring space consumption. Lecture Notes in Computer Science, 592:177– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foong:1992:PPH

[121] Wai-Keong Foong. Or-parallel Prolog with heuristic task distribution. Lecture Notes in Computer Science, 592: 193–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hein:1992:WCS

[122] Hans-Gunther Hein and Manfred Meyer. A WAM compilation scheme. Lecture Notes in Computer Science, 592:201–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henson:1992:SPI

[123] Martin C. Henson. Safe positive induction in the programming logic TK. Lecture Notes in Computer Science, 592: 215–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ivanets:1992:WSP

[124] S. Ivanets, N. Ilinsky, and M. Krylov. WAM specification for parallel execution on SIMD computer. Lecture Notes in Computer Science, 592:232– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janssens:1992:APB

[125] G. Janssens and M. Bruynooghe. On abstracting the procedural behaviour of logic programs. Lecture Notes in Computer Science, 592:240–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kehrer:1992:TEE

[126] Norbert Kehrer and Gustaf Neumann. Treating enhanced entity relationship models in a declarative style. Lecture Notes in Computer Science, 592:263– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitaev:1992:PGR

[127] Eugeny L. Kitaev. Processing of ground regular terms in Prolog. Lecture Notes in Computer Science, 592: 271–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mantsivoda:1992:CFa

[128] Andrei Mantsivoda and Vyacheslav Petukhin. Compiling Flang. Lecture Notes in Computer Science, 592:286– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer:1992:FFD

[129] Manfred Meyer, Hans-Gunther Hein, and Jorg Muller. FIDO: Finite domain consistency techniques in logic programming. Lecture Notes in Computer Science, 592:294–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miglioli:1992:CLA

[130] Pierangelo Miglioli, Ugo Moscato, and Mario Ornaghi. A constructive logic approach to database theory. Lecture Notes in Computer Science, 592:302– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miller:1992:ASL

[131] Dale Miller. Abstract syntax and logic programming. Lecture Notes in Computer Science, 592:322–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neiman:1992:DSG

[132] Vladimir S. Neiman. Deduction search with generalized terms. Lecture Notes in Computer Science, 592:338– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neumann:1992:STP

[133] Gustaf Neumann. A simple transformation from Prolog-written metalevel interpreters into compilers and its implementation. Lecture Notes in Computer Science, 592:349–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parigot:1992:FDA

[134] Michel Parigot. Free deduction: An analysis of “computations” in classical logic. Lecture Notes in Computer Science, 592:361–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pliuskeviciene:1992:GTC

[135] Aida Pliuskeviciene. Gentzen-type calculi for modal logic S4 with Barcan formula. Lecture Notes in Computer Science, 592:381–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pliuskevicius:1992:LFL

[136] Regimantas Pliuskevicius. Logical foundation for logic programming based on first order linear temporal logic. Lecture Notes in Computer Science, 592:391–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Powers:1992:LPP

[137] David M. Powers. Logic programming with pseudo-resolution. Lecture Notes in Computer Science, 592:407– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reynolds:1992:BPD

[138] T. J. Reynolds and P. Kefals. BRAVE: An OR-Parallel dialect of Prolog and its application to artificial intelligence. Lecture Notes in Computer Science, 592:415–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Russo:1992:DDE

[139] Francesco Russo and Mirko Sancassani. A declarative debugging environment for DATALOG. Lecture Notes in Computer Science, 592:433–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakalauskaite:1992:SCFa

[140] Jurate Sakalauskaite. A sequent calculus for a first order linear temporal logic with explicit time. Lecture Notes in Computer Science, 592:442–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sebillot:1992:LBL

[141] P. Sebillot. A logical-based language for feature specification and transmission control. Lecture Notes in Computer Science, 592:452–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tarau:1992:PTW

[142] Paul Tarau. Program transformations and WAM-Support for the compilation of definite metaprograms. Lecture Notes in Computer Science, 592:462– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanderHoek:1992:SCL

[143] Wiebe van der Hoek. Some considerations on the logic psub F D — A logic combining modality and probability. Lecture Notes in Computer Science, 592:474–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voronkov:1992:LPB

[144] Andrei Voronkov. Logic programming with bounded quantifiers. Lecture Notes in Computer Science, 592: 486–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MacGinnes:1992:HOO

[145] S. MacGinnes. How objective is object-oriented analysis. Lecture Notes in Computer Science, 593:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rose:1992:IOA

[146] T. Rose, C. Maltzahn, and M. Jarke. Integrating object and agent worlds. Lecture Notes in Computer Science, 593:17–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borras:1992:FIP

[147] P. Borras, A. Doucet, and P. Pfeffer. A fully integrated programming environment for an object-oriented database. Lecture Notes in Computer Science, 593:33–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jesus:1992:AGD

[148] L. Jesus and R. Carapuca. Automatic generation of documentation for information systems. Lecture Notes in Computer Science, 593:48–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Opdahl:1992:FPE

[149] A. L. Opdahl and A. Solvberg. A framework for performance engineering during information system development. Lecture Notes in Computer Science, 593:65–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poo:1992:FSM

[150] C-C D. Poo. A framework for software maintenance. Lecture Notes in Computer Science, 593:88–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zicari:1992:SOO

[151] R. Zicari, F. Cacace, C. Capelli, A. Galipo, A. Pirovano, and A. Romboli. The SOL object-oriented database language. Lecture Notes in Computer Science, 593:105–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andonoff:1992:IDO

[152] E. Andonoff, C. Sallaberry, and G. Zurfluh. Interactive design of objectoriented databases. Lecture Notes in Computer Science, 593:128–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghosh:1992:CGF

[153] B. C. Ghosh and V. Wuwongse. Conceptual graphs as a framework for deductive object-oriented databases. Lecture Notes in Computer Science, 593: 147–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ip:1992:KBT

[154] S. Ip and T. Holden. A knowledge based technique for the process modelling of information systems: The object life cycle diagram. Lecture Notes in Computer Science, 593:164–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hainaut:1992:DCT

[155] J.-L. Hainaut, M. Cadelli, B. Decuyper, and O. Marchand. Database CASE tool architecture: Principles for flexible design strategies. Lecture Notes in Computer Science, 593:187–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shaw:1992:SKE

[156] M. L. G. Shaw and B. R. Gaines. The synthesis of knowledge engineering and software engineering. Lecture Notes in Computer Science, 593:208–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagelstein:1992:ROD

[157] J. Hagelstein and D. Roelants. Reconciling operational and declarative specifications. Lecture Notes in Computer Science, 593:221–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leonard:1992:NFL

[158] M. Leonard and I. Prince. NelleN: A framework for literate data modelling. Lecture Notes in Computer Science, 593:239–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rolland:1992:NLA

[159] C. Rolland and C. Proix. A natural language approach for requirements engineering. Lecture Notes in Computer Science, 593:257–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fouque:1992:BTS

[160] G. Fouque and C. Vrain. Building a tool for software code analysis: A machine learning approach. Lecture Notes in Computer Science, 593:278– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sutcliffe:1992:SCM

[161] A. Sutcliffe and N. Maiden. Supporting component matching for software reuse. Lecture Notes in Computer Science, 593:290–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Constantopoulos:1992:BSR

[162] P. Constantopoulos and E. Pataki. A browser for software reuse. Lecture Notes in Computer Science, 593:304– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1992:ESE

[163] E. Dubois, P. Du Bois, and A. Rifaut. Elaborating, structuring and expressing formal requirements of composite systems. Lecture Notes in Computer Science, 593:327–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopez:1992:OOO

[164] O. P. Lopez, F. Hayes, and S. Bear. OASIS: An object-oriented specification language. Lecture Notes in Computer Science, 593:348–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

terHofstede:1992:DMC

[165] A. H. M. ter Hofstede, H. A. Proper, and Th. P. van der Weide. Data modelling in complex application domains. Lecture Notes in Computer Science, 593:364–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Budgen:1992:ADP

[166] D. Budgen and G. Friel. Augmenting the design process transformations from abstract design representations. Lecture Notes in Computer Science, 593:378–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dewal:1992:MRA

[167] S. Dewal. A methodology for requirements analysis and evaluation of SDEs. Lecture Notes in Computer Science, 593:394–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reim:1992:OII

[168] F. Reim. Organizational integration of the information system design process. Lecture Notes in Computer Science, 593:410–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dalianis:1992:MVC

[169] H. Dalianis. A method for validating a conceptual model by natural language discourse generation. Lecture Notes in Computer Science, 593:425–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halpin:1992:AVC

[170] T. A. Halpin and J. I. McCormack. Automated validation of conceptual schema constraints. Lecture Notes in Computer Science, 593:445–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fonkam:1992:AES

[171] M. M. Fonkam and W. A. Gray. An approach to eliciting the semantics of relational databases. Lecture Notes in Computer Science, 593:463–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verrijn-Stuart:1992:MII

[172] A. A. Verrijn-Stuart and G. J. Ramackers. Model integration in information planning tools. Lecture Notes in Computer Science, 593:481–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Osmon:1992:IID

[173] P. Osmon and P. Sleat. IDRIS: Interactive design of reactive information systems. Lecture Notes in Computer Science, 593:494–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Comyn-Wattiau:1992:CCI

[174] I. Comyn-Wattiau and M. Bouzeghoub. Constraint confrontation: An important step in view integration. Lecture Notes in Computer Science, 593:507– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brough:1992:MCG

[175] M. Brough. Methods for CASE: A generic framework. Lecture Notes in Computer Science, 593:524–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rossi:1992:MEF

[176] M. Rossi, M. Gustafsson, K. Smolander, L.-A. Johansson, and K. Lyytinen. Metamodeling editor as a front end tool for a CASE shell. Lecture Notes in Computer Science, 593:546–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oquendo:1992:MCE

[177] F. Oquendo, J.-D. Zucker, and P. Griffiths. A Meta-CASE environment for software process-centred CASE environments. Lecture Notes in Computer Science, 593:568–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Casanova:1992:DCM

[178] M. A. Casanova, A. S. Hemerly, and A. L. Furtado. A declarative conceptual modelling language: Description and example applications. Lecture Notes in Computer Science, 593: 589–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Costal:1992:MRA

[179] D. Costal and A. Olive. A method for reasoning about deductive conceptual models of information systems. Lecture Notes in Computer Science, 593: 612–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leone:1992:BQM

[180] N. Leone, A. Mecchia, G. Rossi, and P. Rullo. The basic query machine of the KIWIS system. Lecture Notes in Computer Science, 593:632–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alt:1992:RSG

[181] Helmut Alt and Johannes Blomer. Resemblance and symmetries of geometric patterns. Lecture Notes in Computer Science, 594:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fleischer:1992:STC

[182] Rudolf Fleischer, Otfried Fries, Kurt Mehlhorn, Stefan Meiser, Stefan Naher, Hans Rohnert, Stefan Schirra, Klaus Simon, Athanasios Tsakalidis, and Christian Uhrig. Selected topics from computational geometry, data structures and motion planning. Lecture Notes in Computer Science, 594: 25–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofting:1992:PHD

[183] Franz Hofting, Thomas Lengauer, and Egon Wanke. Processing of hierarchically defined graphs and graph families. Lecture Notes in Computer Science, 594:44–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kriegel:1992:CSA

[184] Hans-Peter Kriegel, Thomas Brinkhoff, and Ralf Schneider. The combination of spatial access methods and computational geometry in geographic database systems. Lecture Notes in Computer Science, 594:70–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kriegel:1992:FEI

[185] Hans-Peter Kriegel, Peter Heep, Stephan Heep, Michael Schiwietz, and Ralf Schneider. A flexible and extensible index manager for spatial database systems. Lecture Notes in Computer Science, 594:87–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kriegel:1992:POD

[186] Hans-Peter Kriegel, Holger Horn, and Michael Schiwietz. The performance of object decomposition techniques for spatial query processing. Lecture Notes in Computer Science, 594:104– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muller:1992:DIS

[187] Heinrich Muller and Jorg Winckler. Distributed image synthesis with breadth-first ray tracing and the ray-Z-buffer. Lecture Notes in Computer Science, 594:124–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nilsson:1992:ROC

[188] Bengt J. Nilsson, Thomas Ottmann, Sven Schuierer, and Christian Icking. Restricted orientation computational geometry. Lecture Notes in Computer Science, 594:148–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Noltemeier:1992:MBT

[189] H. Noltemeier, K. Verbarg, and C. Zirkelbach. Monotonous bisectorsup * trees — A tool for efficient partitioning of complex scenes of geometric objects. Lecture Notes in Computer Science, 594:186–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmeltz:1992:LCS

[190] Bernd Schmeltz. Learning convex sets under uniform distribution. Lecture Notes in Computer Science, 594:204– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Six:1992:SAS

[191] Hans-Werner Six and Peter Widmayer. Spatial access structures for geometric databases. Lecture Notes in Computer Science, 594:214–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Welzl:1992:STL

[192] Emo Welzl. On spanning trees with low crossing numbers. Lecture Notes in Computer Science, 594:233–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dietzfelbinger:1992:HPU

[193] Martin Dietzfelbinger and Friedhelm Meyer auf der Heide. High performance universal hashing, with applications to shared memory simulations. Lecture Notes in Computer Science, 594:250–269, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feldmann:1992:DGT

[194] R. Feldmann, P. Mysliwietz, and B. Monien. Distributed game tree search on a massively parallel system. Lecture Notes in Computer Science, 594:270–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kunde:1992:BSR

[195] Manfred Kunde and Thomas Tensi. Balanced strategies for routing on meshes. Lecture Notes in Computer Science, 594:289–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kutylowski:1992:CBF

[196] Miroslaw Kutylowski. Complexity of Boolean functions on PRAMs — lower bound techniques. Lecture Notes in Computer Science, 594:309–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kroger:1992:EVG

[197] Berthold Kroger and Oliver Vornberger. Enumerative vs. genetic optimization two parallel algorithms for the bin packing problem. Lecture Notes in Computer Science, 594:330–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reischuk:1992:AEM

[198] Rudiger Reischuk and Bernd Schmeltz. Area efficient methods to increase the reliability of circuits. Lecture Notes in Computer Science, 594:363–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levene:1992:NUR

[199] M. Levene. The nested universal relation database model. Lecture Notes in Computer Science, 595:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aronsson:1992:MPT

[200] Martin Aronsson. Methodology and programming techniques in GCLA II. Lecture Notes in Computer Science, 596:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boley:1992:ELP

[201] Harold Boley. Extended logic-plus-functional programming. Lecture Notes in Computer Science, 596:45–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brogi:1992:OLT

[202] Antonio Brogi, Evelina Lamma, and Paola Mello. Open logic theories. Lecture Notes in Computer Science, 596: 73–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eriksson:1992:FVC

[203] Lars-Henrik Eriksson. A finitary version of the calculus of partial inductive definitions. Lecture Notes in Computer Science, 596:89–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Felty:1992:LPA

[204] Amy Felty. A logic programming approach to implementing higher-order term rewriting. Lecture Notes in Computer Science, 596:135–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franzen:1992:OAK

[205] Torkel Franzen, Seif Haridi, and Sverker Janson. An overview of the andorra kernel language. Lecture Notes in Computer Science, 596:163–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hallnas:1992:LCI

[206] Lars Hallnas. Logical and computational invariants of programs. Lecture Notes in Computer Science, 596:181– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hannan:1992:ILC

[207] John Hannan. Implementing lambdacalculus reduction strategies in extended logic programming languages. Lecture Notes in Computer Science, 596:193–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanschke:1992:TRP

[208] Philipp Hanschke. Terminological reasoning and partial inductive definitions. Lecture Notes in Computer Science, 596:221–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kreuger:1992:GID

[209] Per Kreuger. GCLA II A definitional approach to control. Lecture Notes in Computer Science, 596:239–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michaylov:1992:NSS

[210] Spiro Michaylov and Frank Pfenning. Natural semantics and some of its meta-theory in elf. Lecture Notes in Computer Science, 596:299–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reed:1992:NHA

[211] David W. Reed, Donald W. Loveland, and Bruce T. Smith. The near-Horn approach to disjunctive logic programming. Lecture Notes in Computer Science, 596:345–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gusgen:1992:I

[212] Hans Werner Gusgen and J. Hertzberg. Introduction. Lecture Notes in Computer Science, 597:1–4, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gusgen:1992:OW

[213] Hans Werner Gusgen and J. Hertzberg. The office world. Lecture Notes in Computer Science, 597:5–16, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gusgen:1992:PCB

[214] Hans Werner Gusgen and J. Hertzberg. A perspective of constraint-based reasoning: an introductory tutorial, volume 597 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CO-DEN LNCSD9. ISBN 0-387-55510-2 (Berlin), 3-540-55510-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 123 pp. LCCN Q340 .G8 1992.

Gusgen:1992:BA

[215] Hans Werner Gusgen and J. Hertzberg. Backtracking approaches. Lecture Notes in Computer Science, 597:57–66, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gusgen:1992:SFT

[216] Hans Werner Gusgen and J. Hertzberg. From simple filtering to tagging. Lecture Notes in Computer Science, 597: 67–82, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gusgen:1992:AFP

[217] Hans Werner Gusgen and J. Hertzberg. Approaches to filtering in parallel. Lecture Notes in Computer Science, 597: 83–96, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gusgen:1992:OA

[218] Hans Werner Gusgen and J. Hertzberg. Optimization approaches. Lecture Notes in Computer Science, 597:97– 110, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ma:1992:TAP

[219] QingMing Ma and John C. Reynolds. Types, abstraction, and parametric polymorphism, part 2. Lecture Notes in Computer Science, 598:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gateley:1992:CVC

[220] John Gateley and Bruce F. Duba. Callby-value combinatory logic and the lambda-value calculus. Lecture Notes in Computer Science, 598:41–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1992:ODS

[221] Scott F. Smith. From operational to denotational semantics. Lecture Notes in Computer Science, 598:54–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mislove:1992:SLS

[222] Michael W. Mislove and Frank J. Oles. A simple language supporting angelic nondeterminism and parallel composition. Lecture Notes in Computer Science, 598:77–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruce:1992:ETS

[223] Kim B. Bruce. The equivalence of two semantic definitions for inheritance in object-oriented languages. Lecture Notes in Computer Science, 598:102– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nelson:1992:PRF

[224] Neal Nelson. Primitive recursive functionals with dependent types. Lecture Notes in Computer Science, 598:125– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leavens:1992:THR

[225] Gary T. Leavens and Don Pigozzi. Typed homomorphic relations extended with subtypes. Lecture Notes in Computer Science, 598:144–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barwise:1992:ILD

[226] Jon Barwise. Information links in domain theory. Lecture Notes in Computer Science, 598:168–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rutten:1992:NSP

[227] J. J. M. M. Rutten. Nonwellfounded sets and programming language semantics. Lecture Notes in Computer Science, 598:193–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gray:1992:SST

[228] John W. Gray. Simultaneous substitution in the typed lambda calculus. Lecture Notes in Computer Science, 598:207–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barr:1992:HTT

[229] Michael Barr. HSP type theorems in the category of posets. Lecture Notes in Computer Science, 598:221–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jung:1992:DD

[230] Achim Jung, Leonid Libkin, and Hermann Puhlmann. Decomposition of domains. Lecture Notes in Computer Science, 598:235–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huth:1992:CCC

[231] Michael Huth. Cartesian closed categories of domains and the space proj(D). Lecture Notes in Computer Science, 598:259–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heckmann:1992:UPD

[232] Reinhold Heckmann. An upper power domain construction in terms of strongly compact sets. Lecture Notes in Computer Science, 598:272–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wand:1992:CPR

[233] Mitchell Wand. Correctness of procedure representations in higher-order assembly language. Lecture Notes in Computer Science, 598:294–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mercouroff:1992:AAC

[234] Nicolas Mercouroff. An algorithm for analyzing communicating processes. Lecture Notes in Computer Science, 598:312–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brookes:1992:CFP

[235] Stephen Brookes and Shai Geva. Continuous functions and parallel algorithms on concrete data structures. Lecture Notes in Computer Science, 598:326–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloom:1992:TOT

[236] Bard Bloom and Marta Kwiatkowska. Trade-offs in true concurrency: Pomsets and Mazurkiewicz traces. Lecture Notes in Computer Science, 598:350– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aceto:1992:RCN

[237] Luca Aceto. On relating concurrency and nondeterminism. Lecture Notes in Computer Science, 598:376–??, 1992.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kasangian:1992:CTA

[238] Stefano Kasangian. On continuous time agents. Lecture Notes in Computer Science, 598:403–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guo-Qiang:1992:MCC

[239] Zhang Guo-Qiang. A monoidal closed category of event structures. Lecture Notes in Computer Science, 598:426– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rosolini:1992:EMQ

[240] Giuseppe Rosolini. An exper model for quest. Lecture Notes in Computer Science, 598:436–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manes:1992:EIT

[241] Ernest G. Manes. Equations for ifthen-else. Lecture Notes in Computer Science, 598:446–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloom:1992:PCM

[242] Stephen L. Bloom and Zoltan Esik. Program correctness and matricial iteration theories. Lecture Notes in Computer Science, 598:457–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klarlund:1992:LPM

[243] Nils Klarlund. Liminf progress measures. Lecture Notes in Computer Science, 598:477–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Melton:1992:C

[244] Austin Melton, Bernd S. W. Schroeder, and George E. Strecker. Connections. Lecture Notes in Computer Science, 598:492–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allard:1992:UVC

[245] F. Allard. A user’s view of current practice and possibilities. Lecture Notes in Computer Science, 599:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Killin:1992:MMO

[246] J. Killin. The management and maintenance of an operational KADS system development. Lecture Notes in Computer Science, 599:6–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Georges:1992:KET

[247] M. Georges. Knowledge engineering trends in europe. Lecture Notes in Computer Science, 599:7–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morik:1992:AML

[248] K. Morik. Applications of machine learning. Lecture Notes in Computer Science, 599:9–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eriksson:1992:CMA

[249] H. Eriksson and M. A. Musen. Conceptual models for automatic generation of knowledge-acquisition tools. Lecture Notes in Computer Science, 599:14–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greboval:1992:AOC

[250] C. Greboval and G. Kassel. An approach to operationalize conceptual models: The shell aide. Lecture Notes in Computer Science, 599:37–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Linster:1992:LMM

[251] M. Linster. Linking modeling to make sense and modeling to implement systems in an operational modeling environment. Lecture Notes in Computer Science, 599:55–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gappa:1992:CGD

[252] U. Gappa and K. Poeck. Common ground and differences of the KADS and strong-problem-solving-shell approach. Lecture Notes in Computer Science, 599:75–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schreiber:1992:DPS

[253] G. Schreiber, B. Wielinga, and H. Akkermans. Differentiating problem solving methods. Lecture Notes in Computer Science, 599:95–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanHeijst:1992:UGD

[254] G. van Heijst, B. Wielinga, and N. Shadbolt. Using generalised directive models in knowledge acquisition. Lecture Notes in Computer Science, 599:112–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geelen:1992:TFF

[255] P. Geelen, Z. Ruttkay, and J. Treur. Towards a formal framework to compare protocol interpretations and task specifications. Lecture Notes in Computer Science, 599:133–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsujino:1992:KAD

[256] K. Tsujino, V. G. Dabija, and S. Nishida. Knowledge acquisition driven by constructive and interactive induction. Lecture Notes in Computer Science, 599:153–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nedellec:1992:KRU

[257] C. Nedellec and K. Causse. Knowledge refinement using knowledge acquisition and machine learning methods. Lecture Notes in Computer Science, 599:171–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schweiger:1992:GCE

[258] J. Schweiger. Generating configuration expert systems from conceptual specifications of the expert knowledge. Lecture Notes in Computer Science, 599:191–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonker:1992:YAF

[259] W. Jonker and J. W. Spee. Yet another formalisation of KADS conceptual models. Lecture Notes in Computer Science, 599:211–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neubert:1992:KMK

[260] S. Neubert and R. Studer. The KEEP model, a knowledge engineering process model. Lecture Notes in Computer Science, 599:230–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nwana:1992:DDK

[261] H. S. Nwana, R. C. Paton, M. J. R. Shave, and T. J. M. Bench-Capon. Domain-driven knowledge modelling: Mediating and intermediate representations for knowledge acquisition. Lecture Notes in Computer Science, 599: 250–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Portman:1992:PKE

[262] M.-M. Portman and S. M. Easterbrook. PMI: Knowledge elicitation and De Bono’s thinking tools. Lecture Notes in Computer Science, 599:264–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larichev:1992:NAS

[263] O. I. Larichev. A new approach to the solution of expert classification problems. Lecture Notes in Computer Science, 599:283–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dieng:1992:KAE

[264] R. Dieng, A. Giboin, P.-A. Tourtier, and O. Corby. Knowledge acquisition for explainable, multi-expert, knowledge-based design systems. Lecture Notes in Computer Science, 599: 298–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charlet:1992:CMB

[265] J. Charlet, J.-P. Krivine, and C. Reynaud. Causal model-based knowledge acquisition tools: Discussion of experiments. Lecture Notes in Computer Science, 599:318–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maurer:1992:KBM

[266] F. Maurer. Knowledge base maintenance and consistency checking in MOLTKE/HyDi. Lecture Notes in Computer Science, 599:337–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allemang:1992:AKK

[267] D. Allemang and T. E. Rothenfluh. Acquiring knowledge of knowledge acquisition: A self-study of generic tasks. Lecture Notes in Computer Science, 599:353–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Porter:1992:RAD

[268] D. Porter. Reusable analysis and design components for knowledge-based system development. Lecture Notes in Computer Science, 599:373–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manago:1992:ADK

[269] M. Manago, N. Conruyt, and J. le Renard. Acquiring descriptive knowledge for classification and identification. Lecture Notes in Computer Science, 599:392–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmalhofer:1992:IDC

[270] F. Schmalhofer, T. Reinartz, and B. Tschaitschian. Intelligent documentation as a catalyst for developing cooperative knowledge-based systems. Lecture Notes in Computer Science, 599:406–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bradshaw:1992:EAD

[271] J. M. Bradshaw, P. Holm, O. Kipersztok, and T. Nguyen. eQuality: An application of DDucks to process management. Lecture Notes in Computer Science, 599:425–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abadi:1992:OFR

[272] M. Abadi and L. Lamport. An oldfashioned recipe for real time. Lecture Notes in Computer Science, 600:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alur:1992:VAS

[273] R. Alur, C. Courcoubetis, and D. Dill. Verifying automata specifications of probabilistic real-time systems. Lecture Notes in Computer Science, 600:28–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alur:1992:TTA

[274] R. Alur and D. Dill. The theory of timed automata. Lecture Notes in Computer Science, 600:45–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alur:1992:LMR

[275] R. Alur and T. A. Henzinger. Logics and models of real time: A survey. Lecture Notes in Computer Science, 600: 74–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeten:1992:SOR

[276] J. C. M. Baeten and J. A. Bergstra. The state operator in real time process algebra. Lecture Notes in Computer Science, 600:107–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bolognesi:1992:TPA

[277] T. Bolognesi and F. Lucidi. Timed process algebras with urgent interactions and a unique powerful binary operator. Lecture Notes in Computer Science, 600:124–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courtiat:1992:TSB

[278] J. P. Courtiat and M. Diaz. Time in state based formal description techniques for distributed systems. Lecture Notes in Computer Science, 600:149– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emerson:1992:RTM

[279] E. Allen Emerson. Real-time and the mu-calculus. Lecture Notes in Computer Science, 600:176–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halbwachs:1992:IRP

[280] N. Halbwachs and F. Rocheteau. Implementing reactive programs on circuits: A hardware implementation of LUSTRE. Lecture Notes in Computer Science, 600:195–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hansen:1992:SCD

[281] M. R. Hansen and Chaochen Zhou. Semantics and completeness of duration calculus. Lecture Notes in Computer Science, 600:209–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1992:TTS

[282] T. A. Henzinger, Z. Manna, and A. Pnueli. Timed transition systems. Lecture Notes in Computer Science, 600:226–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hooman:1992:CVR

[283] J. Hooman. Compositional verification of real-time systems using extended Hoare triples. Lecture Notes in Computer Science, 600:252–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huizing:1992:SRS

[284] C. Huizing and R. Gerth. Semantics of reactive systems in abstract time. Lecture Notes in Computer Science, 600:291–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joseph:1992:PPP

[285] M. Joseph. Problems, promises and performance: Some questions for realtime system specification. Lecture Notes in Computer Science, 600:315– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klusener:1992:ART

[286] A. S. Klusener. Abstraction in real time process algebra. Lecture Notes in Computer Science, 600:325–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koymans:1992:RTP

[287] R. Koymans. (real) time: A philosophical perspective. Lecture Notes in Computer Science, 600:353–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1992:SAR

[288] I. Lee and R. Gerber. Specification and analysis of resource-bound real-time systems. Lecture Notes in Computer Science, 600:371–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lynch:1992:FBS

[289] N. Lynch and F. Vaandrager. Forward and backward simulations for timing-based systems. Lecture Notes in Computer Science, 600:397–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maler:1992:THS

[290] O. Maler, Z. Manna, and A. Pnueli. From timed to hybrid systems. Lecture Notes in Computer Science, 600: 447–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mok:1992:CID

[291] A. K. Mok. Coping with implementation dependencies in real-time system verification. Lecture Notes in Computer Science, 600:485–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morzenti:1992:VRT

[292] A. Morzenti. Validating real-time systems by executing logic specifications. Lecture Notes in Computer Science, 600:502–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nicollin:1992:OST

[293] X. Nicollin and J. Sifakis. An overview and synthesis on timed process algebras. Lecture Notes in Computer Science, 600:526–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nicollin:1992:ATG

[294] X. Nicollin, J. Sifakis, and S. Yovine. From ATP to timed graphs and hybrid systems. Lecture Notes in Computer Science, 600:549–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ostroff:1992:VSC

[295] J. S. Ostroff. Verification of safety critical systems using TIM/RTTL. Lecture Notes in Computer Science, 600: 573–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schenke:1992:PST

[296] M. Schenke. Predicative specification of timed processes. Lecture Notes in Computer Science, 600:603–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneider:1992:PTP

[297] F. B. Schneider, B. Bloom, and K. Marzullo. Putting time into proof outlines. Lecture Notes in Computer Science, 600:618–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneider:1992:TCT

[298] S. Schneider, J. Davies, D. M. Jackson, G. M. Reed, J. N. Reed, and A. W. Roscoe. Timed CSP: Theory and practice. Lecture Notes in Computer Science, 600:640–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wood:1992:SCM

[299] W. G. Wood. A specification of the cat and mouse problem. Lecture Notes in Computer Science, 600:676–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zwiers:1992:LAR

[300] J. Zwiers. Layering and action refinement for timed systems. Lecture Notes in Computer Science, 600:687– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1992:MSP

[301] Jingsen Chen. Merging and splitting priority queues and deques in parallel. Lecture Notes in Computer Science, 601:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bshouty:1992:LBC

[302] Nader H. Bshouty. Lower bounds for the complexity of functions in a realistic RAM model. Lecture Notes in Computer Science, 601:12–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kenyon:1992:BDT

[303] Claire Kenyon and Valerie King. On Boolean decision trees with faulty nodes. Lecture Notes in Computer Science, 601:24–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golumbic:1992:IGI

[304] Martin Charles Golumbic and Ron Shamir. Interval graphs, interval orders and the consistency of temporal events. Lecture Notes in Computer Science, 601:32–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voronkov:1992:HOF

[305] Andrei Voronkov. Higher order functions in first order theory. Lecture Notes in Computer Science, 601:43–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dosch:1992:RRS

[306] Walter Dosch. Reduction relations in strict applicative languages. Lecture Notes in Computer Science, 601:55–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kortsarz:1992:AAM

[307] Guy Kortsarz and David Peleg. Approximation algorithms for minimum time broadcast. Lecture Notes in Computer Science, 601:67–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yosi:1992:CRN

[308] Ben-Asher Yosi, David Peleg, and Assaf Schuster. The complexity of reconfiguring network models. Lecture Notes in Computer Science, 601:79–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gal:1992:OMD

[309] Shmuel Gal, Yona Hollander, and Alon Itai. Optimal mapping in direct mapped cache environments. Lecture Notes in Computer Science, 601:91–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohen:1992:NAG

[310] Edith Cohen and Nimrod Megiddo. New algorithms for generalized network flows. Lecture Notes in Computer Science, 601:103–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miller:1992:FPR

[311] Victor S. Miller. Factoring polynomials via relation-finding. Lecture Notes in Computer Science, 601:115–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pan:1992:NRI

[312] Victor Pan. New resultant inequalities and complex polynomial factorization. Lecture Notes in Computer Science, 601:122–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Linzer:1992:CST

[313] Elliot Linzer. Can symmetric Toeplitz solvers be strongly stable? Lecture Notes in Computer Science, 601:137– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sadeh:1992:BPC

[314] Ilan Sadeh and Amir Averbuch. Bounds on parallel computation of multivariate polynomials. Lecture Notes in Computer Science, 601:147–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dwork:1992:TLS

[315] Cynthia Dwork, Maurice Herlihy, Serge A. Plotkin, and Orli Waarts. Time-lapse snapshots. Lecture Notes in Computer Science, 601:154–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gawlick:1992:CTM

[316] Rainer Gawlick, Nancy Lynch, and Nir Shavit. Concurrent timestamping made simple. Lecture Notes in Computer Science, 601:171–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Helary:1992:DET

[317] Jean-Michel Helary and Michel Raynal. Distributed evaluation: A tool for constructing distributed detection programs. Lecture Notes in Computer Science, 601:184–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luk-Zilberman:1992:FAT

[318] Evgeny V. Luk-Zilberman. Foundations of asymptotical theory of determinate compact testing. Lecture Notes in Computer Science, 601:195–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartman:1992:OKC

[319] Irith Ben-Arroyo Hartman. Optimal K-colouring and K-nesting of intervals. Lecture Notes in Computer Science, 601:207–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maurer:1992:WHS

[320] H. Maurer. Why hypermedia systems are important. Lecture Notes in Computer Science, 602:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moore:1992:PCC

[321] L. Moore and D. Smith. Project CALC: Calculus as a laboratory course. Lecture Notes in Computer Science, 602: 16–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murray:1992:SAM

[322] J. H. Murray and S. A. Malone. The structures of advanced multimedia learning environments: Reconfiguring space, time, story, and text. Lecture Notes in Computer Science, 602:21–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petruk:1992:APS

[323] M. W. Petruk. Adjusting to the paradigm shift in teaching and learning or: What do I do now? Lecture Notes in Computer Science, 602:34–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shneiderman:1992:ECE

[324] B. Shneiderman. Engagement and construction: Educational strategies for the Post-TV era. Lecture Notes in Computer Science, 602:39–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Woolf:1992:BKB

[325] B. P. Woolf. Building knowledge-based tutors. Lecture Notes in Computer Science, 602:46–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allinson:1992:LSC

[326] L. Allinson. Learning styles and computer-based learning environments. Lecture Notes in Computer Science, 602:61–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anoskey:1992:TGI

[327] A. M. Anoskey and R. Catrambone. Text and graphics in instructional design. Lecture Notes in Computer Science, 602:74–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Badre:1992:APV

[328] A. Badre, M. Beranek, J. M. Morris, and J. Stasko. Assessing program visualization systems as instructional aids. Lecture Notes in Computer Science, 602:87–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balle:1992:LCA

[329] C. Balle, F. Dubreuil, and B. Bouret. At last! A computer aid for the teaching of electronics which answers all the nagging questions about capacitors. Lecture Notes in Computer Science, 602:100–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beltran:1992:OOA

[330] Th. Beltran and A. Peninou. An object-oriented approach to produce educational hypermedia software. Lecture Notes in Computer Science, 602: 111–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertels:1992:CMP

[331] K. Bertels, P. Vanneste, and C. De Backer. A cognitive model of programming knowledge for procedural languages. Lecture Notes in Computer Science, 602:124–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1992:PEE

[332] C. Brown, H. Fell, V. K. Proulx, and R. Rasala. Programming by experimentation and example. Lecture Notes in Computer Science, 602:136– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clement:1992:ECL

[333] D. Clement, C. Vieville, and P. Vilers. An experiment of cooperative learning with hypercard. Lecture Notes in Computer Science, 602:148–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diessel:1992:EDC

[334] Th. Diessel and A. Lehmann. An ITS for engineering domains: Concept, design and application. Lecture Notes in Computer Science, 602:161–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dumslaff:1992:SSM

[335] U. Dumslaff and J. Ebert. Structuring the subject matter. Lecture Notes in Computer Science, 602:174–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Egnatoff:1992:MDD

[336] W. J. Egnatoff. MIDI draw: Designing an impressionistic medium for young musicians, artists, and writers. Lecture Notes in Computer Science, 602:187–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

ElHani:1992:SAI

[337] O. ElHani and G. Gouarderes. Standardized architecture for integrated open courseware. Lecture Notes in Computer Science, 602:198–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Forcheri:1992:FTH

[338] P. Forcheri and M. T. Molfino. Formal techniques in higher education: A proposal. Lecture Notes in Computer Science, 602:212–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Futtersack:1992:QDI

[339] M. Futtersack and J.-M. Labat. QUIZ: A distributed intelligent tutoring system. Lecture Notes in Computer Science, 602:225–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gonzalez:1992:ITS

[340] J. C. Gonzalez, J. J. Sancho, J. M. Carbo, A. Patak, and F. Sanz. Intelligent tutorial system in medicine through an interactive testing program: HyperMIR. Lecture Notes in Computer Science, 602:238–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gunzenhauser:1992:DKB

[341] R. Gunzenhauser and A. Zimmermann. DCE: A knowledge-based tutoring and advisory system: Tutoring strategies and architecture. Lecture Notes in Computer Science, 602:247–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hashim:1992:DSS

[342] R. Hashim and H. Hamburger. Discourse style and situation viewpoint for a conversational language tutor. Lecture Notes in Computer Science, 602: 258–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hayashi:1992:OSC

[343] T. Hayashi and Y. Yano. Open structured CAI system for Kanji learning. Lecture Notes in Computer Science, 602:271–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hesser:1992:SPG

[344] L. A. Hesser, R. P. Hogan, and A. P. Mizell. The sum of the parts is greater than the whole in online graduate education. Lecture Notes in Computer Science, 602:283–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hidaka:1992:DGM

[345] K. Hidaka. Development of GeoBlock: A micro-world for learning and teaching geometry. Lecture Notes in Computer Science, 602:294–??, 1992. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hohmann:1992:SCA

[346] L. Hohmann, M. Guzdial, and E. Soloway. SODA: A computer aided design environment for the doing and learning of software design. Lecture Notes in Computer Science, 602:307– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jackson:1992:GCA

[347] D. Jackson and M. A. Bell. A graphical CAL author language. Lecture Notes in Computer Science, 602:320–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaminski:1992:KBA

[348] D. M. Kaminski. A knowledge base approach to learning to program in Prolog. Lecture Notes in Computer Science, 602:332–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katz:1992:SAC

[349] S. Katz, A. Lesgold, G. Eggan, M. Gordin, and L. Greenberg. Selfadjusting curriculum planning in Sherlock II. Lecture Notes in Computer Science, 602:343–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lawson:1992:IES

[350] Ph. A. Lawson. Integrating an educational simulation into a logic design course. Lecture Notes in Computer Science, 602:356–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1992:AAP

[351] M.-Ch. Lee. An algorithm animation programming environment. Lecture Notes in Computer Science, 602: 367–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1992:AHB

[352] J. Li, A. Rovick, and J. Michael. ABASE: A hypermedia-based tutoring and authoring system. Lecture Notes in Computer Science, 602:380–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lianjing:1992:TDI

[353] H. Lianjing. A tool for developing intelligent tutoring systems. Lecture Notes in Computer Science, 602:391– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lilly:1992:ACM

[354] R. Lilly. Applying computer models of phonological competence to CALL. Lecture Notes in Computer Science, 602:392–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Livingston:1992:VLT

[355] L. A. Livingston. Vigilance in a long-term cognitive computing task: The effects of subject strategy and screen colour on performance. Lecture Notes in Computer Science, 602:405– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1992:MMC

[356] Ch. Lu and D. Frye. Mastering the machine: A comparison of the mouse and touch screen for children’s use of computers. Lecture Notes in Computer Science, 602:417–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marcenac:1992:ASW

[357] P. Marcenac. An authoring system for ITS which is based on a generic level of tutoring strategies. Lecture Notes in Computer Science, 602:428–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michael:1992:UMS

[358] J. Michael, A. Rovick, M. Evens, L. Shim, Ch. Woo, and N. Kim. The uses of multiple student inputs in modeling and lesson planning in CAI and ICAI programs. Lecture Notes in Computer Science, 602:441–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muhlhauser:1992:PNN

[359] M. Muhlhauser and J. Schaper. Project NESTOR: New approaches to cooperative multimedia authoring/learning. Lecture Notes in Computer Science, 602:453–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muldner:1992:DIC

[360] T. Muldner and R. Blondon. Design and implementation of courseware for teaching programming languages. Lecture Notes in Computer Science, 602:466–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muldner:1992:OYA

[361] T. Muldner and M. Elammari. OB-JECTOR: Yet another authoring system. Lecture Notes in Computer Science, 602:478–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nicolson:1992:DES

[362] R. I. Nicolson. Design and evaluation of the SUMIT intelligent tutoring assistant for arithmetic. Lecture Notes in Computer Science, 602:491–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nicolson:1992:SRD

[363] R. I. Nicolson and A. J. Fawcett. Spelling remediation for dyslexic children using the selfspell programs. Lecture Notes in Computer Science, 602: 503–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nicolson:1992:UHC

[364] R. I. Nicolson and A. J. Simpson. Using HyperCard to create a flexible learning package for statistics: Costs, benefits and effectiveness. Lecture Notes in Computer Science, 602:516–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nodenot:1992:ESE

[365] Th. Nodenot. Educational software engineering: A methodology based on cooperative developments. Lecture Notes in Computer Science, 602:529– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okamoto:1992:ECC

[366] R. Okamoto and Y. Yano. Environmental CAL for conversation pattern learning. Lecture Notes in Computer Science, 602:542–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deLaPassardiere:1992:ANT

[367] B. de La Passardiere and A. Dufresne. Adaptive navigational tools for educational hypermedia. Lecture Notes in Computer Science, 602:555–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seminar:1992:TSE

[368] K. B. Seminar and R. N. Robson. A tool for storing and exploring ideas. Lecture Notes in Computer Science, 602:568–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vassileva:1992:DCC

[369] J. Vassileva. Dynamic CAL-Courseware generation within an ITS-Shell architecture. Lecture Notes in Computer Science, 602:581–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilson:1992:CUP

[370] E. Wilson, I. Lindley, and S. Pullen. CALLGUIDE: Using programmable hypertext as a shell for CALL programs. Lecture Notes in Computer Science, 602:592–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Looney:1992:DAD

[371] M. J. Looney and A. O’Brien. Distributed application designed using MASCOT and implemented in Ada. Lecture Notes in Computer Science, 603:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raines:1992:RTA

[372] Gary Raines. Real time Ada in the international space station freedom. Lecture Notes in Computer Science, 603: 9–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1992:MAO

[373] John A. Anderson and John D. Sheffler. Managing Ada object-oriented development. Lecture Notes in Computer Science, 603:20–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wearing:1992:SEA

[374] Alison Wearing. Software engineering, Ada and metrics. Lecture Notes in Computer Science, 603:35–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duell:1992:UAS

[375] Rob Duell, Hugo J. Sebel, and Franklin C. A. de Wit. Using Ada source code generators in a large project. Lecture Notes in Computer Science, 603:47–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zage:1992:DCM

[376] Wayne M. Zage, Dolores M. Zage, Manjari Bhargava, and Dale J. Gaumer. Design and code metrics through a DIANA-Based tool. Lecture Notes in Computer Science, 603:60–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Groeneveld:1992:UAI

[377] Miech Groeneveld. Using Ada in integrating ATC systems. Lecture Notes in Computer Science, 603:72–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:1992:EAS

[378] W. M. Thomas, A. Delis, and V. R. Basili. An evaluation of Ada source code reuse. Lecture Notes in Computer Science, 603:80–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maymir-Ducharme:1992:PER

[379] Fred A. Maymir-Ducharme. Porting embedded real-time Ada software. Lecture Notes in Computer Science, 603: 92–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

delaPuente:1992:REH

[380] Juan A. de la Puente, Juan Zamorano, Alejandro Alonso, and Jose L. Fernandez. Reusable executives for hard realtime systems in Ada. Lecture Notes in Computer Science, 603:104–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burns:1992:DHR

[381] A. Burns and A. J. Wellings. Designing hard real-time systems. Lecture Notes in Computer Science, 603:116– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gobin:1992:RSS

[382] M. Gobin, M. Timmerman, and F. J. A. Gielen. Runtime system support for data-oriented synchronization in Ada-9X. Lecture Notes in Computer Science, 603:128–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brosgol:1992:DAA

[383] Benjamin M. Brosgol, Robert I. Eachus, and David E. Emery. Decimal arithmetic in Ada. Lecture Notes in Computer Science, 603:138–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jingde:1992:TDN

[384] Cheng Jingde. Task dependence net as a representation for concurrent Ada programs. Lecture Notes in Computer Science, 603:150–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lander:1992:DAE

[385] Leslie C. Lander and Sandeep Mitra. Detection and avoidance of elaboration-time problems for multiunit real-time Ada applications. Lecture Notes in Computer Science, 603: 165–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ottens:1992:SMS

[386] Arlet Ottens and Hans Toetenel. Simulation of mosca specifications in Ada. Lecture Notes in Computer Science, 603:182–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Silberg:1992:CRV

[387] Steen Silberg. Considerations with regard to validation of Ada debuggers. Lecture Notes in Computer Science, 603:197–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Treurniet:1992:SAT

[388] W. Treurniet. SWG APSE test support toolset (assessment of the CAIS-A interface set). Lecture Notes in Computer Science, 603:214–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thanh-Nu:1992:CIS

[389] Do Thanh-Nu. Compilation integration: A solution for the challenge of developing and reusing Ada software on different platforms. Lecture Notes in Computer Science, 603:230–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elrad:1992:EWE

[390] Tzilla Elrad, Lee Sungyoung, and Ufuk Verun. Extending working environments for the development of reactive/ adaptive systems with intelligent controls. Lecture Notes in Computer Science, 603:242–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lundberg:1992:PSP

[391] Lars Lundberg. Predicting the speedup of parallel Ada programs. Lecture Notes in Computer Science, 603:257– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hummel:1992:HPA

[392] Susan Flynn Hummel. A highley parallel Ada task scheduler. Lecture Notes in Computer Science, 603:275– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bazalgette:1992:SAT

[393] G. Bazalgette, D. Bekele, C. Bernon, M. Filali, J. M. Rigaud, and A. Sayah. STRAda — an Ada transformation and distribution system. Lecture Notes in Computer Science, 603:287–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mangold:1992:AMP

[394] Karlotto Mangold. AMPATS — A multi-processor Ada tool set. Lecture Notes in Computer Science, 603:300– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bayassi:1992:PUA

[395] M. Bayassi, H. Bitteur, J.-F. Jezequel, and P. Legrain. A practical use of the Ada rendez-vous paradigm in distributed systems. Lecture Notes in Computer Science, 603:312–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lirov:1992:GSA

[396] Y. Lirov. Gaining strategic advantage with real-time distributed artificial intelligence. Lecture Notes in Computer Science, 604:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cremers:1992:IDI

[397] A. B. Cremers, G. Kniesel, T. Lemke, and L. Plumer. Intelligent databases and interoperability. Lecture Notes in Computer Science, 604:15–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fox:1992:TPT

[398] Mark S. Fox. The TOVE project: Towards a common-sense model of the enterprise. Lecture Notes in Computer Science, 604:25–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radig:1992:ADI

[399] B. Radig, W. Eckstein, K. Klotz, T. Messer, and J. Pauli. Automatization in the design of image understanding systems. Lecture Notes in Computer Science, 604:35–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dincbas:1992:CPA

[400] M. Dincbas. Constraint programming — an alternative to expert systems. Lecture Notes in Computer Science, 604:46–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamamoto:1992:CBR

[401] H. Yamamoto and H. Fujimoto. Casebased reasoning in expert systems assisting production line design. Lecture Notes in Computer Science, 604:49–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vanwelkenhuysen:1992:SMB

[402] J. Vanwelkenhuysen. Scaling-up modelbased troubleshooting by exploiting design functionalities. Lecture Notes in Computer Science, 604:59–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clark:1992:AKB

[403] G. Clark, P. Mehta, and T. Thomson. Application of knowledge-based systems to optimised building maintenance management. Lecture Notes in Computer Science, 604:69–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kehl:1992:AMB

[404] W. Kehl, H. Hopfmuller, T. Koussev, and M. Newstead. Application of model-based reasoning to the maintenance of telecommunication networks. Lecture Notes in Computer Science, 604:79–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Newstead:1992:AIM

[405] M. A. Newstead, B. Stahl, and G. Schapeler. Advanced information modelling for integrated network management applications. Lecture Notes in Computer Science, 604:89–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karamouzis:1992:ICB

[406] T. Karamouzis and S. Feyock. An integration of case-based and model-based reasoning and its application to physical system faults. Lecture Notes in Computer Science, 604:100–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becks:1992:APJ

[407] K.-H. Becks, J. Dahm, and F. Seidel. Analysing particle jets with artificial neural networks. Lecture Notes in Computer Science, 604:109–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monostori:1992:CBC

[408] L. Monostori and A. Bothe. Convergence behaviour of connectionist models in large scale diagnostic problems. Lecture Notes in Computer Science, 604:113–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kotterba:1992:PRA

[409] B. Kotterba. Pattern recognition approach to an acoustical quality test of burnt ceramic products. Lecture Notes in Computer Science, 604:123– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muller:1992:ESE

[410] B. Muller. Enhancing software engineering capabilities of PROLOG by object-oriented concepts. Lecture Notes in Computer Science, 604:127– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borden:1992:SDM

[411] A. Borden and U. Cinar. Specifying decision-making processes. Lecture Notes in Computer Science, 604:139– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reynolds:1992:OSR

[412] R. G. Reynolds, J. I. Maletic, and E. Zannoni. Operationalizing software reuse as a problem in inductive learning. Lecture Notes in Computer Science, 604:143–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zeiler:1992:TRC

[413] Ir. W. Zeiler. Towards a real CAD-System using artificial intelligence technology. Lecture Notes in Computer Science, 604:154–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pasternak:1992:WPK

[414] B. Pasternak, G. Gabrielides, and R. Sprengel. WIZ — A prototype for knowledge-based drawing interpretation. Lecture Notes in Computer Science, 604:164–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lekova:1992:ISF

[415] A. Lekova, D. Batanov, and N. Nikolaev. Intelligent system for featurebased modelling of machine parts. Lecture Notes in Computer Science, 604: 174–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vila:1992:IPC

[416] Ll. Vila, C. Sierra, A. B. Martinez, and J. Climent. Intelligent process control by means of expert systems and machine vision. Lecture Notes in Computer Science, 604:185–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schnepf:1992:TGM

[417] U. Schnepf, A. Asteroth, M. S. Fischer, and K. Moller. Tracking and grasping of moving objects — a behaviour-based approach. Lecture Notes in Computer Science, 604:195–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mertsching:1992:PKB

[418] B. Mertsching. PANTER — knowledge based image analysis system for workpiece recognition. Lecture Notes in Computer Science, 604:205–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nonnenmann:1992:KUK

[419] U. Nonnenmann and J. K. Eddy. KITSS: Using knowledge-based software engineering for functional testing. Lecture Notes in Computer Science, 604:215–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Botten:1992:CKB

[420] N. Botten. Complex knowledge base verification using matrices. Lecture Notes in Computer Science, 604:225– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saglietti:1992:STS

[421] F. Saglietti. Structural testing strategies applied to knowledge-based systems. Lecture Notes in Computer Science, 604:236–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barschdorff:1992:RRE

[422] D. Barschdorff, A. Bothe, U. Gartner, and A. Jager. Retraining and redundancy elimination for a condensed nearest neighbour network. Lecture Notes in Computer Science, 604:246– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pagurek:1992:MPD

[423] B. Pagurek, N. Dawes, and R. Kaye. A multiple paradigm diagnostic system for wide area communication networks. Lecture Notes in Computer Science, 604:256–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lackner:1992:RNN

[424] G. Lackner, S. S. Melsheimer, and J. N. Beard. Recursive neural net modeling of a multi-zone tenter frame dryer. Lecture Notes in Computer Science, 604:266–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bayliss:1992:RDM

[425] D. Bayliss, S. Berry, D. Curtis, and B. Cox. Representation of the design to manufacture methodology of the armour protection element of the fibre optic submarine cable within an expert system. Lecture Notes in Computer Science, 604:276–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernardi:1992:PPM

[426] A. Bernardi, C. Klauck, and R. Legleitner. PIM — planning in manufacturing — CAPP using skeletal plans. Lecture Notes in Computer Science, 604: 284–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmidt:1992:VMK

[427] D. Schmidt, J. Haddock, W. A. Wallace, and R. Wright. Visual modelling: A knowledge acquisition method for intelligent process control systems. Lecture Notes in Computer Science, 604: 294–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lehmann:1992:MLC

[428] F. Lehmann, R. Seising, and E. Walther-Klaus. Machine learning in communication nets. Lecture Notes in Computer Science, 604:304–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Specht:1992:SMB

[429] D. Specht and S. Weib. Supporting model-based diagnosis with explanation-based learning and analogical inferences. Lecture Notes in Computer Science, 604:314–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belanche:1992:KBS

[430] Ll. Belanche, M. Sanchez, U. Cortes, and P. Serra. A knowledge-based system for the diagnosis of waste-water treatment plants. Lecture Notes in Computer Science, 604:324–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LiMin:1992:CFL

[431] Fu LiMin. Connectionism for fuzzy learning in rule-based expert systems. Lecture Notes in Computer Science, 604:337–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Qi:1992:ECS

[432] Guan Qi and G. Friedrich. Extending constraint satisfaction problem solving in structural design. Lecture Notes in Computer Science, 604:341–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ukita:1992:MPT

[433] A. Ukita and T. Kitagawa. A modulation package tuning machine applying fuzzy logic. Lecture Notes in Computer Science, 604:351–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramos:1992:TIR

[434] C. Ramos and E. Oliveira. Towards intelligent robotic assemblies. Lecture Notes in Computer Science, 604:361– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kernebeck:1992:APM

[435] U. Kernebeck. Action planning for multiple robots in space. Lecture Notes in Computer Science, 604:371–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schrott:1992:CMC

[436] A. Schrott. A calibration of a mobile camera used for a camera-guided grasping by an eye-in-hand robot. Lecture Notes in Computer Science, 604:381– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirn:1992:SSR

[437] S. Kirn and J. Schneider. STRICT: Selecting the ‘right’ architecture. Lecture Notes in Computer Science, 604: 391–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cervoni:1992:STC

[438] L. Cervoni and F. Rousseaux. Solving temporal constraints satisfaction problems with an object-oriented model. Lecture Notes in Computer Science, 604:401–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arai:1992:IMD

[439] H. Arai, Y. Fukazawa, T. Kadokura, and T. Hasegawa. An integration mechanism for design models in the design environment DATE. Lecture Notes in Computer Science, 604:411–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kopisch:1992:CPA

[440] M. Kopisch and A. Gunther. Configuration of a passenger aircraft cabin based on conceptual hierarchy, constraints and flexible control. Lecture Notes in Computer Science, 604:421– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brinkop:1992:CIM

[441] A. Brinkop and N. Laudwein. Configuration of industrial mixing-machines-development of a knowledge-based system. Lecture Notes in Computer Science, 604:431–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Najmann:1992:TFC

[442] O. Najmann and B. Stein. A theoretical framework for configuration. Lecture Notes in Computer Science, 604: 441–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freisleben:1992:SMP

[443] B. Freisleben. Stock market prediction with backpropagation networks. Lecture Notes in Computer Science, 604: 451–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deGroot:1992:FTS

[444] C. de Groot and D. Wurtz. Forecasting time series with connectionist nets: Applications in statistics, signal processing and economics. Lecture Notes in Computer Science, 604:461–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VandeMerckt:1992:ISL

[445] T. Van de Merckt. ILISCE: A system for learning control heuristics in a scheduling environment. Lecture Notes in Computer Science, 604:471– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weichert:1992:KKB

[446] G. Weichert and R. Lackmann. KUSET — knowledge based user support for electron beam testing. Lecture Notes in Computer Science, 604:481– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burgard:1992:KBD

[447] W. Burgard, S. Luttringhaus-Kappel, and L. Plumer. Knowledge-based design of ergonomic lighting for underground scenarios. Lecture Notes in Computer Science, 604:491–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruning:1992:MAS

[448] R. Bruning, W. Hauenschild, and W. John. A monitoring approach supporting performance analysis of expert systems for the EMC design of printed circuit boards. Lecture Notes in Computer Science, 604:495–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zalik:1992:RGO

[449] B. Zalik, N. Guid, and A. Vesel. Representing geometric objects using constraint description graphs. Lecture Notes in Computer Science, 604:505– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DAloisi:1992:TLR

[450] D. D’Aloisi. A terminological language for representing complex knowledge. Lecture Notes in Computer Science, 604:515–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goss:1992:KRD

[451] S. Goss, I. Wallace, and K. Bluff. Knowledge representation and decision making: A hybrid approach. Lecture Notes in Computer Science, 604:525– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmidt:1992:KAT

[452] G. Schmidt. Knowledge acquisition from text in a complex domain. Lecture Notes in Computer Science, 604: 529–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reid:1992:PPA

[453] R. A. Reid and M. M. Matthews. Parallel parsing of ambiguous languages on hypercube architectures. Lecture Notes in Computer Science, 604:539– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puppe:1992:TKA

[454] F. Puppe and U. Gappa. Towards knowledge acquisition by experts. Lecture Notes in Computer Science, 604:

546–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stefanowski:1992:RSA

[455] J. Stefanowski, R. Nowicki, and R. Slowinski. The rough sets approach to knowledge analysis for classification support in technical diagnostics of mechanical objects. Lecture Notes in Computer Science, 604:556–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagen:1992:BDP

[456] B. W. Hagen. Bi-directional probabilistic assessment. Lecture Notes in Computer Science, 604:566–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nauck:1992:RUU

[457] D. Nauck, F. Klawonn, R. Kruse, and U. Lohs. Reasoning under uncertainty with temporal aspects. Lecture Notes in Computer Science, 604:572– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vale:1992:ESA

[458] Z. A. Vale and A. Machado e. Moura. An expert system approach for power system diagnosis. Lecture Notes in Computer Science, 604:581–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becks:1992:DEE

[459] K.-H. Becks, A. Hemker, J. Ortmann, G. Schlageter, R. Meyer, and A. B. Cremers. DELPHI-EXPERT: An expert system for error diagnosis in high energy physics detectors. Lecture Notes in Computer Science, 604:585–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brockmann:1992:CRT

[460] W. Brockmann. Combining realtime with knowledge processing techniques. Lecture Notes in Computer Science, 604:594–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aiken:1992:TAI

[461] R. M. Aiken and D. Allemang. Tailoring advanced instructional software for AI. Lecture Notes in Computer Science, 604:604–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tabandeh:1992:IUU

[462] A. S. Tabandeh. Interpreting unexpected user activity in an intelligent user interface. Lecture Notes in Computer Science, 604:614–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mikulecka:1992:YAK

[463] J. Mikulecka and P. Mikulecky. Yet another knowledge-based front-end-to-database project. Lecture Notes in Computer Science, 604:622–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plabmeier:1992:FDB

[464] F. Plabmeier, R. Kuke, D. Exner, and K. F. Lehmann. Fault diagnosis based on simulation models. Lecture Notes in Computer Science, 604:626–??, 1992.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ray:1992:RTF

[465] A. K. Ray and R. B. Misra. Realtime fault diagnosis — using occupancy grids and neural network technology. Lecture Notes in Computer Science, 604:636–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurz:1992:ESF

[466] E. Kurz. Expert systems for fault diagnosis on CNC machines. Lecture Notes in Computer Science, 604:641– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jun:1992:DEL

[467] Gu Jun. Design efficient local search algorithms. Lecture Notes in Computer Science, 604:651–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ketteler:1992:DBP

[468] G. Ketteler and M. Lenart. The design of building parts by using knowledge based systems. Lecture Notes in Computer Science, 604:655–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mason:1992:IFK

[469] K. P. Mason. Information fusion in a knowledge-based classification and tracking system. Lecture Notes in Computer Science, 604:666–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quante:1992:NST

[470] F. Quante, H. Kirsch, and M. Ruckhaberle. Numerical and syntactic tools for fusion and diagnosis provided with the TOPMUSS-System. Lecture Notes in Computer Science, 604:676– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Milzner:1992:LPE

[471] K. Milzner. Learning performance estimations in a knowledge based CADenvironment. Lecture Notes in Computer Science, 604:680–??, 1992. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malhotra:1992:OOF

[472] P. Malhotra and R. E. Seviora. Object oriented framework for generating machine understanding of a digital system design. Lecture Notes in Computer Science, 604:690–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skelton:1992:EPC

[473] C. J. Skelton, C. Hammer M. Lopez, M. J. Reeve, P. Townsend, and K. F. Wong. EDS: A parallel computer system for advanced information processing. Lecture Notes in Computer Science, 605:3–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1992:DAD

[474] Y. Liu, A. K. Singh, and R. L. Bagrodia. A decompositional approach to the design of efficient parallel programs. Lecture Notes in Computer Science, 605:21–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katajainen:1992:SEP

[475] J. Katajainen, C. Levcopoulos, and O. Petersson. Space-efficient parallel merging. Lecture Notes in Computer Science, 605:37–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hromkovic:1992:EIN

[476] J. Hromkovic, V. Muller, O. Sykora, and I. Vrto. On embedding interconnection networks into rings of processors. Lecture Notes in Computer Science, 605:53–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glauert:1992:AMP

[477] J. Glauert. Asynchronous mobile processes and graph rewriting. Lecture Notes in Computer Science, 605:63–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Capello:1992:PIN

[478] F. Capello, J.-L. Bechennec, and J-L Giavitto. PTAH: Introduction to a new parallel architecture for highly numeric processing. Lecture Notes in Computer Science, 605:81–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Osmon:1992:TPP

[479] P. Osmon, T. Stiemerling, T. Valsamidis, A. Whitcroft, T. Wilkinson, and N. Williams. The Topsy Project: a position paper. Lecture Notes in Computer Science, 605:97–??, 1992.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feldmann:1992:OAD

[480] R. Feldmann, J. Hromkovic, S. Madhavapeddy, B. Monien, and P. Mysliwietz. Optimal algorithms for dissemination of information in generalized communication modes. Lecture Notes in Computer Science, 605:115–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Das:1992:EPA

[481] S. K. Das and C. C.-Y. Chen. Efficient parallel algorithms on interval graphs. Lecture Notes in Computer Science, 605:131–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collette:1992:SSC

[482] T. Collette, H. Essafi, D. Juvin, and J. Kaiser. SYMPATIX: a SIMD computer performing the low and intermediate levels of image processing. Lecture Notes in Computer Science, 605: 147–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Getov:1992:DPF

[483] V. Getov. 1-dimensional parallel FFT benchmark on SUPRENUM. Lecture Notes in Computer Science, 605:163– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farrag:1992:DFT

[484] A. A. Farrag, R. Dawson, and Q. Yao. On designing fault-tolerant extensions with optimal fanout for complete bipartite networks. Lecture Notes in Computer Science, 605:177–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhong:1992:ASD

[485] X. Zhong and V. M. Lo. Applicationspecific deadlock free wormhole routing on multicomputers. Lecture Notes in Computer Science, 605:193–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1992:SPM

[486] M. Dubois, L. Barroso, Y.-S. Chen, and K. Oner. Scalability problems in multiprocessors with private caches. Lecture Notes in Computer Science, 605: 211–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chavel:1992:PIO

[487] P. Chavel and J. Taboury. Promises and issues in optical computing. Lecture Notes in Computer Science, 605: 231–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bahsoun:1992:MDR

[488] J.-P. Bahsoun and L. Feraud. A model to design reusable parallel software components. Lecture Notes in Computer Science, 605:245–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:1992:ERO

[489] L. Thomas. Extensibility and reuse of object-oriented synchronization components. Lecture Notes in Computer Science, 605:261–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spiliopoulou:1992:UPP

[490] M. Spiliopoulou, M. Hatzopoulos, and C. Vassilakis. Using parallelism and pipeline for the optimisation of join queries. Lecture Notes in Computer Science, 605:279–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marek:1992:PEP

[491] R. Marek and E. Rahm. Performance evaluation of parallel transaction processing in shared nothing database systems. Lecture Notes in Computer Science, 605:295–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bi:1992:EEM

[492] H. Bi. Explicit expression of multidimensional data parallelism. Lecture Notes in Computer Science, 605:313– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jezequel:1992:PMP

[493] J.-M. Jezequel, F. Berghuel, and F. Andre. Programming massively parallel architectures with sequential object oriented languages. Lecture Notes in Computer Science, 605:329–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OBoyle:1992:NPT

[494] M. O’Boyle and G. A. Hedayat. A new program transformation to minimise communication in distributed memory architectures. Lecture Notes in Computer Science, 605:347–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rudalics:1992:DTE

[495] M. Rudalics. Distributed termination enforcement. Lecture Notes in Computer Science, 605:363–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1992:VSA

[496] F. Lin and T. Shih. Verification of systolic architecture designs. Lecture Notes in Computer Science, 605:381– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeVerge:1992:ROA

[497] H. Le Verge. Reduction operators in ALPHA. Lecture Notes in Computer Science, 605:397–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreau:1992:OSP

[498] L. Moreau. An operational semantics for a parallel functional language with continuations. Lecture Notes in Computer Science, 605:415–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Queinnec:1992:CDE

[499] C. Queinnec. A concurrent and distributed extension of scheme. Lecture Notes in Computer Science, 605:431– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robert:1992:NTC

[500] Y. Robert and S. W. Song. New techniques for cycle shrinking. Lecture Notes in Computer Science, 605:449– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jin:1992:LRT

[501] G. Jin and F. Chen. Loop restructuring techniques for thrashing problem. Lecture Notes in Computer Science, 605:465–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freeman:1992:CPM

[502] E. T. Freeman and D. P. Friedman. Characterizing the paralation model using dynamic assignment. Lecture Notes in Computer Science, 605:483– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cox:1992:ISF

[503] S. Cox, S.-Y. Huang, P. Kelly, J. Liu, and F. Taylor. An implementation of static functional process networks. Lecture Notes in Computer Science, 605: 497–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1992:TCP

[504] A. J. Martin. Translating concurrent programs into VLSI chips. Lecture Notes in Computer Science, 605:515– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wise:1992:MBC

[505] M. J. Wise. Message-brokers and communicating Prolog processes. Lecture Notes in Computer Science, 605:535– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mourlas:1992:EPP

[506] C. Mourlas and C. Halatsis. Extensions to a parallel Prolog system to support real-time applications. Lecture Notes in Computer Science, 605:551–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bic:1992:DPC

[507] L. Bic, J. M. A. Roy, and M. Nagel. Declarative programming for conventional MIMD multiprocessors. Lecture Notes in Computer Science, 605:569– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ning:1992:MLS

[508] Q. Ning and G. R. Gao. Minimizing loop storage allocation for an argument-fetching dataflow architecture model. Lecture Notes in Computer Science, 605:585–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ali:1992:PMB

[509] K. A. M. Ali and R. Karlsson. Performance of Muse on the BBN Butterfly TC2000. Lecture Notes in Computer Science, 605:603–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1992:PFP

[510] G. Gupta and V. Santos Costa. Andor parallelism in full Prolog with paged binding arrays. Lecture Notes in Computer Science, 605:617–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duato:1992:IEV

[511] J. Duato. Improving the efficiency of virtual channels with time-dependent selection functions. Lecture Notes in Computer Science, 605:635–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilson:1992:UOC

[512] G. V. Wilson. Using opportunistic combining networks to reduce contention in multicomputers. Lecture Notes in Computer Science, 605:651– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baig:1992:MMM

[513] M. S. Baig, T. A. El-Ghazawi, and N. A. Alexandridis. Mixed-mode multicomputers with load adaptability. Lecture Notes in Computer Science, 605: 669–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hidaka:1992:APM

[514] Y. Hidaka, H. Koike, and H. Tanaka. Architecture of parallel management kernel fo PIE64. Lecture Notes in Computer Science, 605:685–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meybodi:1992:CDS

[515] M. R. Meybodi. Concurrent data structures for hypercube machine. Lecture Notes in Computer Science, 605:703– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Codenotti:1992:RMS

[516] B. Codenotti, M. Leoncini, and G. Resta. Repeated matrix squaring for the parallel solution of linear systems. Lecture Notes in Computer Science, 605:725–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuntz:1992:PEC

[517] J.-M. Kuntz. Performance evaluation of cache memories in tightly coupled multiprocessor systems. Lecture Notes in Computer Science, 605:735– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krishnamoorthy:1992:ESA

[518] S. Krishnamoorthy and A. Choudhary. An evaluation of set-associativity in two-level caches for shared memory multiprocessors. Lecture Notes in Computer Science, 605:751–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuhn:1992:RCV

[519] E. Kuhn and F. Puntigam. Reliable communication in VPL. Lecture Notes in Computer Science, 605:769– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1992:PSS

[520] P. Anderson, D. Bolton, and P. Kelly. Paragon specifications: Structure, analysis and implementation. Lecture Notes in Computer Science, 605:783– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1992:SDM

[521] A. Gupta. Stanford DASH multiprocessor: The hardware and software approach. Lecture Notes in Computer Science, 605:801–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grefen:1992:DAS

[522] P. W. P. J. Grefen and P. M. G. Apers. Dynamic action scheduling in a parallel database system. Lecture Notes in Computer Science, 605:809–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chachaty:1992:CAD

[523] C. Chachaty, P. Borla-Salamet, and M. Ward. A compositional approach for the design of a parallel query processing language. Lecture Notes in Computer Science, 605:825–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cherief:1992:BFB

[524] F. Cherief. Back and forth bisimulations on prime event structures. Lecture Notes in Computer Science, 605: 843–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fanchon:1992:DCP

[525] J. Fanchon. Dynamic concurrent processes. Lecture Notes in Computer Science, 605:859–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carino:1992:EDP

[526] F. Carino and P. Kostamaa. Exegesis of DBC/1012 and P-90 industrial supercomputer database machines. Lecture Notes in Computer Science, 605:877–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lavington:1992:EPP

[527] S. H. Lavington, M. E. Waite, J. Robinson, and N. E. J. Dewhurst. Exploiting parallelism in primitive operations on bulk data types. Lecture Notes in Computer Science, 605:893–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBosschere:1992:CSM

[528] K. De Bosschere and J.-M. Jacquet. Comparative semantics of microlog. Lecture Notes in Computer Science, 605:911–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Singh:1992:FPC

[529] G. Singh and A. J. Bernstein. A framework for parallel composition of protocols. Lecture Notes in Computer Science, 605:927–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gorlatch:1992:PPD

[530] S. Gorlatch. Parallel program development for a recursive numerical algorithm. Lecture Notes in Computer Science, 605:945–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knabe:1992:DPC

[531] F. Knabe. A distributed protocol for channel-based communication with choice. Lecture Notes in Computer Science, 605:947–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damy:1992:RCP

[532] S. Damy and G. R. Perrin. A refinement of communicating processes. Lecture Notes in Computer Science, 605: 949–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caspi:1992:DFT

[533] P. Caspi and A. Girault. Distributing finite transition systems. Lecture Notes in Computer Science, 605:951– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Capobianchi:1992:CPP

[534] R. Capobianchi, R. Guerraoui, A. Lanusse, and R. Roux. Coping with the process proliferation problem in concurrent object-based language implementations. Lecture Notes in Computer Science, 605:953–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gabarro:1992:CFD

[535] J. Gabarro and R. Gavalda. Correctness of flat data parallel algorithms: An axiomatic approach and examples. Lecture Notes in Computer Science, 605:955–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chamski:1992:GME

[536] Z. Chamski. Generating memoryefficient imperative data structures from systolic programs. Lecture Notes in Computer Science, 605:957– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carrara:1992:GTP

[537] J.-P. Carrara, J.-L. Basille, J.-F. Larue, and D. Juvin. Geometric transforms on parallel architecture. Lecture Notes in Computer Science, 605:959–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kacsuk:1992:ICD

[538] P. Kacsuk. Implementing cut in a distributed data driven OR-Parallel Prolog environment. Lecture Notes in Computer Science, 605:961–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deBeaucoudrey:1992:OIP

[539] N. de Beaucoudrey, A. Bellemain, D. Phalippou, P. Chavel, D. Fortin, I. Exman, L. Rudolph, J.-P. Schnell, and J.-P. Pocholle. Optical interconnects for parallel systems: Demonstration of an optical link with multiplequantum-well opto-electronic arrays.

Lecture Notes in Computer Science, 605:963–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1992:PTA

[540] S. F. Liu and M. L. Soffa. Parallel task assignment by graph partitioning. Lecture Notes in Computer Science, 605:965–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klingler:1992:PRL

[541] A. Klingler and V. Penner. Parallelizing recursive logic programs through decomposition. Lecture Notes in Computer Science, 605:967–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Houzet:1992:PAN

[542] D. Houzet, J.-L. Basille, and J.-Y. Latil. Performance analysis of the network of the GFLOPS parallel architecture. Lecture Notes in Computer Science, 605:969–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weihl:1992:PSP

[543] W. Weihl, E. Brewer, A. Colbrook, Ch. Dellarocas, W. Hsieh, A. Joseph, C. Waldspurger, and P. Wang. PRE-LUDE: A system for portable parallel software. Lecture Notes in Computer Science, 605:971–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dimopoulos:1992:RHD

[544] N. J. Dimopoulos, M. Chowdhury, R. Sivakumar, and V. Dimakopoulos.

Routing in hypercycles. deadlock free and backtracking stategies. Lecture Notes in Computer Science, 605:973– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Justo:1992:SPE

[545] G. R. R. Justo and P. H. Welch. Serialisation as a paradigm for the engineering of parallel programs. Lecture Notes in Computer Science, 605:975– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perret:1992:SSP

[546] J. Perret and Ph. Gautron. SIM-CDL: Simulating parallel programming in a sequential environment. Lecture Notes in Computer Science, 605:977– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karjoth:1992:XFM

[547] G. Karjoth. XFSM: A formal model of communicating state machines for implementation specifications. Lecture Notes in Computer Science, 605:979– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nash:1992:XMP

[548] J. M. Nash and P. M. Dew. XPRAM model and programming interface. Lecture Notes in Computer Science, 605: 981–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knuth:1992:AH

[549] D. E. Knuth. Axioms and hulls. Lecture Notes in Computer Science, 606:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wos:1992:IAL

[550] Larry Wos. The impossibility of the automation of logical reasoning (abstract). Lecture Notes in Computer Science, 607:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ammon:1992:APM

[551] Kurt Ammon. Automatic proofs in mathematical logic and analysis. Lecture Notes in Computer Science, 607: 4–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shang-Ching:1992:PGS

[552] Chou Shang-Ching and Gao Xiao-Shan. Proving geometry statements of constructive type. Lecture Notes in Computer Science, 607:20–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hines:1992:CVS

[553] L. M. Hines. The central variable strategy of strive. Lecture Notes in Computer Science, 607:35–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baader:1992:UUD

[554] Franz Baader and Klaus U. Schulz. Unification in the union of disjoint equational theories: Combining decision procedures. Lecture Notes in Computer Science, 607:50–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nipkow:1992:RUL

[555] Tobias Nipkow and Zhenyu Qian. Reduction and unification in lambda calculi with subtypes. Lecture Notes in Computer Science, 607:66–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dougherty:1992:CLA

[556] Daniel J. Dougherty and Patricia Johann. A combinatory logic approach to higher-order E-unification. Lecture Notes in Computer Science, 607:79–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bibel:1992:CU

[557] Wolfgang Bibel, Steffen Holldobler, and Jorg Wurtz. Cycle unification. Lecture Notes in Computer Science, 607: 94–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yelick:1992:PCP

[558] Katherine A. Yelick and Stephen J. Garland. A parallel completion procedure for term rewriting systems. Lecture Notes in Computer Science, 607: 109–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McAllester:1992:GR

[559] David McAllester. Grammar rewriting. Lecture Notes in Computer Science, 607:124–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cichon:1992:PIC

[560] Adam Cichon and Pierre Lescanne. Polynomial interpretations and the complexity of algorithms. Lecture Notes in Computer Science, 607:139– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fegaras:1992:UTC

[561] Leonidas Fegaras, Tim Sheard, and David Stemple. Uniform traversal combinators: Definition, use and properties. Lecture Notes in Computer Science, 607:148–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uribe:1992:SUU

[562] Tomas E. Uribe. Sorted unification using set constraints. Lecture Notes in Computer Science, 607:163–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frisch:1992:AVS

[563] Alan M. Frisch and Anthony G. Cohn. An abstract view of sorted unification. Lecture Notes in Computer Science, 607:178–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boudet:1992:UOS

[564] Alexandre Boudet. Unification in order-sorted algebras with overloading. Lecture Notes in Computer Science, 607:193–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smullyan:1992:PP

[565] Raymond Smullyan. Puzzles and paradoxes. Lecture Notes in Computer Science, 607:208–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCune:1992:EAD

[566] William McCune and Larry Wos. Experiments in automated deduction with condensed detachment. Lecture Notes in Computer Science, 607:209– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Astrachan:1992:CLM

[567] Owen L. Astrachan and Mark E. Stickel. Caching and lemmaizing in model elimination theorem provers. Lecture Notes in Computer Science, 607:224–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Digricoli:1992:LCP

[568] Vincent J. Digricoli and Eugene Kochendorfer. LIM+ challenge problems by RUE hyper-resolution. Lecture Notes in Computer Science, 607:239– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jackson:1992:CPI

[569] Peter Jackson. Computing prime implicates incrementally. Lecture Notes in Computer Science, 607:253–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sutcliffe:1992:LIS

[570] Geoff Sutcliffe. Linear-input subset analysis. Lecture Notes in Computer Science, 607:268–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benhamou:1992:TSS

[571] Belaid Benhamou and Lakhdar Sais. Theoretical study of symmetries in propositional calculus and applications. Lecture Notes in Computer Science, 607:281–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basin:1992:DM

[572] David Basin and Toby Walsh. Difference matching. Lecture Notes in Computer Science, 607:295–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hesketh:1992:UMR

[573] Jane Hesketh, Alan Bundy, and Alan Smaill. Using middle-out reasoning to control the synthesis of tail-recursive programs. Lecture Notes in Computer Science, 607:310–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walsh:1992:UPP

[574] Toby Walsh, Alex Nunes, and Alan Bundy. The use of proof plans to sum series. Lecture Notes in Computer Science, 607:325–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Protzen:1992:DC

[575] Martin Protzen. Disproving conjectures. Lecture Notes in Computer Science, 607:340–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bauer:1992:IBT

[576] Mathias Bauer. An interval-based temporal logic in a multivalued setting. Lecture Notes in Computer Science, 607:355–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fisher:1992:NFF

[577] Michael Fisher. A normal form for first-order temporal formulae. Lecture Notes in Computer Science, 607:370– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caferra:1992:SEN

[578] Ricardo Caferra and Stephane Demri. Semantic entailment in non classical logics based on proofs found in classical logic. Lecture Notes in Computer Science, 607:385–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inoue:1992:ENF

[579] Katsumi Inoue, Miyuki Koshimura, and Ryuzo Hasegawa. Embedding negation as failure into a model generation theorem prover. Lecture Notes in Computer Science, 607:400–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyer:1992:ACP

[580] Robert S. Boyer and Yu Yuan. Automated correctness proofs of machine code programs for a commercial microprocessor. Lecture Notes in Computer Science, 607:416–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hantao:1992:PCR

[581] Zhang Hantao and Hua Xin. Proving the Chinese remainder theorem by the cover set induction. Lecture Notes in Computer Science, 607:431–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Madden:1992:APO

[582] Peter Madden. Automatic program optimization through proof transformation. Lecture Notes in Computer Science, 607:446–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mints:1992:PST

[583] Grigori Mints. Proof search theory and practice in the (former) USSR (tentative). Lecture Notes in Computer Science, 607:461–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bachmair:1992:BPS

[584] Leo Bachmair, Harald Ganzinger, Christopher Lynch, and Wayne Snyder. Basic paramodulation and superposition. Lecture Notes in Computer Science, 607:462–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nieuwenhuis:1992:TPO

[585] Robert Nieuwenhuis and Albert Rubio. Theorem proving with ordering constrained clauses. Lecture Notes in Computer Science, 607:477–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manna:1992:SRR

[586] Zohar Manna and Richard Waldinger. The special-relation rules are incomplete. Lecture Notes in Computer Science, 607:492–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beckert:1992:IMA

[587] Bernhard Beckert and Reiner Hahnle. An improved method for adding equality to free variable semantic tableaux. Lecture Notes in Computer Science, 607:507–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shankar:1992:PSI

[588] N. Shankar. Proof search in the intuitionistic sequent calculus. Lecture Notes in Computer Science, 607:522– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfenning:1992:IMT

[589] Frank Pfenning and Ekkehard Rohwedder. Implementing the meta-theory of deductive systems. Lecture Notes in Computer Science, 607:537–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1992:TBT

[590] Wilfred Z. Chen. Tactic-based theorem proving and knowledge-based forward chaining: An experiment with Nuprl and Ontic. Lecture Notes in Computer Science, 607:552–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farmer:1992:LT

[591] William M. Farmer, Joshua D. Guttman, and F. Javier Thayer. Little theories. Lecture Notes in Computer Science, 607:567–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christian:1992:STC

[592] Jim Christian. Some termination criteria for narrowing and E-narrowing. Lecture Notes in Computer Science, 607:582–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dershowitz:1992:DMC

[593] Nachum Dershowitz, Subrata Mitra, and G. Sivakumar. Decidable matching for convergent systems. Lecture Notes in Computer Science, 607:589– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kesner:1992:FSO

[594] Delia Kesner. Free sequentiality in orthogonal order-sorted rewriting systems with constructors. Lecture Notes in Computer Science, 607:603– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sekar:1992:PEF

[595] R. C. Sekar and I. V. Ramakrishnan. Programming with equations: A framework for lazy parallel evaluation. Lecture Notes in Computer Science, 607:618–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohn:1992:MSL

[596] Anthony G. Cohn. A many sorted logic with possibly empty sorts. Lecture Notes in Computer Science, 607: 633–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voronkov:1992:TPN

[597] Andrei Voronkov. Theorem proving in non-standard logics based on the inverse method. Lecture Notes in Computer Science, 607:648–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vershinin:1992:OML

[598] Konstantine Vershinin and Igor Romanenko. One more logic with uncertainty and resolution principle for it. Lecture Notes in Computer Science, 607:663–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dafa:1992:NDA

[599] Li Dafa. A natural deduction automated theorem proving system. Lecture Notes in Computer Science, 607: 668–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nipkow:1992:I

[600] Tobias Nipkow and Lawrence C. Paulson. Isabelle-91. Lecture Notes in Computer Science, 607:673–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sutcliffe:1992:SGL

[601] Geoff Sutcliffe. The semantically guided linear deduction system. Lecture Notes in Computer Science, 607: 677–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ammon:1992:SS

[602] Kurt Ammon. The SHUNYATA system. Lecture Notes in Computer Science, 607:681–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shang-Ching:1992:GTP

[603] Chou Shang-Ching. A geometry theorem prover for macintoshes. Lecture Notes in Computer Science, 607:686– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xin:1992:FFR

[604] Hua Xin and Zhang Hantao. FRI: Failure-resistant induction in RRL. Lecture Notes in Computer Science, 607:691–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hantao:1992:HHP

[605] Zhang Hantao. Herky: High performance rewriting in RRL. Lecture Notes in Computer Science, 607:696– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farmer:1992:ISD

[606] William M. Farmer, Joshua D. Guttman, and F. Javier Thayer. IMPS: System description. Lecture Notes in Computer Science, 607:701–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alexander:1992:PET

[607] Geoffrey D. Alexander and David A. Plaisted. Proving equality theorems with hyper-linking. Lecture Notes in Computer Science, 607:706–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chirimar:1992:XGI

[608] Jawahar Chirimar, Carl A. Gunter, and Myra VanInwegen. Xpnet: A graphical interface to proof nets with an efficient proof checker. Lecture Notes in Computer Science, 607:711–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barker-Plummer:1992:AND

[609] Dave Barker-Plummer, Sidney C. Bailin, and Andrew S. Merrill. and: Automated natural deduction. Lecture Notes in Computer Science, 607:716– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uribe:1992:OFF

[610] Tomas E. Uribe, Alan M. Frisch, and Michael K. Mitchell. An overview of FRAPPS 2.0: A framework for resolution-based automated proof procedure systems. Lecture Notes in Computer Science, 607:721–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barker-Plummer:1992:GTP

[611] Dave Barker-Plummer and Alex Rothenberg. The GAZER theorem prover. Lecture Notes in Computer Science, 607:726–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lusk:1992:RPT

[612] Ewing L. Lusk, William W. McCune, and John Slaney. ROO: A parallel theorem prover. Lecture Notes in Computer Science, 607:731–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1992:RAF

[613] T. C. Wang and Allen Goldberg. RVF: An automated formal verification system. Lecture Notes in Computer Science, 607:735–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schumann:1992:KPT

[614] Johann M. Ph. Schumann. KPROP — an AND-Parallel theorem prover for propositional logic implemented in KL1. Lecture Notes in Computer Science, 607:740–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blackburn:1992:RIH

[615] K. Blackburn. A report on ICL HOL. Lecture Notes in Computer Science, 607:743–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Owre:1992:PPV

[616] S. Owre, J. M. Rushby, and N. Shankar. PVS: A prototype verification system. Lecture Notes in Computer Science, 607:748–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reif:1992:KSS

[617] Wolfgang Reif. The KIV system: Systematic construction of verified software. Lecture Notes in Computer Science, 607:753–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beckert:1992:TBT

[618] Bernhard Beckert, Stefan Gerberding, Reiner Hahnle, and Werner Kernig. The tableau-based theorem prover a p_{3T} for multiple-valued logics. Lecture Notes in Computer Science, 607:758– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clarke:1992:ATP

[619] Edmund Clarke and Xudong Zhao. Analytica — A theorem prover in Mathematica. Lecture Notes in Computer Science, 607:761–765, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneider:1992:FP

[620] Klaus Schneider, Ramayya Kumar, and Thomas Kropf. The FAUST-Prover. Lecture Notes in Computer Science, 607:766–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Craigen:1992:ESD

[621] Dan Craigen, Sentot Kromodimoeljo, Irwin Meisels, Bill Pase, and Mark Saaltink. Eves system description. Lecture Notes in Computer Science, 607: 771–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hasegawa:1992:MPT

[622] Ryuzo Hasegawa, Miyuki Koshimura, and Hiroshi Fujita. MGTP: A parallel theorem prover based on lazy model generation. Lecture Notes in Computer Science, 607:776–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lusk:1992:BPW

[623] Ewing L. Lusk and Larry Wos. Benchmark problems in which equality plays the major role. Lecture Notes in Computer Science, 607:781–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Randell:1992:CTT

[624] D. A. Randell, A. G. Cohn, and Z. Cui. Computing transivity tables: A challenge for automated theorem provers. Lecture Notes in Computer Science, 607:786–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1992:ITH

[625] J. R. Anderson. Intelligent tutoring and high school mathematics. Lecture Notes in Computer Science, 608:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breuker:1992:GWC

[626] J. Breuker. Generality watching: ITS caught between science and engineering. Lecture Notes in Computer Science, 608:11–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clancey:1992:GMR

[627] W. J. Clancey. Guidon-Manage revisited: A socio-technical systems approach. Lecture Notes in Computer Science, 608:21–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DuBoulay:1992:PEN

[628] B. Du Boulay. Programming environments for novices. Lecture Notes in Computer Science, 608:37–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greer:1992:GCL

[629] J. E. Greer. Granularity and context in learning. Lecture Notes in Computer Science, 608:38–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lesgold:1992:GIT

[630] A. M. Lesgold. Going from intelligent tutors to tools for learning. Lecture Notes in Computer Science, 608:39–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergeron:1992:AMC

[631] A. Bergeron. Assisted mathematics: The case of discrete probabilities. Lecture Notes in Computer Science, 608:40–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhuiyan:1992:LRT

[632] S. Bhuiyan, J. E. Greer, and G. I. McCalla. Learning recursion through the use of a mental model-based programming environment. Lecture Notes in Computer Science, 608:50–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chevallier:1992:SST

[633] R. Chevallier. STUDIA: un système tutoriel intelligent coopératif fondé sur la négociation et sur un modèle dynamique de dialogue. Lecture Notes in Computer Science, 608:58–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frasson:1992:IID

[634] C. Frasson, M. Kaltenbach, J. Gecsei, and J.-Y. Djamen. An iconic intentiondriven ITS environment. Lecture Notes in Computer Science, 608:66–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirashima:1992:PPE

[635] T. Hirashima, A. Kashihara, and J. Toyoda. Providing problem explanation for ITS. Lecture Notes in Computer Science, 608:76–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurup:1992:FAT

[636] M. Kurup, J. E. Greer, and G. I. Mc-Calla. The fawlty article tutor. Lecture Notes in Computer Science, 608:84–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lamontagne:1992:TEG

[637] C. Lamontagne and J. Bourdeau. Towards an epistemology for guided discovery tutoring: The popperian connection. Lecture Notes in Computer Science, 608:92–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merrill:1992:MPV

[638] D. C. Merrill, B. J. Reiser, R. Beekelaar, and A. Hamid. Making process visible: Scaffolding learning with reasoning-congruent representations. Lecture Notes in Computer Science, 608:103–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moore:1992:WSM

[639] J. L. Moore. Winch simulation: Multiple, linked representations of linear functions. Lecture Notes in Computer Science, 608:111–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sime:1992:LEB

[640] J.-A. Sime and R. Leitch. A learning environment based on multiple qualitative models. Lecture Notes in Computer Science, 608:116–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takeuchi:1992:EEE

[641] A. Takeuchi and S. Otsuki. EX-PITS: An experimental environment on ITS. Lecture Notes in Computer Science, 608:124–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yibin:1992:ITS

[642] M. Yibin and L. Jianxiang. Intelligent tutoring system for symbolic calculation. Lecture Notes in Computer Science, 608:132–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1992:CTK

[643] T.-W. Chan. Curriculum tree: A knowledge-based architecture for intelligent tutoring systems. Lecture Notes in Computer Science, 608:140– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choquet:1992:LUD

[644] C. Choquet, T. Mengelle, G. Gouarderes, and B. Malsallez. L’assistance à l’usager dans DIGITEF: un mécanisme réutilisable. Lecture Notes in Computer Science, 608:148–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chouraqui:1992:RPA

[645] E. Chouraqui and C. Inghilterra. Résolution par analogie de problèmes géometriques dans une perspective tutorielle. Lecture Notes in Computer Science, 608:156–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chu:1992:PDA

[646] R. W. Chu and Ch. M. Mitchell. A pedagogical design and architecture for intelligent tutoring and aiding in supervisory control of complex systems. Lecture Notes in Computer Science, 608:164–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Girard:1992:AM

[647] J. Girard, G. Gauthier, and S. Levesque. Une architecture multiagent. Lecture Notes in Computer Science, 608:172– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herzog:1992:EKS

[648] Ch. Herzog. From elementary knowledge schemes towards heuristic expertise — designing an ITS in the field of parallel programming. Lecture Notes in Computer Science, 608:183–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suthers:1992:ASQ

[649] D. D. Suthers. Answering student queries: Functionality and mechanisms. Lecture Notes in Computer Science, 608:191–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yum:1992:IRA

[650] K.-K. Yum and Th. J. Richards. Instruction as reasoning about multiple objectives. Lecture Notes in Computer Science, 608:199–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gavignet:1992:RCG

[651] E. Gavignet and M. Grandbastien. Représentation des connaissances dans un générateur de systèmes d’E.I.A.O. Lecture Notes in Computer Science, 608:209–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khuwaja:1992:KRI

[652] R. A. Khuwaja, M. W. Evens, A. A. Rovick, and J. A. Michael. Knowledge representation for an intelligent tutoring system based on a multilevel causal model. Lecture Notes in Computer Science, 608:217–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sack:1992:KBC

[653] W. Sack. Knowledge base compilation and the language design game. Lecture Notes in Computer Science, 608: 225–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanMarcke:1992:IE

[654] K. Van Marcke. Instructional expertise. Lecture Notes in Computer Science, 608:234–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1992:CEC

[655] T. Chan, Y. S. Chee, and E. L. Lim. COGNITIO: An extended computational theory of cognition. Lecture Notes in Computer Science, 608:244– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daniel:1992:ADS

[656] M. P. Daniel, L. Nicaud, V. Prince, and M. P. Pery-Woodley. Apport du style linguistique a la modelisation cognitive d’un eleve. Lecture Notes in Computer Science, 608:252–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghemri:1992:CFS

[657] L. Ghemri. A cognitive framework for second language error diagnosis. Lecture Notes in Computer Science, 608: 260–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuzmycz:1992:EFB

[658] M. Kuzmycz and G. I. Webb. Evaluation of feature based modelling in subtraction. Lecture Notes in Computer Science, 608:269–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayers:1992:MAC

[659] A. Mayers and B. Lefebvre. Une modelisation de l’architecture cognitive d’un étudiant pour un système tutoriel intelligent. Lecture Notes in Computer Science, 608:277–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwarz:1992:BMC

[660] B. B. Schwarz, A. S. Kohn, and L. B. Resnick. Bootstrapping mental constructions: A learning system about negative numbers. Lecture Notes in Computer Science, 608:286–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mobus:1992:TTG

[661] C. Mobus, K. Pitschke, and O. Schroder. Towards the theory-guided design of help systems for programming and modelling tasks. Lecture Notes in Computer Science, 608:294–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tattersall:1992:NAI

[662] C. Tattersall. A new architecture for intelligent help systems. Lecture Notes in Computer Science, 608:302–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1992:UT

[663] H. Wang and A. Kushniruk. The UNIX tutor. Lecture Notes in Computer Science, 608:317–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allen:1992:TII

[664] R. Allen, C. Desmoulins, and L. Trilling. Tuteurs intelligents et intelligence artificielle: problèmes posés en construction de figures géométriques. Lecture Notes in Computer Science, 608: 325–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baron:1992:GEA

[665] M. Baron and P. Simonnet. Génération d’exercices en algèbre, premières approches dans le cadre du projet APLUSIX. Lecture Notes in Computer Science, 608:335–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gegg-Harrison:1992:AAD

[666] T. S. Gegg-Harrison. ADAPT: Automated debugging in an adaptive Prolog tutor. Lecture Notes in Computer Science, 608:343–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nicaud:1992:RNG

[667] J.-F. Nicaud. Reference network: A genetic model for intelligent tutoring systems. Lecture Notes in Computer Science, 608:351–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lessard:1992:FFM

[668] G. Lessard, M. Levison, E. Girard, and D. Maher. Form, frequency, markedness and strategies in second language performance modelling. Lecture Notes in Computer Science, 608: 360–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Or-Bach:1992:WSB

[669] R. Or-Bach and E. Bar-On. Why should an ITS bother with students’ explanations? Lecture Notes in Computer Science, 608:372–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Recker:1992:SSL

[670] M. M. Recker and P. Pirolli. Student strategies for learning programming from a computational environment. Lecture Notes in Computer Science, 608:382–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1992:ILT

[671] Y. Wang and R. Garigliano. An intelligent language tutoring system for handling errors caused by transfer. Lecture Notes in Computer Science, 608:

395–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brna:1992:ACD

[672] P. Brna and A. Caiger. The application of cognitive diagnosis to the quantitative analysis of simple electrical circuits. Lecture Notes in Computer Science, 608:405–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corbett:1992:SMM

[673] A. T. Corbett and J. R. Anderson. Student modeling and mastery learning in a computer-based programming tutor. Lecture Notes in Computer Science, 608:413–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dion:1992:AMD

[674] P. Dion and R. Lelouche. Application de la méthodologie du traçage de modèle à un environnement d’apprentissage utilisant une stratégie pédagogique non directive. Lecture Notes in Computer Science, 608:421– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frasson:1992:PDN

[675] C. Frasson and D. Ramazani. Prédiction du niveau d’acquisition des connaissances dans la modelisation de l’étudiant. Lecture Notes in Computer Science, 608:435–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1992:IPU

[676] X. Huang and G. I. McCalla. Instructional planning using focus of attention. Lecture Notes in Computer Science, 608:443–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kono:1992:CHS

[677] Y. Kono, M. Ikeda, and R. Mizoguchi. To contradict is human — student modeling of inconsistency. Lecture Notes in Computer Science, 608:451– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marcenac:1992:MOM

[678] P. Marcenac. Mise en oeuvre d’un modèle de l’apprenant générique dans le système EDDI. Lecture Notes in Computer Science, 608:459–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsuda:1992:SMD

[679] N. Matsuda and T. Okamoto. Student model diagnosis for adaptive instruction in ITS. Lecture Notes in Computer Science, 608:467–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nehemie:1992:SAS

[680] P. Nehemie. A systemic approach for student modeling in a multi-agent aided learning environment. Lecture Notes in Computer Science, 608:475– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Talbi:1992:DCA

[681] M. Talbi and M. Joab. Diagnostic cognitif de l’apprenant par apprentissage symbolique. Lecture Notes in Computer Science, 608:483–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Villano:1992:PSM

[682] M. Villano. Probabilistic student models: Bayesian belief networks and knowledge space theory. Lecture Notes in Computer Science, 608:491– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brusilovsky:1992:FIK

[683] P. L. Brusilovsky. A framework for intelligent knowledge sequencing and task sequencing. Lecture Notes in Computer Science, 608:499–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gutwin:1992:UPM

[684] C. Gutwin and G. I. McCalla. The use of pedagogic misrepresentation in tutorial dialogue. Lecture Notes in Computer Science, 608:507–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Labat:1992:PPE

[685] J.-M. Labat, M. Futtersack, and M. Vivet. Planification pédagogique: de l’expertise humaine à sa modelisation dans un STI. Lecture Notes in Computer Science, 608:515–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Major:1992:CSI

[686] N. Major and H. Reichgelt. COCA: A shell for intelligent tutoring systems. Lecture Notes in Computer Science, 608:523–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohan:1992:EIC

[687] P. Mohan, J. E. Greer, and M. Jones. Enhancing the instructional capabilities of intelligent tutoring systems. Lecture Notes in Computer Science, 608: 531–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pirolli:1992:MLS

[688] P. Pirolli and M. Wilson. Measuring learning strategies and understanding: A research framework. Lecture Notes in Computer Science, 608:539–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sebrechts:1992:TTE

[689] M. M. Sebrechts. From testing to training: Evaluating automated diagnosis in statistics and algebra. Lecture Notes in Computer Science, 608:559–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DelSoldato:1992:DRL

[690] T. Del Soldato. Detecting and reacting to the learner’s motivational state. Lecture Notes in Computer Science, 608:567–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aleven:1992:AGE

[691] V. Aleven and K. D. Ashley. Automated generation of examples for a tutorial in case-based argumentation. Lecture Notes in Computer Science, 608:575–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anjaneyulu:1992:ADF

[692] K. S. R. Anjaneyulu and J. R. Anderson. The advantages of data flow diagrams for beginning programming. Lecture Notes in Computer Science, 608: 585–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murray:1992:TTP

[693] T. Murray and B. P. Woolf. Tools for teacher participation in ITS design. Lecture Notes in Computer Science, 608:593–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valley:1992:EES

[694] K. Valley. Explanation in expert system shells: A tool for exploration and learning. Lecture Notes in Computer Science, 608:601–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bell:1992:ITD

[695] B. L. Bell and R. G. Feifer. Intelligent tutoring with dumb software. Lecture Notes in Computer Science, 608:615– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gutstein:1992:UET

[696] E. Gutstein. Using expert tutor knowledge to design a self-improving intelligent tutoring system. Lecture Notes in Computer Science, 608:625–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Winkels:1992:MEE

[697] R. Winkels and J. Breuker. Modeling expertise for educational purposes. Lecture Notes in Computer Science, 608:633–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1992:DLC

[698] T.-W. Chan, I.-L. Chung, R.-G. Ho, W.-J. Hou, and G.-L. Lin. Distributed learning companion system: WEST revisited. Lecture Notes in Computer Science, 608:643–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dillenbourg:1992:PPH

[699] P. Dillenbourg and J. A. Self. People power: A human-computer collaborative learning system. Lecture Notes in Computer Science, 608:651–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diaz-Ilarraza:1992:UMA

[700] A. Diaz-Ilarraza, J. A. Elorriaga, I. Fernandez-Castro, J. Gutierrez-Serrano, and J. A. Vadillo-Zorita. User modeling and architecture in industrial ITSs. Lecture Notes in Computer Science, 608:661–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eggert:1992:CCL

[701] A. A. Eggert, C. H. Middlecamp, and A. T. Jacob. CHEMPROF: The chemical literacy problem. Lecture Notes in Computer Science, 608:669–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McKendree:1992:GTQ

[702] J. McKendree, B. Radlinski, and M. E. Atwood. The Grace Tutor: A qualified success. Lecture Notes in Computer Science, 608:677–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Best:1992:EBR

[703] E. Best. Esprit basic research action 3148 DEMON (design methods based on nets) — aims, scope and achievements. Lecture Notes in Computer Science, 609:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Best:1992:BCN

[704] E. Best, R. Devillers, and J. G. Hall. The box calculus: a new causal algebra with multi-label communication. Lecture Notes in Computer Science, 609: 21–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broy:1992:MFM

[705] M. Broy and Th. Streicher. Modular functional modelling of Petri nets with individual tokens. Lecture Notes in Computer Science, 609:70–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czaja:1992:ISA

[706] I. Czaja, R. J. van Glabbeek, and U. Goltz. Interleaving semantics and action refinement with atomic choice. Lecture Notes in Computer Science, 609:89–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Devillers:1992:MPS

[707] R. Devillers. Maximality preservation and the ST-idea for action refinements. Lecture Notes in Computer Science, 609:108–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fanchon:1992:FNM

[708] J. Fanchon. A fifo-net model for processes with asynchronous communication. Lecture Notes in Computer Science, 609:152–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hopkins:1992:BNA

[709] R. Hopkins, J. Hall, and O. Botti. A basic-net algebra for program semantics and its application to OC-CAM. Lecture Notes in Computer Science, 609:179–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keesmaat:1992:EVS

[710] N. W. Keesmaat and H. C. M. Kleijn. The effect of vector synchronization: residue and loss. Lecture Notes in Computer Science, 609:215–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koutny:1992:MSD

[711] M. Koutny. Modelling systems with dynamic priorities. Lecture Notes in Computer Science, 609:251–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rozoy:1992:DLM

[712] B. Rozoy. On distributed languages and models for concurrency. Lecture Notes in Computer Science, 609:267– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogler:1992:PWV

[713] W. Vogler. Partial words versus processes: a short comparison. Lecture Notes in Computer Science, 609:292– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernardinello:1992:SBN

[714] L. Bernardinello and F. De Cindio. A survey of basic net models and modular net classes. Lecture Notes in Computer Science, 609:304–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Campos:1992:STP

[715] J. Campos and M. Silva. Structural techniques and performance bounds of stochastic Petri net models. Lecture Notes in Computer Science, 609:352– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gastin:1992:SRL

[716] P. Gastin and A. Petit. A survey of recognizable languages with infinite traces. Lecture Notes in Computer Science, 609:392–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pomello:1992:SEN

[717] L. Pomello, G. Rozenberg, and C. Simone. A survey of equivalence notions for net based systems. Lecture Notes in Computer Science, 609:410–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vonMartial:1992:CPA

[718] F. (Frank) von Martial. Coordinating plans of autonomous agents, volume 610 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 0-387-55615-X (New York), 3-540-55615-X (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 246 pp. LCCN Q337 .M36 1992.

Papazoglou:1992:ING

[719] M. P. Papazoglou and J. Zeleznikow. Introduction: The next generation of information systems — from intelligence to distribution and cooperation. Lecture Notes in Computer Science, 611:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Danforth:1992:DOM

[720] S. Danforth and E. Simon. A data and operation model for advanced database systems. Lecture Notes in Computer Science, 611:9–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

George:1992:OOD

[721] R. George, B. P. Buckles, and F. E. Petry. An object-oriented data model to represent uncertainty in coupled artificial intelligence-database systems. Lecture Notes in Computer Science, 611:37–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jelly:1992:CAD

[722] I. E. Jelly and J. P. Gray. Common architectures for databases and knowledge-based systems. Lecture Notes in Computer Science, 611:49–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Debenham:1992:CMK

[723] J. Debenham. The construction of maintainable knowledge bases. Lecture Notes in Computer Science, 611:58–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuen:1992:AQR

[724] H. S. Yuen, S. Ho, and J. Zeleznikow. Adding qualitative reasoning to an organizational database for management decision support. Lecture Notes in Computer Science, 611:79–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stolze:1992:BHC

[725] M. Stolze, M. Gutknecht, and R. Pfeifer. Building human-centred intelligent cooperative information systems with IKEA. Lecture Notes in Computer Science, 611:104–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Richards:1992:DOQ

[726] T. Richards and L. Richards. Database organization for qualitative analysis: The NUDIST system. Lecture Notes in Computer Science, 611:116–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bland:1992:UPE

[727] R. Bland, J. Cowie, T. Kane, C. Rattray, and I. Wilson. Using a Prolog engine to integrate multiple knowledge sources: The KCM/ help-desk project. Lecture Notes in Computer Science, 611:134–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marinos:1992:ROT

[728] L. Marinos and R. A. Smit. From relations to objects: A translation methodology for an object-oriented frontend to RDBMSs. Lecture Notes in Computer Science, 611:148–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vittal:1992:FCA

[729] J. Vittal, B. Silver, W. Frawley, G. Iba, T. Fawcett, S. Dusseault, and J. Doleac. A framework for cooperative adaptable information systems. Lecture Notes in Computer Science, 611:169–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirn:1992:PSF

[730] S. Kirn, A. Scherer, and G. Schlageter. Problem solving in federative environments: The FRESCO concept of cooperative agents. Lecture Notes in Computer Science, 611:185–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernus:1992:HDI

[731] P. Bernus. Heterogeneous database integration architecture based on a conversation theoretic skeleton. Lecture Notes in Computer Science, 611:204– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laufmann:1992:CGD

[732] S. C. Laufmann. Coarse-grained distributed agents for transparent access to remote systems. Lecture Notes in Computer Science, 611:223–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber:1992:FCI

[733] J. C. Weber. A forward-chaining information framework. Lecture Notes in Computer Science, 611:238–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Werkman:1992:UNC

[734] K. J. Werkman. Using negotiation and coordination in multi-agent intelligent cooperative information systems. Lecture Notes in Computer Science, 611:251–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsai:1992:DCA

[735] J. J.-P. Tsai and R.-Y. Sheu. A distributed cooperative agents architecture for software development. Lecture Notes in Computer Science, 611:271– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karagiannis:1992:KSL

[736] D. Karagiannis, F. J. Kurfeb, and H.- W. Schmidt. Knowledge selection in large knowledge bases. Lecture Notes in Computer Science, 611:291–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nierstrasz:1992:TOC

[737] Oscar Nierstrasz. Towards an object calculus. Lecture Notes in Computer Science, 612:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Honda:1992:ACS

[738] Kohei Honda and Mario Tokoro. On asynchronous communication semantics. Lecture Notes in Computer Science, 612:21–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papthomas:1992:UFP

[739] Michael Papthomas. A unifying framework for process calculus semantics of concurrent object-oriented languages. Lecture Notes in Computer Science, 612:53–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolfram:1992:SSF

[740] D. A. Wolfram and Joseph A. Goguen. A sheaf semantics for FOOPS expressions. Lecture Notes in Computer Science, 612:81–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shibayama:1992:SLO

[741] Etsuya Shibayama. Semantic layers of object-based concurrent computing. Lecture Notes in Computer Science, 612:99–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

America:1992:FTP

[742] Pierre America. Formal techniques for parallel object-oriented languages. Lecture Notes in Computer Science, 612: 119–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vasconcelos:1992:TSA

[743] Vasco Vasconcelos and Mario Tokoro. Traces semantics for actor systems. Lecture Notes in Computer Science, 612:141–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andreoli:1992:DPM

[744] Jean-Marc Andreoli, Remo Pareschi, and Marc Bourgois. Dynamic programming as multiagent programming. Lecture Notes in Computer Science, 612: 163–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McHale:1992:SP

[745] Ciaran McHale, Bridget Walsh, Sean Baker, and Alexis Donnelly. Scheduling predicates. Lecture Notes in Computer Science, 612:177–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saleh:1992:CCM

[746] Hayssam Saleh and Philippe Gautron. A concurrency control mechanism for C++ objects. Lecture Notes in Computer Science, 612:195–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsuoka:1992:OOC

[747] Satoshi Matsuoka, Takuo Watanabe, Yuuji Ichisugi, and Akinori Yonezawa. Object-oriented concurrent reflective architectures. Lecture Notes in Computer Science, 612:211–??, 1992. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartmann:1992:ADD

[748] Thorsten Hartmann and Ralf Jungclaus. Abstract description of distributed object systems. Lecture Notes in Computer Science, 612:227– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wegner:1992:DIO

[749] Peter Wegner. Design issues for objectbased concurrency. Lecture Notes in Computer Science, 612:245–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1992:PWO

[750] Anonymous. Panel: What is an object? Lecture Notes in Computer Science, 612:257–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurtz:1992:CFS

[751] Stuart A. Kurtz, John C. Mitchell, and Michael J. O’Donnell. Connecting formal semantics to constructive intuitions. Lecture Notes in Computer Science, 613:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lipton:1992:KSD

[752] James Lipton. Kripke semantics for dependent type theory and realizability interpretations. Lecture Notes in Computer Science, 613:22–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1992:RSC

[753] Scott F. Smith. Reflective semantics of constructive type theory: Preliminary report. Lecture Notes in Computer Science, 613:33–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thompson:1992:SNM

[754] Simon Thompson. Are subsets necessary in Martin-Löf type theory? Lecture Notes in Computer Science, 613: 46–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jianguo:1992:DTB

[755] Lu Jianguo and Xu Jiafu. Development transformation based on higher order type theory. Lecture Notes in Computer Science, 613:58–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murthy:1992:CPP

[756] Chetan R. Murthy. Classical proofs as programs: How, what, and why. Lecture Notes in Computer Science, 613: 71–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Napierala:1992:CTT

[757] Maria Napierala. Classical type theory. Lecture Notes in Computer Science, 613:89–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yong:1992:ACC

[758] Sun Yong. Axiomatization of calculus of constructions. Lecture Notes in Computer Science, 613:108–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Swarup:1992:LVA

[759] Vipin Swarup and Uday S. Reddy. A logical view of assignments. Lecture Notes in Computer Science, 613:131– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fellows:1992:CIG

[760] Michael R. Fellows and Michael A. Langston. Constructivity issues in graph algorithms. Lecture Notes in Computer Science, 613:150–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coquand:1992:CTC

[761] Thierry Coquand. Constructive topology and combinatorics. Lecture Notes in Computer Science, 613:159– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chirimar:1992:ICR

[762] Jawahar Chirimar and Douglas J. Howe. Implementing constructive real analysis: Preliminary report. Lecture Notes in Computer Science, 613:165– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tucker:1992:ESS

[763] J. V. Tucker and J. I. Zucker. Examples of semicomputable sets of real and complex numbers. Lecture Notes in Computer Science, 613:179–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greenleaf:1992:BME

[764] Newcomb Greenleaf. Bringing mathematics education into the algorithmic age. Lecture Notes in Computer Science, 613:199–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Myers:1992:TSC

[765] J. Paul Myers and Ronald E. Prather. The type structure of CAT. Lecture Notes in Computer Science, 613:218– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weihrauch:1992:SPA

[766] Klaus Weihrauch and P. Krolak. A simple and powerful approach for studying constructivity, computability, and complexity. Lecture Notes in Computer Science, 613:228–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanko:1992:WST

[767] J. G. Hanko, E. M. Kuerner, J. D. Northcutt, and G. A. Wall. Workstation support for time-critical applications. Lecture Notes in Computer Science, 614:4–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeffay:1992:KSL

[768] K. Jeffay, D. L. Stone, and F. D. Smith. Kernel support for live digital audio and video. Lecture Notes in Computer Science, 614:10–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mercer:1992:PCP

[769] C. W. Mercer and H. Tokuda. Priority consistency in protocol architectures. Lecture Notes in Computer Science, 614:22–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hehmann:1992:IHA

[770] D. Hehmann, R. G. Herrtwich, W. Schulz, T. Schutt, and R. Steinmetz. Implementing HeiTS architecture and implementation strategy of the Heidelberg High-Speed Transport System. Lecture Notes in Computer Science, 614:33–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coulson:1992:PSD

[771] G. Coulson, F. Garcia, D. Hutchison, and D. Shepherd. Protocol support for distributed multimedia applications. Lecture Notes in Computer Science, 614:45–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruston:1992:ICT

[772] L. Ruston, G. Blair, G. Coulson, and N. Davies. Integrating computing and telecommunications — A tale of two architectures. Lecture Notes in Computer Science, 614:57–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrari:1992:DAD

[773] D. Ferrari. Design and applications of a delay jitter control scheme for packetswitching internetworks. Lecture Notes in Computer Science, 614:72–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Almeida:1992:EES

[774] N. Almeida, J. Cabral, and A. Alves. End-to-end synchronization in packetswitched networks. Lecture Notes in Computer Science, 614:84–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leydekkers:1992:SMD

[775] P. Leydekkers and B. Teunissen. Synchronization of multimedia data streams in open distributed environments. Lecture Notes in Computer Science, 614:94–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bulterman:1992:MSU

[776] D. C. A. Bulterman and R. van Liere. Multimedia synchronization and UNIX. Lecture Notes in Computer Science, 614:108–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Little:1992:SBC

[777] T. D. C. Little and A. Ghafoor. Scheduling of bandwidth-constrained multimedia traffic. Lecture Notes in Computer Science, 614:120–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoepner:1992:PSM

[778] P. Hoepner. Presentation scheduling of multimedia objects and its impact on network and operating system support. Lecture Notes in Computer Science, 614:132–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Field:1992:ACN

[779] B. Field and T. Znati. Alpha-channel — A network level abstraction to support real-time communication. Lecture Notes in Computer Science, 614:148– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Banerjea:1992:RTC

[780] A. Banerjea and B. A. Mah. The real-time channel administration protocol. Lecture Notes in Computer Science, 614:160–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolfinger:1992:CMD

[781] B. Wolfinger and M. Moran. A continuous media data transport service and protocol for real-time communication in high-speed networks. Lecture Notes in Computer Science, 614:171– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gusella:1992:DCM

[782] R. Gusella and M. Maresca. Design considerations for a multimedia network distribution center. Lecture Notes in Computer Science, 614:185– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stevens:1992:NGN

[783] S. M. Stevens. Next generation network and operating system requirements for continuous time media. Lecture Notes in Computer Science, 614: 197–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gutfreund:1992:DIB

[784] S. Gutfreund, J. Diaz-Gonzalez, R. Sasnett, and V. Phuah. Dynamicity issues in broadband network computing. Lecture Notes in Computer Science, 614:209–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cordes:1992:MMS

[785] R. Cordes, D. Wybranietz, and R. Vautz. Managing multimedia sessions on a private broadband communication system. Lecture Notes in Computer Science, 614:217–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1992:TSM

[786] D. P. Anderson and P. Chan. Toolkit support for multiuser audio/video applications. Lecture Notes in Computer Science, 614:230–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Northcutt:1992:SST

[787] J. D. Northcutt and E. M. Kuerner. System support for time-critical applications. Lecture Notes in Computer Science, 614:242–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frey:1992:AFM

[788] S. Frey and D. P. Ingold. An application framework for multimedia communication. Lecture Notes in Computer Science, 614:255–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gibbs:1992:PEM

[789] S. Gibbs, C. Breiteneder, L. Dami, V. de Mey, and D. Tsichritzis. A programming environment for multimedia applications. Lecture Notes in Computer Science, 614:265–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blakowski:1992:TSE

[790] G. Blakowski, J. Hubel, and U. Langrehr. Tools for specifying and executing synchronized multimedia presentations. Lecture Notes in Computer Science, 614:271–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steinmetz:1992:ACM

[791] R. Steinmetz and J. C. Fritzsche. Abstractions for continuous-media programming. Lecture Notes in Computer Science, 614:283–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanko:1992:IMS

[792] J. G. Hanko, D. Berry, T. Jacobs, and D. Steinberg. Integrated multimedia at SUN microsystems. Lecture Notes in Computer Science, 614:300–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meibner:1992:AAM

[793] K. Meibner. Architectural aspects of multimedia CD-I integration in UNIX/ X-Windows workstations. Lecture Notes in Computer Science, 614:314– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lamparter:1992:XMT

[794] B. Lamparter and W. Effelsberg. Xmovie transmission and presentation of digital movies under X. Lecture Notes in Computer Science, 614:328– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rosenberg:1992:PMD

[795] J. Rosenberg, G. Cruz, and T. Judd. Presenting multimedia documents over a digital network. Lecture Notes in Computer Science, 614:346–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hinz:1992:CCP

[796] R. Hinz. The CCITT communication protocol for videophone teleconferencing equipment. Lecture Notes in Computer Science, 614:358–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zolzer:1992:FBD

[797] U. Zolzer and N. Kalff. FDDI-Based digital audio interfaces. Lecture Notes in Computer Science, 614:368– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cousin:1992:DVT

[798] B. Cousin. Digital video transmission and the FDDI token ring protocol. Lecture Notes in Computer Science, 614:375–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biersack:1992:PSF

[799] E. W. Biersack. A performance study of forward error correction in ATM networks. Lecture Notes in Computer Science, 614:391–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1992:WW

[800] Anonymous. Workshop wrap-up. Lecture Notes in Computer Science, 614: 403–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trappl:1992:AIP

[801] R. Trappl. AI: Introduction, paradigms, applications (including CBR), impacts, visions. Lecture Notes in Computer Science, 617:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Havel:1992:AIC

[802] I. M. Havel. Artificial intelligence and connectionism: Some philosophical implications. Lecture Notes in Computer Science, 617:25–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gilbert:1992:LRI

[803] D. R. Gilbert and C. J. Hogger. Logic for representing and implementing knowledge about system behaviour. Lecture Notes in Computer Science, 617:42–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stepankova:1992:PST

[804] O. Stepankova and P. Stepanek. Prolog: A step towards the future of programming. Lecture Notes in Computer Science, 617:50–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kriwaczek:1992:ICL

[805] F. Kriwaczek. An introduction to constraint logic programming. Lecture Notes in Computer Science, 617:82–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eder:1992:LD

[806] J. Eder. Logic and databases. Lecture Notes in Computer Science, 617:95–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kubat:1992:IML

[807] M. Kubat. Introduction to machine learning. Lecture Notes in Computer Science, 617:104–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brazdil:1992:AIL

[808] P. B. Brazdil. Approaches to inductive logic programming. Lecture Notes in Computer Science, 617:139–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moscatelli:1992:AML

[809] S. Moscatelli and Y. Kodratoff. Advanced machine learning techniques for computer vision. Lecture Notes in Computer Science, 617:161–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steel:1992:NCT

[810] S. Steel. Notes on current trends in AI planning. Lecture Notes in Computer Science, 617:198–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelleher:1992:ARM

[811] G. Kelleher. The application of reason maintenance systems in planning and scheduling. Lecture Notes in Computer Science, 617:208–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lazansky:1992:PAP

[812] J. Lazansky. Practical applications of planning tasks. Lecture Notes in Computer Science, 617:238–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klir:1992:RUM

[813] G. J. Klir. The role of uncertainty measures and principles in AI. Lecture Notes in Computer Science, 617: 245–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jirousek:1992:IPM

[814] R. Jirousek. Introduction to probabilistic methods of knowledge representation and processing. Lecture Notes in Computer Science, 617:255–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hajek:1992:BF

[815] P. Hajek and D. Harmanec. On belief functions. Lecture Notes in Computer Science, 617:286–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kouba:1992:DAU

[816] Z. Kouba. Data analysis and uncertainty processing. Lecture Notes in Computer Science, 617:308–??, 1992.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marik:1992:SAK

[817] V. Marik and T. Vlcek. Some aspects of knowledge engineering. Lecture Notes in Computer Science, 617: 316–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Findler:1992:AKA

[818] N. V. Findler. An automatic knowledge acquisition tool. Lecture Notes in Computer Science, 617:338–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doran:1992:DAA

[819] J. Doran. Distributed AI and its applications. Lecture Notes in Computer Science, 617:368–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cuena:1992:ASG

[820] J. Cuena. Architectures for second generation knowledge based systems. Lecture Notes in Computer Science, 617:373–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stepankova:1992:IQR

[821] O. Stepankova. An introduction to qualitative reasoning. Lecture Notes in Computer Science, 617:404–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mozetic:1992:MBD

[822] I. Mozetic. Model-based diagnosis: An overview. Lecture Notes in Computer Science, 617:419–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bratko:1992:DSS

[823] I. Bratko. Dynamic system simulation with qualitative differential equations. Lecture Notes in Computer Science, 617:431–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dorffner:1992:INP

[824] G. Dorffner. An introduction to neurocomputing and its possible role in AI. Lecture Notes in Computer Science, 617:440–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Edmonds:1992:IPD

[825] E. A. Edmonds and B. S. Murray. Intelligent presentation and dialogue management. Lecture Notes in Computer Science, 617:465–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hajicova:1992:LAN

[826] E. Hajicova. Linguistic aspects of natural language processing. Lecture Notes in Computer Science, 617:477– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brodie:1992:PDC

[827] M. L. Brodie. The promise of distributed computing and the challenges of legacy systems. Lecture Notes in Computer Science, 618:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gradwell:1992:OOR

[828] D. J. L. Gradwell. Object-oriented requirements capture and analysis — the Orca project. Lecture Notes in Computer Science, 618:29–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monk:1992:MVC

[829] S. R. Monk and I. Sommerville. A model for versioning of classes in object-oriented databases. Lecture Notes in Computer Science, 618:42–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tuv:1992:SMH

[830] E. Tuv, A. Poulovassilis, and M. Levene. A storage manager for the hypernode model. Lecture Notes in Computer Science, 618:59–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saake:1992:VFI

[831] G. Saake and R. Jungclaus. Views and formal implementation in threelevel schema architecture for dynamic objects. Lecture Notes in Computer Science, 618:78–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holman:1992:MCS

[832] A. Holman. The Meiko computing surface: A parallel and scalable open systems platform for Oracle. Lecture Notes in Computer Science, 618:96–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Page:1992:SPD

[833] J. Page. A study of a parallel database machine and its performance — the NCR/Teradata DBC/1012. Lecture Notes in Computer Science, 618: 115–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Unwalla:1992:CLM

[834] M. Unwalla and J. Kerridge. Control of a large massively parallel database machine using SQL catalogue extensions, and a DSDL in preference to an operating system. Lecture Notes in Computer Science, 618:138–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sutton:1992:IML

[835] D. R. Sutton and P. J. H. King. Integration of modal logic and the functional data model. Lecture Notes in Computer Science, 618:156–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kay:1992:RAM

[836] M. H. Kay, P. J. Rivett, and T. J. Walters. The raleigh activity model: Integrating versions, concurrency, and access control. Lecture Notes in Computer Science, 618:175–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cooper:1992:GDM

[837] R. Cooper and Z. Qin. A graphical data modelling program with constraint specification and management. Lecture Notes in Computer Science, 618:192–??, 1992. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Qutaishat:1992:AMS

[838] M. A. Qutaishat, N. J. Fiddian, and W. A. Gray. Association merging in a schema meta-integration system for a heterogeneous object-oriented database environment. Lecture Notes in Computer Science, 618:209–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diaz:1992:GAR

[839] O. Diaz and S. M. Embury. Generating active rules from high-level specifications. Lecture Notes in Computer Science, 618:227–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fonkam:1992:EIC

[840] M. M. Fonkam and W. A. Gray. Employing integrity constraints for query modification and intensional answer generation in multi-database systems. Lecture Notes in Computer Science, 618:244–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orlowska:1992:AAR

[841] Ewa Orlowska. Algebraic aspects of the relational knowledge representation: Modal relation algebras. Lecture Notes in Computer Science, 619:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schumacher:1992:LM

[842] Christian Schumacher. A logic for memory. Lecture Notes in Computer Science, 619:23–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morreau:1992:APP

[843] Michael Morreau. Actions with preconditions and postconditions. Lecture Notes in Computer Science, 619:46–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gehne:1992:TCW

[844] Jurgen Gehne. Testclasses and closed world assumptions for non-Horn theories. Lecture Notes in Computer Science, 619:56–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pearce:1992:RNI

[845] David Pearce. Reasoning with negative information, II: Hard negation, strong negation and logic programs. Lecture Notes in Computer Science, 619:63–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1992:LAS

[846] Gerd Wagner. Lindenbaum-algebraic semantics of logic programs. Lecture Notes in Computer Science, 619:80–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuhrmann:1992:CLC

[847] Andre Fuhrmann. Conditional logics and cumulative logics. Lecture Notes in Computer Science, 619:92–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herre:1992:SNR

[848] Heinrich Herre. Semantics of nonmonotonic reasoning in logic programming. Lecture Notes in Computer Science, 619:104–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wansing:1992:FTH

[849] Heinrich Wansing. Formulas-as-types for a hierarchy of sublogics of intuitionistic propositional logic. Lecture Notes in Computer Science, 619:125– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schroeder-Heister:1992:CEL

[850] Peter Schroeder-Heister. Cut elimination for logics with definitional reflection. Lecture Notes in Computer Science, 619:146–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Archangelsky:1992:MLL

[851] D. A. Archangelsky and M. A. Taitslin. Modal linear logic. Lecture Notes in Computer Science, 620:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baliga:1992:MLH

[852] G. Baliga, J. Case, S. Jain, and M. Suraj. Machine learning of higher order programs. Lecture Notes in Computer Science, 620:9–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beigel:1992:QAV

[853] R. Beigel, M. Kummer, and F. Stephan. Quantifying the amount of verboseness. Lecture Notes in Computer Science, 620:21–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benton:1992:SLP

[854] P. N. Benton. Strictness logic and polymorphic invariance. Lecture Notes in Computer Science, 620:33–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1992:PLN

[855] A. L. Brown, S. Mantha, and T. Wakayama. Preference logics and non-monotonicity in logic programming. Lecture Notes in Computer Science, 620:45–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calo:1992:EFG

[856] A. Calo and J. A. Makowsky. The Ehrenfeucht-Fraisse games for transitive closure. Lecture Notes in Computer Science, 620:57–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cerans:1992:FFI

[857] K. Cerans. Feasibility of finite and infinite paths in data dependent programs. Lecture Notes in Computer Science, 620:69–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1992:IMR

[858] L. Chen. An interleaving model for real-time systems. Lecture Notes in Computer Science, 620:81–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christensen:1992:LCA

[859] S. Christensen. A logical characterization of asynchronously communicating agents. Lecture Notes in Computer Science, 620:93–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deGroote:1992:DCP

[860] P. de Groote. Denotations for classical proofs — preliminary results. Lecture Notes in Computer Science, 620: 105–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dershowitz:1992:OAL

[861] N. Dershowitz and E. M. Reingold. Ordinal arithmetic with list structures. Lecture Notes in Computer Science, 620:117–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Edalat:1992:CC

[862] A. Edalat. Continuous I-categories. Lecture Notes in Computer Science, 620:127–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fitting:1992:MVN

[863] M. Fitting. Many-valued nonmonotonic modal logics. Lecture Notes in Computer Science, 620:139– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galmiche:1992:ADA

[864] D. Galmiche and G. Perrier. Automated deduction in additive and multiplicative linear logic. Lecture Notes in Computer Science, 620:151–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gamatie:1992:ISF

[865] B. Gamatie. Intensionally stable functions. Lecture Notes in Computer Science, 620:163–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1992:CPT

[866] A. Gupta. A constructive proof that trees are well-quasi-ordered under minors. Lecture Notes in Computer Science, 620:174–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hemachandra:1992:BRT

[867] L. A. Hemachandra, S. Jain, and N. K. Vereshchagin. Banishing robust Turing completeness. Lecture Notes in Computer Science, 620:186–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirokawa:1992:BFB

[868] S. Hirokawa. Balanced formulas, BCK-Minimal formulas and their proofs. Lecture Notes in Computer Science, 620:198–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoofman:1992:NSM

[869] R. Hoofman. Non-stable models of linear logic. Lecture Notes in Computer Science, 620:209–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

King:1992:OOC

[870] A. King and P. Soper. Ordering optimizations for concurrent logic programs. Lecture Notes in Computer Science, 620:221–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knijnenburg:1992:CIP

[871] P. M. W. Knijnenburg and F. Nordemann. A categorical interpretation of partial function logic and Hoare logic. Lecture Notes in Computer Science, 620:229–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kossovsky:1992:PCC

[872] N. K. Kossovsky and A. B. Prokhoroff. The polynomial complexity of conjunctive normal form satisfiability, when the number of conjunctions and negations is limited. Lecture Notes in Computer Science, 620:241–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kotov:1992:TLC

[873] S. V. Kotov. Typed lambda-calculus with recursive definitions. Lecture Notes in Computer Science, 620:246– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lano:1992:STF

[874] K. Lano. Set theoretic foundations for fuzzy set theory, and their applications. Lecture Notes in Computer Science, 620:258–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lesske:1992:CSA

[875] F. Lesske. Constructive specifications of abstract data types using temporal logic. Lecture Notes in Computer Science, 620:269–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1992:IBM

[876] R. Li and J. Carmo. An intervalbased modal logic for system specification. Lecture Notes in Computer Science, 620:281–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luo:1992:UTD

[877] Z. Luo. A unifying theory of dependent types: The schematic approach. Lecture Notes in Computer Science, 620:293–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makarov:1992:MMS

[878] V. P. Makarov. MSL — A mathematical specification language. Lecture Notes in Computer Science, 620:305– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Megrelis:1992:PAO

[879] A. Megrelis. Partial algebra + ordersorted algebra = galactic algebra. Lecture Notes in Computer Science, 620: 314–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Momigliano:1992:MNH

[880] A. Momigliano. Minimal negation and hereditary harrop formulae. Lecture Notes in Computer Science, 620:326– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mullins:1992:KAR

[881] J. Mullins. Kleene automata and recursion theory. Lecture Notes in Computer Science, 620:336–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mycroft:1992:IPT

[882] A. Mycroft. Incremental polymorphic type checking with update. Lecture Notes in Computer Science, 620:347– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plaza:1992:OLO

[883] J. A. Plaza. Operators on lattices of omega-Herbrand interpretations. Lecture Notes in Computer Science, 620: 358–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pliuskeviciene:1992:SCP

[884] A. Pliuskeviciene. Sequential calculus for proving the properties of regular programs. Lecture Notes in Computer Science, 620:370–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pliuskevicius:1992:CSC

[885] R. Pliuskevicius. Complete sequential calculi for the first order symmetrical linear temporal logic with until and since. Lecture Notes in Computer Science, 620:382–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rabinovich:1992:NME

[886] A. Rabinovich. Non modularity and expressibility for nets of relations. Lecture Notes in Computer Science, 620: 394–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reif:1992:CGM

[887] W. Reif. Correctness of generic modules. Lecture Notes in Computer Science, 620:406–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sadi:1992:PCS

[888] K. Sadi and F. Meunier. An andparallelism cooperative scheme for full Prolog interpreters on a Transputer-Based architecture. Lecture Notes in Computer Science, 620:418–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakalauskaite:1992:SCFb

[889] J. Sakalauskaite. A sequent calculus for a first order linear temporal logic with equality. Lecture Notes in Computer Science, 620:430–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schlingloff:1992:EPM

[890] B.-H. Schlingloff. On the expressive power of modal logics on trees. Lecture Notes in Computer Science, 620: 441–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shilov:1992:PDL

[891] N. V. Shilov. Propositional dynamic logic with fixed points: Algorithmic tools for verification of finite state machines. Lecture Notes in Computer Science, 620:452–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spreen:1992:EOC

[892] D. Spreen. Effective operators and continuity revisited. Lecture Notes in Computer Science, 620:459–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stewart:1992:LCB

[893] I. A. Stewart. Logical characterizations of bounded query classes I: Logspace oracle machines. Lecture Notes in Computer Science, 620:470–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tiuryn:1992:SEC

[894] J. Tiuryn. Solving equational constraints in polymorphic types. Lecture Notes in Computer Science, 620:480– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tupailo:1992:GSN

[895] S. Tupailo. Gentzen-style and Novikovstyle cut-elimination. Lecture Notes in Computer Science, 620:493–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanderHoek:1992:GME

[896] W. van der Hoek and J.-J. Ch. Meyer. Graded modalities in epistemic logic. Lecture Notes in Computer Science, 620:503–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerbessiotis:1992:DBS

[897] A. V. Gerbessiotis and L. G. Valiant. Direct bulk-synchronous parallel algorithms (invited). Lecture Notes in Computer Science, 621:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freivalds:1992:MLI

[898] R. Freivalds and C. H. Smith. Memory limited inductive inference machines. Lecture Notes in Computer Science, 621:19–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fich:1992:RSI

[899] F. Fich, M. Kowaluk, K. Loryś, M. Kutylowski, and P. Ragde. Retrieval of scattered information by EREW, CREW and CRCW PRAMs. Lecture Notes in Computer Science, 621:30–41, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Razborov:1992:SDT

[900] A. A. Razborov. On small depth threshold circuits (invited). Lecture Notes in Computer Science, 621:42–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pippenger:1992:EAS

[901] N. Pippenger. An elementary approach to some analytic asymptotics. Lecture Notes in Computer Science, 621:53–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czumaj:1992:OPA

[902] A. Czumaj. An optimal parallel algorithm for computing a near-optimal order of matrix multiplications. Lecture Notes in Computer Science, 621:62–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kortsarz:1992:GSS

[903] G. Kortsarz and D. Peleg. Generating sparse 2-Spanners. Lecture Notes in Computer Science, 621:73–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Awerbuch:1992:LDG

[904] B. Awerbuch, B. Berger, L. Cowen, and D. Peleg. Low-diameter graph decomposition is in NC. Lecture Notes in Computer Science, 621:83–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

He:1992:PAC

[905] X. He. Parallel algorithm for cograph recognition with applications. Lecture Notes in Computer Science, 621:94–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghosh:1992:PAA

[906] S. K. Ghosh and A. Maheshwari. Parallel algorithms for all minimum link paths and link center problems. Lecture Notes in Computer Science, 621:

106–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaufmann:1992:OMP

[907] M. Kaufmann and J. F. Sibeyn. Optimal multi-packet routing on the torus. Lecture Notes in Computer Science, 621:118–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinotti:1992:PAP

[908] M. C. Pinotti and G. Pucci. Parallel algorithms for priority queue operations. Lecture Notes in Computer Science, 621:130–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dietz:1992:HCP

[909] P. F. Dietz. Heap construction in the parallel comparison tree model. Lecture Notes in Computer Science, 621: 140–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyar:1992:ERC

[910] J. Boyar and K. S. Larsen. Efficient rebalancing of chromatic search trees. Lecture Notes in Computer Science, 621:151–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jakoby:1992:CSP

[911] A. Jakoby and R. Reischuk. The complexity of scheduling problems with communication delays for trees. Lecture Notes in Computer Science, 621: 165–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

dAmore:1992:LUP

[912] F. d’Amore and V. Liberatore. The list update problem and the retrieval of sets. Lecture Notes in Computer Science, 621:178–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cunto:1992:GTB

[913] W. Cunto and V. Yriarte. GKD-Trees: Binary trees that combine multi-dimensional data handling, node size, and fringe reorganization. Lecture Notes in Computer Science, 621: 192–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sen:1992:FCS

[914] S. Sen. Fractional cascading simplified. Lecture Notes in Computer Science, 621:212–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giammarresi:1992:DCP

[915] D. Giammarresi and G. F. Italiano. Dynamic 2- and 3-Connectivity on planar graphs. Lecture Notes in Computer Science, 621:221–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hershberger:1992:FDC

[916] J. Hershberger, M. Rauch, and S. Suri. Fully dynamic 2-Connectivity in planar graphs. Lecture Notes in Computer Science, 621:233–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McDiarmid:1992:NIN

[917] C. McDiarmid, B. Reed, A. Schrijver, and B. Shepherd. Non-interfering network flows. Lecture Notes in Computer Science, 621:245–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kant:1992:TPG

[918] G. Kant and H. L. Bodlaender. Triangulating planar graphs while minimizing the maximum degree. Lecture Notes in Computer Science, 621:258– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertolazzi:1992:HDS

[919] P. Bertolazzi, R. F. Cohen, G. Di Battista, R. Tamassia, and I. G. Tollis. How to draw a series-parallel digraph. Lecture Notes in Computer Science, 621:272–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Furer:1992:CRG

[920] M. Furer and C. R. Subramanian. Coloring random graphs. Lecture Notes in Computer Science, 621:284–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kloks:1992:TSK

[921] T. Kloks and H. Bodlaender. Testing superperfection of kappa-trees. Lecture Notes in Computer Science, 621:292– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez-Baca:1992:PPG

[922] D. Fernandez-Baca and G. Slutzki. Parametric problems on graphs of bounded tree-width. Lecture Notes in Computer Science, 621:304–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gonnet:1992:ETD

[923] G. Gonnet. Efficient two-dimensional searching (invited). Lecture Notes in Computer Science, 621:317–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chew:1992:IGP

[924] L. P. Chew and K. Kedem. Improvements on geometric pattern matching problems. Lecture Notes in Computer Science, 621:318–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joseph:1992:DDS

[925] D. Joseph, J. Meidanis, and P. Tiwari. Determining DNA sequence similarity using maximum independent set algorithms for interval graphs. Lecture Notes in Computer Science, 621:326– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Welzl:1992:NRL

[926] E. Welzl. New results on linear programming and related problems (invited). Lecture Notes in Computer Science, 621:338–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golin:1992:DCP

[927] M. J. Golin. Dynamic closest pairs — A probabilistic approach. Lecture Notes in Computer Science, 621:340– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deBerg:1992:TTD

[928] M. de Berg, M. van Kreveld, and J. Snoeyink. Two- and threedimensional point location in rectangular subdivisions. Lecture Notes in Computer Science, 621:352–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chazelle:1992:DBN

[929] B. Chazelle and L. Palios. Decomposing the boundary of a nonconvex polyhedron. Lecture Notes in Computer Science, 621:364–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hershberger:1992:CPM

[930] J. Hershberger and J. Snoeyink. Convex polygons made from few lines and convex decompositions of polyhedra. Lecture Notes in Computer Science, 621:376–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenhof:1992:MVM

[931] H.-P. Lenhof and M. Smid. Maintaining the visibility map of spheres while moving the viewpoint on a circle at infinity. Lecture Notes in Computer Science, 621:388–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albers:1992:VDM

[932] G. Albers and T. Roos. Voronoi diagrams of moving points in higher dimensional spaces. Lecture Notes in Computer Science, 621:399–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katajainen:1992:SMS

[933] J. Katajainen and T. Pasanen. Sorting multisets stably in minimum space. Lecture Notes in Computer Science, 621:410–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petersson:1992:FAS

[934] O. Petersson and A. Moffat. A framework for adaptive sorting. Lecture Notes in Computer Science, 621:422– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vob:1992:F

[935] A. Vob. Foreword. Lecture Notes in Computer Science, 622:V–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmalhofer:1992:RBK

[936] F. Schmalhofer. Relations between knowledge engineering and cognition. Lecture Notes in Computer Science, 622:3–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dallemagne:1992:MAP

[937] G. Dallemagne, G. Klinker, D. Marques, J. McDermott, and D. Tung. Making application programming more worthwhile. Lecture Notes in Computer Science, 622:6–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manago:1992:UIT

[938] M. Manago and N. Conruyt. Using information technology to solve real world problems. Lecture Notes in Computer Science, 622:23–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shadbolt:1992:FFF

[939] N. Shadbolt. Facts, fantasies and frameworks: The design of a knowledge acquisition workbench. Lecture Notes in Computer Science, 622:39–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfeifer:1992:MEB

[940] R. Pfeifer, T. Rothenfluh, M. Stolze, and F. Steiner. Mapping expert behavior onto task-level frameworks: The need for “eco-pragmatic” approaches to knowledge engineering. Lecture Notes in Computer Science, 622:59–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fensel:1992:KAI

[941] D. Fensel. Knowledge acquisition and the interpretative paradigm. Lecture Notes in Computer Science, 622:78–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janetzko:1992:CBR

[942] D. Janetzko and G. Strube. Case-based reasoning and model-based knowledge acquisition. Lecture Notes in Computer Science, 622:99–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmalhofer:1992:RPH

[943] F. Schmalhofer, Ch. Globig, and J. Thoben. The refitting of plans by a human expert. Lecture Notes in Computer Science, 622:115–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergmann:1992:KAG

[944] R. Bergmann. Knowledge acquisition by generating skeletal plans from real world cases. Lecture Notes in Computer Science, 622:125–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Branskat:1992:KAC

[945] S. Branskat. Knowledge acquisition from cases. Lecture Notes in Computer Science, 622:134–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reimann:1992:TEC

[946] P. Reimann and T. J. Shult. Transforming examples into cases. Lecture Notes in Computer Science, 622:139– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Althoff:1992:CBR

[947] K.-D. Althoff and S. Web. Casebased reasoning and expert system development. Lecture Notes in Computer Science, 622:146–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strube:1992:RCS

[948] G. Strube. The role of cognitive science in knowledge engineering. Lecture Notes in Computer Science, 622:161– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schlenker:1992:KAE

[949] B. Schlenker and Th. Wetter. Knowledge acquisition as an empirically based modelling activity. Lecture Notes in Computer Science, 622:175– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Linster:1992:SPM

[950] M. Linster. Shifting positions: Moving from a cognitive science point of view to a knowledge engineering stance. Lecture Notes in Computer Science, 622:183–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puppe:1992:TQE

[951] F. Puppe and U. Gappa. Two questions from expert system developers to cognitive scientists. Lecture Notes in Computer Science, 622:190–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Woodward:1992:CBK

[952] J. B. Woodward, M. L. G. Shaw, and B. R. Gaines. The cognitive basis of knowledge engineering. Lecture Notes in Computer Science, 622:194– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wetter:1992:CAS

[953] Th. Wetter. A comparative assessment of selected approaches in the focal area of knowledge engineering and cognition. Lecture Notes in Computer Science, 622:225–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1992:PIK

[954] M. Li and P. M. B. Vitanyi. Philosophical issues in KolmogorovComplexity. Lecture Notes in Computer Science, 623:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Straubing:1992:CCE

[955] H. Straubing. Circuit complexity and the expressive power of generalized first-order formulas. Lecture Notes in Computer Science, 623:16–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeSantis:1992:OMS

[956] A. De Santis, G. Persiano, and M. Yung. One-message statistical zeroknowledge proofs and space-bounded verifier. Lecture Notes in Computer Science, 623:28–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keranen:1992:ASA

[957] V. Keranen. Abelian squares are avoidable on 4 letters. Lecture Notes in Computer Science, 623:41–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karhumaki:1992:PST

[958] J. Karhumaki, W. Plandowski, and W. Rytter. Polynomial size test sets for context-free languages. Lecture Notes in Computer Science, 623:53–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nishida:1992:QDO

[959] T. Y. Nishida. Quasi-deterministic OL systems. Lecture Notes in Computer Science, 623:65–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buntrock:1992:GCS

[960] G. Buntrock and K. Lorys. On growing context-sensitive languages. Lecture Notes in Computer Science, 623:77–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shallit:1992:NSL

[961] J. Shallit. Numeration systems, linear recurrences, and regular sets. Lecture Notes in Computer Science, 623:89–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krob:1992:EPR

[962] D. Krob. The equality problem for rational series with multiplicities in the tropical semiring is undecidable. Lecture Notes in Computer Science, 623: 101–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clerbout:1992:SCR

[963] M. Clerbout, M. Latteux, Y. Roos, and W. Zielonka. Semi-commutations and rational expressions. Lecture Notes in Computer Science, 623:113–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ibarra:1992:NRC

[964] O. H. Ibarra and N. Q. Tran. New results concerning synchronized finite automata. Lecture Notes in Computer Science, 623:126–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelfriet:1992:GNF

[965] J. Engelfriet. A Greibach Normal Form for context-free graph grammars. Lecture Notes in Computer Science, 623:138–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peladeau:1992:RGD

[966] P. Peladeau and A. Podelski. On reverse and general definite tree languages. Lecture Notes in Computer Science, 623:150–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arvind:1992:RSL

[967] V. Arvind, Y. Han, L. Hemachandra, J. Kobler, A. Lozano, M. Mundhenk, M. Ogiwara, U. Schoning, R. Silvestri, and T. Thierauf. Reductions to sets of low information content. Lecture Notes in Computer Science, 623:162– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sheu:1992:LHH

[968] M.-J. Sheu and T. J. Long. UP and the low and high hierarchies: A relativized separation. Lecture Notes in Computer Science, 623:174–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flajolet:1992:AAA

[969] Ph. Flajolet. Analytic analysis of algorithms. Lecture Notes in Computer Science, 623:186–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirschenhofer:1992:HCQ

[970] P. Kirschenhofer, H. Prodinger, and W. Szpankowski. How to count quickly and accurately. Lecture Notes in Computer Science, 623:211–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmid:1992:ACL

[971] U. Schmid. The average CRI-Length of a tree collision resolution algorithm in presence of multiplicity-dependent capture effects. Lecture Notes in Computer Science, 623:223–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dietzfelbinger:1992:PHF

[972] M. Dietzfelbinger, J. Gil, Y. Matias, and N. Pippenger. Polynomial hash functions are reliable. Lecture Notes in Computer Science, 623:235–246, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sekar:1992:APM

[973] R. C. Sekar, R. Ramesh, and I. V. Ramakrishnan. Adaptive pattern matching. Lecture Notes in Computer Science, 623:247–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mansour:1992:RIA

[974] Y. Mansour. Randomized interpolation and approximation of sparse polynomials. Lecture Notes in Computer Science, 623:261–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodlaender:1992:TSA

[975] H. L. Bodlaender, M. R. Fellows, and T. J. Warnow. Two strikes against perfect phylogeny. Lecture Notes in Computer Science, 623:273–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1992:DSD

[976] G.-Q. Zhang. Disjunctive systems and L-domains. Lecture Notes in Computer Science, 623:284–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Apostolico:1992:OPA

[977] A. Apostolico, D. Breslauer, and Z. Galil. Optimal parallel algorithms for periods, palindromes and squares. Lecture Notes in Computer Science, 623:296–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broder:1992:NPT

[978] A. Z. Broder, A. M. Frieze, E. Shamir, and E. Upfal. Near-perfect token distribution. Lecture Notes in Computer Science, 623:308–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagerup:1992:FIM

[979] T. Hagerup and M. Kutylowski. Fast integer merging on the EREW PRAM. Lecture Notes in Computer Science, 623:318–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khuller:1992:AAG

[980] S. Khuller and R. Thurimella. Approximation algorithms for graph augmentation. Lecture Notes in Computer Science, 623:330–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Westbrook:1992:FIP

[981] J. Westbrook. Fast incremental planarity testing. Lecture Notes in Computer Science, 623:342–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LaPoutre:1992:MTC

[982] J. A. La Poutre. Maintenance of triconnected components of graphs. Lecture Notes in Computer Science, 623: 354–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vazirani:1992:SCT

[983] V. V. Vazirani and M. Yannakakis. Suboptimal cuts: Their enumeration, weight and number. Lecture Notes in Computer Science, 623:366–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buchberger:1992:GBI

[984] B. Buchberger. Gröbner bases: An introduction. Lecture Notes in Computer Science, 623:378–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bundgen:1992:BAT

[985] R. Bundgen. Buchberger’s algorithm: The term rewriter’s point of view. Lecture Notes in Computer Science, 623: 380–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Comon:1992:CRS

[986] H. Comon. Completion of rewrite systems with membership constraints. Lecture Notes in Computer Science, 623:392–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rote:1992:NMB

[987] G. Rote. A new metric between polygons, and how to compute it. Lecture Notes in Computer Science, 623:404– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paterson:1992:NNG

[988] M. S. Paterson and F. F. Yao. On nearest-neighbor graphs. Lecture Notes in Computer Science, 623:416– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matousek:1992:TEM

[989] J. Matousek and R. Seidel. A tail estimate for Mulmuley’s segment intersection algorithm. Lecture Notes in Computer Science, 623:427–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chazelle:1992:LBC

[990] B. Chazelle and B. Rosenberg. Lower bounds on the complexity of simplex range reporting on a pointer machine. Lecture Notes in Computer Science, 623:439–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolaitis:1992:ILC

[991] P. G. Kolaitis and M. Y. Vardi. Infinitary logic for computer science. Lecture Notes in Computer Science, 623: 450–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1992:CTP

[992] E. Chang, Z. Manna, and A. Pnueli. Characterization of temporal property classes. Lecture Notes in Computer Science, 623:474–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ong:1992:LLC

[993] C.-H. L. Ong. Lazy lambda calculus: Theories, models and local structure characterization. Lecture Notes in Computer Science, 623:487–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laenens:1992:LPS

[994] E. Laenens, D. Vermeir, and C. Zaniolo. Logic programming semantics made easy. Lecture Notes in Computer Science, 623:499–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Debray:1992:CDA

[995] S. K. Debray. On the complexity of dataflow analysis of logic programs. Lecture Notes in Computer Science, 623:509–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cortesi:1992:CAI

[996] A. Cortesi, G. File, and W. Winsborough. Comparison of abstract interpretations. Lecture Notes in Computer Science, 623:521–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Phoa:1992:PCS

[997] W. Phoa and M. Fourman. A proposed categorical semantics for pure ML. Lecture Notes in Computer Science, 623:533–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1992:WGD

[998] Th. A. Henzinger, Z. Manna, and A. Pnueli. What good are digital clocks? Lecture Notes in Computer Science, 623:545–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moller:1992:BAT

[999] F. Moller and Ch. Tofts. Behavioral abstraction in TCCS. Lecture Notes in Computer Science, 623:559–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1992:TPN

[1000] C. Brown and D. Gurr. Timing Petri nets categorically. Lecture Notes in Computer Science, 623:571–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gastin:1992:ACA

[1001] P. Gastin and A. Petit. Asynchronous cellular automata for infinite traces. Lecture Notes in Computer Science, 623:583–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoogers:1992:TSP

[1002] P. W. Hoogers, H. C. M. Kleijn, and P. S. Thiagarajan. A trace semantics for Petri nets. Lecture Notes in Computer Science, 623:595–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogler:1992:ACP

[1003] W. Vogler. Asynchronous communication of Petri nets and the refinement of transitions. Lecture Notes in Computer Science, 623:605–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montanari:1992:PAL

[1004] U. Montanari and D. Yankelevich. A parametric approach to localities. Lecture Notes in Computer Science, 623: 617–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Degano:1992:PT

[1005] P. Degano and C. Priami. Proved trees. Lecture Notes in Computer Science, 623:629–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Olderog:1992:IBL

[1006] E.-R. Olderog. Interfaces between languages for communicating systems. Lecture Notes in Computer Science, 623:641–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sannella:1992:TFD

[1007] D. Sannella and A. Tarlecki. Toward formal development of programs from algebraic specifications: Modeltheoretic foundations. Lecture Notes in Computer Science, 623:656–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fix:1992:PCU

[1008] L. Fix, N. Francez, and O. Grumberg. Program composition via unification. Lecture Notes in Computer Science, 623:672–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Milner:1992:BB

[1009] R. Milner and D. Sangiorgi. Barbed bisimulation. Lecture Notes in Computer Science, 623:685–??, 1992. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rabinovich:1992:CEB

[1010] A. Rabinovich. Checking equivalences between concurrent systems of finite agents. Lecture Notes in Computer Science, 623:696–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cleaveland:1992:TPP

[1011] R. Cleaveland, S. A. Smolka, and A. Zwarico. Testing preorders for probabilistic processes. Lecture Notes in Computer Science, 623:708–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Przymusinska:1992:SCP

[1012] Halina Przymusinska, Teodor Przymusinski, and Hirohisa Seki. Soundness and completeness of partial deductions for well-founded semantics. Lecture Notes in Computer Science, 624:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holldobler:1992:DPF

[1013] Steffen Holldobler. On deductive planning and the frame problem. Lecture Notes in Computer Science, 624:13–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harland:1992:RFC

[1014] James Harland and David Pym. On resolution in fragments of classical linear logic. Lecture Notes in Computer Science, 624:30–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galmiche:1992:PAP

[1015] Didier Galmiche and Guy Perrier. A procedure for automatic proof nets construction. Lecture Notes in Computer Science, 624:42–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bowen:1992:FLI

[1016] James Bowen and Dennis Bahler. Free logic and infinite constraint networks. Lecture Notes in Computer Science, 624:54–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wuthrich:1992:TPK

[1017] Beat Wuthrich. Towards probabilistic knowledge bases. Lecture Notes in Computer Science, 624:66–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bryant:1992:TLG

[1018] Barrett R. Bryant and Aiqin Pan. Two-level grammar: A functional/ logic query language for database and knowledge-base systems. Lecture Notes in Computer Science, 624:78–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freitag:1992:EDD

[1019] Burkhard Freitag. Extending deductive database languages by embedded implications. Lecture Notes in Computer Science, 624:84–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lusk:1992:CRL

[1020] Ewing L. Lusk. Controlling redundancy in large search spaces. Lecture Notes in Computer Science, 624:96–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baaz:1992:RMV

[1021] Matthias Baaz and Christian G. Fermuller. Resolution for many-valued logics. Lecture Notes in Computer Science, 624:107–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baumgartner:1992:OTR

[1022] Peter Baumgartner. An ordered theory resolution calculus. Lecture Notes in Computer Science, 624:119–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCune:1992:AAD

[1023] William McCune and Larry Wos. Application of automated deduction to the search for single axioms for exponent groups. Lecture Notes in Computer Science, 624:131–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hai-Ping:1992:ELB

[1024] Ko Hai-Ping and Mark E. Nadel. Elementary lower bounds for the lengths of refutations. Lecture Notes in Computer Science, 624:137–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Egly:1992:SPQ

[1025] Uwe Egly. Shortening proofs by quantifier introduction. Lecture Notes in Computer Science, 624:148–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Millroth:1992:RCN

[1026] Hakan Millroth. Reform compilation for nonlinear recursion. Lecture Notes in Computer Science, 624:160– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neumerkel:1992:PIF

[1027] Ulrich Neumerkel. Pruning infinite failure branches in programs with occurcheck. Lecture Notes in Computer Science, 624:172–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ireland:1992:UPC

[1028] Andrew Ireland. The use of planning critics in mechanizing inductive proofs. Lecture Notes in Computer Science, 624:178–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parigot:1992:LCA

[1029] Michel Parigot. lambdamu-calculus: An algorithmic interpretation of classical natural deduction. Lecture Notes in Computer Science, 624:190–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

BoydelaTour:1992:BPA

[1030] Thierry Boy de la Tour and Christoph Kreitz. Building proofs by analogy via the Curry–Howard isomorphism. Lecture Notes in Computer Science, 624:202–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baker:1992:UCO

[1031] Siani Baker, Andrew Ireland, and Alan Smaill. On the use of the constructive omega-rule within automated deduction. Lecture Notes in Computer Science, 624:214–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ertel:1992:PTP

[1032] Wolfgang Ertel. OR-Parallel theorem proving with random competition. Lecture Notes in Computer Science, 624:226–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suttner:1992:PCM

[1033] Christian B. Suttner. Parallel computation of multiple sets-of-support. Lecture Notes in Computer Science, 624: 238–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hausman:1992:TUA

[1034] Bogumil Hausman. Towards using the andorra kernel language for industrial real-time applications. Lecture Notes in Computer Science, 624:250–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ringeissen:1992:UCE

[1035] Christophe Ringeissen. Unification in a combination of equational theories with shared constants and its application to primal algebras. Lecture Notes in Computer Science, 624:261–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bachmair:1992:NCR

[1036] Leo Bachmair and Harald Ganzinger.

Non-clausal resolution and superposition with selection and redundancy criteria. Lecture Notes in Computer Science, 624:273–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gramlich:1992:RIW

[1037] Bernhard Gramlich. Relating innermost, weak, uniform and modular termination of term rewriting systems. Lecture Notes in Computer Science, 624:285–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gabbrielli:1992:TSS

[1038] Maurizio Gabbrielli, Giorgio Levi, and Daniele Turi. A two step semantics for logic programs with negation. Lecture Notes in Computer Science, 624:297–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baral:1992:GNF

[1039] Chitta Baral. Generalized negation as failure and semantics of normal disjunctive logic programs. Lecture Notes in Computer Science, 624:309– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bai:1992:GMT

[1040] Mino Bai and Howard A. Blair. General model theoretic semantics for higherorder Horn logic programming. Lecture Notes in Computer Science, 624: 320–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1992:DDD

[1041] Jose Alberto Fernandez and Jack Minker. Disjunctive deductive databases. Lecture Notes in Computer Science, 624:332–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voinov:1992:NCO

[1042] Alexander Voinov. Netlog — A concept oriented logic programming language. Lecture Notes in Computer Science, 624:357–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischer:1992:PFE

[1043] Michael Fischer and Richard Owens.

From the past to the future: Executing temporal logic programs. Lecture Notes in Computer Science, 624:369– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walther:1992:CIA

[1044] Christoph Walther. Computing induction axioms. Lecture Notes in Computer Science, 624:381–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antimirov:1992:CEE

[1045] Valentin Antimirov and Anatoli Degtyarev. Consistency of equational enrichments. Lecture Notes in Computer Science, 624:393–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Curzon:1992:PLV

[1046] Paul Curzon. A programming logic for a verified structured assembly language. Lecture Notes in Computer Science, 624:403–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salzer:1992:UIS

[1047] Gernot Salzer. The unification of infinite sets of terms and its applications. Lecture Notes in Computer Science, 624:409–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kohlhase:1992:UOS

[1048] Michael Kohlhase. Unification in ordersorted type theory. Lecture Notes in Computer Science, 624:421–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sattler-Klein:1992:ICS

[1049] Andrea Sattler-Klein. Infinite, canonical string rewriting systems generated by completion. Lecture Notes in Computer Science, 624:433–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alexandre:1992:SSL

[1050] Francis Alexandre, Khaled Bsaies, Jean Pierre Finance, and Alain Quere. Spes: A system for logic program transformation. Lecture Notes in Computer Science, 624:445–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andreoli:1992:LOL

[1051] Jean-Marc Andreoli and Remo Pareschi. Linear objects: A logic framework for open system programming. Lecture Notes in Computer Science, 624:448– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bauer:1992:IIS

[1052] Bernhard Bauer and Rolf Hennicker.

ISAR: An interactive system for algebraic implementation proofs. Lecture Notes in Computer Science, 624:451– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beeson:1992:MCS

[1053] Michael Beeson. Mathpert: Computer support for learning algebra, trig and calculus. Lecture Notes in Computer Science, 624:454–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bocca:1992:MPD

[1054] Jorge Bocca, Michael Dahmen, and Michael Freeston. MegaLog — A platform for developing knowledge base management systems. Lecture Notes in Computer Science, 624:457–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouhoula:1992:SAT

[1055] Adel Bouhoula, Emmanuel Kounalis, and Michael Rusinowitch. SPIKE, an automatic theorem prover. Lecture Notes in Computer Science, 624:460– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dafa:1992:ATL

[1056] Li Dafa. An application to teaching in logic course of ATP based on natural deduction. Lecture Notes in Computer Science, 624:463–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dawson:1992:GLE

[1057] Mark Dawson. A generic logic environment. Lecture Notes in Computer Science, 624:466–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dorochevsky:1992:EPP

[1058] Michel Dorochevsky, Li Liang-Liang, Mike Reeve, Kees Schuerman, and Andre Veron. ElipSys. A parallel programming system based on logic. Lecture Notes in Computer Science, 624:469– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ducasse:1992:OHL

[1059] Mireille Ducasse. Opium — A high level debugging environment. Lecture Notes in Computer Science, 624:472– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fraus:1992:ITP

[1060] Ulrich Fraus and Heinrich Hussmann. An inductive theorem prover based on narrowing. Lecture Notes in Computer Science, 624:475–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaasterland:1992:CAS

[1061] Terry Gaasterland, Parke Godfrey, Jack Minker, and Lev Novik. A cooperative answering system. Lecture Notes in Computer Science, 624:478– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leach:1992:MPT

[1062] Javier Leach and Susana Nieva. MIZ-PR: A theorem prover for polymorphic and recursive functions. Lecture Notes in Computer Science, 624:481– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manoury:1992:PPL

[1063] Pascal Manoury, Michel Parigot, and Marianne Simonot. ProPre. A programming language with proofs. Lecture Notes in Computer Science, 624: 484–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1992:FWI

[1064] Julio Garcia Martin and Juan Jose Moreno-Navarro. FRIENDLY-WAM: An interactive tool to understand the compilation of Prolog. Lecture Notes in Computer Science, 624:487–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meier:1992:SBP

[1065] Micha Meier. SEPIA: A basis for Prolog extensions. Lecture Notes in Computer Science, 624:490–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nilsson:1992:EDS

[1066] Hans Nilsson. The external database in SICStus Prolog. Lecture Notes in Computer Science, 624:493–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Noye:1992:KSS

[1067] Jacques Noye. The KCM system: Speeding-up logic programming through hardware support. Lecture Notes in Computer Science, 624:496– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Romanenko:1992:LW

[1068] Igor Romanenko. Logician’s workbench. Lecture Notes in Computer Science, 624:499–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sawamura:1992:EGR

[1069] Hajime Sawamura, Toshiro Minami, and Kyoko Ohashi. EUODHILOS: A general reasoning system for a variety of logics. Lecture Notes in Computer Science, 624:501–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vielle:1992:EVS

[1070] Laurent Vielle, Petra Bayer, Volker Kuchenhoff, Alexandre Lefebvre, and Rainer Manthey. The EKS-VI system. Lecture Notes in Computer Science, 624:504–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wallace:1992:CP

[1071] Mark Wallace and Thierry Le Provost. CHIP and Propia. Lecture Notes in Computer Science, 624:507–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogler:1992:I

[1072] W. (Walter) Vogler. Introduction. Lecture Notes in Computer Science, 625: 1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogler:1992:PNT

[1073] W. (Walter) Vogler. Petri nets and their semantics. Lecture Notes in Computer Science, 625:9–32, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogler:1992:PCD

[1074] W. (Walter) Vogler. Parallel composition and deadlocking. Lecture Notes in Computer Science, 625:33–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogler:1992:BPR

[1075] W. (Walter) Vogler. Behaviour preserving refinement of places and transitions. Lecture Notes in Computer Science, 625:75–130, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogler:1992:ARI

[1076] W. (Walter) Vogler. Action refinement and interval words. Lecture Notes in Computer Science, 625:131–182, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogler:1992:ARB

[1077] W. (Walter) Vogler. Action refinement and bisimulation. Lecture Notes in Computer Science, 625:183–216, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogler:1992:POS

[1078] W. (Walter) Vogler. Partial order semantics for nets with capacities. Lecture Notes in Computer Science, 625:217–238, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bargury:1992:EPT

[1079] Y. Bargury and J. Makowsky. The expressive power of transitive closure and 2-way multihead automata. Lecture Notes in Computer Science, 626:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beierle:1992:CPW

[1080] Ch. Beierle and E. Borger. Correctness proof for the WAM with types. Lecture Notes in Computer Science, 626:15–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Best:1992:MCP

[1081] E. Best and J. Esparza. Model checking of persistent Petri nets. Lecture Notes in Computer Science, 626:35–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chirimar:1992:PTD

[1082] J. Chirimar and J. Lipton. Provability in TBLL: A decision procedure. Lecture Notes in Computer Science, 626:53–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dahlhaus:1992:HIF

[1083] E. Dahlhaus. How to implement first order formulas in local memory machine models. Lecture Notes in Computer Science, 626:68–??, 1992.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feferman:1992:NAA

[1084] S. Feferman. A new approach to abstract data types II. computation on ADTs as ordinary computation. Lecture Notes in Computer Science, 626: 79–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernando:1992:PRS

[1085] T. Fernando. A primitive recursive set theory and AFA: On the logical complexity of the largest bisimulation. Lecture Notes in Computer Science, 626:96–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flum:1992:BT

[1086] J. Flum. On bounded theories. Lecture Notes in Computer Science, 626: 111–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goerdt:1992:CPP

[1087] A. Goerdt. The cutting plane proof system with bounded degree of falsity. Lecture Notes in Computer Science, 626:119–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreno:1992:DVD

[1088] J. C. Gonzalez Moreno, M. T. Hortala Gonzalez, and M. Rodriguez Artalejo. Denotational versus declarative semantics for functional programming. Lecture Notes in Computer Science, 626:134–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gradel:1992:TCL

[1089] E. Gradel. On transitive closure logic. Lecture Notes in Computer Science, 626:149–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heusch:1992:SAP

[1090] P. Heusch and E. Speckenmeyer. Some aspects of the probabilistic behavior of variants of resolution. Lecture Notes in Computer Science, 626:164–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirshfeld:1992:SQR

[1091] J. Hirshfeld. Safe queries in relational databases with functions. Lecture Notes in Computer Science, 626: 173–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hooker:1992:LIP

[1092] J. N. Hooker. Logical inference and polyhedral projection. Lecture Notes in Computer Science, 626:184–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hosli:1992:SL

[1093] B. Hosli. Stable logic. Lecture Notes in Computer Science, 626:201–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rao:1992:TMP

[1094] M. R. K. Krishna Rao, D. Kapur, and R. K. Shyamasundar. A transformational methodology for proving termination of logic programs. Lecture Notes in Computer Science, 626:213– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lehmann:1992:PL

[1095] D. Lehmann. Plausibility logic. Lecture Notes in Computer Science, 626: 227–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leiss:1992:TKA

[1096] H. Leiß. Towards Kleene algebra with recursion. Lecture Notes in Computer Science, 626:242–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meinke:1992:ESA

[1097] K. Meinke. Equational specification of abstract types and combinators. Lecture Notes in Computer Science, 626: 257–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mundici:1992:NFI

[1098] D. Mundici. Normal forms in infinitevalued logic: The case of one variable. Lecture Notes in Computer Science, 626:272–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oguztuzun:1992:FFO

[1099] H. Oguztuzun. A fragment of first order logic adequate for observation equivalence. Lecture Notes in Computer Science, 626:278–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinchinat:1992:OPC

[1100] S. Pinchinat. Ordinal processes in comparative concurrency semantics. Lecture Notes in Computer Science, 626: 293–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deLavalette:1992:LSM

[1101] G. R. Renardel de Lavalette. Logical semantics of modularisation. Lecture Notes in Computer Science, 626:306– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmerl:1992:CEP

[1102] U. R. Schmerl. A cut-elimination procedure designed for evaluating proofs as programs. Lecture Notes in Computer Science, 626:316–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwichtenberg:1992:MCP

[1103] H. Schwichtenberg. Minimal from classical proofs. Lecture Notes in Computer Science, 626:326–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seibert:1992:QHW

[1104] S. Seibert. Quantifier hierarchies over word relations. Lecture Notes in Computer Science, 626:329–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steffen:1992:CRD

[1105] E. Steffen. Complexity results for the default-and the autoepistemic logic. Lecture Notes in Computer Science, 626:339–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stewart:1992:CNP

[1106] I. A. Stewart. On completeness for NP via projection translations. Lecture Notes in Computer Science, 626: 353–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thistle:1992:COA

[1107] J. G. Thistle and W. M. Wonham.

Control of omega-automata, Church’s problem, and the emptiness problem for tree omega-automata. Lecture Notes in Computer Science, 626:367– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Troelstra:1992:CTR

[1108] A. S. Troelstra. Comparing the theory of representations and constructive mathematics. Lecture Notes in Computer Science, 626:382–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tyszkiewicz:1992:IQT

[1109] J. Tyszkiewicz. Infinitary queries and their asymptotic probabilities I: Properties definable in transitive closure logic. Lecture Notes in Computer Science, 626:396–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voronkov:1992:CPS

[1110] A. Voronkov. On completeness of program synthesis systems. Lecture Notes in Computer Science, 626:411– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weiermann:1992:PTT

[1111] A. Weiermann. Proving termination for term rewriting systems. Lecture Notes in Computer Science, 626:419– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pustejovsky:1992:I

[1112] J. Pustejovsky. Introduction. Lecture Notes in Computer Science, 627:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nirenburg:1992:SDO

[1113] S. Nirenburg and L. Levin. Syntaxdriven and ontology-driven lexical semantics. Lecture Notes in Computer Science, 627:5–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer:1992:KMT

[1114] I. Meyer. Knowledge management for terminology-intensive applications: Needs and tools. Lecture Notes in Computer Science, 627:21–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sowa:1992:LSL

[1115] J. Sowa. Logical structures in the lexicon. Lecture Notes in Computer Science, 627:39–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1992:CML

[1116] J. Martin. Conventional metaphor and the lexicon. Lecture Notes in Computer Science, 627:61–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burkert:1992:RSK

[1117] G. Burkert and P. Forster. Representation of semantic knowledge with term subsumption languages. Lecture Notes in Computer Science, 627:75–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ostler:1992:PMS

[1118] N. Ostler and B. T. S. Atkins. Predictable meaning shift: Some linguistic properties of lexical implication rules. Lecture Notes in Computer Science, 627:87–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Copestake:1992:LOU

[1119] A. Copestake and T. Briscoe. Lexical operations in a unification-based framework. Lecture Notes in Computer Science, 627:101–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anick:1992:LSL

[1120] P. Anick and S. Bergler. Lexical structures for linguistic inference. Lecture Notes in Computer Science, 627:121– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meijs:1992:MWK

[1121] W. Meijs and P. Vossen. In so many words: Knowledge as a lexical phenomenon. Lecture Notes in Computer Science, 627:137–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Macpherson:1992:RLW

[1122] M. Macpherson. Redefining the “level” of the “word”. Lecture Notes in Computer Science, 627:155–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

White:1992:LWK

[1123] J. S. White. Lexical and world knowledge: Theoretical and applied viewpoints. Lecture Notes in Computer Science, 627:171–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herweg:1992:ART

[1124] M. Herweg. Aspectual requirements of temporal connectives: Evidence for a two-level approach to semantics. Lecture Notes in Computer Science, 627: 185–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerstl:1992:MIL

[1125] P. Gerstl. A model for the interaction of lexical and non-lexical knowledge in the determination of word meaning. Lecture Notes in Computer Science, 627:201–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Evens:1992:LE

[1126] M. Evens et al. For the lexicon that has everything. Lecture Notes in Computer Science, 627:219–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calzolari:1992:ARS

[1127] N. Calzolari. Acquiring and representing semantic information in a lexical knowledge base. Lecture Notes in Computer Science, 627:235–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palmer:1992:GLR

[1128] M. Palmer. General lexical representation for an effect predicate. Lecture Notes in Computer Science, 627:245– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dahlgren:1992:ASL

[1129] K. Dahlgren. The autonomy of shallow lexical knowledge. Lecture Notes in Computer Science, 627:255–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Door:1992:TLK

[1130] B. J. Door. A two-level knowledge representation for machine translation: Lexical semantics and tense/ aspect. Lecture Notes in Computer Science, 627:269–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Onyshkevych:1992:LOT

[1131] B. A. Onyshkevych and S. Nirenburg. Lexicon, ontology, and text meaning. Lecture Notes in Computer Science, 627:289–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsukawa:1992:DCD

[1132] T. Matsukawa and E. Yokota. Development of the concept dictionary — implementation of lexical knowledge. Lecture Notes in Computer Science, 627:305–319, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mercer:1992:PDR

[1133] R. Mercer. Presuppositions and default reasoning: A study in lexical pragmatics. Lecture Notes in Computer Science, 627:321–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vosselman:1992:CVM

[1134] G. (George) Vosselman. Computer vision and matching. Lecture Notes in Computer Science, 628:1–4, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vosselman:1992:CMM

[1135] G. (George) Vosselman. A classification of matching methods. Lecture Notes in Computer Science, 628:5–34, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vosselman:1992:FDR

[1136] G. (George) Vosselman. Formal description of relational matching. Lecture Notes in Computer Science, 628: 35–44, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vosselman:1992:PDC

[1137] G. (George) Vosselman. Problem definition and contributions of the thesis. Lecture Notes in Computer Science, 628:45–50, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vosselman:1992:ITS

[1138] G. (George) Vosselman. Information theory: selected topics. Lecture Notes in Computer Science, 628:51–66, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vosselman:1992:EMB

[1139] G. (George) Vosselman. Evaluation of mappings between relational descriptions. Lecture Notes in Computer Science, 628:67–87, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vosselman:1992:TSM

[1140] G. (George) Vosselman. Tree search methods and heuristics. Lecture Notes in Computer Science, 628:88– 111, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vosselman:1992:RIM

[1141] G. (George) Vosselman. Relational image and model description. Lecture Notes in Computer Science, 628:112– 122, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vosselman:1992:EFO

[1142] G. (George) Vosselman. Evaluation function for object location. Lecture Notes in Computer Science, 628:123– 143, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vosselman:1992:SPT

[1143] G. (George) Vosselman. Strategy and performance of the tree search for object location. Lecture Notes in Computer Science, 628:144–162, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vosselman:1992:MIB

[1144] G. (George) Vosselman. Mutual information between a continuous signal and a discretized noisy observation. Lecture Notes in Computer Science, 628:181–184, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vosselman:1992:DCP

[1145] G. (George) Vosselman. Distribution of the coordinates of points on a sphere. Lecture Notes in Computer Science, 628:185–186, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vosselman:1992:CPD

[1146] G. (George) Vosselman. Conditional probability density function of the image line length. Lecture Notes in Computer Science, 628:187–188, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agarwal:1992:RSS

[1147] P. K. Agarwal and J. Matousek. On range searching with semialgebraic sets. Lecture Notes in Computer Science, 629:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diaz:1992:GLP

[1148] J. Diaz. Graph layout problems. Lecture Notes in Computer Science, 629: 14–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lange:1992:PRR

[1149] K.-J. Lange, P. Rossmanith, and W. Rytter. Parallel recognition and ranking of context-free languages. Lecture Notes in Computer Science, 629: 24–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mihail:1992:ECP

[1150] M. Mihail. On the expansion of combinatorial polytopes. Lecture Notes in Computer Science, 629:37–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orponen:1992:NNC

[1151] P. Orponen. Neural networks and complexity theory. Lecture Notes in Computer Science, 629:50–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tucker:1992:TCS

[1152] J. V. Tucker and J. I. Zucker. Theory of computation over stream algebras, and its applications. Lecture Notes in Computer Science, 629:62–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vishkin:1992:MPAa

[1153] U. Vishkin. Methods in parallel algorithmics. Lecture Notes in Computer Science, 629:81–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watanabe:1992:CSD

[1154] O. Watanabe. On the complexity of small description and related topics. Lecture Notes in Computer Science, 629:82–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiedermann:1992:WPM

[1155] J. Wiedermann. Weak parallel machines: a new class of physically feasible parallel machine models. Lecture Notes in Computer Science, 629:95–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wigderson:1992:CGC

[1156] A. Wigderson. The complexity of graph connectivity. Lecture Notes in Computer Science, 629:112–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bast:1992:PPD

[1157] H. Bast, M. Dietzfelbinger, and T. Hagerup. A perfect parallel dictionary. Lecture Notes in Computer Science, 629:133–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becker:1992:SRT

[1158] B. Becker and J. Hartmann. Some remarks on the test complexity of iterative logic arrays. Lecture Notes in Computer Science, 629:142–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burtschick:1992:DSR

[1159] H.-J. Burtschick and A. Hoene. The degree structure of 1-L reductions. Lecture Notes in Computer Science, 629: 153–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cai:1992:PPA

[1160] J. Cai, L. A. Hemachandra, and J. Vyskoc. Promise problems and access to unambiguous computation. Lecture Notes in Computer Science, 629: 162–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chari:1992:CIC

[1161] S. Chari, D. Ranjan, and P. Rohatgi.

On the complexity of incremental computation. Lecture Notes in Computer Science, 629:172–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choffrut:1992:RTC

[1162] Ch. Choffrut and M. Goldwurm. Rational transductions and complexity of counting problems. Lecture Notes in Computer Science, 629:181–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Comon:1992:NEE

[1163] H. Comon and M. Fernandez. Negation elimination in equational formulae. Lecture Notes in Computer Science, 629:191–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Costa:1992:OI

[1164] J. F. Costa, A. Sernadas, C. Sernadas, and H.-D. Ehrich. Object interaction. Lecture Notes in Computer Science, 629:200–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Curien:1992:SNS

[1165] P.-L. Curien, T. Hardin, and A. Rios. Strong normalization of substitutions. Lecture Notes in Computer Science, 629:209–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daley:1992:PPL

[1166] R. Daley and B. Kalyanasundaram. Probabilistic and pluralistic learners with mind changes. Lecture Notes in Computer Science, 629:218–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damm:1992:PCI

[1167] C. Damm, M. Holzer, and K.-J. Lange. Parallel complexity of iterated morphisms and the arithmetic of small numbers. Lecture Notes in Computer Science, 629:227–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Derencourt:1992:CPW

[1168] D. Derencourt, J. Karhumaki, M. Latteux, and A. Terlutte. On computational power of weighted finite automata. Lecture Notes in Computer Science, 629:236–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feldmann:1992:SEN

[1169] R. Feldmann and P. Mysliwietz. The shuffle exchange network has a Hamiltonian path. Lecture Notes in Computer Science, 629:246–??, 1992. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gastin:1992:PPC

[1170] P. Gastin and A. Petit. Poset properties of complex traces. Lecture Notes in Computer Science, 629:255–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goerdt:1992:TU

[1171] A. Goerdt. A treshold for unsatisfiability. Lecture Notes in Computer Science, 629:264–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gold:1992:DSP

[1172] R. Gold. Dataflow semantics for Petri nets. Lecture Notes in Computer Science, 629:275–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guessarian:1992:ABS

[1173] I. Guessarian and M. Veloso-Peixoto.

About boundedness for some DAT-ALOG and DATALOGsup neg programs. Lecture Notes in Computer Science, 629:284–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagerup:1992:MSS

[1174] T. Hagerup and O. Petersson. Merging and sorting strings in parallel. Lecture Notes in Computer Science, 629: 298–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jancar:1992:CCF

[1175] P. Jancar, F. Mraz, and M. Platek. Characterization of context-free languages by erasing automata. Lecture Notes in Computer Science, 629:307– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kari:1992:IDW

[1176] L. Kari. Insertion and deletion of words: Determinism and reversibility. Lecture Notes in Computer Science, 629:315–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kascak:1992:SUO

[1177] F. Kascak. Small universal one-state linear operator algorithm. Lecture Notes in Computer Science, 629:327– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laneve:1992:MCP

[1178] C. Laneve and U. Montanari. Mobility in the CC-Paradigm. Lecture Notes in Computer Science, 629:336–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lange:1992:EPI

[1179] K.-J. Lange and P. Rossmanith. The emptiness problem for intersections of regular languages. Lecture Notes in Computer Science, 629:346–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leung:1992:FAL

[1180] H. Leung. On finite automata with limited nondeterminism. Lecture Notes in Computer Science, 629:355–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Litovsky:1992:DCL

[1181] I. Litovsky, Y. Metivier, and E. Sopena.

Definitions and comparisons of local computations on graphs. Lecture Notes in Computer Science, 629:364– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1992:EUU

[1182] B. Martin. Efficient unidimensional universal cellular automaton. Lecture Notes in Computer Science, 629:374– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maruyama:1992:ITW

[1183] O. Maruyama and S. Miyano. Inferring a tree from walks. Lecture Notes in Computer Science, 629:383–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayordomo:1992:AES

[1184] E. Mayordomo. Almost every set in exponential time is P-bi-immune. Lecture Notes in Computer Science, 629: 392–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murphy:1992:FSO

[1185] D. Murphy and A. Poigne. A functorial semantics for observed concurrency. Lecture Notes in Computer Science, 629:401–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ochmanski:1992:MCS

[1186] E. Ochmanski. Modelling concurrency with semi-commutations. Lecture Notes in Computer Science, 629: 412–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pedersen:1992:DPC

[1187] J. Pedersen. Decision problems for cellular automata and their semigroups. Lecture Notes in Computer Science, 629:421–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinna:1992:NE

[1188] G. M. Pinna and A. Poigne. On the nature of events. Lecture Notes in Computer Science, 629:430–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Preilowski:1992:NPA

[1189] W. Preilowski, E. Dahlhaus, and G. Wechsung. New parallel algorithms for convex hull and triangulation in 3dimensional space. Lecture Notes in Computer Science, 629:442–??, 1992.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Przymusinski:1992:TSC

[1190] T. C. Przymusinski. Two simple characterizations of well-founded semantics. Lecture Notes in Computer Science, 629:451–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramesh:1992:FAS

[1191] S. Ramesh. Fully abstract semantics for higher order communicating systems. Lecture Notes in Computer Science, 629:463–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reimen:1992:STA

[1192] N. Reimen. Superposable trellis automata. Lecture Notes in Computer Science, 629:472–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roos:1992:MPH

[1193] T. Roos and G. Albers. Maintaining proximity in higher dimensional spaces. Lecture Notes in Computer Science, 629:483–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Szilard:1992:CRL

[1194] A. Szilard, S. Yu, K. Zhang, and J. Shallit. Characterizing regular languages with polynomial densities. Lecture Notes in Computer Science, 629: 494–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1992:SSC

[1195] D. Wang. A strategy for speedingup the computation of characteristic sets. Lecture Notes in Computer Science, 629:504–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wrathall:1992:ORT

[1196] C. Wrathall, V. Diekert, and F. Otto. One-rule trace-rewriting systems and confluence. Lecture Notes in Computer Science, 629:511–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Milner:1992:ILP

[1197] R. Milner. Invited lecture: The polyadic pi-calculus. Lecture Notes in Computer Science, 630:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boreale:1992:TEM

[1198] M. Boreale and R. De Nicola. Testing equivalences for mobile processes. Lecture Notes in Computer Science, 630: 2–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jategaonkar:1992:TEP

[1199] L. Jategaonkar and A. Meyer. Testing equivalence for Petri nets with action refinement. Lecture Notes in Computer Science, 630:17–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sangiorgi:1992:PWB

[1200] D. Sangiorgi and R. Milner. The problem of “weak bisimulation up to”. Lecture Notes in Computer Science, 630: 32–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brinksma:1992:UFM

[1201] E. Brinksma. On the uniqueness of fixpoints modulo observation congruence. Lecture Notes in Computer Science, 630:47–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Groote:1992:VPS

[1202] J. F. Groote and F. Moller. Verification of parallel systems via decomposition. Lecture Notes in Computer Science, 630:62–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerth:1992:IRR

[1203] R. Gerth, R. Kuiper, and J. Segers.

Interface refinement in reactive systems. Lecture Notes in Computer Science, 630:77–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hennessy:1992:CTP

[1204] M. Hennessy. Concurrent testing of processes. Lecture Notes in Computer Science, 630:94–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boudol:1992:TPL

[1205] G. Boudol, I. Castellani, M. Hennessy, and A. Kiehn. A theory of processes with localities. Lecture Notes in Computer Science, 630:108–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burkart:1992:MCC

[1206] O. Burkart and B. Steffen. Model checking for context-free processes. Lecture Notes in Computer Science, 630:123–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christensen:1992:BED

[1207] S. Christensen, H. Huttel, and Colin Stirling. Bisimulation equivalence is decidable for all context-free processes. Lecture Notes in Computer Science, 630:138–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christensen:1992:DBD

[1208] S. Christensen. Distributed bisimularity is decidable for a class of infinite state-space processes. Lecture Notes in Computer Science, 630:148–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pnueli:1992:ITH

[1209] A. Pnueli. Invited talk: How vital is liveness? verifying timing properties of reactive and hybrid systems. Lecture Notes in Computer Science, 630:162– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stomp:1992:PSP

[1210] F. A. Stomp. Preserving specific properties in program development. Lecture Notes in Computer Science, 630: 176–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peled:1992:SSG

[1211] D. Peled. Sometimes “some” is as good as “all”. Lecture Notes in Computer Science, 630:192–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaivola:1992:WCS

[1212] R. Kaivola and A. Valmari. The weakest compositional semantic equivalence preserving nexttime-less linear temporal logic. Lecture Notes in Computer Science, 630:207–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goltz:1992:PTL

[1213] U. Goltz, R. Kuiper, and W. Penczek.

Propositional temporal logics and equivalences. Lecture Notes in Computer Science, 630:222–??, 1992. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pratt:1992:ITD

[1214] V. R. Pratt. Invited talk: The duality of time and information. Lecture Notes in Computer Science, 630:237– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goubault:1992:HHD

[1215] E. Goubault and T. P. Jensen. Homology of higher dimensional automata. Lecture Notes in Computer Science, 630:254–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rensink:1992:PC

[1216] A. Rensink. Posets for configurations! Lecture Notes in Computer Science, 630:269–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meseguer:1992:SPN

[1217] J. Meseguer, U. Montanari, and V. Sassone. On the semantics of Petri nets. Lecture Notes in Computer Science, 630:286–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Badouel:1992:SOS

[1218] E. Badouel and P. Darondeau. Structural operational specifications and trace automata. Lecture Notes in Computer Science, 630:302–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lynch:1992:ITM

[1219] N. Lynch, B. Lampson, and J. Søgaard-Andersen. Invited talk: At-most-once message delivery: A case study in algorithm verification. Lecture Notes in Computer Science, 630:317–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reingold:1992:GAP

[1220] N. Reingold, D.-W. Wang, and L. D. Zuck. Games I/O automata play. Lecture Notes in Computer Science, 630: 325–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alur:1992:MTT

[1221] R. Alur, C. Courcoubetis, N. Halbwachs, D. Dill, and H. Wong-Toi. Minimization of timed transition systems. Lecture Notes in Computer Science, 630:340–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davies:1992:UCV

[1222] J. Davies and S. Schneider. Using CSP to verify a timed protocol over a fair medium. Lecture Notes in Computer Science, 630:355–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weinberg:1992:TER

[1223] H. B. Weinberg and L. D. Zuck. Timed ethernet: Real-time formal specification of ethernet. Lecture Notes in Computer Science, 630:370–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karjoth:1992:ILS

[1224] G. Karjoth. Implementing LOTOS specifications by communicating state machines. Lecture Notes in Computer Science, 630:386–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeten:1992:ITD

[1225] J. C. M. Baeten and J. A. Bergstra.

Invited talk: Discrete time process algebra. Lecture Notes in Computer Science, 630:401–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klusener:1992:SST

[1226] S. Klusener. The silent step in time. Lecture Notes in Computer Science, 630:421–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vaandrager:1992:ATT

[1227] F. Vaandrager and N. Lynch. Action transducers and timed automata. Lecture Notes in Computer Science, 630: 436–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsen:1992:CVP

[1228] K. G. Larsen and A. Skou. Compositional verification of probabilistic processes. Lecture Notes in Computer Science, 630:456–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeten:1992:APP

[1229] J. C. M. Baeten, J. A. Bergstra, and S. A. Smolka. Axiomatizing probabilistic processes: ACP with generative probabilities. Lecture Notes in Computer Science, 630:472–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shapiro:1992:EAC

[1230] E. Shapiro. Embeddings among concurrent programming languages. Lecture Notes in Computer Science, 630: 486–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rabinovich:1992:LTL

[1231] A. Rabinovich. Logic of trace languages. Lecture Notes in Computer Science, 630:504–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parrow:1992:MSV

[1232] J. Parrow and P. Sjodin. Multiway synchronization verified with coupled simulation. Lecture Notes in Computer Science, 630:518–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

German:1992:PGM

[1233] S. M. German. Programming in a general model of synchronization. Lecture Notes in Computer Science, 630:534– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maraninchi:1992:OCS

[1234] F. Maraninchi. Operational and compositional semantics of synchronous automaton compositions. Lecture Notes in Computer Science, 630:550– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agha:1992:TTA

[1235] G. Agha, I. A. Mason, S. Smith, and C. Talcott. Towards a theory of actor computation. Lecture Notes in Computer Science, 630:565–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanus:1992:ILI

[1236] M. Hanus. Invited lecture: Improving control of logic programs by using functional logic languages. Lecture Notes in Computer Science, 631:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuchen:1992:IPI

[1237] H. Kuchen, J. J. Moreno-Navarro, and M. V. Hermenegildo. Independent AND-Parallel implementation of narrowing. Lecture Notes in Computer Science, 631:24–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dorochevsky:1992:BTG

[1238] M. Dorochevsky and A. Veron. Binding techniques and garbage collection for OR-Parallel CLP systems. Lecture Notes in Computer Science, 631:39–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gomard:1992:PAL

[1239] C. K. Gomard and P. Sestoft. Path analysis for lazy data structures. Lecture Notes in Computer Science, 631: 54–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Apt:1992:WOC

[1240] K. R. Apt and A. Pellegrini. Why the occur-check is not a problem. Lecture Notes in Computer Science, 631:69–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Attali:1992:IEN

[1241] I. Attali, J. Chazarain, and S. Gilette.

Incremental evaluation of natural semantics specifications. Lecture Notes in Computer Science, 631:87–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barthelemy:1992:SOP

[1242] F. Barthelemy and E. Villemonte de la Clergerie. Subsumption-oriented pushdown automata. Lecture Notes in Computer Science, 631:100–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kahrs:1992:UUL

[1243] S. Kahrs. Unlimp, uniqueness as a leitmotiv for implementation. Lecture Notes in Computer Science, 631:115– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pennings:1992:UCF

[1244] M. Pennings, D. Swierstra, and H. Vogt. Using cached functions and constructors for incremental attribute evaluation. Lecture Notes in Computer Science, 631:130–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rosendahl:1992:SAA

[1245] M. Rosendahl. Strictness analysis for attribute grammars. Lecture Notes in Computer Science, 631:145–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saidi:1992:CDT

[1246] S. Saidi and J.-F. Boulicaut. Checking and debugging of two-level grammars. Lecture Notes in Computer Science, 631:158–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bossi:1992:SF

[1247] A. Bossi, N. Cocco, and S. Etalle. On safe folding. Lecture Notes in Computer Science, 631:172–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amtoft:1992:UFT

[1248] T. Amtoft. Unfold/fold transformations preserving termination properties. Lecture Notes in Computer Science, 631:187–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alexandre:1992:TTL

[1249] F. Alexandre. A technique for transforming logic programs by fold-unfold transformations. Lecture Notes in Computer Science, 631:202–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henrard:1992:FED

[1250] J. Henrard and B. Le Charlier. FOLON: An environment for declarative construction of logic programs. Lecture Notes in Computer Science, 631:217–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hans:1992:CIS

[1251] W. Hans. A complete indexing scheme for WAM-Based abstract machines. Lecture Notes in Computer Science, 631:232–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krall:1992:FPB

[1252] A. Krall and T. Berger. Fast Prolog with a VAM_{1p} based Prolog compiler. Lecture Notes in Computer Science, 631:245–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holzbaur:1992:MVA

[1253] C. Holzbaur. Metastructures vs. attributed variables in the context of extensible unification. Lecture Notes in Computer Science, 631:260–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cousot:1992:ILC

[1254] P. Cousot and R. Cousot. Invited lecture: Comparing the Galois connection and widening/ narrowing approaches to abstract interpretation. Lecture Notes in Computer Science, 631:269– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verschaetse:1992:DLS

[1255] K. Verschaetse and D. De Schreye.

Derivation of linear size relations by abstract interpretation. Lecture Notes in Computer Science, 631:296–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Englebert:1992:GAI

[1256] V. Englebert, B. Le Charlier, D. Roland, and P. Van Hentenryck. Generic abstract interpretation algorithms for prolog: Two optimization techniques and their experimental evaluation. Lecture Notes in Computer Science, 631: 311–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoa:1992:BIH

[1257] A. Hui Bon Hoa. A bottom-up interpreter for a higher-order logic programming language. Lecture Notes in Computer Science, 631:326–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muck:1992:CEC

[1258] A. Muck. CAMEL: An extension of the categorical abstract machine to compile functional/logic programs. Lecture Notes in Computer Science, 631:341– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hans:1992:ILE

[1259] W. Hans, R. Loogen, and S. Winkler. On the interaction of lazy evaluation and backtracking. Lecture Notes in Computer Science, 631:355–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamkar:1992:IDS

[1260] M. Kamkar, N. Shahmehri, and P. Fritzson. Interprocedural dynamic slicing. Lecture Notes in Computer Science, 631:370–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nilsson:1992:ADL

[1261] H. Nilsson and P. Fritzson. Algorithmic debugging of lazy functional languages. Lecture Notes in Computer Science, 631:385–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ducasse:1992:GTQ

[1262] M. Ducasse. A general trace query mechanism based on Prolog. Lecture Notes in Computer Science, 631:400– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plaza:1992:FDL

[1263] J. A. Plaza. Fully declarative logic programming. Lecture Notes in Computer Science, 631:415–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lindstrom:1992:OLS

[1264] G. Lindstrom, J. Maluszynski, and T. Ogi. Our LIPS are sealed: Interfacing functional and logic programming systems. Lecture Notes in Computer Science, 631:428–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alpuente:1992:AII

[1265] M. Alpuente, M. Falaschi, and F. Manzo. Analyses of inconsistency for incremental equational logic programming. Lecture Notes in Computer Science, 631:443–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rebelsky:1992:TIL

[1266] S. A. Rebelsky. I/O trees and interactive lazy functional programming. Lecture Notes in Computer Science, 631: 458–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamel:1992:UEL

[1267] L. H. Hamel. UCG-E: An equational logic programming system. Lecture Notes in Computer Science, 631:473– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cattrall:1992:RPS

[1268] D. Cattrall and C. Runciman. A relational programming system with inferred representations. Lecture Notes in Computer Science, 631:475–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moura:1992:IAS

[1269] H. Moura. An implementation of action semantics (summary). Lecture Notes in Computer Science, 631:477–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tarau:1992:BCP

[1270] P. Tarau. BinProlog: a continuation passing style Prolog engine. Lecture Notes in Computer Science, 631:479– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thiemann:1992:LLT

[1271] P. Thiemann. LaToKi: A language toolkit for bottom-up evaluation of functional programs. Lecture Notes in Computer Science, 631:481–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reus:1992:IHO

[1272] B. Reus. Implementing higher-order functions in an algebraic specification language with narrowing. Lecture Notes in Computer Science, 631:483– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vilo:1992:IAG

[1273] J. Vilo. Implementing attribute grammars by computational models. Lecture Notes in Computer Science, 631: 485–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1992:PLA

[1274] C. M. Li and P. Y. Gloess. ProQuery:

Logical access to an OODB. Lecture Notes in Computer Science, 631:487– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Solnon:1992:IIR

[1275] C. Solnon and M. Rueher. Inference of inheritance relationships from Prolog programs: a system developed with Prolog III. Lecture Notes in Computer Science, 631:489–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barth:1992:CPM

[1276] P. Barth. CLP(PB), A metainterpreter in CLP(R). Lecture Notes in Computer Science, 631:491–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ait-Kaci:1992:OOO

[1277] H. Ait-Kaci. Outline of an objectoriented calculus of higher type (invited talk). Lecture Notes in Computer Science, 632:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrig:1992:HLR

[1278] H. Ehrig and F. Parisi-Presicce. Highlevel-replacement systems for equational algebraic specifications. Lecture Notes in Computer Science, 632:3–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lescanne:1992:TRS

[1279] P. Lescanne. Termination of rewrite systems by elementary interpretations. Lecture Notes in Computer Science, 632:21–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gnaedig:1992:TOS

[1280] I. Gnaedig. Termination of order-sorted rewriting. Lecture Notes in Computer Science, 632:37–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gramlich:1992:GSC

[1281] B. Gramlich. Generalized sufficient conditions for modular termination of rewriting. Lecture Notes in Computer Science, 632:53–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Apt:1992:TFO

[1282] K. R. Apt, E. Marchiori, and C. Palamidessi. A theory of first-order built-in’s of Prolog. Lecture Notes in Computer Science, 632:69–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gabbrielli:1992:FSP

[1283] M. Gabbrielli and M. C. Meo. Fixpoint semantics for partial computed answer substitutions and call patterns. Lecture Notes in Computer Science, 632:84–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbuti:1992:OSP

[1284] R. Barbuti, M. Codish, R. Giacobazzi, and M. Maher. Oracle semantics for Prolog. Lecture Notes in Computer Science, 632:100–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hermann:1992:RBP

[1285] M. Hermann. On the relation between primitive recursion, schematization and divergence. Lecture Notes in Computer Science, 632:115–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffmann:1992:TRS

[1286] B. Hoffmann. Term rewriting with sharing and memoization. Lecture Notes in Computer Science, 632:128– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoy:1992:DT

[1287] S. Antoy. Definitional trees. Lecture Notes in Computer Science, 632:143– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meseguer:1992:MLP

[1288] J. Meseguer. Multiparadigm logic programming (invited talk). Lecture Notes in Computer Science, 632:158– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoon:1992:NLR

[1289] Hong Hoon. Non-linear real constraints in constraint logic programming. Lecture Notes in Computer Science, 632:201–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopez-Fraguas:1992:GSC

[1290] F.-J. Lopez-Fraguas. A general scheme for constraint functional logic programming. Lecture Notes in Computer Science, 632:213–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanus:1992:IRN

[1291] M. Hanus. Incremental rewriting in narrowing derivations. Lecture Notes in Computer Science, 632:228–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Middeldorp:1992:CCR

[1292] A. Middeldorp and E. Hamoen. Counterexamples to completeness results for basic narrowing (extended abstract).

Lecture Notes in Computer Science, 632:244–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Echahed:1992:UNS

[1293] R. Echahed. Uniform narrowing strategies. Lecture Notes in Computer Science, 632:259–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lysne:1992:PCC

[1294] O. Lysne. Proof by consistency in constructive systems with final algebra semantics. Lecture Notes in Computer Science, 632:276–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bundgen:1992:FAG

[1295] R. Bundgen and H. Eckhardt. A fast algorithm for ground normal form analysis. Lecture Notes in Computer Science, 632:291–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hardin:1992:ECL

[1296] T. Hardin. Eta-conversion for the languages of explicit substitutions. Lecture Notes in Computer Science, 632: 306–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

King:1992:SAC

[1297] A. King and P. Soper. Serialisation analysis of concurrent logic programs. Lecture Notes in Computer Science, 632:322–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giovini:1992:ITP

[1298] A. Giovini, F. Morando, and A. Capani. Implementation of a toolset for prototyping algebraic specifications of concurrent systems. Lecture Notes in Computer Science, 632:335–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laneve:1992:APE

[1299] C. Laneve and U. Montanari. Axiomatizing permutation equivalence in the lambda-calculus. Lecture Notes in Computer Science, 632:350–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maher:1992:CVL

[1300] M. Maher. A CLP view of logic programming (invited talk). Lecture Notes in Computer Science, 632:364– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aravindan:1992:PDL

[1301] C. Aravindan and Minh Dung Phan. Partial deduction of logic programs w.r.t. well-founded semantics. Lecture Notes in Computer Science, 632:384– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balbiani:1992:FLP

[1302] P. Balbiani. The finiteness of logic programming derivations. Lecture Notes in Computer Science, 632:403–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bachmair:1992:TPH

[1303] L. Bachmair, H. Ganzinger, and U. Waldmann. Theorem proving for hierarchic first-order theories. Lecture Notes in Computer Science, 632:420– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Socher-Ambrosius:1992:GOS

[1304] R. Socher-Ambrosius. A goal oriented strategy based on completion. Lecture Notes in Computer Science, 632:435– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boudet:1992:SET

[1305] A. Boudet and E. Contejean. On nsyntactic equational theories. Lecture Notes in Computer Science, 632:446– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vakarelov:1992:MTA

[1306] D. Vakarelov. A modal theory of arrows. arrow logics I. Lecture Notes in Computer Science, 633:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MacNish:1992:KMS

[1307] C. MacNish. Knowledge without modality: A simplified framework for chronological ignorance. Lecture Notes in Computer Science, 633:25–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pliuskevicius:1992:DCS

[1308] R. Pliuskevicius. Design complete sequential calculus for continuous fixpoint temporal logic. Lecture Notes in Computer Science, 633:36–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muskens:1992:LOC

[1309] R. Muskens. Logical omniscience and classical logic. Lecture Notes in Computer Science, 633:52–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwast:1992:WIT

[1310] K. L. Kwast and S. van Denneheuvel. Weak implication: Theory and applications. Lecture Notes in Computer Science, 633:65–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Royer:1992:DIR

[1311] V. Royer and J. J. Quantz. Deriving inference rules for terminological logics. Lecture Notes in Computer Science, 633:84–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fronhofer:1992:LPL

[1312] B. Fronhofer. Linear proofs and linear logic. Lecture Notes in Computer Science, 633:106–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weydert:1992:RRA

[1313] E. Weydert. Relevance and revision — about generalizing syntax-based belief revision. Lecture Notes in Computer Science, 633:126–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rott:1992:MBC

[1314] H. Rott. Modellings for belief change:

Base contraction, multiple contraction, and epistemic entrenchment. Lecture Notes in Computer Science, 633:139– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Froidevaux:1992:FDL

[1315] C. Froidevaux and J. Mengin. A framework for default logics. Lecture Notes in Computer Science, 633:154–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hunter:1992:CPN

[1316] A. Hunter. A conceptualization of preferences in non-monotonic proof theory. Lecture Notes in Computer Science, 633:174–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vreeswijk:1992:RDA

[1317] G. Vreeswijk. Reasoning with defeasible arguments: Examples and applications. Lecture Notes in Computer Science, 633:189–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Besnard:1992:ADG

[1318] Ph. Besnard and E. Gregoire. About deductive generalization. Lecture Notes in Computer Science, 633:212– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernando:1992:TSD

[1319] T. Fernando. Transition systems and dynamic semantics. Lecture Notes in Computer Science, 633:232–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mircheva:1992:DSI

[1320] M. S. Mircheva. Declarative semantics for inconsistent database programs. Lecture Notes in Computer Science, 633:252–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bittel:1992:TBT

[1321] O. Bittel. Tableau-based theorem proving and synthesis of lambda-terms in the intuitionistic logic. Lecture Notes in Computer Science, 633:262–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goltz:1992:CTS

[1322] H.-J. Goltz. A constructive type system based on data terms. Lecture Notes in Computer Science, 633:279–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zabel:1992:ORP

[1323] N. Zabel. An ordered resolution and paramodulation calculus for finite many-valued logics. Lecture Notes in Computer Science, 633:304–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prehofer:1992:ECL

[1324] C. Prehofer. An efficient constraint language for polymorphic order-sorted resolution. Lecture Notes in Computer Science, 633:319–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pereira:1992:DTW

[1325] L. M. Pereira, J. J. Alferes, and J. N. Aparicio. Default theory for well founded semantics with explicit negation. Lecture Notes in Computer Science, 633:339–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Furbach:1992:CAD

[1326] U. Furbach. Computing answers for disjunctive logic programs. Lecture Notes in Computer Science, 633:357– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Witteveen:1992:ELP

[1327] C. Witteveen. Expanding logic programs. Lecture Notes in Computer Science, 633:373–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herre:1992:DLP

[1328] H. Herre and D. Pearce. Disjunctive logic programming, constructivity and strong negation. Lecture Notes in Computer Science, 633:391–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhatt:1992:TFS

[1329] S. N. Bhatt, F. R. K. Chung, F. T.

Leighton, and A. L. Rosenberg. Tolerating faults in synchronization networks. Lecture Notes in Computer Science, 634:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turdik:1992:IH

[1330] P. Turdik. On incomplete hypercubes. Lecture Notes in Computer Science, 634:13–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamanaka:1992:RNH

[1331] N. Hamanaka, J. Nakagoshi, and T. Tanaka. Reducing network hardware quantity by employing multiprocessor cluster structure in distributed memory parallel processors. Lecture Notes in Computer Science, 634:25–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ziavras:1992:CMR

[1332] S. G. Ziavras. Connection machine results for pyramid embedding algorithms. Lecture Notes in Computer Science, 634:31–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hatz:1992:INB

[1333] V. Hatz. Interconnection networks based on block designs. Lecture Notes in Computer Science, 634:37–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

David:1992:PMC

[1334] V. David, C. Fraboul, J. Y. Rousselot, and P. Siron. Partitioning and mapping communication graphs on a modular reconfigurable parallel architecture. Lecture Notes in Computer Science, 634:43–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Munthe-Kaas:1992:GSE

[1335] H. Munthe-Kaas. Generalized shuffleexchange networks. Lecture Notes in Computer Science, 634:49–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leu:1992:ERM

[1336] E. Leu and A. Shiper. Execution replay: A mechanism for integrating a visualization tool with a symbolic debugger. Lecture Notes in Computer Science, 634:55–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bauch:1992:SMD

[1337] A. Bauch, T. Kosch, E. Maehle, and W. Oberloer. The software-monitor DELTA-T and its use for performance measurements of some farming variants on the Multi-Transputer system DAMP. Lecture Notes in Computer Science, 634:67–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bemmerl:1992:VMP

[1338] T. Bemmerl and P. Braun. Visualization of message passing parallel programs. Lecture Notes in Computer Science, 634:79–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mansour:1992:PPO

[1339] N. Mansour and G. C. Fox. Parallel physical optimization algorithms for data mapping. Lecture Notes in Computer Science, 634:91–??, 1992.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becher:1992:PMP

[1340] J. D. Becher and K. L. Beck. Profiling on a massively parallel computer. Lecture Notes in Computer Science, 634:97–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hersch:1992:MMV

[1341] R. D. Hersch, B. Tonelli, and B. Krummenacher. A multiprocessor multiwindow visualization subsystem. Lecture Notes in Computer Science, 634:103– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beranek:1992:DRD

[1342] A. Beranek. Data race detection based on execution replay for parallel applications. Lecture Notes in Computer Science, 634:109–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adamo:1992:CPE

[1343] J. M. Adamo, C. Bonello, and L. Trejo. The CNET programming environment: An overview. Lecture Notes in Computer Science, 634:115–??, 1992. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lemke:1992:PCV

[1344] M. Lemke and D. Quinlan. P++, a C++ virtual shared grids based programming environment for architectureindependent development of structured grid applications. Lecture Notes in Computer Science, 634:121–??, 1992.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arvind:1992:DCR

[1345] D. K. Arvind and J. Knight. Detection of concurrency-related errors in joyce. Lecture Notes in Computer Science, 634:127–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lavault:1992:AED

[1346] C. Lavault. Analysis of an efficient distributed algorithm for mutual exclusion (average-case analysis of path reversal). Lecture Notes in Computer Science, 634:133–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schepers:1992:IPD

[1347] J. Schepers. Invariance properties in distributed systems. Lecture Notes in Computer Science, 634:145–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makhaniok:1992:SPP

[1348] M. Makhaniok and R. Manner. Synchronization of parallel processes in distributed systems. Lecture Notes in Computer Science, 634:157–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moukeli:1992:SPC

[1349] P. Moukeli. Statistical probabilistic clock synchronization algorithm. Lecture Notes in Computer Science, 634: 163–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nowinski:1992:SAM

[1350] W. L. Nowinski. A SIMD architecture for medical imaging. Lecture Notes in Computer Science, 634:169–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1992:CIP

[1351] R. Lin and S. Olariu. Computing the inner product on reconfigurable buses with shift switching. Lecture Notes in Computer Science, 634:181–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1992:NSA

[1352] S. P. S. Lam. A novel sorting array processor. Lecture Notes in Computer Science, 634:193–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horton:1992:TPS

[1353] G. Horton, R. Knirsch, and H. Vollath. The time-parallel solution of parabolic partial differencial equations using the frequency-filtering method. Lecture Notes in Computer Science, 634:205– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Griebel:1992:CTP

[1354] M. Griebel, W. Huber, U. Rude, and T. Stortkuhl. The combination technique for parallel sparse-grid-preconditioning or, solution of PDEs on workstation networks. Lecture Notes in Computer Science, 634:217–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schafer:1992:CDM

[1355] M. Schafer, M. M. Gutzmann, and M. Schwehm. Comparing the DAP, Meiko and Suprenum with a fluid dynamic benchmark. Lecture Notes in Computer Science, 634:229–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kabakchiev:1992:PDA

[1356] C. A. Kabakchiev and V. P. Behar.

Parallel detection algorithm of radar signals. Lecture Notes in Computer Science, 634:241–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andonov:1992:ELS

[1357] R. Andonov and P. Quinton. Efficient linear systolic array for the knapsack problem. Lecture Notes in Computer Science, 634:247–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xue:1992:LRA

[1358] J. Xue. On the loading, recovery and access of stationary data in systolic arrays. Lecture Notes in Computer Science, 634:259–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Myoupo:1992:DML

[1359] J.-F. Myoupo and A.-C. Fabret. Designing modular linear systolic arrays using dependence graph regular partitions. Lecture Notes in Computer Science, 634:265–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lang:1992:RSB

[1360] B. Lang. Reducing symmetric banded matrices to tridiagonal form — A comparison of a new parallel algorithm with two serial algorithms on the IPSC/860. Lecture Notes in Computer Science, 634:271–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilkinson:1992:IBI

[1361] B. Wilkinson and L. S. Mulholland. An implementation of the BLAS on the i860: A RISC approach to software for RISC devices. Lecture Notes in Computer Science, 634:283–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charrier:1992:PMP

[1362] P. Charrier and J. Roman. Partitioning and mapping for parallel nested dissection on distributed memory architectures. Lecture Notes in Computer Science, 634:295–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsao:1992:AST

[1363] N.-K. Tsao. On the accuracy of solving triangular systems in Parallel-III. Lecture Notes in Computer Science, 634:307–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amestoy:1992:LAC

[1364] P. R. Amestoy, M. J. Dayde, I. S. Duff, and P. Morere. Linear algebra calculations on the BBN TC2000. Lecture Notes in Computer Science, 634:319– 330, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1992:PHA

[1365] G. Chen, H. B. Keller, S. H. Lui, and B. Roux. Parallel homotopy algorithm for large sparse generalized eigenvalue problems: Application to hydrodynamic stability analysis. Lecture Notes in Computer Science, 634:331– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stpiczynski:1992:PAS

[1366] P. Stpiczynski. Parallel algorithms for solving linear recurrence systems. Lecture Notes in Computer Science, 634: 343–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bar-On:1992:NPF

[1367] I. Bar-On and O. Munk. A new parallel factorization A = DDsup t BC for band symmetric positive definite matrices. Lecture Notes in Computer Science, 634:349–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waring:1992:CER

[1368] L. C. Waring and M. Clint. Computation of the eigenvalues of real symmetric matrices using a processor farm. Lecture Notes in Computer Science, 634:355–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bull:1992:NPA

[1369] J. M. Bull and T. L. Freeman. Numerical performance of an asynchronous Jacobi iteration. Lecture Notes in Computer Science, 634:361–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Milovanovic:1992:MIA

[1370] E. I. Milovanovic, I. Z. Milovanovic, and M. K. Stojcev. Matrix inversion algorithm for linear array processor. Lecture Notes in Computer Science, 634:367–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petiton:1992:MPP

[1371] S. Petiton and C. Weill-Duflos. Massively parallel preconditioners for the sparse conjugate gradient method. Lecture Notes in Computer Science, 634: 373–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:1992:BAO

[1372] S. J. Thomas. A block algorithm for orthogonalization in elliptic norms. Lecture Notes in Computer Science, 634: 379–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deCampos:1992:APA

[1373] G. Lino de Campos. Asynchronous polycyclic architecture. Lecture Notes in Computer Science, 634:387–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bird:1992:CTM

[1374] P. L. Bird, N. P. Topham, and S. Manoharan. A comparison of two memory models for high performance computers. Lecture Notes in Computer Science, 634:399–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruighaver:1992:DMA

[1375] A. B. Ruighaver. A decoupled multicomputer architecture with optical full interconnection. Lecture Notes in Computer Science, 634:405–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mzoughi:1992:VHS

[1376] A. Mzoughi, M. Lalam, and D. Litaize.

Very high speed vectorial processors using serial multiport memory as data memory. Lecture Notes in Computer Science, 634:411–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keane:1992:RMP

[1377] J. A. Keane and K. R. Mayes. Resource management on a packet-based parallel graph reduction machine. Lecture Notes in Computer Science, 634: 417–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Govindarajan:1992:LCM

[1378] R. Govindarajan and S. S. Nemawarkar. A large context multithreaded architecture. Lecture Notes in Computer Science, 634:423–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hellwagner:1992:PER

[1379] H. Hellwagner. On the practical efficiency of randomized shared memory. Lecture Notes in Computer Science, 634:429–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lahjomri:1992:KSV

[1380] Z. Lahjomri and T. Priol. KOAN: A shared virtual memory for the iPSC/2 hypercube. Lecture Notes in Computer Science, 634:441–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murer:1992:SDS

[1381] S. Murer and P. Farber. A scalable distributed shared memory. Lecture Notes in Computer Science, 634:453– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roch:1992:CPL

[1382] J. L. Roch, A. Vermeerbergen, and G. Villard. Cost prediction for load balancing: Application to algebraic computations. Lecture Notes in Computer Science, 634:467–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shteiman:1992:EAL

[1383] K. Shteiman, D. Feitelson, L. Rudolph, and I. Exman. Envelopes in adaptive local queues for MIMD load balancing. Lecture Notes in Computer Science, 634:479–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tarnvik:1992:DPT

[1384] E. Tarnvik. Dynamo — A portable tool for dynamic load balancing on distributed memory multicomputers. Lecture Notes in Computer Science, 634: 485–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levaire:1992:SCD

[1385] J.-L. Levaire. Semantics and compilation of the data-parallel switch statement. Lecture Notes in Computer Science, 634:491–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Essafi:1992:KPF

[1386] H. Essafi, M. Pic, and D. Juvin. Kproject — first step: To improve data manipulations and representations on parallel computers. Lecture Notes in Computer Science, 634:503–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muchnick:1992:FCP

[1387] V. B. Muchnick and A. V. Shafarenko.

F-code: A portable software platform for data-parallel languages. Lecture Notes in Computer Science, 634:509– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gengler:1992:PBF

[1388] M. Gengler and G. Coray. A parallel best-first B and B with synchronization phases. Lecture Notes in Computer Science, 634:515–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1992:TUT

[1389] H. Wu. On tests of uniform tree circuits. Lecture Notes in Computer Science, 634:527–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1992:IPS

[1390] J. Chen and C. Levcopoulos. Improved parallel sorting of presorted sequences. Lecture Notes in Computer Science, 634:539–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duboux:1992:MDM

[1391] T. Duboux, A. Ferreira, and M. Gastaldo. MIMD dictionary machines: From theory to practice. Lecture Notes in Computer Science, 634:545– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klingspor:1992:PSO

[1392] F. Klingspor, D. Luhofer, and T. Rottke. Parallel searching for 3D-objects. Lecture Notes in Computer Science, 634:551–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roch:1992:PGL

[1393] J. L. Roch and G. Villard. Parallel gcd and lattice basis reduction. Lecture Notes in Computer Science, 634: 557–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akl:1992:PMS

[1394] S. G. Akl and K. Qiu. Parallel minimum spanning forest algorithms on the star and pancake interconnection networks. Lecture Notes in Computer Science, 634:565–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Embrechts:1992:PAD

[1395] H. Embrechts and D. Roose. Parallel algorithms for the distance transformation. Lecture Notes in Computer Science, 634:571–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanKleij:1992:MSD

[1396] R. van Kleij and F. Kuijper. A multiprocessor system for displaying quadric CSG models. Lecture Notes in Computer Science, 634:583–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Olariu:1992:EIC

[1397] S. Olariu, J. L. Schwing, and J. Zhang.

Efficient image computations on reconfigurable meshes. Lecture Notes in Computer Science, 634:589–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levaire:1992:MPT

[1398] J. L. Levaire and S. Ubeda. A multiscale parallel thinning algorithm. Lecture Notes in Computer Science, 634: 595–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerasoulis:1992:SSP

[1399] A. Gerasoulis and T. Yang. Static scheduling of parallel programs for message passing architectures. Lecture Notes in Computer Science, 634:601– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanDongen:1992:PTM

[1400] V. H. Van Dongen, R. G. Gao, and Q. Ning. A polynomial time method for optimal software pipelining. Lecture Notes in Computer Science, 634: 613–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gasperoni:1992:SLP

[1401] F. Gasperoni and U. Schwiegelshohn. Scheduling loops on parallel processors: A simple algorithm with close to optimum performance. Lecture Notes in Computer Science, 634:625–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kushner:1992:APG

[1402] E. J. Kushner. Automatic parallelization of grid-based applications for the iPSC/860. Lecture Notes in Computer Science, 634:637–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

David:1992:SAP

[1403] P. David. Semantic analysis for parallelizing C. Lecture Notes in Computer Science, 634:647–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yeh:1992:TSA

[1404] H.-J. Yeh. A task scheduling algorithm for the parallel expression evaluation in a reconfigurable fully digit on-line network. Lecture Notes in Computer Science, 634:653–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barnett:1992:UCN

[1405] L. M. Barnett and C. Lengauer. Unimodularity considered non-essential. Lecture Notes in Computer Science, 634:659–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gutzmann:1992:PSS

[1406] M. M. Gutzmann and K. Steffan.

PEPSIM-ST: A simulator tool for benchmarking. Lecture Notes in Computer Science, 634:665–??, 1992. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pozo:1992:PMS

[1407] R. Pozo. Performance modeling of sparse matrix methods for distributed memory architectures. Lecture Notes in Computer Science, 634:677–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schnekenburger:1992:PTA

[1408] T. Schnekenburger, M. Friedrich, A. Weininger, and T. Schoen. ParSim: A tool for the analysis of parallel and distributed programs. Lecture Notes in Computer Science, 634:689–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berthou:1992:EEP

[1409] J.-Y. Berthou and P. Klein. Estimating the effective performance of program parallelization on shared memory MIMD multiprocessors. Lecture Notes in Computer Science, 634:701– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duprat:1992:SPF

[1410] J. Duprat and M. Fiallos Aguilar. On the simulation of pipelining of fully digit on-line floating-point adder networks on massively parallel computers. Lecture Notes in Computer Science, 634:707–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tout:1992:PES

[1411] K. R. Tout and D. J. Evans. A parallel expert system using a backward chaining strategy. Lecture Notes in Computer Science, 634:713–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paugam-Moisy:1992:OSC

[1412] H. Paugam-Moisy. Optimal speedup conditions for a parallel back-propagation algorithm. Lecture Notes in Computer Science, 634:719–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bourzoufi:1992:TBB

[1413] H. Bourzoufi, G. Goncalves, and B. Toursel. A time-bounded binding access with low task-switching cost in an OR-Parallel Prolog. Lecture Notes in Computer Science, 634:725– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ungerer:1992:TSL

[1414] T. Ungerer and E. Zehendner. Threads and subinstruction level parallelism in a dataflow architecture. Lecture Notes in Computer Science, 634:731– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonchev:1992:HDA

[1415] B. Bonchev and M. Iliev. A hybrid dataflow architecture with multiple tokens. Lecture Notes in Computer Science, 634:737–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kusakabe:1992:PCS

[1416] S. Kusakabe, T. Hoshide, R. I.

Taniguchi, and M. Amamiya. Parallelism control scheme in a dataflow architecture. Lecture Notes in Computer Science, 634:743–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Unger:1992:PNB

[1417] H. Unger and K. Ben Achour. A Petri-net-based method to the design of parallel programs for a multiprocessor system. Lecture Notes in Computer Science, 634:749–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rabhi:1992:PSS

[1418] F. A. Rabhi. A parallel structure for static iterative transformation algorithms. Lecture Notes in Computer Science, 634:755–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyle:1992:CNM

[1419] J. M. Boyle, M. Clint, S. Fitzpatrick, and T. J. Harmer. The construction of numerical mathematical software for the AMT DAP by program transformation. Lecture Notes in Computer Science, 634:761–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gannon:1992:UOO

[1420] D. Gannon, J. K. Lee, and S. Narayana. On using object-oriented parallel programming to build distributed algebraic abstractions. Lecture Notes in Computer Science, 634:769–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demian:1992:ISO

[1421] V. Demian and J.-C. Mignot. Implementation of the self-organizing feature map on parallel computers. Lecture Notes in Computer Science, 634:775– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuchen:1992:CTD

[1422] H. Kuchen, H. Stoltze, and F. Lucking. Candela — A topology description language. Lecture Notes in Computer Science, 634:777–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shukhman:1992:PGS

[1423] B. Shukhman. Parallelism in generation of sequences for Monte Carlo simulation. Lecture Notes in Computer Science, 634:779–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pin:1992:DPE

[1424] C. Pin and E. V. Krishnamurthy. Data parallel evaluation-interpolation algorithm for solving functional matrix equations. Lecture Notes in Computer Science, 634:781–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohring:1992:DTE

[1425] S. Ohring. Dynamic tree embeddings into de Bruijn graphs. Lecture Notes in Computer Science, 634:783–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berrendorf:1992:MAS

[1426] R. Berrendorf. Memory access in shared virtual memory. Lecture Notes in Computer Science, 634:785– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weston:1992:PIJ

[1427] J. S. Weston, M. Clint, and C. W.

Bleakney. Parallel implementations of Jacobi’s algorithm for the eigensolution of large matrices using array processors. Lecture Notes in Computer Science, 634:787–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schutz:1992:PAV

[1428] B. Schutz and R. Rauscher. A parallel architecture for a VLSI-Hardware-Realization of a numerical stable variant of the simplex-method. Lecture Notes in Computer Science, 634:789– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1992:GPC

[1429] L. Zhang, M. Galindo, D. Marquie, and Y. Raynaud. Generating physical channel access protocol data unit of CCSDS recommendation by use of the Transputers and OCCAM. Lecture Notes in Computer Science, 634:791–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arbab:1992:MCI

[1430] F. Arbab, I. Herman, and P. Spilling. Manifold: Concepts and implementation. Lecture Notes in Computer Science, 634:793–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eglin-Leclerc:1992:HCS

[1431] M. C. Eglin-Leclerc, J. Julliand, and G. R. Perrin. How to compile systems of recurrence equations into networks of communicating processes. Lecture Notes in Computer Science, 634:795– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bahi:1992:PAS

[1432] M. Bahi. Parallel algorithms simulation applied to differential-algebraic systems and electrical network application. Lecture Notes in Computer Science, 634:797–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mahjoub:1992:OPS

[1433] Z. Mahjoub and F. Karoui-Sahtout. Optimal performances and scheduling for parallel algorithms with equal cost tasks. Lecture Notes in Computer Science, 634:799–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tan:1992:LBN

[1434] G. Tan and W.-N. Chin. Load balancing in a neighbourhood-based multiprocessor. Lecture Notes in Computer Science, 634:801–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crookes:1992:DBI

[1435] D. Crookes, A. P. McHale, and N. Beney. A DAP-Based implementation of a portable parallel image processing machine. Lecture Notes in Computer Science, 634:803–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gabarro:1992:DPA

[1436] J. Gabarro. The design of a parallel algorithm to solve the word problem for the free partially commutative groups. Lecture Notes in Computer Science, 634:805–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vajtersic:1992:VMP

[1437] M. Vajtersic. A VLSI multigrid Poisson solver amenable to biharmonic equation. Lecture Notes in Computer Science, 634:807–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kulkarni:1992:MIC

[1438] P. J. Kulkarni, N. R. Phadnis, and V. Udpikar. Monochrome image coding using hierarchical closed loop vector quantizer on a multiprocessing system. Lecture Notes in Computer Science, 634:809–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srinivas:1992:SFP

[1439] S. Srinivas and K. Dybvig. Superscalar floating-point vector computation in scheme. Lecture Notes in Computer Science, 634:811–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Datta:1992:EPA

[1440] A. K. Datta and R. K. Sen. An efficient parallel algorithm for maximal matching. Lecture Notes in Computer Science, 634:813–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adamo:1992:MAD

[1441] J. M. Adamo and N. Alhafez. Minimal, adaptive and deadlock-free routing for multiprocessors. Lecture Notes in Computer Science, 634:815–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desprez:1992:TSR

[1442] F. Desprez and B. Tourancheau. A theoretical study of reconfigurability for numerical algorithms on a reconfigurable network. Lecture Notes in Computer Science, 634:817–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fron:1992:MSU

[1443] A. Fron. Multiprocessor simulation using object-oriented programming. Lecture Notes in Computer Science, 634: 819–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stojcev:1992:RNC

[1444] M. K. Stojcev, I. Z. Milovanovic, E. I. Milovanovic, and G. S. Jovanovic. Real-number codes for faulttolerant matrix inversion on processor arrays. Lecture Notes in Computer Science, 634:821–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer:1992:AMP

[1445] J. W. Meyer. Automatic mapping of parallel programs onto processor networks. Lecture Notes in Computer Science, 634:823–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zaversek:1992:DAT

[1446] P. Zaversek and P. Kolbezen. Dynamic allocation on the Transputer network. Lecture Notes in Computer Science, 634:825–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arbenz:1992:BIS

[1447] P. Arbenz and M. Oettli. Block implementations of the symmetric QR and Jacobi algorithms. Lecture Notes in Computer Science, 634:827–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knirsch:1992:IPA

[1448] R. Knirsch and K. Hofmann. An improved parallel algorithm for the solution of molecular dynamics problems on MIMD multiprocessors. Lecture Notes in Computer Science, 634:829– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leite:1992:CMI

[1449] N. J. Leite and G. Bertrand. Computational models for image processing: Towards a concise parallel language. Lecture Notes in Computer Science, 634:831–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1992:BFC

[1450] J. Wu and E. B. Fernandez. Broadcasting in faulty cube-connected-cycles with minimum recovery time. Lecture Notes in Computer Science, 634:833– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Talbi:1992:PRM

[1451] E. G. Talbi, P. Bessiere, J. M. Ahuactzin, and E. Mazer. Parallel robot motion planning in a dynamic environment. Lecture Notes in Computer Science, 634:835–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mackenzie:1992:HBI

[1452] L. M. Mackenzie, M. Ould-Khaoua, R. J. Sutherland, and T. Kelly. A hypergraph-based interconnection network for large multicomputers. Lecture Notes in Computer Science, 634: 837–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grammatikakis:1992:GPR

[1453] M. D. Grammatikakis and J.-S. Jwo. Greedy permutation routing on caylay graphs. Lecture Notes in Computer Science, 634:839–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adamo:1992:EUC

[1454] J. M. Adamo and L. Trejo. Experimenting upon the CG algorithm within the CNET programming environment. Lecture Notes in Computer Science, 634:841–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barth:1992:ABM

[1455] D. Barth. An algorithm of broadcasting in the mesh of trees. Lecture Notes in Computer Science, 634:843–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1992:CEP

[1456] L. Zhou. Complexity estimation in the PIPS parallel programming environment. Lecture Notes in Computer Science, 634:845–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anisimov:1992:GPN

[1457] N. Anisimov, A. Kovalenko, P. Postupalski, and A. Simanchuk. A graphical Petri net based editor for a visualization of distributed and parallel systems. Lecture Notes in Computer Science, 634:847–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freisleben:1992:APD

[1458] B. Freisleben and T. Kielmann. Automatic parallelization of divide-and-conquer algorithms. Lecture Notes in Computer Science, 634:849–??, 1992.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conradi:1992:TRF

[1459] R. Conradi, C. Fernstrom, A. Fuggetta, and B. Snowdon. Towards a reference framework for process concepts. Lecture Notes in Computer Science, 635: 3–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Starke:1992:SCI

[1460] G. Starke and M. von der Beeck. SA/ CM/IM for process modelling. Lecture Notes in Computer Science, 635:18–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benali:1992:SPM

[1461] K. Benali and J. C. Derniame. Software process modelling: What, who and when. Lecture Notes in Computer Science, 635:21–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Devarenne:1992:NPE

[1462] A. Devarenne and C. Ozanne. The need of a process engineering method. Lecture Notes in Computer Science, 635: 26–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sommerville:1992:I

[1463] I. Sommerville. Introduction. Lecture Notes in Computer Science, 635:33–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lonchamp:1992:SSA

[1464] J. Lonchamp. Supporting social activities of software process. Lecture Notes in Computer Science, 635:34–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sommerville:1992:USP

[1465] I. Sommerville and T. Rodden. Understanding the software process as a social process. Lecture Notes in Computer Science, 635:55–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engels:1992:SCB

[1466] G. Engels and L. P. J. Groenewegen. Specification of coordinated behaviour in the software development process. Lecture Notes in Computer Science, 635:58–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rockwell:1992:SDC

[1467] R. Rockwell. Software development is a communication process. Lecture Notes in Computer Science, 635:61–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernstrom:1992:I

[1468] C. Fernstrom. Introduction. Lecture Notes in Computer Science, 635:65–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bandinelli:1992:PES

[1469] S. Bandinelli, A. Fuggetta, C. Ghezzi, and S. Grigolli. Process enactment in SPADE. Lecture Notes in Computer Science, 635:67–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambriola:1992:OAT

[1470] V. Ambriola and C. Montangero. OIKOS at the age of three. Lecture Notes in Computer Science, 635:84–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emmerich:1992:SDP

[1471] W. Emmerich, W. Schafer, and J. Welsh. Suitable database for processcentered environments do not yet exists. Lecture Notes in Computer Science, 635:94–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaccheri:1992:IRE

[1472] L. Jaccheri and S. Gai. Initial requirements for E3: An environment for experimenting and evolving software process. Lecture Notes in Computer Science, 635:99–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schafer:1992:I

[1473] W. Schafer. Introduction. Lecture Notes in Computer Science, 635:105– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sa:1992:IFS

[1474] J. Sa and B. Warboys. Integrating a formal specification method with PML. Lecture Notes in Computer Science, 635:106–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffmann:1992:MDS

[1475] C. Hoffmann, B. Kramer, and B. Dinler. Multiparadigm description of system development processes. Lecture Notes in Computer Science, 635:123– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greenwood:1992:UCS

[1476] R. M. Greenwood. Using CSP and system dynamics as process engineering tools. Lecture Notes in Computer Science, 635:138–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakagawa:1992:PVP

[1477] A. T. Nakagawa. Process versus product, abstraction and formalism: A personal perspective. Lecture Notes in Computer Science, 635:146–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belkhatir:1992:ORS

[1478] N. Belkhatir and W. L. Melo. The object role software process model. Lecture Notes in Computer Science, 635: 150–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bourdon:1992:ISC

[1479] M. Bourdon. Impact of the supported coordination levels on process modelling. Lecture Notes in Computer Science, 635:153–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Estublier:1992:RSP

[1480] J. Estublier. Reusing software processes. Lecture Notes in Computer Science, 635:156–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Favaro:1992:PME

[1481] J. Favaro. Process modelling at the European Space Agency. Lecture Notes in Computer Science, 635:159–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Warboys:1992:I

[1482] B. Warboys. Introduction. Lecture Notes in Computer Science, 635:165– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conradi:1992:DUI

[1483] R. Conradi, M. L. Jaccheri, C. Mazzi, M. N. Nguyen, and A. Aarsten. Design, use and implementation of SPELL, a language for software process modelling and evolution. Lecture Notes in Computer Science, 635:167–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Snowdon:1992:EPC

[1484] R. A. Snowdon. An example of process change. Lecture Notes in Computer Science, 635:178–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gruhn:1992:EFN

[1485] V. Gruhn and R. Jegelka. An evaluation of FUNSOFT nets. Lecture Notes in Computer Science, 635:196– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arbaoui:1992:PDM

[1486] S. Arbaoui, S. Mouline, F. Oquendo, and G. Tassart. PEACE: Describing and managing evolving knowledge in the software process. Lecture Notes in Computer Science, 635:215–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gruhn:1992:SPV

[1487] V. Gruhn and A. Saalmann. Software process validation based on FUN-SOFT nets. Lecture Notes in Computer Science, 635:223–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kawalek:1992:PMC

[1488] P. J. Kawalek. The process modelling cookbook: Orientation, description and experience. Lecture Notes in Computer Science, 635:227–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galle:1992:APM

[1489] J. Galle. Applying process modelling. Lecture Notes in Computer Science, 635:230–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taylor:1992:USM

[1490] I. C. Taylor. Using SPL to model ISO 9000. Lecture Notes in Computer Science, 635:237–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keller:1992:CPM

[1491] R. K. Keller and N. H. Madhavji. A comprehensive process model for discussing and recording scientific papers. Lecture Notes in Computer Science, 635:245–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kowalski:1992:TPL

[1492] Robert Kowalski. Theory and practice in logic programming. Lecture Notes in Computer Science, 636:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fruhwirth:1992:CLP

[1493] Thom Fruhwirth, Alexander Herold, Volker Kuchenhoff, Thierry Le Provost, Pierre Lim, Eric Monfroy, and Mark Wallace. Constraint logic programming. Lecture Notes in Computer Science, 636:3–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

dAndrea:1992:SOA

[1494] Michel d’Andrea. Scheduling and optimisation in the automobile industry. Lecture Notes in Computer Science, 636:36–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Evans:1992:FSU

[1495] Owen V. Evans. Factory scheduling using finite domains. Lecture Notes in Computer Science, 636:45–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gailly:1992:PPA

[1496] Pierre-Joseph Gailly, Sylvie Bescos, Wolfgang Krautter, and Christophe Bisiere. The prince project and its applications. Lecture Notes in Computer Science, 636:54–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsur:1992:GID

[1497] Shalom Tsur. A (gentle) introduction to deductive databases. Lecture Notes in Computer Science, 636:64–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beierle:1992:KBP

[1498] Christoph Beierle. Knowledge based PPS applications in PROTOS-L. Lecture Notes in Computer Science, 636: 80–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chiopris:1992:SBE

[1499] Carlo Chiopris. The SECReTS banking expert system from phase 1 to phase 2. Lecture Notes in Computer Science, 636:91–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fox:1992:LEC

[1500] John Fox. Logic engineering and clinical dilemmas. Lecture Notes in Computer Science, 636:100–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freeman:1992:KBA

[1501] Edward H. Freeman. A knowledgebased approach to strategic planning.

Lecture Notes in Computer Science, 636:109–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plumer:1992:ESM

[1502] Lutz Plumer. Expert systems in mining. Lecture Notes in Computer Science, 636:118–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hess:1992:NFL

[1503] Michael Hess. Natural and formal language processing. Lecture Notes in Computer Science, 636:127–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dahl:1992:PNL

[1504] Deborah A. Dahl. Pundit — natural language interfaces. Lecture Notes in Computer Science, 636:176–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grossi:1992:EDM

[1505] Thomas Grossi, Didier Bronisz, and Francois Jean-Marie. The ESTEAM-316 dialogue manager. Lecture Notes in Computer Science, 636:186–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kowalski:1992:LLP

[1506] Robert Kowalski. Legislation as logic programs. Lecture Notes in Computer Science, 636:203–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pletat:1992:KRN

[1507] Udo Pletat. Knowledge representation for natural language processing. Lecture Notes in Computer Science, 636: 231–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reintjes:1992:STV

[1508] Peter Reintjes. A set of tools for VHDL design. Lecture Notes in Computer Science, 636:242–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bundy:1992:TNR

[1509] Alan Bundy. Tutorial notes: Reasoning about logic programs. Lecture Notes in Computer Science, 636:252–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ed-Dbali:1992:SFS

[1510] Abdel Ali Ed-Dbali and Pierre Deransart. Software formal specification by logic programming: The example of standard Prolog. Lecture Notes in Computer Science, 636:278–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breuer:1992:ACP

[1511] Peter Breuer. The art of computer un-programming: Reverse engineering in Prolog. Lecture Notes in Computer Science, 636:290–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ducasse:1992:OAD

[1512] Mireille Ducasse. OPIUM — an advanced debugging system. Lecture Notes in Computer Science, 636:303– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baj:1992:ATP

[1513] Fabio Baj and Michael Rosner. Automatic theorem proving within the portable AI lab. Lecture Notes in Computer Science, 636:313–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilson:1992:UGC

[1514] Paul R. Wilson. Uniprocessor garbage collection techniques. Lecture Notes in Computer Science, 637:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdullahi:1992:CSD

[1515] S. E. Abdullahi, E. E. Miranda, and G. A. Ringwood. Collection schemes for distributed garbage. Lecture Notes in Computer Science, 637:43–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bekkers:1992:DMM

[1516] Y. Bekkers, O. Ridoux, and L. Ungaro. Dynamic memory management for sequential logic programming languages. Lecture Notes in Computer Science, 637:82–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juul:1992:CRG

[1517] N. C. Juul and E. Jul. Comprehensive and robust garbage collection in a distributed system. Lecture Notes in Computer Science, 637:103–??, 1992.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plainfosse:1992:EFT

[1518] D. Plainfosse and M. Shapiro. Experience with a fault-tolerant garbage collector in a distributed Lisp system. Lecture Notes in Computer Science, 637:116–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Venkatasubramanian:1992:SDG

[1519] N. Venkatasubramanian, G. Agha, and C. Talcott. Scalable distributed garbage collection for systems of active objects. Lecture Notes in Computer Science, 637:134–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puaut:1992:DGC

[1520] I. Puaut. Distributed garbage collection of active objects with no global synchronisation. Lecture Notes in Computer Science, 637:148–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Langendoen:1992:MMP

[1521] K. G. Langendoen, H. L. Muller, and W. G. Vree. Memory management for parallel tasks in shared memory. Lecture Notes in Computer Science, 637: 165–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeSergent:1992:IMT

[1522] T. Le Sergent and B. Berthomieu. Incremental multi-threaded garbage collection on virtually shared memory architectures. Lecture Notes in Computer Science, 637:179–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seward:1992:GGC

[1523] J. Seward. Generational garbage collection for lazy graph reduction. Lecture Notes in Computer Science, 637: 200–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chailloux:1992:CGC

[1524] E. Chailloux. A conservative garbage collector with ambiguous roots for static typechecking languages. Lecture Notes in Computer Science, 637:218– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mateu:1992:EIC

[1525] L. Mateu. An efficient implementation for coroutines. Lecture Notes in Computer Science, 637:230–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heck:1992:IAF

[1526] B. C. Heck and D. S. Wise. An implementation of an applicative file system. Lecture Notes in Computer Science, 637:248–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duvvuru:1992:CTM

[1527] S. Duvvuru, R. Sundararajan, E. Tick, A. V. S. Sastry, L. Hansen, and X. Zhong. A compile-time memoryreuse scheme for concurrent logic programs. Lecture Notes in Computer Science, 637:264–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hayes:1992:FCI

[1528] B. Hayes. Finalization in the collector interface. Lecture Notes in Computer Science, 637:277–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Edelson:1992:PCG

[1529] D. R. Edelson. Precompiling C++ for garbage collection. Lecture Notes in Computer Science, 637:299–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Samples:1992:GCC

[1530] A. D. Samples. Garbage collectioncooperative C++. Lecture Notes in Computer Science, 637:315–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pique:1992:DRC

[1531] J. F. Pique. Dynamic revision of choice points during garbage collection in Prolog [II/III]. Lecture Notes in Computer Science, 637:330–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tarau:1992:EMM

[1532] P. Tarau. Ecological memory management in a continuation passing Prolog engine. Lecture Notes in Computer Science, 637:344–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nettles:1992:RBI

[1533] S. Nettles, J. O’Toole, D. Pierce, and N. Haines. Replication-based incremental copying collection. Lecture Notes in Computer Science, 637:357– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolodner:1992:AIG

[1534] E. K. Kolodner and W. E. Weihl. Atomic incremental garbage collection. Lecture Notes in Computer Science, 637:365–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hudson:1992:ICM

[1535] R. L. Hudson and J. E. B. Moss. Incremental collection of mature objects. Lecture Notes in Computer Science, 637:388–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1992:OTD

[1536] M. S. Lam, P. R. Wilson, and T. G. Moher. Object type directed garbage collection to improve locality. Lecture Notes in Computer Science, 637:404– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delacour:1992:ARI

[1537] V. Delacour. Allocation regions and implementation contracts. Lecture Notes in Computer Science, 637:426– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rojemo:1992:CGG

[1538] N. Rojemo. A concurrent generational garbage collector for a parallel graph reducer. Lecture Notes in Computer Science, 637:440–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weemeeuw:1992:GCA

[1539] P. Weemeeuw and B. Demoen.

Garbage collection in aurora: An overview. Lecture Notes in Computer Science, 637:454–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merrall:1992:CGC

[1540] S. C. Merrall and J. A. Padget. Collections and garbage collection. Lecture Notes in Computer Science, 637: 473–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuasa:1992:MMG

[1541] T. Yuasa. Memory management and garbage collection of an extended Common Lisp system for massively parallel SIMD architecture. Lecture Notes in Computer Science, 637:490–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baker:1992:NFT

[1542] H. G. Baker. NREVERSAL of fortune — the thermodynamics of garbage collection. Lecture Notes in Computer Science, 637:507–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rocha:1992:NNT

[1543] A. F. Rocha. Neural nets: A theory for brains and machines. Lecture Notes in Computer Science, 638:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golledge:1992:DPU

[1544] Reginald G. Golledge. Do people understand spatial concepts: The case of first-order primitives. Lecture Notes in Computer Science, 639:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Snodgrass:1992:TD

[1545] Richard T. Snodgrass. Temporal databases. Lecture Notes in Computer Science, 639:22–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Couclelis:1992:PMO

[1546] Helen Couclelis. People manipulate objects (but cultivate fields): Beyond the raster-vector debate in GIS. Lecture Notes in Computer Science, 639:65–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hough:1992:TSE

[1547] Robbin R. Hough. Time and space: An economic model. Lecture Notes in Computer Science, 639:78–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takeuchi:1992:CLP

[1548] Keiichi Takeuchi. The changing language of and persisting patterns in the urban design of edo/tokyo. Lecture Notes in Computer Science, 639:97–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guttenberg:1992:TBT

[1549] Albert Z. Guttenberg. Toward a behavioral theory of regionalization. Lecture Notes in Computer Science, 639:

110–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Falcidieno:1992:DMP

[1550] Bianca Falcidieno, Caterina Pienovi, and Michela Spagnuolo. Descriptive modeling and prescriptive modeling in spatial data handling. Lecture Notes in Computer Science, 639:122–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montello:1992:GEK

[1551] Daniel R. Montello. The geometry of environmental knowledge. Lecture Notes in Computer Science, 639:136– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papadias:1992:SRU

[1552] Dimitris Papadias and Timos Sellis.

Spatial reasoning using symbolic arrays. Lecture Notes in Computer Science, 639:153–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freksa:1992:UOI

[1553] Christian Freksa. Using orientation information for qualitative spatial reasoning. Lecture Notes in Computer Science, 639:162–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jungert:1992:OPV

[1554] Erland Jungert. The observer’s point of view: An extension of symbolic projections. Lecture Notes in Computer Science, 639:179–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Egenhofer:1992:RAG

[1555] Max J. Egenhofer and Khaled K. Al-Taha. Reasoning about gradual changes of topological relationships. Lecture Notes in Computer Science, 639:196–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gold:1992:MN

[1556] Christopher M. Gold. The meaning of “neighbour”. Lecture Notes in Computer Science, 639:220–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeFloriani:1992:HTB

[1557] Leila De Floriani and Enrico Puppo. A hierarchical triangle-based model for terrain description. Lecture Notes in Computer Science, 639:236–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hadzilacos:1992:MET

[1558] Thanasis Hadzilacos and Nectaria Tryfona. A model for expressing topological integrity constraints in geographic databases. Lecture Notes in Computer Science, 639:252–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fotheringham:1992:ESI

[1559] A. Stewart Fotheringham and Andrew Curtis. Encoding spatial information: The evidence for hierarchical processing. Lecture Notes in Computer Science, 639:269–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freundschuh:1992:TRB

[1560] Scott M. Freundschuh. Is there a relationship between spatial cognition and environmental patterns? Lecture Notes in Computer Science, 639:288– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mark:1992:CIG

[1561] David M. Mark. Counter-intuitive geographic ‘facts’: Clues for spatial reasoning at geographic scales. Lecture Notes in Computer Science, 639:305– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Futch:1992:SLR

[1562] Shaun Futch, David N. Chin, Matthew McGranaghan, and Jinn-Guey Lay. Spatial-linguistic reasoning in LEI (locality and elevation interpreter). Lecture Notes in Computer Science, 639: 318–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lindholm:1992:UMI

[1563] Mikko Lindholm and Tapani Sarjakoski. User models and information theory in the design of a query interface for GIS. Lecture Notes in Computer Science, 639:328–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Timpf:1992:CMW

[1564] Sabine Timpf, Gary S. Volta, David W.

Pollock, and Max J. Egenhofer. A conceptual model of wayfinding using multiple levels of abstraction. Lecture Notes in Computer Science, 639:348– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hiraki:1992:TAS

[1565] Kazuo Hiraki and Yuichiro Anzai. Towards acquiring spatio-temporal knowledge from sensor data. Lecture Notes in Computer Science, 639:368– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbanente:1992:AAK

[1566] A. Barbanente, D. Borri, F. Esposito, P. Leo, G. Maciocco, and F. Selicato. Automatically acquiring knowledge by digital maps in artificial intelligence planning techniques. Lecture Notes in Computer Science, 639:379–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Whigham:1992:MIG

[1567] P. A. Whigham, R. I. McKay, and J. R. Davis. Machine induction of geospatial knowledge. Lecture Notes in Computer Science, 639:402–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paladino:1992:TQG

[1568] Ombretta Paladino. Treatment of qualitative geographic information in monitoring environmental pollution. Lecture Notes in Computer Science, 639: 418–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1992:KA

[1569] Anonymous. Keynote address. Lecture Notes in Computer Science, 640:1–2, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Humphrey:1992:TDS

[1570] Watts S. Humphrey. Toward a discipline for software engineering. Lecture Notes in Computer Science, 640:3–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hall:1992:SE

[1571] Anthony Hall. Is software engineering? Lecture Notes in Computer Science, 640:5–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gibbs:1992:SSE

[1572] Norm E. Gibbs. Session 1: Software engineering education in the 1990s. Lecture Notes in Computer Science, 640: 9–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grant:1992:SEE

[1573] Douglas D. Grant. Software engineering education in the 1990s– the way forward. Lecture Notes in Computer Science, 640:11–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garlan:1992:ECA

[1574] David Garlan, Mary Shaw, Chris Okasaki, Curtis M. Scott, and Roy F. Swonger. Experience with a course on architectures for software systems. Lecture Notes in Computer Science, 640:

23–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shepard:1992:TRD

[1575] Terry Shepard and Dan Hoffman. On teaching the rational design process. Lecture Notes in Computer Science, 640:44–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coulter:1992:SUP

[1576] Neal S. Coulter. Session 2: University perspective on industry-oriented courses and software reuse. Lecture Notes in Computer Science, 640:63–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruegge:1992:TIO

[1577] Bernd Bruegge. Teaching an industryoriented software engineering course. Lecture Notes in Computer Science, 640:65–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Offutt:1992:IRR

[1578] A. Jefferson Offutt and Roland H. Untch. Integrating research, reuse, and integration into software engineering courses. Lecture Notes in Computer Science, 640:88–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sindre:1992:SRE

[1579] Guttorm Sindre, Even-Andre Karlsson, and Tor Stalhane. Software reuse in an education perspective. Lecture Notes in Computer Science, 640:99–??, 1992.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carpenter:1992:SPD

[1580] Maribeth B. Carpenter. Session 3:

Panel discussion. Lecture Notes in Computer Science, 640:115–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lulofs:1992:IRA

[1581] Timothy J. Lulofs. Industry requirements for academic and continuing education programs in software engineering. Lecture Notes in Computer Science, 640:117–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pierce:1992:SFM

[1582] Keith R. Pierce. Session 4: Formal methods and engineering instruction. Lecture Notes in Computer Science, 640:119–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lutz:1992:FME

[1583] Michael J. Lutz. Formal methods and the engineering paradigm. Lecture Notes in Computer Science, 640:121– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garlan:1992:FMS

[1584] David Garlan. Formal methods for software engineers: Tradeoffs in curriculum design. Lecture Notes in Computer Science, 640:131–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lai:1992:TPE

[1585] Richard Lai. Teaching protocol engineering in honours year. Lecture Notes in Computer Science, 640:143– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mead:1992:SPD

[1586] Nancy R. Mead. Session 5: Panel discussion. Lecture Notes in Computer Science, 640:157–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lutz:1992:AWT

[1587] Peter H. Lutz. Across the wire: Teaching software engineering at a distance. Lecture Notes in Computer Science, 640:159–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ford:1992:SPD

[1588] Gary Ford. Session 6: Panel discussion. Lecture Notes in Computer Science, 640:161–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gotterbarn:1992:ILP

[1589] Donald Gotterbarn. Issues in licensing professional software engineers. Lecture Notes in Computer Science, 640: 163–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diaz-Herrera:1992:SEP

[1590] Jorge Diaz-Herrera. Session 7: Educating practitioners. Lecture Notes in Computer Science, 640:165–??, 1992.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luginbuhl:1992:TSV

[1591] David R. Luginbuhl and James E.

Cardow. Teaching software verification and validation to software practitioners. Lecture Notes in Computer Science, 640:167–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haikala:1992:CES

[1592] Ilkka J. Haikala and Jukka Marijarvi.

(continuing) education of software professionals. Lecture Notes in Computer Science, 640:180–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taulbee:1992:SEP

[1593] George M. Taulbee. The software engineering — patent law interface: A practitioner’s view. Lecture Notes in Computer Science, 640:194–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vonMayrhauser:1992:SIS

[1594] Anneliese von Mayrhauser. Session 8:

Integrating software engineering with computer science. Lecture Notes in Computer Science, 640:245–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Epley:1992:JML

[1595] Donn Epley. A joint master’s level software engineering subtrack. Lecture Notes in Computer Science, 640:247– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hooper:1992:PSE

[1596] James W. Hooper. Planning for software engineering education within a computer science framework at marshall university. Lecture Notes in Computer Science, 640:257–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tewari:1992:IOO

[1597] Raj Tewari and Frank Friedman. Integrating object-oriented software engineering in the computer science curriculum. Lecture Notes in Computer Science, 640:270–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pierce:1992:SPP

[1598] Keith R. Pierce. Session 9: Product and practice models for the engineering of software. Lecture Notes in Computer Science, 640:285–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DIppolito:1992:PES

[1599] Richard S. D’Ippolito, Kenneth J. Lee, Charles P. Plinta, Jeffrey A. Stewart, Paul D. Bailor, and Thomas C. Hartrum. Putting the engineering into software engineering. Lecture Notes in Computer Science, 640:287–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pesante:1992:SPD

[1600] Linda Hutz Pesante. Session 10: Panel discussion. Lecture Notes in Computer Science, 640:299–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1992:CET

[1601] A. Winsor Brown. Continuing education and training for software process improvement. Lecture Notes in Computer Science, 640:301–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Powel:1992:SPT

[1602] Karen L. Powel. Software process training: A formal and informal approach at McDonnell Douglas Electronic Systems Company. Lecture Notes in Computer Science, 640:308–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coulter:1992:SLP

[1603] Neal S. Coulter. Session 11: Laboratory and project experiences. Lecture Notes in Computer Science, 640:313– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lively:1992:USE

[1604] William M. Lively and Mark Lease.

Undergraduate software engineering laboratory at Texas A&M University. Lecture Notes in Computer Science, 640:315–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kantipudi:1992:SEC

[1605] Manmahesh Kantipudi, James S.

Collofello, Ken W. Collier, and Scott Medeiros. Software engineering course projects: Failures and recommendations. Lecture Notes in Computer Science, 640:324–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grau:1992:UIE

[1606] J. Kaye Grau and Norman Wilde. Use of the individual exchange project model in an undergraduate software engineering laboratory. Lecture Notes in Computer Science, 640:339–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carpenter:1992:SCE

[1607] Maribeth B. Carpenter. Session 12: Corporate experience. Lecture Notes in Computer Science, 640:349–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pavlik:1992:CSE

[1608] Rich Pavlik and Kathy Jacobs. Creating a software engineering training program in a level I organization. Lecture Notes in Computer Science, 640: 351–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coker:1992:CSE

[1609] Sandi Coker, Bud Glick, Lavon Green, and Anneliese von Mayrhauser. Corporate software engineering education for six sigma: Course development and assessment of success. Lecture Notes in Computer Science, 640:360–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Linger:1992:ICS

[1610] Richard C. Linger and R. Alan Spangler. The IBM Cleanroom Software Engineering Technology Transfer Program. Lecture Notes in Computer Science, 640:380–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christel:1992:EIV

[1611] Michael G. Christel. Experiences with an interactive video code inspection laboratory. Lecture Notes in Computer Science, 640:395–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ford:1992:SFP

[1612] Gary Ford. Session 13: Funding, practica, and principles. Lecture Notes in Computer Science, 640:413–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lidtke:1992:PCF

[1613] Doris K. Lidtke and Caroline E. Wardle. Panel: Current funding opportunities through the National Science Foundation. Lecture Notes in Computer Science, 640:415–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Junk:1992:ISE

[1614] William Junk and Paul Oman. The influence of software engineering paradigms on individual and team project results. Lecture Notes in Computer Science, 640:417–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berztiss:1992:EPS

[1615] Alfs T. Berztiss. Engineering principles and software engineering. Lecture Notes in Computer Science, 640:437– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grosch:1992:TAT

[1616] Josef Grosch. Transformation of attributed trees using pattern matching. Lecture Notes in Computer Science, 641:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galvez:1992:GLP

[1617] Jose Fortes Galvez. Generating LR(1) parsers of small size. Lecture Notes in Computer Science, 641:16–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Melichar:1992:SDT

[1618] Borivoj Melichar. Syntax directed translation with LR parsing. Lecture Notes in Computer Science, 641:30–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muller:1992:ADT

[1619] Karel Muller. Attribute-directed topdown parsing. Lecture Notes in Computer Science, 641:37–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baum:1992:AKM

[1620] Beate Baum. Another kind of modular attribute grammars. Lecture Notes in Computer Science, 641:44–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gyimothy:1992:IGE

[1621] Tibor Gyimothy, Zoltan Alexin, and Robert Szucs. Integrated graphic environment to develop applications based on attribute grammars. Lecture Notes in Computer Science, 641:51–??, 1992.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poetzsch-Heffter:1992:IHL

[1622] Arnd Poetzsch-Heffter. Implementing high-level identification specifications. Lecture Notes in Computer Science, 641:59–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abmann:1992:ASS

[1623] Werner Abmann. Another solution of scoping problems in symbol tables. Lecture Notes in Computer Science, 641:66–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Samples:1992:CIA

[1624] A. Dain Samples. Compiler implementation of ADTs using profile data. Lecture Notes in Computer Science, 641: 73–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Riedewald:1992:LLD

[1625] Gunter Riedewald. The LDL — language development laboratory. Lecture Notes in Computer Science, 641:88–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1992:AAS

[1626] Deryck F. Brown, Hermano Moura, and David A. Watt. ACTRESS: An action semantics directed compiler generator. Lecture Notes in Computer Science, 641:95–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber:1992:CFC

[1627] Christian Weber. Creation of a family of compilers and runtime environments by combining reusable components. Lecture Notes in Computer Science, 641:110–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knoop:1992:ICT

[1628] Jens Knoop and Bernhard Steffen. The interprocedural coincidence theorem. Lecture Notes in Computer Science, 641:125–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buth:1992:PCC

[1629] Bettina Buth, K.-H. Buth, M. Franzle, B. von Karger, Y. Lakhneche, K. Langmaack, and M. Muller-Olm. Provably correct compiler development and implementation. Lecture Notes in Computer Science, 641:141–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sudholt:1992:IDF

[1630] Mario Sudholt and Christoph Steigner. On interprocedural data flow analysis for object-oriented languages. Lecture Notes in Computer Science, 641:156– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emmelmann:1992:TCC

[1631] Helmut Emmelmann. Testing completeness of code selector specifications. Lecture Notes in Computer Science, 641:163–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hendren:1992:RAF

[1632] Laurie J. Hendren, Guang R. Gao, Erik R. Altman, and Chandrika Mukerji. A register allocation framework based on hierarchical cyclic interval graphs. Lecture Notes in Computer Science, 641:176–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duesterwald:1992:RPI

[1633] Evelyn Duesterwald, Rajiv Gupta, and Mary Lou Soffa. Register pipelining: An integrated approach to register allocation for scalar and subscripted variables. Lecture Notes in Computer Science, 641:192–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ertl:1992:ISC

[1634] M. Anton Ertl and Andreas Krall.

Instruction scheduling for complex pipelines. Lecture Notes in Computer Science, 641:207–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoogerbrugge:1992:CSP

[1635] Jan Hoogerbrugge and Henk Corporaal. Comparing software pipelining for an operation-triggered and a transporttriggered architecture. Lecture Notes in Computer Science, 641:219–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Griesemer:1992:SID

[1636] Robert Griesemer. Scheduling instructions by direct placement. Lecture Notes in Computer Science, 641:229– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vitek:1992:CTA

[1637] Jan Vitek, R. Nigel Horspool, and James S. Uhl. Compile-time analysis of object-oriented programs. Lecture Notes in Computer Science, 641:236– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersen:1992:PEC

[1638] Lars Ole Andersen. Partial evaluation of C and automatic compiler generation. Lecture Notes in Computer Science, 641:251–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pettersson:1992:TPM

[1639] Mikael Pettersson. A term patternmatch compiler inspired by finite automata theory. Lecture Notes in Computer Science, 641:258–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knopp:1992:IPP

[1640] Jurgen Knopp. Improving the performance of parallel LISP by compile time analysis. Lecture Notes in Computer Science, 641:271–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Langendoen:1992:FCG

[1641] Koen Langendoen and Pieter H. Hartel. FCG: A code generator for lazy functional languages. Lecture Notes in Computer Science, 641:278–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mantsivoda:1992:CFc

[1642] Andrei Mantsivoda and Vyacheslav Petukhin. Compiling Flang. Lecture Notes in Computer Science, 641:297– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Viklund:1992:IOH

[1643] Lars Viklund, Johan Herber, and Peter Fritzson. The implementation of ObjectMath — A high-level programming environment for scientific computing. Lecture Notes in Computer Science, 641:312–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Case:1992:RSK

[1644] John Case, Dayanand S. Rajan, and Anil M. Shende. Representing the spatial / kinematic domain and lattice computers. Lecture Notes in Computer Science, 642:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhixiang:1992:SCA

[1645] Chen Zhixiang and Wolfgang Maass.

A solution of the credit assignment problem in the case of learning rectangles. Lecture Notes in Computer Science, 642:26–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grefenstette:1992:LDS

[1646] John J. Grefenstette. Learning decision strategies with genetic algorithms. Lecture Notes in Computer Science, 642:35–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lavrac:1992:BKD

[1647] Nada Lavrac and Saso Dzeroski. Background knowledge and declarative bias in inductive concept learning. Lecture Notes in Computer Science, 642:51–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiehagen:1992:TMI

[1648] Rolf Wiehagen and Thomas Zeugmann. Too much information can be too much for learning efficiently. Lecture Notes in Computer Science, 642: 72–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ammon:1992:SEL

[1649] Kurt Ammon. Some experiments with a learning procedure. Lecture Notes in Computer Science, 642:87–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Apsitis:1992:UIC

[1650] Kalvis Apsitis, Rusins Freivalds, Martins Krikis, Raimonds Simanovskis, and Juris Smotrovs. Unions of identifiable classes of total recursive functions. Lecture Notes in Computer Science, 642:99–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baliga:1992:LMS

[1651] Ganesh Baliga, Sanjay Jain, and Arun Sharma. Learning from multiple sources of inaccurate data. Lecture Notes in Computer Science, 642:108– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Case:1992:SSL

[1652] John Case, Chen Keh-Jiann, and Sanjay Jain. Strong separation of learning classes. Lecture Notes in Computer Science, 642:129–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohen:1992:DGA

[1653] William W. Cohen. Desiderata for generalization-to-N algorithms. Lecture Notes in Computer Science, 642: 140–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daley:1992:PPP

[1654] Robert P. Daley, Bala Kalyanasundaram, and Mahendran Velauthapillai. The power of probabilism in Popperian FINite learning. Lecture Notes in Computer Science, 642:151–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flach:1992:AVF

[1655] Peter A. Flach. An analysis of various forms of ‘jumping to conclusions’. Lecture Notes in Computer Science, 642:170–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freivalds:1992:IIA

[1656] Rusins Freivalds and Achim G. Hoffmann. An inductive inference approach to classification. Lecture Notes in Computer Science, 642:187–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gasarch:1992:AQV

[1657] William I. Gasarch and Mahendran Velauthapillai. Asking questions versus verifiability. Lecture Notes in Computer Science, 642:197–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Indurkhya:1992:PAC

[1658] Bipin Indurkhya. Predictive analogy and cognition. Lecture Notes in Computer Science, 642:214–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kinber:1992:LCR

[1659] Efim Kinber. Learning a class of regular expressions via restricted subset queries. Lecture Notes in Computer Science, 642:232–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lange:1992:UAM

[1660] Steffen Lange and Thomas Zeugmann. A unifying approach to monotonic language learning on informant. Lecture Notes in Computer Science, 642:244– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mukouchi:1992:CFI

[1661] Yasuhito Mukouchi. Characterization of finite identification. Lecture Notes in Computer Science, 642:260–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OHara:1992:MRP

[1662] Scott O’Hara. A model of the redescription process in the context of geometric proportional analogy problems. Lecture Notes in Computer Science, 642:268–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Provost:1992:ISE

[1663] Foster John Provost and Bruce G.

Buchanan. Inductive strengthening: The effects of a simple heuristic for restricting hypothesis space search. Lecture Notes in Computer Science, 642: 294–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takada:1992:IDS

[1664] Yuji Takada and Rani Siromoney. On identifying DNA splicing systems from examples. Lecture Notes in Computer Science, 642:305–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:GI

[1665] A. (Annegret) Habel. General introduction. Lecture Notes in Computer Science, 643:1–4, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:HH

[1666] A. (Annegret) Habel. Hyperedges and hypergraphs. Lecture Notes in Computer Science, 643:5–12, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:RHH

[1667] A. (Annegret) Habel. Replacement of hyperedges by hypergraphs. Lecture Notes in Computer Science, 643:13–16, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:HRG

[1668] A. (Annegret) Habel. Hyperedgereplacement grammars. Lecture Notes in Computer Science, 643:17–23, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:MH

[1669] A. (Annegret) Habel. Modifications of HRG’s. Lecture Notes in Computer Science, 643:24–30, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:RC

[1670] A. (Annegret) Habel. Related concepts. Lecture Notes in Computer Science, 643:31–41, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:BN

[1671] A. (Annegret) Habel. Bibliographic note. Lecture Notes in Computer Science, 643:42, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:PD

[1672] A. (Annegret) Habel. Parallel derivations. Lecture Notes in Computer Science, 643:43–44, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:CFL

[1673] A. (Annegret) Habel. A contextfreeness lemma. Lecture Notes in Computer Science, 643:45–50, 1992.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:DT

[1674] A. (Annegret) Habel. Derivation trees. Lecture Notes in Computer Science, 643:51–52, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:BNb

[1675] A. (Annegret) Habel. Bibliographic note. Lecture Notes in Computer Science, 643:53–54, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:CP

[1676] A. (Annegret) Habel. Closure properties. Lecture Notes in Computer Science, 643:55–62, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:KTC

[1677] A. (Annegret) Habel. A Kleene-type characterization. Lecture Notes in Computer Science, 643:63–66, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:FPT

[1678] A. (Annegret) Habel. A fixed-point theorem. Lecture Notes in Computer Science, 643:67–68, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:BNc

[1679] A. (Annegret) Habel. Bibliographic note. Lecture Notes in Computer Science, 643:69–70, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:SH

[1680] A. (Annegret) Habel. Simplifications of HRG’s. Lecture Notes in Computer Science, 643:71–77, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:PLH

[1681] A. (Annegret) Habel. A pumping lemma for HRL’s. Lecture Notes in Computer Science, 643:78–86, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:APL

[1682] A. (Annegret) Habel. Applications of the pumping lemma. Lecture Notes in Computer Science, 643:87–93, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:PTH

[1683] A. (Annegret) Habel. Parikh’s theorem for HRL’s. Lecture Notes in Computer Science, 643:94–95, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:BNd

[1684] A. (Annegret) Habel. Bibliographic note. Lecture Notes in Computer Science, 643:96, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:GGG

[1685] A. (Annegret) Habel. Graphgenerating grammars of small order. Lecture Notes in Computer Science, 643:97–100, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:GGP

[1686] A. (Annegret) Habel. The graphgenerating power of HRG’s. Lecture Notes in Computer Science, 643:101– 106, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:SGG

[1687] A. (Annegret) Habel. String-graph-generating grammars of small order. Lecture Notes in Computer Science, 643:107–109, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:SGGb

[1688] A. (Annegret) Habel. The stringgraph-generating power of HRG’s. Lecture Notes in Computer Science, 643: 110–115, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:BNe

[1689] A. (Annegret) Habel. Bibliographic note. Lecture Notes in Computer Science, 643:116, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:SCG

[1690] A. (Annegret) Habel. Some compatible graph-theoretic properties. Lecture Notes in Computer Science, 643:117– 127, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:GVC

[1691] A. (Annegret) Habel. A general view of compatible predicates. Lecture Notes in Computer Science, 643:128–132, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:EAG

[1692] A. (Annegret) Habel. Efficient analysis of graph properties. Lecture Notes in Computer Science, 643:133–135, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:MGT

[1693] A. (Annegret) Habel. A metatheorem for graph-theoretic decision problems. Lecture Notes in Computer Science, 643:136–138, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:FTH

[1694] A. (Annegret) Habel. A filter theorem for HRL’s. Lecture Notes in Computer Science, 643:139–140, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:NCG

[1695] A. (Annegret) Habel. Non-compatible graph-theoretic properties. Lecture Notes in Computer Science, 643:141, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:UGT

[1696] A. (Annegret) Habel. Unsolvable graph-theoretic decision problems. Lecture Notes in Computer Science, 643: 142, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:RR

[1697] A. (Annegret) Habel. Related research. Lecture Notes in Computer Science, 643:143, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:BNf

[1698] A. (Annegret) Habel. Bibliographic note. Lecture Notes in Computer Science, 643:144, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:SCGb

[1699] A. (Annegret) Habel. Some compatible graph-theoretic functions. Lecture Notes in Computer Science, 643:145– 155, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:GVCb

[1700] A. (Annegret) Habel. A general view of compatible functions. Lecture Notes in Computer Science, 643:156–160, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:MBP

[1701] A. (Annegret) Habel. A metatheorem for boundedness problems. Lecture Notes in Computer Science, 643:161– 166, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:RRb

[1702] A. (Annegret) Habel. Related research. Lecture Notes in Computer Science, 643:167–166, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:UBP

[1703] A. (Annegret) Habel. Unsolvable boundedness problems. Lecture Notes in Computer Science, 643:167, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:BNg

[1704] A. (Annegret) Habel. Bibliographic note. Lecture Notes in Computer Science, 643:168, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:HRGb

[1705] A. (Annegret) Habel. Hypergraphreplacement grammars. Lecture Notes in Computer Science, 643:169–177, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:HRGc

[1706] A. (Annegret) Habel. Hypergraphreplacement grammars with application conditions. Lecture Notes in Computer Science, 643:178–181, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:PHR

[1707] A. (Annegret) Habel. Parallel hyperedge-replacement grammars. Lecture Notes in Computer Science, 643: 182, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:FGG

[1708] A. (Annegret) Habel. Figuregenerating grammars based on hyperedge replacement. Lecture Notes in Computer Science, 643:183–191, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:BNh

[1709] A. (Annegret) Habel. Bibliographic note. Lecture Notes in Computer Science, 643:192, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:C

[1710] A. (Annegret) Habel. Conclusion. Lecture Notes in Computer Science, 643: 193–194, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1992:B

[1711] A. (Annegret) Habel. Bibliography.

Lecture Notes in Computer Science, 643:195–206, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Szpankowski:1992:PAG

[1712] Wojciech Szpankowski. Probabilistic analysis of generalized suffix trees. Lecture Notes in Computer Science, 644: 1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Regnier:1992:LAS

[1713] Mireille Regnier. A language approach to string searching evaluation. Lecture Notes in Computer Science, 644:15–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atallah:1992:PMM

[1714] Mikhail J. Atallah, Philippe Jacquet, and Wojciech Szpankowski. Pattern matching with mismatches: A probabilistic analysis and a randomized algorithm. Lecture Notes in Computer Science, 644:27–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jong:1992:FMK

[1715] Yong Kim Jong and John Shawe-Taylor. Fast multiple keyword searching. Lecture Notes in Computer Science, 644:41–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacobson:1992:HIC

[1716] Guy Jacobson and Vo Kiem-Phong. Heaviest increasing/common subsequence problems. Lecture Notes in Computer Science, 644:52–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knight:1992:ARE

[1717] James R. Knight and Eugene W. Myers. Approximate regular expression pattern matching with concave gap penalties. Lecture Notes in Computer Science, 644:66–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pevzner:1992:MLC

[1718] Pavel A. Pevzner and Michael S. Waterman. Matrix longest common subsequence problem, duality and Hilbert bases. Lecture Notes in Computer Science, 644:77–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chia-Hsiang:1992:RED

[1719] Chang Chia-Hsiang and Robert Paige.

From regular expressions to DFA’s using compressed NFA’s. Lecture Notes in Computer Science, 644:88–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischetti:1992:IPO

[1720] Vincent A. Fischetti, Gad M. Landau, Jeanette P. Schmidt, and Peter H. Sellers. Identifying periodic occurrences of a template with applications to protein structure. Lecture Notes in Computer Science, 644:109–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sankoff:1992:EDG

[1721] David Sankoff. Edit distance for genome comparison based on non-local operations. Lecture Notes in Computer Science, 644:118–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischer:1992:DSM

[1722] Daniel Fischer, Ruth Nussinov, and Haim J. Wolfson. 3-D substructure matching in protein molecules. Lecture Notes in Computer Science, 644: 133–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaizhong:1992:FSP

[1723] Zhang Kaizhong, Dennis Shasha, and Jason T. L. Wang. Fast serial and parallel algorithms for approximate tree matching with VLDC’s. Lecture Notes in Computer Science, 644:148– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kilpelainen:1992:GTM

[1724] Pekka Kilpelainen and Heikki Mannila. Grammatical tree matching. Lecture Notes in Computer Science, 644:159– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1992:TEC

[1725] William I. Chang and Jordan Lampe. Theoretical and empirical comparisons of approximate string matching algorithms. Lecture Notes in Computer Science, 644:172–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeza-Yates:1992:FPA

[1726] Ricardo A. Baeza-Yates and Chris H. Perleberg. Fast and practical approximate string matching. Lecture Notes in Computer Science, 644:182–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Revuz:1992:DTC

[1727] Dominique Revuz and Marc Zipstein. DZ: A text compression algorithm for natural languages. Lecture Notes in Computer Science, 644:190–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pevzner:1992:MAG

[1728] Pavel A. Pevzner. Multiple alignment with guaranteed error bounds and communication cost. Lecture Notes in Computer Science, 644:202–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Irving:1992:TAL

[1729] Robert W. Irving and Campbell B.

Fraser. Two algorithms for the longest common subsequence of three (or more) strings. Lecture Notes in Computer Science, 644:211–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hui:1992:CSS

[1730] Lucas C. K. Hui. Color set size problem with applications to string matching. Lecture Notes in Computer Science, 644:227–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mehta:1992:CDC

[1731] Dinesh P. Mehta and Sartaj Sahni.

Computing display conflicts in string and circular string visualization. Lecture Notes in Computer Science, 644: 241–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amir:1992:ERD

[1732] Amihood Amir, Martin Farach, and Yossi Matias. Efficient randomized dictionary matching algorithms. Lecture Notes in Computer Science, 644:259– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Idury:1992:DDM

[1733] Ramana M. Idury and Alejandro A.

Schaffer. Dynamic dictionary matching with failure functions. Lecture Notes in Computer Science, 644:273–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1992:VO

[1734] P. P. Chen. ER vs. OO. Lecture Notes in Computer Science, 645:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Navathe:1992:NTY

[1735] S. B. Navathe. The next ten years of modeling, methodologies, and tools. Lecture Notes in Computer Science, 645:3–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thalheim:1992:FCC

[1736] B. Thalheim. Fundamentals of cardinality constraints. Lecture Notes in Computer Science, 645:7–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1992:ECC

[1737] J. Zhou and P. Baumann. Evaluation of complex cardinality constraints. Lecture Notes in Computer Science, 645:24–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kappel:1992:LRI

[1738] G. Kappel and M. Schrefl. Local referential integrity. Lecture Notes in Computer Science, 645:41–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauh:1992:ETC

[1739] O. Rauh and E. Stickel. Entity tree clustering — A method for simplifying ER designs. Lecture Notes in Computer Science, 645:62–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hainaut:1992:TSM

[1740] J. L. Hainaut. A temporal statistical model for entity-relationship schemas. Lecture Notes in Computer Science, 645:79–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Song:1992:SSR

[1741] W. W. Song, P. Johannesson, and J. A.

Bubenko. Semantic similarity relations in schema integration. Lecture Notes in Computer Science, 645:97–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castano:1992:CRC

[1742] S. Castano, V. De Antonellis, and B. Zonta. Classifying and reusing conceptual schemas. Lecture Notes in Computer Science, 645:121–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scheer:1992:EDM

[1743] A.-W. Scheer. Embedding data modelling in a general architecture for integrated information systems. Lecture Notes in Computer Science, 645:139– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buitelaar:1992:ULI

[1744] P. Buitelaar and R. van de Riet. The use of a lexicon to interpret ER diagrams: A LIKE project. Lecture Notes in Computer Science, 645:162– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garcia:1992:ERB

[1745] H.-M. Chen Garcia and O. Liu R. Sheng. An entity-relationship-based methodology for distributed database design: An integrated approach towards combined logical and distribution designs. Lecture Notes in Computer Science, 645:178–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gandhi:1992:SBD

[1746] M. Gandhi and E. L. Robertson. A specification-based data model. Lecture Notes in Computer Science, 645: 194–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pirri:1992:DDD

[1747] F. Pirri and C. Pizzuti. Data dictionary design: A logic programming approach. Lecture Notes in Computer Science, 645:210–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miura:1992:NQV

[1748] T. Miura. Nesting quantification in a visual data manipulation language. Lecture Notes in Computer Science, 645:226–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kehrer:1992:EPE

[1749] N. Kehrer and G. Neumann. An EER prototyping environment and its implementation in a datalog language. Lecture Notes in Computer Science, 645: 243–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ling:1992:TER

[1750] T.-W. Ling and M.-L. Lee. A theory for entity-relationship view updates. Lecture Notes in Computer Science, 645:262–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herzig:1992:TCD

[1751] R. Herzig and M. Gogolla. Transforming conceptual data models into an object model. Lecture Notes in Computer Science, 645:280–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tresch:1992:MOM

[1752] M. Tresch and M. H. Scholl. Meta object management and its application to database evolution. Lecture Notes in Computer Science, 645:299–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breiteneder:1992:MAV

[1753] C. Breiteneder, S. Gibbs, and D. Tsichritzis. Modelling of audio/ video data. Lecture Notes in Computer Science, 645:322–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Auddino:1992:SVI

[1754] A. Auddino, Y. Dennebouy, Y. Dupont, E. Fontana, S. Spaccapietra, and Z. Tari. SUPER-Visual interaction with an object-based ER model. Lecture Notes in Computer Science, 645: 340–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bono:1992:NLR

[1755] G. Bono and P. Ficorilli. Natural language restatement of queries expressed in a graphical language. Lecture Notes in Computer Science, 645: 357–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elmasri:1992:TQL

[1756] R. Elmasri and V. Kouramajian. A temporal query language based on conceptual entities and roles. Lecture Notes in Computer Science, 645:375– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tari:1992:DOO

[1757] Z. Tari. On the design of objectoriented databases. Lecture Notes in Computer Science, 645:389–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halper:1992:PRO

[1758] M. Halper, J. Geller, and Y. Perl. ’part’ relations for object-oriented databases. Lecture Notes in Computer Science, 645:406–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bratsberg:1992:UCE

[1759] S. E. Bratsberg. Unified class evolution by object-oriented views. Lecture Notes in Computer Science, 645:423– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katona:1992:CAR

[1760] Gyula O. H. Katona. Combinatorial and algebraic results for database relations. Lecture Notes in Computer Science, 646:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1992:SDD

[1761] Jose Alberto Fernandez and Jack Minker. Semantics of disjunctive deductive databases. Lecture Notes in Computer Science, 646:21–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsalenko:1992:DTR

[1762] Michail Sh. Tsalenko. Database theory in Russia (1979–1991). Lecture Notes in Computer Science, 646:51–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demetrovics:1992:CDA

[1763] Janos Demetrovics, L. Ronyai, and nam Son Hua. On the composition and decomposition of attributes and tuples. Lecture Notes in Computer Science, 646:71–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kivinen:1992:ADI

[1764] Jyrki Kivinen and Heikki Mannila. Approximate dependency inference from relations. Lecture Notes in Computer Science, 646:86–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tuijn:1992:VDD

[1765] Chris Tuijn and Marc Gyssens. Views and decompositions of databases from a categorical perspective. Lecture Notes in Computer Science, 646:99–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abiteboul:1992:CIL

[1766] Serge Abiteboul, Moshe Vardi, and Victor Vianu. Computing with infinitary logic. Lecture Notes in Computer Science, 646:113–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grumbach:1992:QLC

[1767] Stephane Grumbach and Christophe Tollu. Query languages with counters. Lecture Notes in Computer Science, 646:124–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breazu-Tannen:1992:NEQ

[1768] Val Breazu-Tannen, Peter Buneman, and Limsoon Wong. Naturally embedded query languages. Lecture Notes in Computer Science, 646:140–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fabret:1992:OIC

[1769] Françoise Fabret, Mireille Regnier, and Eric Simon. Optimizing incremental computation of datalog programs with non-deterministic semantics. Lecture Notes in Computer Science, 646:155– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abiteboul:1992:OAD

[1770] Serge Abiteboul and Allen Van Gelder.

Optimizing active databases using the split technique. Lecture Notes in Computer Science, 646:171–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Finger:1992:UAI

[1771] Marcelo Finger and Dov M. Gabbay. Updating atomic information in labelled database systems. Lecture Notes in Computer Science, 646:188– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Man:1992:CSS

[1772] Han Wong Man and D. Agrawal.

Context-specific synchronization for atomic data types. Lecture Notes in Computer Science, 646:201–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fekete:1992:HAN

[1773] Alan Fekete, Nancy Lynch, and William E. Weihl. Hybrid atomicity for nested transactions. Lecture Notes in Computer Science, 646:216–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuo:1992:MVD

[1774] Dean Kuo. Model and verification of a data manager based on ARIES. Lecture Notes in Computer Science, 646: 231–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barback:1992:MII

[1775] Monica D. Barback, Jorge Lobo, and James J. Lu. Minimizing indefinite information in disjunctive deductive databases. Lecture Notes in Computer Science, 646:246–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gire:1992:WFS

[1776] Françoise Gire. Well-founded semantics and stable semantics of semi-strict programs. Lecture Notes in Computer Science, 646:261–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papadimitriou:1992:FED

[1777] Christos H. Papadimitriou and Martha Sideri. On finding extensions of default theories. Lecture Notes in Computer Science, 646:276–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dong:1992:IED

[1778] Guozhu Dong and Rodney Topor. Incremental evaluation of datalog queries. Lecture Notes in Computer Science, 646:282–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feder:1992:DUE

[1779] Tomas Feder and Yatin Saraiya. Decidability and undecidability of equivalence for linear datalog with applications to normal-form optimizations. Lecture Notes in Computer Science, 646:297–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grumbach:1992:PDT

[1780] Stephane Grumbach. A paradox in database theory. Lecture Notes in Computer Science, 646:312–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VandenBussche:1992:HFS

[1781] Jan Van den Bussche and Dirk Van Gucht. A hierarchy of faithful set creation in pure OODBs. Lecture Notes in Computer Science, 646:326–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schewe:1992:IGC

[1782] Klaus-Dieter Schewe, Joachim W.

Schmidt, and Ingrid Wetzel. Identification, genericity and consistency in object-oriented databases. Lecture Notes in Computer Science, 646:341– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beneventano:1992:SCO

[1783] Domenico Beneventano and Sonia Bergamaschi. Subsumption for complex object data models. Lecture Notes in Computer Science, 646:357– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duzi:1992:SIC

[1784] Marie Duzi. Semantic information connected with data. Lecture Notes in Computer Science, 646:376–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonatti:1992:DFS

[1785] Piero Bonatti, Sarit Kraus, and V. S. Subrahmanian. Declarative foundations of secure deductive databases. Lecture Notes in Computer Science, 646:391–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jakobsson:1992:MVL

[1786] Hakan Jakobsson. On materializing views and on-line queries. Lecture Notes in Computer Science, 646:407– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinchuk:1992:MDS

[1787] Anatoly P. Pinchuk and Konstantin V. Shvachko. Maintaining dictionaries: Space-saving modifications of B- trees. Lecture Notes in Computer Science, 646:421–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arni:1992:STM

[1788] Natraj Arni, Sergio Greco, and Domenico Sacca. Set-term matching in logic programming. Lecture Notes in Computer Science, 646:436–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bull:1992:TSO

[1789] John A. Bull, Gong Li, and Karen R. Sollins. Towards security in an open systems federation. Lecture Notes in Computer Science, 648:3–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelter:1992:TLA

[1790] Udo Kelter. Type-level access controls for distributed structurally objectoriented database systems. Lecture Notes in Computer Science, 648:21–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kessler:1992:CWM

[1791] Volker Kessler. On the Chinese wall model. Lecture Notes in Computer Science, 648:41–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jingsha:1992:FMA

[1792] He Jingsha and Virgil D. Gligor. Formal methods and automated tool for timing-channel identification in TCB source code. Lecture Notes in Computer Science, 648:57–??, 1992. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toussaint:1992:SSI

[1793] Marie-Jeanne Toussaint. Separating the specification and implementation phases in cryptology. Lecture Notes in Computer Science, 648:77–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1992:FSS

[1794] Andrew J. I. Jones and Marek Sergot. Formal specification of security requirements using the theory of normative positions. Lecture Notes in Computer Science, 648:103–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desmedt:1992:BTC

[1795] Yvo Desmedt. Breaking the traditional computer security barriers. Lecture Notes in Computer Science, 648:125– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hauser:1992:VMA

[1796] Ralf C. Hauser and E. Stewart Lee. Verification and modelling of authentication protocols. Lecture Notes in Computer Science, 648:141–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Molva:1992:KAK

[1797] Refik Molva, Gene Tsudik, Els Van Herreweghen, and Stefano Zatti. KryptoKnight authentication and key distribution system. Lecture Notes in Computer Science, 648:155–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tarah:1992:AMC

[1798] Anas Tarah and Christian Huitema. Associating metrics to certification paths. Lecture Notes in Computer Science, 648:175–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fabre:1992:OOV

[1799] Jean-Charles Fabre and Brian Randell. An object-oriented view of fragmented data processing for fault and intrusion tolerance in distributed systems. Lecture Notes in Computer Science, 648:193–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1992:DTI

[1800] Michael Wang and Andrzej Goscinski. The development and testing of the identity-based conference key distribution system for the RHODOS distributed system. Lecture Notes in Computer Science, 648:209–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsudik:1992:PES

[1801] Gene Tsudik. Policy enforcement in stub autonomous domains. Lecture Notes in Computer Science, 648:229– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwok-Yan:1992:FAA

[1802] Lam Kwok-Yan and Dieter Gollmann.

Freshness assurance of authentication protocols. Lecture Notes in Computer Science, 648:261–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyd:1992:FFA

[1803] Colin Boyd. A formal framework for authentication. Lecture Notes in Computer Science, 648:273–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwok-Yan:1992:TAD

[1804] Lam Kwok-Yan and Thomas Beth.

Timely authentication in distributed systems. Lecture Notes in Computer Science, 648:293–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sandhu:1992:PCS

[1805] Ravi S. Sandhu and Sushil Jajodia. Polyinstantiation for cover stories. Lecture Notes in Computer Science, 648: 307–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kang:1992:TPM

[1806] Iwen E. Kang and Thomas F. Keefe. On transaction processing for multilevel secure replicated databases. Lecture Notes in Computer Science, 648: 329–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pernul:1992:SCP

[1807] Gunther Pernul. Security constraint processing in multilevel secure AMAC schemata. Lecture Notes in Computer Science, 648:349–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

dAusbourg:1992:MMM

[1808] Bruno d’Ausbourg and Jean-Henri Llareus. Msup 2 S: A machine for multilevel security. Lecture Notes in Computer Science, 648:373–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calas:1992:GMD

[1809] Christel Calas. GDoM: a multilevel document manager. Lecture Notes in Computer Science, 648:393–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1992:USG

[1810] Ross J. Anderson. UEPS — A second generation electronic wallet. Lecture Notes in Computer Science, 648: 411–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daemen:1992:HDM

[1811] Joan Daemen, Rene Govaerts, and Joos Vandewalle. A hardware design model for cryptographic algorithms. Lecture Notes in Computer Science, 648:419–434, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habra:1992:ASA

[1812] Naji Habra, Baudouin Le Charlier, Abdelaziz Mounji, and Isabelle Mathieu. ASAX: Software architecture and rulebased language for universal audit trail analysis. Lecture Notes in Computer Science, 648:435–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gardenfors:1992:BRV

[1813] P. Gardenfors. Belief revision: A vademecum. Lecture Notes in Computer Science, 649:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DelCerro:1992:MPT

[1814] L. Farinas Del Cerro and A. Herzig.

Meta-programming through intensional deduction: Some examples. Lecture Notes in Computer Science, 649: 11–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konolige:1992:AAM

[1815] K. Konolige. An autoepistemic analysis of metalevel reasoning in logic programming. Lecture Notes in Computer Science, 649:26–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Komorowski:1992:IPD

[1816] J. Komorowski. An introduction to partial deduction. Lecture Notes in Computer Science, 649:49–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeSchreye:1992:TTL

[1817] D. De Schreye and K. Verschaetse. Tutorial on termination of logic programs. Lecture Notes in Computer Science, 649:70–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanHarmelen:1992:DNR

[1818] F. van Harmelen. Definable naming relations in metalevel systems. Lecture Notes in Computer Science, 649:89–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brogi:1992:MML

[1819] A. Brogi, P. Mancarella, D. Pedreschi, and F. Turini. Meta for modularizing logic programming. Lecture Notes in Computer Science, 649:105–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Debray:1992:COL

[1820] S. K. Debray. Compiler optimizations for lowlevel redundancy elimination: An application of metalevel Prolog primitives. Lecture Notes in Computer Science, 649:120–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Costantini:1992:RAM

[1821] S. Costantini, P. Dell’Acqua, and G. Lanzarone. Reflective agents in metalogic programming. Lecture Notes in Computer Science, 649:135–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cervesato:1992:LMP

[1822] I. Cervesato and G. F. Rossi. Logic meta-programming facilities in log. Lecture Notes in Computer Science, 649:148–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bahgat:1992:PDH

[1823] R. Bahgat. The Pandora deadlock handler metalevel relation. Lecture Notes in Computer Science, 649:162– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benkerimi:1992:OOP

[1824] K. Benkerimi and P. M. Hill. Objectoriented programming in godel: An experiment. Lecture Notes in Computer Science, 649:177–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeSchreye:1992:SLH

[1825] D. De Schreye and B. Martens. A sensible least Herbrand semantics for untyped vanilla meta-programming and its extension to a limited form of amalgamation. Lecture Notes in Computer Science, 649:192–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christiansen:1992:CRM

[1826] H. Christiansen. A complete resolution method for logical meta-programming languages. Lecture Notes in Computer Science, 649:205–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonatti:1992:MTS

[1827] P. Bonatti. Model theoretic semantics for demo. Lecture Notes in Computer Science, 649:220–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giunchiglia:1992:HML

[1828] F. Giunchiglia, L. Serafini, and A. Simpson. Hierarchical meta-logics: Intuitions, proof-theory, and semantics. Lecture Notes in Computer Science, 649:235–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiggins:1992:NCA

[1829] G. A. Wiggins. Negation and control in automatically generated logic programs. Lecture Notes in Computer Science, 649:250–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bossi:1992:TNP

[1830] A. Bossi, N. Cocco, and S. Etalle. Transforming normal programs by replacement. Lecture Notes in Computer Science, 649:265–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Traff:1992:MPR

[1831] J. L. Traff and S. D. Prestwich. Metaprogramming for reordering literals in deductive databases. Lecture Notes in Computer Science, 649:280–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruynooghe:1992:PNO

[1832] M. Bruynooghe and G. Janssens. Propagation: A new operation in a framework for abstract interpretation of logic programs. Lecture Notes in Computer Science, 649:294–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mesnard:1992:CQP

[1833] F. Mesnard and J.-G. Ganascia.

CLP(Q) for proving interargument relations. Lecture Notes in Computer Science, 649:308–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamfelt:1992:RFM

[1834] A. Hamfelt and A. Hansson. Representation of fragmentary multilayered knowledge. Lecture Notes in Computer Science, 649:321–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grabowski:1992:MCA

[1835] J. Grabowski. Metaprograms for change, assumptions, objects, and inheritance. Lecture Notes in Computer Science, 649:336–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vishkin:1992:MPAb

[1836] U. Vishkin. Methods in parallel algorithmics and who may need to know them? Lecture Notes in Computer Science, 650:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1992:RPA

[1837] D. T. Lee. Rectilinear paths among rectilinear obstacles. Lecture Notes in Computer Science, 650:5–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1992:LTA

[1838] M. H. Chen and S. L. Lee. Linear time algorithms for k-cutwidth problem. Lecture Notes in Computer Science, 650:21–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watanabe:1992:ECA

[1839] T. Watanabe, T. Mashita, and S. Taoka. The k-edge-connectivity augmentation problem of weighted graphs. Lecture Notes in Computer Science, 650:31–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patkar:1992:PLP

[1840] S. Patkar and H. Narayanan. Principal lattice of partitions of submodular functions on graphs: Fast algorithms for principal partition and generic rigidity. Lecture Notes in Computer Science, 650:41–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hwang:1992:ASS

[1841] R. Z. Hwang and R. C. T. Lee. The application of the searching over separators strategy to solve some NP-Complete problems on planar graphs. Lecture Notes in Computer Science, 650:51–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halldorsson:1992:PLG

[1842] M. Halldorsson. Parallel and on-line graph coloring algorithms. Lecture Notes in Computer Science, 650:61–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsumoto:1992:CAR

[1843] T. Matsumoto. Competitive analysis of the round robin algorithm. Lecture Notes in Computer Science, 650:71–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chien:1992:CAL

[1844] B. C. Chien, R. J. Chen, and W. P.

Yang. Competitive analysis of the on-line algorithms for multiple stacks systems. Lecture Notes in Computer Science, 650:78–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lai:1992:SAA

[1845] T. W. Lai. Self-adjusting augmented search trees. Lecture Notes in Computer Science, 650:88–??, 1992. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gonzalez:1992:ACM

[1846] T. F. Gonzalez and T. Murayama. Algorithms for a class of min-cut and max-cut problem. Lecture Notes in Computer Science, 650:97–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ho:1992:ARO

[1847] J. M. Ho, M. T. Ko, T. H. Ma, and T. Y. Sung. Algorithms for rectilinear optimal multicast tree problem. Lecture Notes in Computer Science, 650:106–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kloks:1992:ATP

[1848] T. Kloks and H. Bodlaender. Approximating treewidth and pathwidth of some classes of perfect graphs. Lecture Notes in Computer Science, 650: 116–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ueno:1992:GSC

[1849] S. Ueno, M. Yamazaki, and Y. Kajitani. Graph spanners and connectivity. Lecture Notes in Computer Science, 650:126–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berkman:1992:RRM

[1850] O. Berkman, Y. Matias, and U. Vishkin. Randomized range-maxima in nearly-constant parallel time. Lecture Notes in Computer Science, 650:135– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Han:1992:FTB

[1851] Y. Han, Y. Igarashi, K. Kanai, and K. Miura. Fault-tolerant broadcasting in binary jumping networks. Lecture Notes in Computer Science, 650:145– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iwama:1992:RPM

[1852] K. Iwama, E. Miyano, and Y. Kambayashi. Routing problems on the mesh of buses. Lecture Notes in Computer Science, 650:155–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jimbo:1992:SNL

[1853] S. Jimbo and A. Maruoka. Selection networks with 8n logsub 2 n size and O(log n) depth. Lecture Notes in Computer Science, 650:165–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Book:1992:ROP

[1854] R. Book. Relativizations of the P = ?NP and other problems: Some developments in structural complexity theory. Lecture Notes in Computer Science, 650:175–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paterson:1992:BCC

[1855] M. Paterson. Boolean circuit complexity (abstract). Lecture Notes in Computer Science, 650:187–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Symvonis:1992:SSP

[1856] A. Symvonis and S. Tragoudas. Searching a solid pseudo 3-Sided orthoconvex grid. Lecture Notes in Computer Science, 650:188–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dendris:1992:EPA

[1857] N. D. Dendris, I. A. Kalafatis, and L. M. Kirousis. An efficient parallel algorithm for geometrically characterizing drawings of a class of 3-D objects. Lecture Notes in Computer Science, 650:198–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sugihara:1992:TCA

[1858] K. Sugihara. Topologically consistent algorithms related to convex polyhedra. Lecture Notes in Computer Science, 650:209–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choi:1992:CRV

[1859] S. H. Choi, S. Y. Shin, and K. Y. Chwa. Characterizing and recognizing visibility graphs of funnel-shaped polygons. Lecture Notes in Computer Science, 650:219–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Itoh:1992:CCN

[1860] T. Itoh and K. Horikawa. On the complexity of composite numbers. Lecture Notes in Computer Science, 650:229– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kobayashi:1992:MID

[1861] K. Kobayashi. On malign input distributions for algorithms. Lecture Notes in Computer Science, 650:239– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arvind:1992:LCS

[1862] V. Arvind, J. Kobler, and M. Mundhenk. Lowness and the complexity of sparse and tally descriptions. Lecture Notes in Computer Science, 650:249– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1992:HIS

[1863] J. Wang and J. Belanger. Honest iteration schemes of randomizing algorithms. Lecture Notes in Computer Science, 650:259–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1992:AVC

[1864] H. S. Lee and R. C. Chang. Approximating vertices of a convex polygon with grid points in the polygon. Lecture Notes in Computer Science, 650: 269–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akutsu:1992:ADG

[1865] T. Akutsu. Algorithms for determining the geometrical congruity in two and three dimensions. Lecture Notes in Computer Science, 650:279–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jennings:1992:RAC

[1866] E. Jennings and A. Lingas. On the relationships among constrained geometric structures. Lecture Notes in Computer Science, 650:289–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Madlener:1992:GSC

[1867] K. Madlener, F. Otto, and A. Sattler-Klein. Generating small convergent systems can be extremely hard. Lecture Notes in Computer Science, 650: 299–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ogawa:1992:CTR

[1868] M. Ogawa. Chew’s theorem revisited — uniquely normalizing property of nonlinear term rewriting systems. Lecture Notes in Computer Science, 650:309–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bolignano:1992:HOC

[1869] D. Bolignano and M. Debabi. Higher order communicating processes with value-passing, assignment and return of results. Lecture Notes in Computer Science, 650:319–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pijls:1992:SIG

[1870] W. Pijls and A. de Bruin. Searching informed game trees. Lecture Notes in Computer Science, 650:332–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iri:1992:HGR

[1871] M. Iri. How to generate realistic sample problems for network optimization. Lecture Notes in Computer Science, 650:342–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martello:1992:GAP

[1872] S. Martello and P. Toth. Generalized assignment problems. Lecture Notes in Computer Science, 650:351–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guevremont:1992:REL

[1873] E. Guevremont and J. Snoeyink. Recognizing an envelope of lines in linear time. Lecture Notes in Computer Science, 650:370–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1992:APC

[1874] W. S. Chan and F. Chin. Approximation of polygonal curves with minimum number of line segments. Lecture Notes in Computer Science, 650: 378–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sarrafzadeh:1992:WKK

[1875] M. Sarrafzadeh, D. Wagner, F. Wagner, and K. Weihe. Wiring knock-knee layouts: A global approach. Lecture Notes in Computer Science, 650:388– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takahashi:1992:AFN

[1876] J. Takahashi, H. Suzuki, and T. Nishizeki. Algorithms for finding noncrossing paths with minimum total length in plane graphs. Lecture Notes in Computer Science, 650:400– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Longpre:1992:SIP

[1877] L. Longpre and O. Watanabe. On symmetry of information and polynomial time invertibility. Lecture Notes in Computer Science, 650:410–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beigel:1992:PAC

[1878] R. Beigel, J. Tarui, and S. Toda. On probabilistic ACC circuits with an exact-threshold output gate. Lecture Notes in Computer Science, 650:420– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurosawa:1992:CSI

[1879] K. Kurosawa and O. Watanabe. Computational and statistical indistinguishabilities. Lecture Notes in Computer Science, 650:430–??, 1992. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fu:1992:SDN

[1880] B. Fu and H. Z. Li. On symmetric differences of NP-Hard sets with weakly-P- selective sets. Lecture Notes in Computer Science, 650:439–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sarrafzadeh:1992:RTA

[1881] M. Sarrafzadeh and D. T. Lee. Restricted track assignment with applications. Lecture Notes in Computer Science, 650:449–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hsu:1992:STC

[1882] W. L. Hsu. A simple test for the consecutive ones property. Lecture Notes in Computer Science, 650:459–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hakata:1992:LCS

[1883] K. Hakata and H. Imai. The longest common subsequence problem for small alphabet size between many strings. Lecture Notes in Computer Science, 650:469–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1992:IDP

[1884] T. W. Lam and K. H. Lee. The implicit dictionary problem revisited. Lecture Notes in Computer Science, 650: 479–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reinhardt:1992:SPW

[1885] K. Reinhardt. Sorting in-place with a worst case complexity of n log n — 1.3n + O(log n) comparisons and ϵn log n + O(1) transports. Lecture Notes in Computer Science, 650:489– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moffat:1992:SMF

[1886] A. Moffat, O. Petersson, and N. C. Wormald. Sorting and/by merging finger trees. Lecture Notes in Computer Science, 650:499–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koymans:1992:SMP

[1887] Ron Koymans. Specifying message passing and time-critical systems with temporal logic. Lecture Notes in Computer Science, 651:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pnueli:1992:SSR

[1888] A. Pnueli. System specification and refinement in temporal logic. Lecture Notes in Computer Science, 652:1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dam:1992:FPB

[1889] M. Dam. Fixed points of Buchi automata. Lecture Notes in Computer Science, 652:39–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramakrishna:1992:ATD

[1890] Y. S. Ramakrishna, L. K. Dillon, L. E. Moser, P. M. Melliar-Smith, and G. Kutty. An automata-theoretic decision procedure for future interval logic. Lecture Notes in Computer Science, 652:51–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Granger:1992:IRS

[1891] P. Granger. Improving the results of static analyses of programs by local decreasing iterations. Lecture Notes in Computer Science, 652:68–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agarwal:1992:IPL

[1892] P. K. Agarwal and M. van Kreveld. Implicit point location in arrangements of line segments with an application to motion planning. Lecture Notes in Computer Science, 652:80–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jadhav:1992:OAI

[1893] S. Jadhav, A. Mukhopadhyay, and B. Bhattacharya. An optimal algorithm for the intersection radius of a set of convex polygons. Lecture Notes in Computer Science, 652:92–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levcopoulos:1992:CST

[1894] C. Levcopoulos and A. Lingas. Csensitive triangulations approximate the MinMax length triangulation. Lecture Notes in Computer Science, 652: 104–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buhrman:1992:SCA

[1895] H. Buhrman and S. Homer. Superpolynomial circuits, almost sparse oracles and the exponential hierarchy. Lecture Notes in Computer Science, 652:116– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuler:1992:SAC

[1896] R. Schuler and T. Yamakami. Structural average case complexity. Lecture Notes in Computer Science, 652:128– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arvind:1992:BTT

[1897] V. Arvind, J. Kobler, and M. Mundhenk. On bounded truth-table, conjunctive, and randomized reductions to sparse sets. Lecture Notes in Computer Science, 652:140–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ganesan:1992:OWF

[1898] K. Ganesan. One-way functions and isomorphism conjecture. Lecture Notes in Computer Science, 652:152– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agarwala:1992:SLD

[1899] R. Agarwala and D. Fernandez-Baca. Solving the Lagrangian dual when the number of constraints is fixed. Lecture Notes in Computer Science, 652:164– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lakshmanan:1992:SQA

[1900] L. V. S. Lakshmanan and D. A. Nonen. Superfiniteness of query answers in deductive databases: An automatatheoretic approach. Lecture Notes in Computer Science, 652:176–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steinbach:1992:PPP

[1901] J. Steinbach. Proving polynomials positive. Lecture Notes in Computer Science, 652:191–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sundararajan:1992:AIS

[1902] R. Sundararajan and J. S. Conery.

An abstract interpretation scheme for groundness, freeness, and sharing analysis of logic programs. Lecture Notes in Computer Science, 652:203–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monsuez:1992:PTA

[1903] B. Monsuez. Polymorphic typing by abstract interpretation. Lecture Notes in Computer Science, 652:217– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huet:1992:GSL

[1904] G. Huet. The gallina specification language: A case study. Lecture Notes in Computer Science, 652:229–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reif:1992:VLS

[1905] W. Reif. Verification of large software systems. Lecture Notes in Computer Science, 652:241–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garg:1992:DUP

[1906] V. K. Garg and B. Waldecker. Detection of unstable predicates in distributed programs. Lecture Notes in Computer Science, 652:253–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patkar:1992:FSR

[1907] S. Patkar and H. Narayanan. Fast sequential and randomized parallel algorithms for rigidity and approximate min k-cut. Lecture Notes in Computer Science, 652:265–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ravi:1992:ATL

[1908] R. Ravi, B. Raghavachari, and P. Klein. Approximation through local optimality: Designing networks with small degree. Lecture Notes in Computer Science, 652:279–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abramsky:1992:GFC

[1909] S. Abramsky. Games and full completeness for multiplicative linear logic (extended abstract). Lecture Notes in Computer Science, 652:291–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godskesen:1992:RTC

[1910] J. C. Godskesen and K. G. Larsen. Real-time calculi and expansion theorems. Lecture Notes in Computer Science, 652:302–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caucal:1992:BBC

[1911] D. Caucal. Branching bisimulation for context-free processes. Lecture Notes in Computer Science, 652:316–??, 1992.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mukund:1992:CLA

[1912] M. Mukund and M. Nielsen. CCS, locations and asynchronous transition systems. Lecture Notes in Computer Science, 652:328–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christoff:1992:RAS

[1913] L. Christoff and I. Christoff. Reasoning about safety and liveness properties for probabilistic processes. Lecture Notes in Computer Science, 652:342– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muthukrishnan:1992:SMU

[1914] S. Muthukrishnan and H. Ramesh.

String matching under a general matching relation. Lecture Notes in Computer Science, 652:356–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martel:1992:CCW

[1915] C. Martel and R. Subramonian. On the complexity of certified write all algorithms. Lecture Notes in Computer Science, 652:368–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Munro:1992:SRO

[1916] J. I. Munro and V. Raman. Selection from read-only memory and sorting with optimum data movement. Lecture Notes in Computer Science, 652: 380–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ravikumar:1992:SOW

[1917] B. Ravikumar. Some observations on 2-Way probabilistic finite automata. Lecture Notes in Computer Science, 652: 392–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mehlhorn:1992:RDA

[1918] K. Mehlhorn. Recent developments in algorithms for the maximum flow problem (abstract). Lecture Notes in Computer Science, 652:404–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lions:1992:ICW

[1919] J. L. Lions. Inaugural conference: World mathematical year 2,000 and computer sciences. Lecture Notes in Computer Science, 653:3–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rabin:1992:DPC

[1920] M. O. Rabin. Dependable parallel computing by randomization. Lecture Notes in Computer Science, 653:19–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Randell:1992:SD

[1921] B. Randell. System dependability. Lecture Notes in Computer Science, 653: 21–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saltzer:1992:TNL

[1922] J. H. Saltzer. Technology, networks, and the library of the year 2,000. Lecture Notes in Computer Science, 653: 51–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seitz:1992:MCE

[1923] C. L. Seitz. Mosaic C: an experimental fine-grain multicomputer. Lecture Notes in Computer Science, 653:69–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ullman:1992:NFD

[1924] J. D. Ullman. New frontiers in database system research. Lecture Notes in Computer Science, 653:87–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Constable:1992:FTS

[1925] R. L. Constable. Formal theories and software systems: fundamental connections between computer science and logic. Lecture Notes in Computer Science, 653:105–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manna:1992:TC

[1926] Z. Manna and A. Pnueli. Time for concurrency. Lecture Notes in Computer Science, 653:129–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Preparata:1992:HPC

[1927] F. P. Preparata and G. Bilardi. Horizons of parallel computation. Lecture Notes in Computer Science, 653:155– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sethi:1992:CSV

[1928] R. Sethi, R. Campbell, S. Dorward, A. Iyengar, C. Kalmanek, G. Murakami, Shieh Ce-Kuen, and Tan See-Mong. Control software for virtualcircuit switches: call processing. Lecture Notes in Computer Science, 653: 175–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brachman:1992:WKR

[1929] R. J. Brachman. What is knowledge representation, and where is it going? Lecture Notes in Computer Science, 653:189–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carroll:1992:CDS

[1930] J. M. Carroll. Creating a design science of human-computer interaction. Lecture Notes in Computer Science, 653: 205–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brady:1992:SR

[1931] J. M. Brady. Sensing robots. Lecture Notes in Computer Science, 653:219– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koenderink:1992:FBP

[1932] J. J. Koenderink. Fundamentals of bicentric perspective. Lecture Notes in Computer Science, 653:233–??, 1992.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Netravali:1992:DHT

[1933] A. N. Netravali. Digital HDTV: a technical challenge. Lecture Notes in Computer Science, 653:253–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Astrom:1992:AC

[1934] K. J. Astrom. Autonomous control. Lecture Notes in Computer Science, 653:267–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brockett:1992:ADC

[1935] R. W. Brockett. Analog and digital computing. Lecture Notes in Computer Science, 653:279–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fleming:1992:SCL

[1936] W. H. Fleming. Stochastic control and large deviations. Lecture Notes in Computer Science, 653:291–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sussmann:1992:DGM

[1937] H. J. Sussmann. Differential-geometric methods: a powerful set of new tools for optimal control. Lecture Notes in Computer Science, 653:301–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Varaiya:1992:CVA

[1938] P. Varaiya. Coordinating vehicles in an automated highway. Lecture Notes in Computer Science, 653:315–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Willsky:1992:OCS

[1939] A. S. Willsky. Opportunities and challenges in signal processing and analysis. Lecture Notes in Computer Science, 653:331–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wong:1992:NCS

[1940] E. Wong. Neural computing and stochastic optimization. Lecture Notes in Computer Science, 653:339–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brezzi:1992:SGM

[1941] F. Brezzi, C. Baiocchi, and L. D. Marini. Stabilization of Galerkin methods and applications to domain decomposition. Lecture Notes in Computer Science, 653:345–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hughes:1992:EIS

[1942] T. J. R. Hughes and Z. Johan. An efficient implementation of the spectral partitioning algorithm on connection machine systems. Lecture Notes in Computer Science, 653:357–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charot:1992:EHT

[1943] F. Charot, P. Frison, E. Gautrin, D. Lavenier, P. Quinton, and C. Wagner. From equations to hardware: Towards systematic mapping of algorithms onto parallel architectures. Lecture Notes in Computer Science, 654: 1–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fukushima:1992:VPR

[1944] K. Fukushima. Visual pattern recognition with neural networks. Lecture Notes in Computer Science, 654:16–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weng:1992:ORS

[1945] J. Weng, T. S. Huang, and N. Ahuja.

Object recognition by a self-organizing neural network which grows adaptively (abstract). Lecture Notes in Computer Science, 654:32–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1992:IPR

[1946] P. S. P. Wang. Intelligent pattern recognition and applications (summary). Lecture Notes in Computer Science, 654:34–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aizawa:1992:PCG

[1947] K. Aizawa and A. Nakamura. Pathcontrolled graph grammars for syntactic pattern recognition. Lecture Notes in Computer Science, 654:37–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cantoni:1992:DOM

[1948] V. Cantoni, L. Cinque, C. Guerra, S. Levialdi, and L. Lombardi. Describing objects by a multi-resolution syntactic approach. Lecture Notes in Computer Science, 654:54–??, 1992.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaudhary:1992:PMO

[1949] V. Chaudhary, K. Kamath, P. Arunachalam, and J. K. Aggarwal. Parallel manipulations of octrees and quadtrees. Lecture Notes in Computer Science, 654:69–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coutelle:1992:VRO

[1950] C. Coutelle, C. Fortunel, M. Eccher, and B. Zavidovique. Vision reflex operators. Lecture Notes in Computer Science, 654:87–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crochemore:1992:NTD

[1951] M. Crochemore and W. Rytter. Note on two-dimensional pattern matching by optimal parallel algorithms. Lecture Notes in Computer Science, 654: 100–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graham:1992:ELD

[1952] P. Graham, S. S. Iyengar, and S.-Q. Zheng. An efficient line drawing algorithm for parallel machines. Lecture Notes in Computer Science, 654:113– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inoue:1992:CRP

[1953] K. Inoue and I. Takanami. A characterization of recognizable picture languages. Lecture Notes in Computer Science, 654:133–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ito:1992:CLS

[1954] A. Ito, K. Inoue, and I. Takanami.

Constant leaf-size hierarchy of twodimensional alternating Turing machines. Lecture Notes in Computer Science, 654:144–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kakusho:1992:SRE

[1955] K. Kakusho, S. Dan, N. Abe, and T. Kitahashi. Shape recovery and error correction based on hypothetical constraints by parallel network for energy minimization. Lecture Notes in Computer Science, 654:159–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kobara:1992:UGP

[1956] K. Kobara, T. Kanaoka, K. Munechika, Y. Hamamoto, and S. Tomita. Use of gradated patterns in an associative neural memory for invariant pattern recognition. Lecture Notes in Computer Science, 654:179–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laroche:1992:CSP

[1957] P. Laroche, M. Nivat, and A. Saoudi.

Context-sensitivity of puzzle grammars. Lecture Notes in Computer Science, 654:195–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morita:1992:PGP

[1958] K. Morita and S. Ueno. Parallel generation and parsing of array languages using reversible cellular automata. Lecture Notes in Computer Science, 654:213–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saoudi:1992:PRM

[1959] A. Saoudi. Parallel recognition of multidimensional images using regular tree grammars. Lecture Notes in Computer Science, 654:231–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saoudi:1992:OPA

[1960] A. Saoudi and M. Nivat. Optimal parallel algorithms for multidimensional image template matching and pattern matching. Lecture Notes in Computer Science, 654:240–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siromoney:1992:LRP

[1961] R. Siromoney, L. Mathew, K. G. Subramanian, and V. R. Dare. Learning of recognizable picture languages. Lecture Notes in Computer Science, 654: 247–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siromoney:1992:CDS

[1962] R. Siromoney, K. G. Subramanian, and V. R. Dare. Circular DNA and splicing systems. Lecture Notes in Computer Science, 654:260–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takada:1992:LEE

[1963] Y. Takada. Learning even equal matrix languages based on control sets. Lecture Notes in Computer Science, 654: 274–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thang:1992:PDP

[1964] G. V. Thang. Parallel dynamic programming algorithms for image recognition based on its contour. Lecture Notes in Computer Science, 654:290– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1992:NMP

[1965] J.-X. Wang, M. Nakamura, T. Jimbo, and M. Umeno. Neural model for pattern recognition. Lecture Notes in Computer Science, 654:296–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1992:TDS

[1966] P. S. P. Wang. Three-dimensional sequential/parallel universal array grammars and object pattern analysis. Lecture Notes in Computer Science, 654: 305–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).